A Bibliography of Publications in Chance

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

14 August 2024 Version 1.27

Title word cross-reference

#10 [1168]. #11 [1185]. #12 [1200]. #13 [1215]. #14 [1235]. #15 [1251]. #16 [1268]. #17 [1286]. #18 [1304]. #19 [1316]. #20 [1331]. #21 [1348]. #22 [1366]. #23 [1380]. #5 [1086]. #6 [1103]. #7 [1119]. #8 [1135]. #9 [1153].

.406 [137]. 1 [444]. 11 [1238]. 16 [444]. 2 [38]. 4 × 4 [1118]. 5 × 5 [1134]. ^{(}R) [706]. k [1491]. P [1435, 1852, 1823]. R [781]. s [1337]. U [781].

*omics [997].

-After-Touchdown [570, 1590]. -D [38]. -means [1491]. -Seed [444]. -Values [1823, 1435]. -Year [1238].

0 [1247]. 0-300-16969-8 [1247].

1 [1249, 44, 313]. 1-4200-7941-7 [1249]. 1-for-45 [1588]. 10 [1614]. 12 [1465, 1459, 1453]. 13 [999]. 15 [401]. 15th [716]. 19 [1723, 1687, 1677, 1823, 1715, 1712, 1686, 1681]. 1918 [1009]. 1988 [85]. 1990 [4, 82]. 1996 [373].

2 [67]. 2.0 [519, 1330]. 2000 [580, 643, 596, 551, 526]. 2002 [667, 865, 782]. 2004 [947, 948]. 2006 [999]. 2007 [1025]. 2008 [1132]. 2009 [1109]. 2010 [1194, 1150]. 2016 [1596, 1589]. 2023 [1777, 1839]. 21st [1844, 474]. 2370724 [948]. 2nd [1769].

3100 [86].

4.5 [487]. ’41 [137]. 47th [80]. 4th [1345].

587 [693].

60 [58, 60].

7 [1249]. 7.0 [470].

8 [1247]. 88 [593].

9 [1250, 1248]. 978 [1250, 1248]. 978-0-472-05007-9 [1248]. 978-1-58488-711-9 [1250]. 9th [1343].

AAA [1375]. Ability [587]. Able [99]. Abnormal [433]. Absolute [417]. Abundance [1767]. Abuse [765, 302]. Abuses [107]. Academia [231]. Academic [678, 1369]. Academics [1784]. Academiks [236]. Acceptance [550]. Access [907, 1843, 1293, 1226, 98]. Accidents [943]. Accompaniment [793]. According [73]. Account [1655]. Accountable [1716]. Accounting [531, 1064, 1020, 661]. Accuracy [1364, 1026, 1238, 1230]. Accurate [536, 482, 1292]. Accurately [426]. Achievements [1850]. Achilles [1296]. Across [1276, 1463, 644]. Acrostic [40, 861, 938, 1164]. Act [782, 1370]. Actigraphy [1209]. Action [1792, 1842, 1724]. Activities [1529]. ActivStats [519]. Actually [890]. Actuarial [1481]. Adapting [886, 1793]. Adaptive [1646, 1580]. Added [1177]. Address [1530]. Addressing [1078, 1533]. Adjective [20, 21]. Adjusted [356, 51, 84, 94, 74, 95, 134, 140, 169, 279]. Adjusting [1537, 203]. Adjustment [643, 674, 705, 643]. Adjustments [97]. administered [1600]. Administration [1784]. Administrative [308]. Admissions [593, 1620, 708]. Ado [115]. Adolescents [554]. Advance [474, 1341]. Advanced [1364]. Advances [178]. Advancing [1531, 1647]. Advantage [316, 1195]. Adventure [838]. Adverb [20, 21].

Adversarial [1579]. Advisor [23, 110]. Advisory [1500, 1798]. Affair [438, 423]. Affect [1324, 907, 315]. affecting [226]. Affymetrix [900]. Afraid [855]. Africa [1194]. Africans [1497]. After [1370, 570, 1590]. Aftermath [422]. Again [84, 94, 1178, 1295]. against [884, 912, 820]. Age [576, 1544, 81, 1787]. Agencies [246]. Agency [1397]. Ages [903]. Aggregated [1141, 1240]. Agile [733]. Aging [903]. Agreement [390]. Ahoy [497]. AI [1848]. Aid [659]. Aids [269, 105, 759, 579, 36]. AIQ [1627]. Air [81, 769, 698]. Airbags [818, 913, 819, 817, 914, 963]. Album [557]. Alcohol [959]. Alcohol-Related [959]. Alexa [1641]. Algorithm} [1390]. Algorithms [186, 1425, 1692]. Allegation [515]. Allegations [91]. Alleged [609]. Allocation [886]. almost [1165]. Along [1486, 1728]. Alpine [1486]. Already [950, 185]. Also [399]. Alternative [774, 208]. Altitude [1851]. altius [555]. Always [387]. am [734]. Amateur [100]. Ambiguity [1360]. America [1324, 50]. American [1777, 584, 814, 1792, 1773, 1798, 1312, 414]. Americans [814]. Amidst [98]. Amnio [960]. Among [317, 1533, 390, 765, 190]. Amy [1789]. Anaheim [393]. Analyses [799, 1621, 1809, 21]. Analysis [1353, 1162, 1714, 1515, 1694, 1579, 1679, 1235, 1251, 137, 699, 455, 716, 1591, 788, 18, 1640, 1531, 23, 1362, 1057, 1256, 905, 115, 783, 311, 1195, 1339, 1372, 1719, 1762, 754, 1771, 1113, 1231, 1266, 1616, 1660, 1710, 573, 585, 600, 838, 831, 1739, 903, 1082, 692, 1602, 1618]. Analyst [1332, 882, 64]. Analytic [733]. Analytics [1511, 1254, 1795, 1440]. Analytics-Highlights [1511]. Analyze [1404, 172]. Analyzed [1359]. Analyzing [512, 1194, 1780, 901, 1225, 741, 1240, 1800, 480]. Anathem [1234]. Anatomy [666, 1751]. Ancestry [933]. Ancient [1179, 16, 659, 1585]. Anderson [679].

Andrei [918]. Andrew [1738]. Andrus [1397]. Angels [889]. Animal [1485]. Animals [832, 391]. Animate [1326]. Annie [1477]. Announcements [1320, 402]. Announcing [1068]. Annual [405]. Anomalies [452]. Anova [836, 835]. Answer [99]. Answering [192]. Answers [628, 1241, 32, 232]. Antarctic [1556]. Anthrax [1110, 668]. Anti [1521]. Anti-isomorphisms [1521]. Antibiotic [1043]. Anticipated [1289]. Anticipations [16]. Antiquity [824]. Any [1799]. Anyway [1636]. AP [1456, 1755, 1455]. Apartment [609]. Appearance [1822]. Appetite [393]. Apple [418, 269]. Application [716, 1050, 1055, 733, 1020]. Application-Specific [733]. Applications [1163, 1093]. Applied [1515, 156, 392, 1282, 1575, 1560, 1318, 781]. Applies [921]. Applying [1723, 1383, 578]. Approach [1730, 1363, 919, 1651, 1685, 1233]. Approaches [904, 1657, 140, 1003, 1754, 780]. Approval [361]. Approved [669]. Apps [1600]. Arabidopsis [998]. arbitrage [556]. archaeology [578]. Area [474, 517]. Areas [1243]. Arguments [884]. ARIMA [1638]. Aristotle [481]. Arizona [1728]. Arrivals [1814]. Ars [404]. Art [1264, 1548, 1293, 1807, 1471, 1191, 1213, 1226, 1243, 1311, 1447, 1390, 1214, 1192, 1864]. Arthur [637]. Article [1185]. Artifact [820]. Artistic [1850]. Arts [1359]. ASA [1534, 1737, 1511, 1409, 343]. Aside [279]. Asking [834]. Aspect [853]. Aspects [822, 1564, 1755]. Aspirin [24]. Assembled [1490]. Assess [1559, 1621]. Assessing [886, 1683, 1198, 1816, 540, 1272]. Assessment [1723, 1733, 985, 1462, 111, 1053]. Assessments [959]. Assigning [970]. Assistant [1740]. Assisted [679]. Association [1792, 1777, 1773, 1798].

Astronomical [1640]. Astrostatistics [1633, 1022]. Athletes [765, 315]. Athletic [541, 767, 209]. Atlantic [1497]. Atmospheric [1649, 1078]. ATP [1520]. Attack [464]. Attraction [1387]. Attribution [716, 1554, 718]. atypical [1383]. Audits [1144, 1143, 1145]. Augsburg [480]. Author [917, 715, 726]. Authors [815, 828, 844, 862, 878, 895, 909, 923, 925, 940, 955, 968, 980, 995, 1014, 1031, 1047, 1061, 1074, 1089, 1106, 1122, 1138, 1156, 1171, 1188, 1203, 1218]. Authorship [716, 718, 1704]. Automated [1092]. Automating [573]. Automation [4]. Availability [921]. Available [1844, 733, 1517]. Average [137, 257, 808, 1238]. Averages [305]. Averaging [808]. Avoiding [1401]. Award [707, 930]. Awarded [1515]. Away [1636, 1670]. Awful [130].

B [1800, 1547]. Babbage [159]. Babe [1499]. Babies [1810]. Baby [1334, 1477, 241]. Back [1194, 1502, 1330, 84, 94]. Backlash [1388]. Bad [1101]. Badges [1388]. Badly [1281]. Bailey [1847]. Ball [1444, 1402, 339]. Balls [1400]. Bankers [1318]. Banning [1206]. Bar [173, 1821, 1816, 1133, 731, 745, 787, 772, 758, 796, 837, 811]. Barriers [1467]. Base [1769]. Baseball [550, 587, 876, 133, 1656, 257, 264, 501, 54, 63, 192, 467, 1227, 1261, 1313, 306, 369, 934, 1482, 871, 582, 1066, 657, 1495, 382, 976, 932]. Based [1700, 474, 18, 885, 1151, 708, 1274, 1233, 1208]. Basketball [508, 1780, 63, 1038, 705, 1039, 444, 1258, 1382, 29]. BasketballAnalyzeR [1780]. Battles [730]. Battleship [1307]. Battlestar [1083]. Bayes [520, 960, 1449, 1865, 1283, 1788, 1809, 632]. Bayesball [41]. Bayesian [1364, 106, 1477, 1362, 398, 1355, 887, 1266, 1267, 1265, 1660, 1810, 899, 1335, 1176, 625].

Bayesians [117]. BCS [722]. Be [1017, 1215, 1404, 1403, 1361, 1325, 1337, 195, 1197, 1280, 369, 1183, 569, 945, 1471, 426, 46, 1028, 777, 1018, 1037, 669, 99]. Beach [595]. Bear [834]. Bears [1114]. Beat [1307, 444, 1101]. Beating [442]. Beatles [1242, 1606]. Beauty [1626]. Because [542]. Becoming [44, 67]. Bed [186]. Before [1467, 1314, 1460, 1040, 499]. Beginning [877]. Begins [95]. Behavior [425, 1740]. Behind [902, 965]. Being [1361, 1478, 1733]. Belief [1862, 1504]. Believe [61, 1768]. Belizaire [145, 348, 393]. Bell [317, 377]. Bellman [1431]. Belly [1180]. Benchmark [200]. Benefit [1645]. Benford [1760, 1715, 1761]. bent [1078]. bent-cable [1078]. Beowulfs [662]. Berger [1386]. Berk [1597]. Bernoulli [1747]. Bertsch [1241, 1247]. Best [1017, 1148, 759, 1444, 376, 210, 657, 463, 1018]. Best-Ball [1444]. Bestseller [126]. Bet [1257, 1193]. Bets [108]. Better [628, 295, 294, 952, 1195, 761, 1019]. Betting [263, 255]. Bettors [263]. Between [1369, 1373, 1285, 1453, 739, 1037, 947, 948, 561, 98, 231]. Beware [1349]. Beyond [1645, 230, 198]. Bias [1716, 658, 1533, 820]. Biases [424, 371]. Bible [432]. Big [262, 1507, 840, 1293, 1239, 1279, 1294, 1259, 1359, 1322, 1310, 1339, 1372, 1388, 1402, 1433, 1446, 1468, 1494, 1570, 1612, 1625, 1442, 1629, 1766]. Binary [1580]. Binning [874]. Binomial [1852]. Bioinformatics [744]. Biological [912]. Biology [1298, 679]. Biomarker [1647]. Biomarker-Driven [1647]. Biomarkers [1070]. Biomedical [1602, 1507, 1758, 1442]. Biometrics [757]. Biostatistics [886, 1813, 1363, 1114, 1443, 1273, 1274, 1306]. Biosurveillance [912]. Bioterrorism [711]. Bird [621]. Birth [170, 324, 402, 1547, 1663, 610]. Birthday [1036]. Birthdays [1036]. Bits [1514]. BK [688]. BK-Plot [688]. Black [292, 762, 1508, 599, 700, 889]. Blackjack [1197, 1371]. Blame [1244]. Blaming [227]. blanks [1019]. Blind [1478]. Blindside [1223]. Block [327]. Blockbusters [568]. Blood [545, 742, 1197, 185, 1042]. Blue [370]. Blunders [1574]. BMDP [330]. BMI [840]. Board [71, 104, 154, 181, 237, 290, 333, 367, 437, 492, 537, 590, 649, 702, 751, 801, 802, 816, 829, 845, 863, 879, 896, 910, 924, 939, 954, 967, 979, 994, 1013, 1030, 1046, 1060, 1073, 1088, 1105, 1121, 1137, 1155, 1170, 1187, 1202, 1217, 1406]. Boca [1250]. Bodies [193]. Body [158, 453]. Bois [1800, 1547]. Bolts [177]. Bond [1384, 1608]. Book [1611, 716, 1249, 1362, 234, 1264, 1250, 1232, 47, 69, 102, 151, 179, 202, 217, 252, 1231, 1247, 1365, 1504, 1503, 1616, 1614, 1615, 1613, 1709, 1710, 1354, 267, 283, 307, 326, 341, 351, 362, 383, 397, 411, 429, 447, 458, 473, 488, 525, 496, 510, 583, 547, 559, 571, 598, 613, 627, 642, 672, 684, 696, 732, 710, 720, 748, 1248, 789]. Books [1510, 270, 1182, 1309, 1133]. Bookshelf [430, 412, 489, 445, 525, 583, 672, 732]. Boom [241]. Boomsma [142]. Bordeaux [322]. Border [1728]. Borders [1726]. Boston [1399]. Both [201]. Bottlings [831]. Bowl [807]. Bowling [1210]. Box [1508, 599, 1622]. Boxplot [1642]. Boys [636]. Bracket [1611]. Brain [788, 977, 937]. Breaking [1400]. Breast [772, 422]. Brève [1523]. Bricks [265]. Bride [387]. Bridesmaid [387]. Bridge [683, 809]. Brief [1514, 500]. Bring [65]. Bringing [1818, 1652]. British [500, 1225]. Broken [369]. Bronx [265]. Brookline [1097]. Brown [292, 1701]. Browsers [1600]. Budget [998, 1844]. Budgets [876]. Bug [301]. BUGS [1362, 398]. Buhler [1]. Build [1291]. Building [1070, 501, 1758, 733, 761, 1306, 1727, 1598, 1165]. Buildings [265]. Bulldog [1406]. Bullet [787, 883]. Bumps [808]. Bureau [1774, 1841, 858]. Burtin [1043, 1068, 1056]. Bus [328]. Bush [731, 738]. Business [790, 823, 1785, 1440]. Businesses [1730]. Busman [106]. butterflies [1028]. Buy [1370, 1658]. Buzz [632]. byproducts [1166].

C [867]. C. [1855]. Cabernets [87]. CABL [884]. cable [1078]. Calculating [159, 1097, 161]. Calculations [1744]. California [87]. Call [1531, 1671]. called [166]. Calling [160, 1442]. Campaign [1505]. Campanile [404, 446, 485, 476, 462, 520, 505]. Campanologia [404]. Can [1534, 999, 295, 294, 1131, 1687, 759, 1403, 1602, 1832, 1670, 314, 1244, 1280, 1370, 369, 1183, 1816, 542, 1319, 209, 1471, 342, 890, 1830, 426, 1434, 256, 46, 183, 1101, 1341, 177]. Canada [441]. Canadian [1430, 626]. Cancer [740, 893, 375, 965, 1026, 1070, 223, 422, 742, 1758, 224, 625]. Candidates [360]. Canonical [191]. Can’t [921, 1383]. Capacity [1727]. Capital [42]. Capitol [674]. Car [389]. Card [358]. Cardano [1289]. Cards [1222]. Care [141, 759, 907, 1102]. Career [766, 112, 1466, 1273, 1147, 1713, 582, 1341, 1318, 1021]. Careers [1511, 1112]. Careful [1818]. caricatures [639]. Carl [1398]. Carlo [1868, 11, 586, 1249]. Carnegie [1738]. Cars [262, 469]. cars.com [1652]. Cartoon [1420, 1422, 1379, 343]. Casanova [1808]. Case [275, 1477, 258, 531, 32, 811, 1397, 370, 1597, 405, 717, 1288, 694, 637, 1844, 1485, 1147, 1724, 1113, 883, 1240, 469, 1517]. Case-Control [1147]. Casella [1309]. Cases [772, 211]. Casino [1257, 823]. Cat [762, 1117]. Catalyst [1457]. Catastrophes [254]. Catch [483, 970]. Catch-22 [970]. Catching [225]. Categorical [1353]. Cauchy [986]. Causal [1854, 1610, 607, 1426, 1451]. Cause [375, 1319, 1532, 1354]. Cautionary [337, 317]. CD [1250]. CD-ROM [1250].

Celebrating [1415, 50, 1839]. Celebration [1340, 1056]. Cell [1206, 832]. Census [356, 212, 308, 643, 1774, 230, 51, 84, 94, 74, 95, 134, 140, 169, 279, 1844, 1841, 1150, 536, 1311, 858, 580, 4, 387, 464, 643, 674, 82, 596, 551, 175, 248, 730, 526]. Census-Taking [212]. Censuses [1352]. Centenary [1056]. Centennial [946]. Center [1649, 1273, 1758, 625]. Central [1657, 985]. centric [1468]. centuries [1865]. Century [1844, 474]. CEO [1796]. Certainty [933]. cetinkaya [1536]. cetinkaya-Rundel [1536]. Chain [1249, 1868, 586]. Chains [38]. Challenge [419, 1544, 1564, 666]. Challenger [1128, 1095]. Challenges [758, 1057, 1465, 1259, 1530]. Challenging [1125, 1528, 1656]. Champion [1666, 1403]. Championships [891]. Chance [135, 189, 205, 221, 239, 261, 274, 299, 321, 495, 509, 524, 539, 553, 566, 577, 592, 604, 620, 635, 651, 664, 677, 690, 703, 714, 727, 752, 764, 778, 791, 1148, 1404, 1260, 1323, 1340, 944, 1490, 1610, 1448, 1623, 467, 1559, 1280, 1469, 1459, 1373, 118, 1387, 616, 1736, 984, 1419, 1023, 1603, 245, 1492, 1360, 1240, 1295, 1312, 1536, 1164, 973, 481, 1357, 1619, 975, 166, 1807, 34, 56, 88, 153, 165, 235, 288, 331, 430, 435, 434, 445, 491, 535, 1043, 731, 745, 787, 772, 758, 796]. Chance [837, 811, 1614, 972, 858, 906]. Chances [28, 1768]. Change [1733, 1558, 340, 1796, 1794, 1555, 1554, 9, 1797, 6, 1798, 1557, 1457]. Changed [1329, 1028]. Changepoint [1803]. Changes [1818, 1276, 1406, 903]. Changing [1380, 1793, 1690]. Chaos [792, 158]. Chapman [1249, 1250]. Chaps [719]. Characteristics [1197, 324]. Charles [159, 471, 713]. Chart [173, 946, 1181, 641]. Charybdis [561]. Chatting [83]. Cheating [91, 168, 1281]. Checkerboards [1490]. Checking [1371, 497]. Cherry [718]. Chess [1679]. Chicago [349, 348]. Chicken [1822]. Chief [1774]. Child [1277, 1184].

Childhood [1547, 1020]. Children [1521, 759, 284]. Chips [497]. chlorofluorocarbons [1078]. Choice [168, 350]. Cholesterol [58]. Choosing [1645]. Chronographic [1825]. Cigarette [1278]. Cinema [1384, 1608]. Circles [1130]. Circular [1436]. Circumstances [1714]. Cities [1732]. Citius [555]. City [187, 889, 480]. Civil [717]. Claim [831]. Claims [875, 1371]. Clash [174]. Class [1652, 820]. Classical [1521, 1289, 1130]. Classification [31, 1610, 1093, 1102]. Classroom [1404, 1260, 1323, 1340, 1448, 1623, 1559, 729, 1469, 1459, 1373, 1387, 1419, 1603, 1492, 1360, 1240, 1295, 1312, 1536, 1686, 1755]. Cleaning [1654]. Clear [688, 663]. Clearly [1832]. Clemens [1065, 1021]. Clients [1369]. Climate [1558, 1550, 1552, 1793, 1796, 1794, 1555, 1554, 1553, 1797, 1798, 1795, 1557]. Climatology [1803]. Clinical [1502, 796, 421, 860, 875, 1085, 1646, 1648, 1098, 711, 625, 1804, 1680]. Closer [1818, 1334]. Clouds [201]. CLT [722]. Cluster [662, 1143]. Clustering [1640, 1491, 905]. Clusters [1637, 1639]. Co [1733]. Co-worker [1733]. Coach [612, 1655]. Coaching [1513, 654, 821, 1254, 655]. Cocktail [1330]. Code [1571, 1594, 1865]. Codes [454]. Coefficients [1149]. Cogitation [1348]. Cognitive [118, 62]. Coin [1261]. Coincidence [1164]. Colandr [1729]. Cold [798, 29]. Cold-Foot [798]. Collaboration [1463]. collaborative [997, 1410]. Collecting [1125]. collection [1125]. College [927, 508, 1467, 722, 766, 513, 1460, 1464, 1080, 1207, 1039, 313, 444, 618, 1583, 942, 1443, 1406]. Colleges [771]. Cøllisiøns [1206]. Colon [893]. Colonoscopy [893]. Colorectal [893]. Column [846, 864, 1482, 1481, 346]. Combinations [1490]. Combinatoria [404].

Combinatorial [564]. Combine [300, 1570]. Come [1408]. Comedies [1646]. Comes [1599, 659]. Comic [1079]. Coming [349, 348, 236]. Comment [867, 868, 681, 888, 884, 638, 836, 848, 930, 622, 639, 850, 849, 541, 851, 594, 542, 606, 607, 916, 948, 852, 21, 839]. Commentary [212, 230, 914, 963, 1008, 1132]. Comments [77, 1118, 1134]. Commerce [134]. Commercial [594]. Commissar [1391]. Committee [1798, 1534]. Committees [860, 1500]. Communicate [919]. Communicating [7]. Communication [825, 1532, 1748]. Communications [1374]. Communities [1822, 1490]. Community [1723, 1775, 1312, 1463]. companies [1594]. Company [1502]. Comparative [904, 1354]. Compared [511]. Comparing [269]. Comparison [1600, 900, 522, 1130]. Comparisons [1358, 1336]. Compass [79]. Compensation [1003, 1004]. Compete [1291]. Competent [44, 67]. Competing [843, 1258]. Competitions [546, 172]. Competitive [1096]. Complementary [774]. Complete [1503, 1866]. Complex [609, 892, 1019]. Complexity [825]. Complicated [1052]. Component [1183]. Components [533]. Composition [461]. Compression [278]. Computation} [1392]. Computational [250, 794, 1265, 1685, 1616]. Computer [1626, 1351, 989]. Computerized [711]. Computers [132]. Computing [1363, 177, 200, 215, 233, 250, 270, 286, 310, 416, 1322, 504, 519, 533, 549, 562, 574, 586, 601, 616, 630, 645, 329, 347, 359, 386, 399, 418, 433, 451, 662, 673, 685, 698, 711, 721, 733, 761, 776, 1462]. Concave [876]. concentration [1078]. Conception [1117]. Concepts [630]. Concern [1078]. Concerned [1740]. Concomitant [335]. Concrete [1404]. Concussions [1651]. Conditional [1183]. Conduct [1277, 1369]. Conference [31]. Confidence [1852]. Confidential [779, 1548, 782].

Confidentiality [1293, 784, 1213, 785, 781]. conflict [1103]. Confronting [708]. Congress [749]. Congressional [999, 354]. Conjecture [10]. Connecting [1860]. Conquer [1570]. Conquers [1038]. Consequences [1620]. Conserving [1335]. Consider [1466, 1054]. Considering [952, 1471]. Consistency [266]. Consolation [873]. Constitution [124]. Consulting [1005, 1163, 1369, 1004, 1560, 243, 1003]. Consumer [686, 952]. Consumerism [729]. consumption [1129]. Contact [1681]. Contacts [993, 1012, 1029, 1045, 1059, 1072, 1087, 1104, 1120, 1136, 1154, 1169, 1186, 1201, 1216, 1236]. Contemporary [1324]. Content [606]. Contest [1043, 1832, 1112, 1128, 1068]. Contests [1822]. Context [1829, 1471]. Contrarian [987]. Contrary [1753]. Contribute [327]. Contribution [1149]. Contributions [1353, 1338, 1416, 1352]. Control [129, 49, 1259, 926, 1147, 339, 248]. Controlling [1007, 1008]. Controversies [580]. Controversy [138, 1865]. Conversation [1467, 1701, 1774, 978, 1777, 858, 906, 1831]. Conversions [1181]. Convex [876]. Coping [254]. Copying [515]. Corner [861, 813, 827, 842, 877, 894, 908, 922, 938, 953, 966, 992, 1011, 1010]. Coronary [1151]. Coronavirus [1683]. Correction [384, 773, 1016]. Corrections [396]. Correlations [1834, 1285]. Correspondence [1162, 1714, 1771]. Cosmology [1634]. Cost [297, 1144, 335, 1660, 1801]. Cost-Effective [335]. Cost-Effectiveness [297, 1660]. Costs [248, 1867]. Could [669]. Count [779, 1569]. Counterterrorism [756, 753]. Counting [1544, 1197, 1001, 606, 603, 193, 469]. Countries [1723]. Counts [621, 799]. County [1760, 916]. County-Level [1760].

Couriers [225]. Course

[1696, 1323, 678, 729, 1858, 166, 618, 1799]. Courses [1448, 1560, 1731]. Coursework [1813]. Court [139, 84, 94, 1493, 887, 174, 561, 615, 356, 1430]. Courtroom [1479]. Courts [25, 1565]. covariate [1019]. Coverage [1150]. COVID [1723, 1687, 1677, 1823, 1715, 1732, 1712, 1686, 1761, 1681]. COVID-19 [1723, 1687, 1677, 1823, 1715, 1712, 1686, 1681]. Cox [96, 1515]. CPS [353]. cracked [1865]. Crash [1705]. Crashes [870, 1717, 1146, 807]. Craze [563]. CRC [1249, 1250]. Creating [1844]. Creation [368]. Creativity [1007, 1008]. Creators [1502]. Credibility [614]. Credit [865]. Crichton [245]. Crime [173, 661, 1631]. Crimes [1720, 692]. Criminal [511, 1481]. criticism [638]. Criticisms [1342]. Criticize [1450]. Critics [831]. Crockett [1667]. Crop [528]. Crossing [694]. Crossword [1664, 1659]. Crowdedness [986]. Crowdfunding [1700]. Crowdsourcing [1625]. Crown [575]. Crucial [1374]. Crunching [201, 364, 346, 385, 398, 416, 460, 470, 487, 287, 271, 330, 311, 216, 251, 178, 68, 45, 101, 150, 346]. Cry [1177]. Cryptic [303, 861, 813, 827, 842, 894, 908, 922, 938, 953, 992, 1011]. Cucumbers [1778]. Cult [1867, 1248]. Culture [1271, 812, 174]. Cumulative [153, 235, 288, 331, 435, 491, 535]. Cup} [1385, 1194, 1025, 440, 918]. Curated [1844]. Curbstoning [797]. Cure [942, 987]. Curious [668]. Current [1358, 246]. Currents [1753]. Curriculum [1456, 1758, 1463, 1453, 1583, 1457]. CURV [1860]. Curve [317, 392, 1526, 377]. Curveball [1400]. Cut [44, 67]. CV [1388]. Cybersecurity [754, 1816]. Cyberspace [388]. Cycle [1499]. Cylon [1083]. Cynthia [1666].

D [38]. Daddy [1334]. Dallas [459].

Damned [1741, 33]. Damsel [1084]. Dangerous [1682]. Dangers [1525]. Dare [1535]. Daring [380]. Darts [408]. Darwin [471]. Data [1353, 173, 1043, 1125, 1725, 1722, 1694, 740, 1817, 1404, 212, 1674, 1340, 1654, 427, 686, 724, 1640, 1757, 1531, 1702, 1209, 998, 23, 1543, 19, 1805, 1832, 20, 1458, 779, 1780, 1507, 1228, 1342, 1315, 1670, 460, 860, 1026, 1256, 422, 1563, 1610, 514, 1448, 1765, 303, 1693, 783, 1734, 1621, 1548, 1844, 1293, 1279, 1259, 1322, 1310, 1339, 1372, 1388, 1433, 1468, 1570, 1612, 1625, 1719, 1254, 1689, 1715, 754, 1652, 1785, 1816, 645, 1718, 1691, 1716, 1651, 594, 1442, 386, 780, 1113, 1266, 1377, 1572, 1661, 1471, 711, 1582, 1360, 948, 1533, 1462, 1226, 1351]. Data [1690, 883, 1240, 1686, 1800, 480, 1434, 256, 46, 1727, 1761, 573, 585, 600, 838, 958, 1390, 1483, 98, 1530, 1748, 1707, 692, 781, 1394, 68, 865, 1383, 1636, 1710, 1843]. Data-centric [1468]. Data-Driven [212, 1651]. Database [685]. Databases [1026]. Dating [1387, 1027]. David [96, 1515, 867]. Davy [1667]. Day [1714, 24, 610]. Days [1523]. Dead [413]. Deadlines [1319]. Deal [840]. Death [206, 1687, 1324, 837, 945, 700, 1451, 641]. Debate [643, 14, 1288]. Debia$ed [1141]. Decades [1528]. Deceiving [1669]. Decennial [387, 82]. Deception [1483]. Decide [313, 1594]. Decides [356]. Deciding [293]. Decision [960, 1430, 64, 1809, 570, 1590, 1274]. Decision-Theoretic [1274]. Decisionmaking [962]. Decisions [643, 271, 350, 1833, 976, 1335, 561, 1054, 1445]. Decomposition [573]. DECstation [86]. Deepwater [1491]. Defeating [1483]. Defeats [394]. Defects [170]. Defense [652, 891, 799, 1025, 1577, 405, 1691]. Defensive [538]. Definite [1185]. Defuse [1320]. Defying [1593]. DeGroot [72]. Delays [448]. Delbrück [1126]. Delegates [970]. Deliberate [1274]. Delight [393]. della [156]. Deming [1805]. Democrats [926]. Demographer [1774]. Demonstrating [1620]. Dendrology [178]. Denny’s [1492]. Deploying [1729]. depth [1777]. Derby [1588]. Derivation [934]. Describes [159]. Describing [950]. Design [406, 921, 1057, 59, 1024, 625, 499, 1374, 1864, 1264]. Designing [513, 624, 335]. Desk [68]. Desperately [770]. Detect [899, 1281]. Detecting [1486, 1578]. Detection [1554, 830]. Deter [767]. Deterioration [265]. Determine [1035]. Determining [766, 686]. Developing [579, 1116]. Development [1644, 1529, 1443, 1459, 1727, 1547]. Developments [1581, 1648]. Deviation [417]. Devised [1277]. Devises [193]. Devolution [419]. Diaconis [1290]. Diagnosis [1055]. Dice [1092]. Dickens [1856]. Did [505, 1650, 1641, 926, 1314]. Dietary [875]. Difference [262, 282, 831]. Differences [661, 866, 1127, 988, 540, 1181, 989]. Different [1210]. Differential [1347, 1790, 1336, 469, 248]. Difficult [203]. Difficulty [514]. Diffusion [736]. Digit [432]. Digital [132]. Digitized [1359]. Dilemma [1697, 1754]. DiMaggio [137, 1067, 1178]. Dimensional [741, 30]. Dining [42, 459]. Dinner [1822]. Dionne [1777]. Dipsea [1787]. Directive [401]. Director [858]. Directors [643]. Dirt [545]. Dirty [1471]. Disability [1052]. Disabled [561]. Disagreement [390]. Disagreements [1437]. Disaster [340]. Discernible [415]. Disclosure [781, 1471, 1243]. Discoveries [1854, 679, 1442]. Discovery [1792, 1842, 644]. Discrepancies [947, 948]. Discrimination [1528, 758, 882]. Discuss [855]. Discussing [1321]. Discussion [818, 1545, 819, 738, 1409]. Disease [1578, 1151]. disenrollment [1102]. Disparities [1049, 890]. Dispensable [144, 143]. Dispersion [1657]. Display [1043]. Displays [1257, 148, 199, 958, 1007, 1132, 1008]. Disputing [415]. Dissemination [1693, 780]. Distance [1238]. Distances [1636]. Distinctive [1455]. Distinguish [1131]. Distinguishing [1572]. Distress [1084]. Distribution [1180, 809, 1036, 668, 1126, 1115]. Distributions [1757, 1250, 1863]. District [999]. District-13 [999]. Diversity [427, 1702, 291, 1804]. Divide [1570, 1100]. Divining [282, 1739]. Division [1851]. DNA [92, 275, 93, 1476, 885, 90, 933]. Do [1520, 295, 294, 617, 263, 327, 843, 1435, 300, 554, 1755, 52, 807, 897, 431, 240, 1314, 248]. Doctor [929]. Doctoral [1758]. Document [1480]. DoD [1580]. Does [92, 506, 891, 1588, 375, 1263, 371, 907, 223, 1195, 110, 695, 767, 636, 1096, 768, 1206]. Doesn’t [929]. Dog [1131]. Doing [771]. Don [1371]. Done [1503, 1178, 1866]. Don’t [1215, 614, 1005, 1246, 1166, 1756, 1520]. Doping [1397]. Doubles [853]. Dough [534]. Down [22, 1569, 1054, 1865, 399]. Drab [1672]. Draft [617, 766, 1591, 1509]. Drafting [766]. Drawing [574]. Drawn [945]. Dread [1377]. Drive [534]. Driven [212, 1647, 1829, 1651]. Drivers [567]. Driving [1238, 889, 389, 1206]. Drop [1603]. Drug [1644, 1430, 190, 1648, 767, 228, 225, 892]. Drug-Testing [767]. Drugs [1116, 854]. Drunk [389]. Dubourg [486]. Due [1565]. dummies [1861, 1365]. During [1009, 806, 1754, 1712, 687, 943]. Duty [1714]. Dynamic [1161, 30]. Dynamics [1700, 790].

E-Books [270]. E-Journals [270]. Each [1332, 1183]. Eagles [197]. ear [1153]. Early [1704, 1133, 499, 1126]. Earned [257]. Earth [1320, 1551, 1609, 124]. Easier [928]. East [1125]. Easy [1450, 603, 1445]. EasyPlot [150]. Ecological [1486, 1490, 211, 1488, 1557]. Econometrics [1327]. Economic [1730, 8, 1712, 78]. Economic-Impact [1730]. Ecosystem [1487]. Ecotones [1486]. Ed [1806]. Eddy [975, 1357]. Edgar [679]. Edge [177, 200, 215, 233, 250, 270, 286, 310, 504, 519, 533, 549, 562, 574, 586, 601, 616, 630, 645, 329, 347, 359, 386, 399, 418, 433, 451, 662, 673, 685, 698, 711, 721, 733, 761, 776, 1354]. Edgeworth [471]. Editing [347]. Edition [1343, 1345, 1709, 1710, 1344, 1249, 1298, 1250, 1247, 1769, 1248]. editor [1175, 163, 344, 477, 490, 493, 508, 494, 523, 552, 619, 634, 653, 665, 725, 911, 1017, 1075, 1090, 1107, 1123, 1157, 1172, 1189, 1204, 1219, 1174, 1063, 1205, 1173, 1190, 1220, 1237, 1253, 1269, 1287, 1305, 1317, 1333, 1350, 220, 273, 424, 1367, 1381, 1395, 1407, 1427, 1439, 1452, 1461, 1460, 1473, 1484, 1496, 1506, 1516, 1527, 1538, 1549, 1561, 1576, 1586, 1595, 1604, 1617, 1632, 1643, 346, 164, 298, 1438, 1252, 260, 457, 188, 238, 981, 996, 1015, 1032, 1010, 1048, 1062, 1076, 1091, 1108, 1124, 1140, 1158, 846, 864, 941, 956, 969, 689, 957, 415]. Editor [926, 538, 1482, 591, 1653, 1665, 1676, 1688, 1699, 1711, 1721, 1735, 1749, 1759, 1772, 1781, 803, 1077, 880, 219, 320, 738, 403, 650, 737, 1481, 633, 334, 479, 565]. Editorial [1550, 991, 890, 970, 71, 104, 154, 181, 237, 290, 333, 367, 437, 492, 537, 590, 649, 702, 751, 801, 802, 816, 829, 845, 863, 879, 896, 910, 924, 939, 954, 967, 979, 994, 1013, 1030, 1046, 1060, 1073, 1088, 1105, 1121, 1137, 1155, 1170, 1187, 1202, 1217]. Editors [35, 57, 89, 204, 804, 136, 1791, 1802, 1815, 1826, 1838, 1849]. Education [1467, 1454, 297, 1275, 811, 1142, 594, 1471, 1274]. Educational [1698]. Educators [1465]. EEG [1093]. Effect [798, 593, 1430, 394, 558, 1066, 1354, 943, 874]. Effective [335, 1532]. Effectiveness [297, 821, 1660]. Effects [258, 1794, 527, 594, 805, 607, 871, 256, 1787, 1631]. Efficiency [484, 782]. Efficient [1619]. Efficiently [183]. effort [755]. EIS [99]. Election [671, 1000, 1144, 182, 1139, 1724, 1145, 1596, 971, 947, 1109, 1146, 948, 85, 797]. Elections [1760, 1708]. Electoral [1115, 942]. Electric [60]. Electrical [58]. electricity [1129]. Electronic [23, 912, 658]. Elegance [148]. Elementary [1597]. Elevation [871]. Elite [541]. Elite-Level [541]. Elo [1679]. Embracing [1558, 1360]. emerged [1865]. Emerging [1680]. Emery [1701]. Emma [1825]. Empathy [1374]. Emphasis [1271]. Emphasize [1755]. Empire [516]. Empirical [1283, 820]. Employment [790, 97, 52, 1631]. Employs [1135]. Enabled [1277]. End [81, 877, 1509, 1040]. Endangered [1335]. Endeavors [1738]. Energy [1161]. Enforcement [469]. Engage [1822, 972]. Engine [453, 585, 1796]. Engineering [1515, 1113]. Engineers [1343]. Engines [159, 600]. English [1249, 1250, 1247, 1248]. Enhance [213]. Enhancing [272, 1804]. Enigma [1865]. Enmity [1033]. Enough [1168, 1522]. Enrich [213]. Entanglement [1783]. Entangling [1099]. Enterprise [1844]. Environment [1161]. Environmental [125, 1486, 531, 335, 1739]. Environments [416]. Epi [287]. Epic [1596, 1303]. Epidemic [105, 1677, 1255, 36]. Epidemics [806]. Epidemiologic [772, 1717]. Epidemiology [1299, 111]. EPL [1593]. EPO [767]. Eponymy [1314]. Equal [1764]. Equate [1588]. Equations [1347, 1329]. Equipment [695]. Equitable [597]. Era [419, 1805, 1293]. Errata [478]. Error [230, 377, 1745, 1867]. Errors [1574, 1020, 1064]. Escape [1718]. Escaped [945]. Escaping [1483]. Esséen [885]. Essentials [1684]. Establishing [1271]. Estimands [1645]. Estimate [1687, 1545, 1542, 1830, 1540]. Estimates [474, 843, 1358]. Estimating [258, 1543, 265, 1636, 1541, 1182, 1210, 558]. Estimation [1828, 1327, 1283, 1767, 1739, 1544]. Ethical [1848, 1740, 1833]. Ethicist [1361]. Ethics [1229, 1228, 1263, 1297, 1246, 1278, 1361, 1325, 1342, 1315, 1375, 1391, 1437, 1445, 1522, 1098, 1799, 1445]. Ethnic [259, 820]. Ethnicity [353, 427]. Euler [499]. Eulerian [499]. Europe [644]. European [1311, 1547]. Evaluate [594, 1177]. Evaluating [390, 1227]. Evaluation [799, 109]. Evaluations [1323, 678, 1150]. Evans [869]. Even [1412]. Event [1183]. Events [1555, 405, 254, 767, 1752]. Ever [112, 369, 657, 444, 1028]. Everglades [728]. Every [99]. Everyday [27, 114]. Everything [841]. Everywhere [1310]. Evidence [91, 593, 1324, 1862, 787, 772, 811, 1437, 1683, 1151, 168, 1166, 1479, 1724, 1405, 1504, 1628, 224, 883, 820, 1505, 989]. Evidence-Based [1151]. Evolution [832, 1024, 855, 526, 1126]. Evolutionary [679]. Evolving [133]. Exam [91, 1455, 1272]. Examination [1162, 515, 1480, 1475]. Examining [1371]. Example [892, 936, 1374, 461, 1598]. Examples [1654, 1378, 1726, 799]. exceptional [1053]. Excerpt [1611]. Excitements [1818]. Executive [493, 988, 650, 633]. Exhalation [1859]. Exhibition [1069]. Exist [506, 1096]. Exit [771, 947, 948]. Exoplanets [1638]. Expanding [173]. Expect [82]. Expectation [1081]. Expectations [1856]. Expected [1658]. Expedition [1389]. Experience [27, 1542]. Experiences [988]. Experiment [902, 1718, 1598, 1845]. Experimental [406, 109, 499]. Experimentation [978, 982, 605, 606, 607]. Experiments [250, 11, 1582]. Expert [25, 1208, 244]. Explaining [947, 948]. Exploitation [1540, 1544]. Exploration [1792, 1842, 1480, 1159]. Exploratory [250, 1256, 1734, 573, 585, 600, 838]. Explore [23]. Exploring [1260, 1552, 20, 1848, 742, 1387, 1761].

Expository [919]. Exposure [870, 223]. Expression [905]. Extra [768]. Extract [1602]. Extraordinary [1371, 1599]. Extraterrestrials [83]. Extreme [624, 1555, 1639]. Eye [194]. Eyeballing [1196].

F [975, 1357]. Fab [1672]. Fable [318]. Face [1642, 1622]. Fact [1371, 1572]. Fact-Checking [1371]. Factors [226, 986, 1809]. Facts [29]. Fail [1596, 1565]. Failings [1041]. faint [1125]. Fair [1211, 130, 1565, 108, 1787, 1619]. Fairness [886, 680, 242, 681]. Faking [514]. Fallacies [920]. Fallacy [1751, 1747]. False [857, 1605, 1717, 990]. False-Positive [1717]. Falsified [422]. Fame [932, 550]. Familiar [734]. Familiarity [1705]. Families [284]. Family [1589, 1255, 636]. Famous [1384, 1608]. Far [1636]. Farewell [1380]. Farm [391]. Farming [1582]. Fascinating [1243]. Fast [1793, 1673]. Fast-Changing [1793]. Faster [567]. Fasting [943]. Fatal [1319, 807]. Fatalities [866]. Fault [227]. Faulting [184]. Favorite [1225]. Favorites [1385]. FDA [1517, 1500]. Feast [196]. Features [1455]. Feb [1248]. Federal [400, 327, 340, 474, 502, 560, 614, 697, 735, 1776, 246, 1352, 65, 43, 98]. Feel [617]. Fellow [380]. Female [1764, 1851]. Ferguson [486]. Fermat [1100]. Ferocious [631]. Fibonacci [157]. Fiction [1572, 245, 859]. Field [521, 1531, 1470, 210, 626]. Fields [58, 60]. Fienberg [1354, 1353, 1356, 1519, 975, 1357]. FIFA [1194]. Fighting [1545]. Figure [1119, 667, 281]. Filling [1019, 557]. Final [1022, 255]. Finally [1385]. Finance [811, 1099, 1318]. Find [1730]. Finding [740, 873, 874]. Findings [857, 1605, 1559, 919, 1704]. Finds [1638]. Finger [1027]. Fingerprinting [92, 93, 90]. Fingerprints [1475]. Finishing [1589]. Finland [1568]. Finney [914, 963, 913].

Firearms [1563]. Fired [777]. Firing [558]. Firings [1113]. Firm [886]. First [73, 826, 1591, 814, 357, 316, 873, 1199, 1248, 1249, 1250, 1247]. First-Round [873]. First-Year [1591]. Fish [160]. Fisher [75, 77, 1671, 76, 1034, 499]. Fit [1263]. Fitting [417, 1250, 1863]. Five [1404, 502, 1545, 155, 569, 1391]. Five-Year [1391]. Fix [1750]. Fixed [1144]. Fixed-Cost [1144]. Fixed-Risk [1144]. Flagstick [1697]. Flawed [1179, 139]. flaws [1594]. Fletcher [1245]. Flight [693]. Flipbooks [1698]. Florence [1814]. Florida [595, 999, 803, 1145]. Flow [1472]. Flu [1677]. Fly [197]. Focused [1731]. Follow [907, 865]. Follow-up [907, 865]. following [807]. Follows [1390]. Food [1131]. Foolish [531]. Foot [798]. Football [927, 722, 513, 1080, 266, 313, 468, 1181, 626]. Forbes [406]. Force [1737]. Forecasting [306]. Forefather [156]. Foreign [865]. Forensic [884, 787, 881, 1480, 1475, 1479, 1474, 1809, 883, 1739]. Forensics [1428]. Forget [1005]. Formal [1695]. Former [1774, 858, 692]. Formula [934]. Formulas [735]. fortius [555]. Fortunes [833]. Forum [1792, 1842, 1797, 1839]. Forum-Statistics [1797]. Fostering [785, 231]. Found [160]. Foundations [1425, 1233]. Four [808, 357, 1827, 1765, 554, 1003]. Fourth [1450, 1344, 1054]. fourth-down [1054]. Fractals [792, 616]. Framework [1732]. Francis [1250, 471, 644, 859]. Francisco [196]. Franklin [916]. Fraud [421, 1109]. Free [406, 1758]. Free-Standing [1758]. Freedle [820]. Freely [733]. Frequency [658]. Fresh [882, 1811]. Friend [1471]. Frontier [739, 1022, 872]. Frontiersman [1667]. Frustrated [1001]. Full [140]. Full-scale [140]. Fun [342, 247]. Functional [788, 1209, 1294, 1372]. Fund [1335]. Funded [1534]. Funding [735]. Further [884, 989]. Fuss [278]. Future [1819, 1558, 8, 1796, 1502, 1009, 1556, 1456, 5, 9, 7, 1475, 6, 780, 1409]. Fuzzy [323].

G [171]. Gain [1692]. GAISE [1583]. Galactica [1083]. Galaxies [1637]. Galaxy [1636]. Galton [471, 1176, 644, 859, 951]. Gambler [1371, 1751, 222]. Gamblers [1257]. Gambling [1081, 824, 1038]. Game [734, 876, 167, 357, 501, 930, 1280, 1402, 658, 1039, 1037, 1040]. Game-Theoretic [930]. Games [1818, 950, 1337, 306, 1736, 1707]. Gaming [823, 1390]. Gap [1418]. Garbage [256]. Garbology [507]. Garden [1733]. Gas [1341]. Gastronomic [145, 146]. Gastronomie [80]. Geary [867]. Geert [1518]. Gel [741]. Gelman [1252]. Gender [1418, 1127, 1564, 988, 1544, 989]. Gene [905]. Genealogy [471]. General [511, 1227, 110]. Generated [38]. Generation [1689, 241, 1419]. Generators [215]. Genes [317, 965]. Genetic [904, 1374, 903]. Genetics [1494]. Genius [1373]. Genocide [1567]. Genomics [739, 904, 1447, 1820]. Genomish [897]. Gentleman [3]. Geometry [521, 659, 338]. geomnet [1497]. Geophysical [121]. George [1309]. German [1589]. Germany [945]. Gerrymandering [1750]. Get [1670, 1610, 300, 116]. Gets [1512, 858, 1594]. Getson [330]. Getting [1776, 55, 1406]. Gibbs [398]. Girls [636]. GIS [733]. Give [610]. Giving [1027]. Glimpse [780]. Global [1543, 872, 1648, 1540, 1125, 1078]. Globally [1830]. Globular [1637]. Glossary [898]. Go [1054, 1102, 1493, 1554, 1517, 558, 468, 357, 316]. Goal [521, 1450, 210, 626]. Goalie [440, 612, 1655]. Goes [347, 399, 1594]. Going [1006, 567]. Gold [667, 1602]. Golden [81]. Golf [1697, 1592, 1444, 597, 1195, 108, 1587, 1037]. Good [1722, 1257, 455, 1323, 1179, 1025, 200, 130, 1734, 210, 1196, 1731, 1726, 2, 1101]. Goodness [1028, 1044, 1058, 1071, 1086, 1103, 1119, 1135, 1153, 1168, 1185, 1200, 1215, 1235, 1251, 1268, 1286, 1304, 1316, 1331, 1348, 1366, 1380]. Google [1260]. Gore [731, 738]. Governed [1148]. Government [8, 28, 1773, 231]. governments [1594]. Governor [1164]. Govys [1775]. GPS [1277]. GPS-Enabled [1277]. Grabovsky [1526]. Grace [148]. Gradebook [1389]. Grades [678, 729]. Gradients [1486]. Grading [603]. Graduate [1256, 1620, 1411]. Grail [1703]. Grant [300]. Granted [614]. Grants [327]. Graph [1017, 1133, 336, 1101, 1018]. Graphic [1836, 1043, 1423, 1568, 629, 644, 1007, 1132, 1748, 1068, 1008]. Graphical [1389, 679, 359, 148, 232, 402, 486, 532, 713, 1663, 1159, 1547]. Graphics [1128, 2, 1547, 30, 216]. Graphs [311, 285, 892]. Grapple [1690]. Grating [323]. Gray [292, 1481]. Great [712, 1221, 369, 323, 1599, 1326, 1614, 319, 1009, 113]. Greater [1673]. Greatest [378]. Greek [824]. Gretzky [113]. Grosses [568]. Group [97, 1250, 945, 1083]. Grow [241]. Growing [1243]. Grows [881]. Guarantee [497]. Guatemalan [1208]. guesses [1019]. Guesstimation [1330]. Guest [212, 230, 890]. Guidance [1341]. Guide [1420, 1423, 1421, 280, 1455, 180, 1866, 1503]. Guidelines [243, 1583]. Guilty [990]. Gummy [1114]. Gun [405, 538]. Guns [1324].

H [72]. Haalvelmo [78]. Hal [1784]. Hall [550, 1249, 1250, 119, 1782, 932]. Hall/CRC [1249]. Hand [1148, 1507, 223, 61, 224, 1206, 29, 62, 1441]. Hand-Held [1206]. Handbook [1868, 1863, 1249, 1250]. Handgun [745]. Handicap [597, 1195]. Handicapping [1444]. Handicaps [1787]. Handle [1766]. Hands [809]. Happened [190, 15]. Happens [1411, 332]. Happiness [1426]. Hard [1342, 1833, 1445]. Hardcover [1249, 1250, 1247, 1248]. Harder [1450]. Hardware [86, 177]. Harmony [66]. Harnessing [1599]. Harold [121, 122, 124, 120, 123]. Hate [1720]. Haven [190]. Having [636, 203]. Haystack [901]. Headlines [965]. Health [1725, 125, 1624, 1558, 746, 759, 774, 788, 810, 893, 1006, 840, 860, 875, 907, 921, 937, 952, 965, 1009, 1026, 1041, 1057, 1070, 1085, 1102, 1116, 1151, 1198, 1277, 872, 1049, 60, 350, 890, 1560, 1779, 1857, 147, 1820, 1743, 365]. Healthcare [1405, 1132]. Healthy [1020, 1064, 1049]. Heart [1151, 1125]. Heavy [394]. Height [415]. Heights [363]. Held [1206]. Help [1534, 993, 1012, 1029, 1045, 1059, 1072, 1087, 1104, 1120, 1136, 1154, 1169, 1186, 1201, 1216, 1236, 295, 294, 759, 770, 1559, 1383, 391, 152, 768, 1177]. Helping [1730, 1794]. Here [617, 746, 759, 774, 788, 810, 893, 840, 860, 875, 907, 921, 937, 952, 965, 1009, 1026, 1041, 1057, 1070, 1085, 1102, 1116, 1151, 1198, 1277, 1383, 1498, 55]. Hertz [58]. Hexagonal [1434]. Hey [1129]. hide [1594]. Hierarchical [836, 899]. High [1070, 391, 1462, 1457, 989]. High-Producing [391]. High-School [1462]. Highlights [1511]. Hill [674]. Hired [777]. Historical [1482, 1002, 480, 1547]. History [516, 105, 1650, 500, 124, 1659, 1133, 499, 1568, 1748]. Hit [1499]. Hite [17]. Hits [1279]. Hitter [1147]. Hitters [506, 264]. Hitting [466, 691, 1396, 1337, 1499, 1166]. HIV [190]. Hockey [484, 293, 1338, 1207, 409]. Hogan [1774]. Hold [1378]. Holiday [430, 445, 106]. Holidays [505]. Holy [1703]. Home [466, 1396, 506, 952, 1207, 1102]. Home-Run [466, 506]. Homeland [652]. homes [1102]. Homicides [1162]. Homme [191]. Homology [1634]. Honest [522]. Honesty [1522]. Hope [323, 1820]. Horizon [1491]. Hormone [1151]. Horse [1230]. Horses [1225]. Hospitals [472]. Hot [61, 440, 395, 29, 62, 1441, 581]. Hotez [1842]. Hour [100]. Household [1844]. Howard [1429, 1774]. Hub [1311]. Hubbard [838]. Human [1828, 1083, 1729, 365, 1544, 158, 1827, 1642, 1564, 1565, 1829, 1562, 453, 350, 1566, 1376, 1622, 1208]. Humans [183]. Humor [1514]. Hundred [155, 1523]. hunted [1865]. Hurrah [700]. Hybrid [1623]. Hype [323, 1820]. Hypergeometric [668]. Hypothesis [1521, 63, 668, 1763, 1779]. Hypothetical [1009]. Hz [60].

Ibry [1314]. Ice [944, 1556, 1207, 1441]. Iceberg [12]. IDEA [1839, 1797]. Ideas [1266, 1326, 1614]. Idées [736]. Identification [885]. Identifying [904, 1639]. Ides [1349]. If [374, 1766, 921, 669, 1291, 1445, 1028]. Ignorance [1324, 1084]. II [1460, 1098, 585]. III [600]. Illinois [1079]. Illness [937]. Illuminations [1405]. Illusion [62, 1782]. Illustrated [1471]. Illustrating [1019]. Illustration [859]. Imagery [830]. Images [38, 788, 741, 338]. Imagination [55]. Imaging [977, 937, 1294]. Imagining [1494]. Immigration [268]. Impact [1223, 799, 1730, 1320, 1009, 847, 1851, 1471, 865, 1712, 148, 1739, 848, 850, 1019, 849, 851, 852]. Impacts [1558, 1557]. Imperfect [1552]. Implant [772]. Implementing [782]. Implementor [99]. Implicit [1533, 285]. Importance [1654, 1238, 1243, 1143, 1557]. Impress [563]. Improve [1405, 209]. Improved [927, 707, 930]. Improvements [1672]. Improving [1026, 1254, 276, 958, 1007, 1008, 1611]. Imputation [1110, 1026, 1020, 1064]. imputing [1019]. In-depth [1777]. Incalculable [247]. Inclusion [1702]. Inconsistencies [358]. Incorporating [1813]. Increase [1813, 223, 314, 807]. Increasing [1725]. Incremental [1672]. Independent [1615]. Index [153, 235, 288, 331, 435, 434, 491, 535, 588, 647, 701, 750, 800, 686, 1400]. Indian [1732]. Indicators [661]. Indigenous [1728]. Individual [1223, 324]. Industrial [1819]. Industry [138, 745, 28, 1225]. Inequality [986, 1752]. Inequality-How [1752]. Inextricable [1275]. Infant [1162]. Infer [935]. Inference [1854, 1610, 1327, 1283, 1328, 1176, 1426, 1451, 1086, 398]. Inferences [125, 765, 1445]. Influences [1472, 324]. Influencing [1792, 1842, 1412]. Influenza [1009, 1198, 806]. Info [287]. Information [792, 1805, 1297, 314, 335, 1479, 782, 1517]. informative [1445]. Informing [1577]. Infovis [1568]. Ingestation [1652]. Ingredients [1416]. Initial [886, 1387]. Initiative [1702, 1792]. Injectors [190]. Ink [1670]. Innate [866]. Innovations [697, 679]. Inquiry [1230]. Inside-Out [1152]. Insider [1455]. Insight [606]. Insights [1832, 1582]. Insignificant [1603, 936]. Inspecting [606]. Inspired [1112]. Instinct [350]. Institutes [1306]. Institutions [676]. Instructional [1718]. instructions [289]. Instruments [1552]. Insurance [907, 528, 629]. Integrated [783]. Integrating [1799]. Integration [1448, 783]. Integrity [1841, 883]. Intelligence [637, 1308, 638]. Intelligent [1024]. Intent [467]. Intentions [1001]. Interact [1369]. Interactions [231]. Interactive [1254]. Intercomparison [1553]. Interdisciplinary [1414]. Interest [1213, 43]. Intergalactic [1635]. Interim [74]. Intermediate [167]. International [253, 1515, 1397]. Internet [286, 364]. Interpretable [1666]. Interpretation [1476]. Interpretations [795]. Interpreting [682]. Interval [1852]. Interventions [1009, 1712]. Interview [1847, 1356, 1386, 1429, 1784, 1842, 1410, 1006, 1518, 946, 1010, 1398, 1290, 1536, 498, 975, 1357]. Interviewer [371]. Intro [1652].

Introduce [1397, 246]. Introducing [1423]. Introduction [1422, 1362, 132, 715, 1294, 1023, 1379, 662, 1378]. Introductory [1696, 1448, 1273, 1274, 1583]. Inventions [544, 486]. Inverse [121]. Investigation [693, 1667, 13, 532, 1079]. Involving [469]. Iowa [1144]. IQ [317, 433, 797]. Iran [1109]. Irregularities [595, 372]. ISBN [1249, 1250, 1247, 1248]. ISBN-10 [1249, 1247]. ISBN-13 [1250, 1248]. ISI [80]. Island [1560]. Isles [1225]. Isn’t [841]. Isolation [1598]. isomorphisms [1521]. Issue [628, 1722, 1633, 1689, 1309, 803]. Issues [912, 1273, 1480]. Item [1127, 1117]. items [1019].

Jabine [1010]. Jack [655]. James

[1384, 1608]. Janet [1410]. Japan [129, 15]. Java [504]. Jay [1321]. Jeffreys [121, 122, 123, 124, 120]. Jelly [1180]. Jensen [1482, 639, 637]. Jeopardy [255, 1159]. Jersey [206]. Jesse [1431]. Jewels [713]. Jim [1386]. Job [295, 294, 724, 1776]. Jobs [1867]. Joe [1178]. John [215]. Johnson [412, 1371]. Jokinen [1053]. Joseph [713, 1245]. Journal [1843]. Journals [448, 270, 1766, 218, 1082]. Journey [1825]. JSM [881]. Judge [1366, 95]. Judging [425, 395]. Judgment [279]. Judicial [140]. Judith [906]. Juiced [699]. Jules [1314]. Junkets [253]. Jury [666, 461]. Jussi [1053]. Just [1636, 856, 542, 55, 785, 1166]. Justice [1470, 1435, 1449, 1493, 1512, 1537, 1569, 1584, 276, 1481, 1594, 1867, 1208].

Kadane} [1321]. Kaplan [655]. Katherine [1853]. Kaufmann [1]. Keep [1729]. Keeping [746]. Kelley [548]. Kendall [171]. Kentucky [1588]. Kettering [96]. Kickers [521, 210]. Kickstarter [1700]. Kids [563]. Kill [1584]. Killed [1714]. Kind [257]. Kindergartners [1307]. Kinematics [1639]. Know [400, 1551, 1609, 1408, 52, 1409, 929].

Knowledge [19, 368]. Kotz [412]. Kullback [1713]. Kurashi [249]. Kyrgios [1403].

L.A. [328]. Labor [724]. Lament [1764]. Lanarkshire [1845]. Land [145, 146]. Landscape [1848, 1690]. Language [1249, 1728, 1250, 1247, 1248]. Large [1259, 1635, 1283]. Large-Scale [1635, 1283]. Larger [1183]. Last [648, 564, 118]. LATE [1854]. Later [502]. Latest [1391]. Latin [1118, 1134, 13, 499]. Laureate [78]. Law [1099, 305, 1760, 1715, 1314]. Laws [1289, 1752, 1761]. Layperson [280]. Lead [787, 1039, 883]. Leaderboard [1520]. Leadership [1529]. Leading [529]. League [550, 1337, 871, 376, 409]. Learn [82, 1434]. Learned [1696, 1729, 1542, 551]. Learner [1822]. Learning [1666, 1404, 1624, 1055, 1730, 1009, 771, 1393, 1572, 1685, 1274, 1669, 1483, 1769]. Least [1337, 417, 1130]. Least-Absolute-Deviation [417]. left [1165]. Legacy [713]. Legal [469]. Legendary [185]. Legislation [258]. Leibniz [1782]. Leicester [1593]. Leisure [687]. Leland [1770]. Length [1242, 1606]. Leninist [476]. Lens [1677]. Leone [1277]. Less [1817]. Lesser [254]. Lesson [551]. Lessons [1723, 1696, 1729, 731, 1114, 1542, 606]. Let [9, 1517, 1628, 436]. Letter [163, 204, 493, 508, 619, 911, 1017, 1172, 1189, 1174, 1205, 1173, 1190, 1220, 1237, 1253, 1269, 1287, 1305, 1317, 1333, 1350, 424, 136, 1367, 1381, 1395, 1407, 1427, 1439, 1452, 1461, 1473, 1484, 1496, 1506, 1516, 1527, 1538, 1549, 1561, 1576, 1586, 1595, 1604, 1617, 1632, 1643, 1175, 298, 1438, 457, 188, 981, 996, 1015, 1032, 1048, 1062, 1076, 1091, 1108, 1124, 1140, 1158, 941, 956, 969, 415, 926, 1791, 1802, 1815, 1826, 1838, 1849, 538, 1653, 1665, 1676, 1688, 1699, 1711, 1721, 1735, 1749, 1759, 1772, 1781, 803, 1077, 658, 880].

Letter [219, 320, 738, 403, 650, 633, 334, 479, 1845, 1075, 1090, 1107, 1123, 1157, 1204, 1219]. Letter-Frequency [658]. Letters [35, 57, 89, 344, 477, 490, 494, 523, 552, 634, 653, 665, 725, 1063, 220, 273, 164, 1252, 260, 238, 689, 957, 591, 565, 717]. Letting [1554]. Level [1760, 1280, 541, 988, 1034, 1623, 989]. Leveling [1470]. Liars [1741]. lied [581]. Lies [1574, 33, 33]. Life [1324, 1345, 1560, 629, 1390, 1020, 1064]. Lifelengths [529]. Light [1040]. Lighter [48, 70, 103, 152, 180, 236, 203, 218, 272, 253, 289, 312, 332, 343, 355, 366, 388, 436, 420, 465, 507]. Lights [1129]. Like [1445, 228, 1572, 428, 542, 389, 1483]. Likelihood [1332]. Likely [1593]. Lilliput [318]. lime [687]. Limit [985]. Limitation [1243]. Limitations [1734]. Limited [1163]. Limiting [1143]. Line [1714, 417, 413, 1133]. Linearity [1508]. Linemen [1223]. Lines [694, 1373, 1167]. Lining [176]. Link [1275, 224]. Linkage [756, 685]. linked [1082]. Linking [170]. List [430, 445, 126]. Listen [645, 1756]. Listicle [1603]. Literacy [359, 989]. Literary [446]. Literate [630]. Literature [1521, 1602, 719, 1704]. Litigious [212]. Little [1084]. Livermore [1431]. Lives [1867]. Livestock [392]. Living [977, 1844]. Load [986]. Locating [1491]. Location [935]. Loci [899]. Logic [323]. Logical [1268, 271, 820]. Logicomix [1303]. Logistic [1445, 1592]. Long [983, 413, 723]. longer [555]. Longitudinal [594, 1078]. Look [1477, 814, 1502, 913, 1618, 1755, 743, 914, 1334, 882, 1384, 1608, 1021, 1067, 1129, 963]. Looked [201]. Looking [1323, 1194, 1042, 1720]. Loops [966]. Lorenz [392]. Lose [1385]. Loser [826]. Losers [568]. Loss [978, 934, 1692]. Losses [873]. Lost [388]. Lottery [410, 1658, 1808, 1368, 500, 1763]. Lotto [485, 130, 439, 443]. Love [1568]. Low [1636]. Low-Resolution [1636]. Lucia [1597]. Luck [602, 1789, 1768, 1587, 1037]. Lucky [210]. Lunacy [1601]. Lung [1070, 223, 224]. Luria [1126].

M&M [370]. Machine [1666, 1729, 1055, 1730, 1393, 1669]. Machinery [837]. Machines [1669]. Made [663]. Madness [987, 1038]. Magic [1326, 1135]. Magical [1326]. Magnetic [788, 58, 1294, 60]. Magnetism [1485]. Mailrooms [668]. Mailtrap [761]. Maintaining [1729, 419]. Major [474, 799, 1443, 1337, 871, 550]. Make [342, 1335, 892]. Makes [1575, 498]. Making [779, 1765, 1413, 350, 1833, 1295, 199, 688, 1054, 284, 1548]. Mammal [704]. Man [675, 191, 498]. Management [1805, 1448, 7, 339]. Manager [558]. Managers [1227]. Managing [724, 39, 325]. Manchester [376]. Mandatory [1430]. Manga [1421]. Mania [485]. Map [935, 781]. Mappers [79]. Mapping [1621, 1635]. Maps [1245]. Marathon [226, 1589, 527, 1399]. March [987, 1038, 1349]. Marey [1314]. Marital [66]. Market [107, 834, 611, 1141, 518]. Marketing [406, 669]. Markets [724, 1292]. Markov [1249, 38, 1868, 586, 918, 612, 1655]. Marks [369]. Marriage [1351]. Marry [112]. Marxist [476]. Mary [868, 1481]. Maryland [1565]. Masonry [265]. Mass [740, 743, 744]. Masses [1624, 688]. Master [207, 736]. Mastermind [186]. Masterpieces [1825]. Masters [1592]. Match [1474]. Matches [1444]. Matching [1036]. Matchmaker [1351]. Maternal [806]. Math [1817, 962, 1418, 821, 719, 193, 1457, 247]. Mathematical [1326, 1751, 1463]. Mathematics [1326, 1626]. Matloff [1264]. Matter [1577, 371, 695, 705, 785]. Matters [1101]. Maurice [171]. Maximizing [1658]. May [1017, 1249, 1247, 1018]. Maybe [1822, 1334]. MCAT [1272]. McGrayne [1247, 1241]. Me [1402]. Mean [1215, 1361, 1334, 1191, 431, 1053]. Means [611, 329, 354, 1491]. Measure [886, 230, 291, 890, 360, 355]. Measured [1280]. Measurement [353, 743, 1150]. Measures [1645, 1723, 1552, 1657, 835]. Measuring [1223, 484, 621, 1862, 870, 1564, 1338, 1049, 1541, 1767, 1504]. Medals [667]. Meddling [373]. Media [295, 294, 1413, 958]. Medical [1271, 1364, 1055, 746, 1263, 1273, 1093, 1405, 1660, 685, 1230, 1778]. Medicare [892]. Medicine [1515, 1275, 774, 1151, 1276, 1099, 589, 1820]. Medics [659]. Meet [1666, 1562, 662]. Meeting [343]. Meets [194, 1440]. Mega [1763]. Memoriam [1, 72, 1770]. Memories [971]. Memory [1487]. Men [413, 1382]. Mendel [874]. Mental [937]. Mere [1002]. Messy [1674]. Meta [1602, 1332, 1618, 115]. Meta-Analysis [115, 1602, 1618]. Meta-Analyst [1332]. Metabolomic [998]. Metabolomics [739, 997]. Method [1723, 1836, 847, 542, 883, 820, 848, 850, 849, 851, 852]. Methodology [1272]. Methods [1364, 1055, 1640, 1476, 1545, 1577, 1228, 1347, 1803, 1327, 1566, 1283, 1616, 1767, 900, 1680, 1383, 1615]. Metric [1080]. Metrics [1520, 1656]. Metropolitan [517]. Meyer [868, 913]. Mice [1388]. Michael [245]. Michigan [1248]. Mick [1165]. Microarray [902]. Microarrays [900]. Microdata [1693]. Might [1337, 313]. Migration [621, 1728]. Mike [466]. Mileage [1341]. Military [1581]. Milk [428, 1845]. Milky [1639]. Millennium [560]. Million [554]. Millions [405, 1763]. Milosevic [692]. Minard [713]. Mine [1536]. Mining [460]. Minority [97, 315]. Minus [1304]. Miracle [1097]. Mirage [991]. Mirandola [156]. Mirrors [1501]. Mis [1182]. Mis-Shelved [1182]. Misattribution [1704]. Misbehaving [1149]. Misclassification [1430]. Misconceptions [1508]. Misconduct [421]. misleading [639]. Misreporting [1026]. Misrepresenting [1670]. Missed [626]. Missile [830]. Missing [1026, 1490, 1094, 1390]. Misspent [49]. Mistakes [1319]. Misuse [823, 887, 1739]. Misuses [107, 637]. Mitigate [1794]. Mitigating [915]. Mixed [1566]. Mixing [1180]. Miyagi [1254]. MLB [646, 1591, 1147]. Mobile [1600, 1277]. Modalities [1494]. Modality [1294]. Model [1515, 521, 759, 258, 474, 18, 952, 1592, 1623, 1553, 1221, 1078]. Model-Based [474, 18]. Modeling [550, 587, 1581, 250, 1009, 1116, 794, 1469, 376, 541, 1652, 1489, 1267, 1392, 1051, 85, 1712, 1463]. Modelling [1298]. Models [836, 1682, 1288, 899, 1177, 1488]. Modern [1695, 675, 1476, 1546, 1543, 1541, 1542, 1327, 659, 1830, 2, 1583]. Modification [1655]. Modified [927]. Molenaar [142]. Molenberghs [1518]. Möller [885]. MoMath [1432]. Momentum [394, 1096]. Mommy [1334]. Monitoring [860, 1139, 1732, 147]. Monopoly [530]. Monte [1249, 1868, 11, 586]. Monty [119, 1782, 1703]. MOOCs [1459]. Mood [605]. Moral [1245]. Morris [72, 1398]. Mortality [1277, 278, 1601]. Most [857, 1605, 1413, 1376, 1384, 1608]. Mother [838]. Motivate [785]. Motivations [1734]. Moto [249]. Motorcycle [1601]. Move [170]. Moveat} [347]. Moves [357]. Movie [390, 568]. Moving [1507]. Moyen [191]. Mozart [1101]. MR [948]. Mr. [1671, 1254, 974]. MR2149936 [963]. MR2370724 [949]. Much [1193, 49, 115, 183, 768, 1212, 177]. Multi [779]. Multi-Way [779]. Multinational [759]. Multinomial [1180]. Multiple [1162, 1110, 1854, 1828, 1026, 1337, 168, 1544, 1020, 1064]. Multiple-Choice [168]. Multiplicity [1644, 1259]. Multispecies [902]. Multivariate [716, 406]. Murder [302]. Murders [1324]. Music [462, 792, 1700, 59]. Musical [794, 795, 793]. Musing [1406]. Musings [34, 56, 88, 135, 165, 189, 205, 221, 239, 299, 321, 495, 509, 524, 539, 553, 566, 577, 592, 604, 620, 635, 651, 664, 677, 690, 703, 714, 727, 752, 764, 778, 791, 261, 274]. Must [374]. My [505, 563, 575, 55]. Mysteries [1000]. Mysterious [370]. Myth [58, 405]. Myths [339].

N=1 [1256]. NAEP [599]. Naked [1377]. Name [309]. Napkin [1330]. narrative [1547]. Nation [291, 26, 147]. National [419, 661, 500, 1542, 1132, 1006, 1649, 594, 409]. Nationality [409]. Nations [676, 838]. Natural [1598, 1066]. Nature [812, 1301]. Navigate [561]. NBA [766, 1513, 1040, 1149]. NCAA [1611, 315, 1851, 987, 1382]. Nearness [600]. Necessarily [1588]. Need [628, 770, 110, 1498]. Needed [1449, 961, 224]. Needles [901]. Needs [400, 686, 310]. Neglected [731, 803]. NELS [594, 593]. Neopythagorean [1657]. Net [230, 1444]. Netflix [1111]. Netherlands [1544]. Network [754, 1082]. Networks [904, 1335]. Neumann [215]. Neuroimaging [1294]. Never [387, 1499]. Nevius [1806]. News [295, 294]. Newspapers [809]. Next [474, 1442, 1146, 1492, 997]. Nexus [1224]. NFL [521, 511, 1450, 210, 1509, 1651]. NHL [1434]. Nick [1403]. Nigeria [1724]. Night [117, 971]. Nightingale [1224, 1814]. NISS [1417]. No [1017, 1401, 1316, 1796, 401, 134, 264, 861, 813, 827, 842, 894, 908, 922, 938, 953, 992, 1011, 1147, 1028, 1018, 555]. No-Hitter [1147]. No-Hitters [264]. Nobel [78, 336]. Nobody [775]. Node [1354]. Nomograms [545]. Non [1430, 1712]. Non-Drug [1430]. Non-Pharmaceutical [1712]. Noninferiority [1085]. Nonparametric [1695, 1358]. NonRANDom [407]. Nontraditional [718]. Nonuniform [1036]. Normal [451, 951]. Norman [1264]. Northern [296]. Nostalgia [1656]. Note [406, 1460, 1430]. Nothing [1262, 1028]. Notice [831]. Notices [32]. Nozer [1846]. NSABP [438, 423]. Nuclear [240]. Null [1554, 63, 1779, 1164]. Number [215, 1629, 313]. Numbers [788, 450, 1302, 157, 382, 17, 760, 73, 1544]. Numerical [1616, 1231]. Numerics [1837]. Nursing [952, 1102]. Nurture [812]. Nurturing [1460]. Nuts [177].

O [1214, 1192, 1548, 1293, 1471, 1191, 1213, 1226, 1243, 1311, 1447]. Oasis [991]. obesity [1020]. Obey [1166]. Obituaries [1631]. Objectives [147]. Observational [1050, 1336]. Observations [1004]. Obtain [1342]. Occult [1385]. Occur [1650]. Ocean [1491, 1488]. Oct [1250]. Octopus [1385]. Oddities [719]. Odds [263, 1067, 93, 1470, 1435, 1449, 1493, 1512, 1537, 1569, 1584, 1593, 83, 1097, 1225, 276, 1481, 482, 498, 1594]. Ode [1563]. Odyssey [520]. Off [304, 1129, 944]. Offense [1025]. Offensive [1223, 587, 1066]. Offerings [886]. Office [374]. Officers [1714]. Official [1564, 947, 948]. Ohio [916]. Oil [1491]. Okay [61]. Old [1295, 973, 838]. Olympic [1589, 373, 555]. Olympics [667, 767, 854]. OMB [427, 401]. Omega [562]. Omics [1531]. Once [1168]. Oncology [1647, 1680]. One [675, 712, 759, 1856, 1523, 1133, 1812, 1835, 1153]. Online [1696]. Only [841, 831, 1445]. Ontario [1368]. Open [1840, 1228, 1843, 1841, 1839, 1639]. Open-Access [1843]. Openings [724]. Openness [1271]. Operating [385]. Operational [799]. Opinion [58, 90, 303, 551, 576, 589, 632, 603, 618, 36, 48]. Opinions [1127]. Opportunities [1467, 1417, 1259]. Opportunity [340]. Opposition [1316]. Oprah [933]. Optimal [1179, 1094]. Optimism [1385]. options [556]. Orbits [38]. Order [789, 158, 456, 324, 615]. Ordinal [1592]. Oriented [1698]. Origin [215, 124, 251]. Orlando [304]. Orleans [22]. Orphaned [759]. orthogonal [1130]. Oscars [856]. Other [1729, 1041, 1499, 1453, 1153, 719, 567, 339]. Ottawa [475, 441]. Ottoman [516]. Our [661, 291, 1634, 1335]. Ours [958]. Outbreaks [1578]. Outcome [1235, 1251, 1698]. Outcome-Oriented [1698]. Outcomes [167, 1210, 959, 426]. Outperform [240]. Outreach [1702]. OutSMARTing [1678]. Ovarian [742]. Overall [1382]. Overestimates [405]. Overhead [199]. Overtime [1450, 293]. Overview [1050, 1829, 1150]. Owl [1224]. Oz [716].

Pâté [1131]. Pace [746]. Package [1497]. Packages [2]. Page [71, 104, 154, 181, 237, 290, 333, 367, 437, 492, 537, 590, 649, 702, 751, 801]. Pages [1401, 506, 521, 534, 550, 563, 575, 587, 602, 617, 631, 646, 660, 675, 687, 699, 712, 722, 734, 762, 777, 789, 798, 812, 841, 876, 891, 920, 1450, 950, 964, 1025, 1513, 1227, 1244, 1261, 1280, 1313, 1338, 1482, 382, 395, 408, 426, 444, 463, 468, 482, 1434, 1249, 1250, 1247, 1248]. Paired [921]. Palm [595]. Pamela [330]. Pandemic [1730, 1683, 1009, 1754, 1712]. Panel [1358, 1409]. Papers [31, 1112, 116]. Paradigm [1404]. Paradox [281, 964, 467, 1184, 611, 1002, 548, 688, 461]. Paradoxes [1328]. Parallax [1818]. Paralyzed [1507]. Parameters [795]. Paris [80]. Parity [133, 1482]. Parker [641]. Part [1548, 445, 1460, 44, 67]. Partially [1752]. Participation [1531, 541, 767, 1724, 784, 1787]. Partnering [420]. Party [1854]. Pascal [1100]. Past [971, 1009, 1456, 1762, 1475, 1409, 713, 760, 643]. Pastrami [3]. Patent [694]. Patents [544]. Path [825]. Paths [1525]. Pathways [1725]. Patient [1116]. Pattern [1382]. Patterns [1486, 644]. Pay [1764, 882, 831]. PCs [385]. Peek [1432, 364]. Peer [1560]. Penman [1667]. Penny [531, 557]. Pentium [301, 300]. People [1131, 1361, 1049, 1756, 1547]. Perceive [928, 1813]. Percentage [871]. Perception [1041, 784]. Perceptions [988]. Peremptory [758]. Perfect [1474, 775]. Performance [786, 226, 1656, 1851, 527, 541, 1509, 209, 1066, 272, 1053]. performances [555]. Perhaps [921, 929]. Persi [1290]. Persistent [1634]. Personal [201, 364, 346, 385, 398, 416, 460, 470, 487, 287, 271, 330, 311, 216, 251, 178, 68, 45, 101, 150, 858, 346]. Personalities [529]. Perspective [1111, 759, 560, 731, 1811, 268, 1274, 111, 1393]. Perspectives [808, 421, 1415]. Pertaining [1273]. PES [95]. Peter [1842]. Petiet [1314]. PGA [983, 1238]. Pharmaceutical [1712]. Pharmacogenetics [1057]. Pharmacogenomics [1057]. Phase [1098]. Philatelic [1118, 1134, 1023]. Philosophical [156]. Phishing [961]. Phone [389, 1206]. Photometric [1636]. Phylakopi [578]. Physical [124, 1285]. Physicist [1821]. Physics} [481]. pi [162, 161]. Picked [1818]. Pickett [717]. Picking [718]. Pico [156]. Picture [1239, 1279, 1294, 1259, 1359, 1322, 1310, 1339, 1372, 1388, 1402, 1433, 1446, 1468, 1494, 1570, 1612, 1625, 33]. Pictures [1069]. Picturing [328]. Pie [418]. Pies [1167]. Pillars [1573]. Pioneer [1056]. Pioneering [1502]. Pipeline [1306]. PIPES [475]. Pitchers} [1400]. Pitching [264, 1147, 1021]. Pitfalls [1661]. Pixel [1850, 1807]. Place [42, 826]. Placebo [669, 670]. Placekicks [455]. Placement [1225]. Places [952]. Plan [892, 526, 1391]. Planning [902]. Plant [998, 1487]. Plate [1663]. Platforms [1600]. Platykurtic [951]. Play [374, 130, 1655]. Played [950]. Player [766, 1591, 1337, 1338, 1195]. Players [511, 456, 1149]. Playfair [380, 352, 381, 486]. Playing [1470, 316]. Playoff [513, 463]. Playoffs [1261, 918]. Plays [1101]. Pleasure [1510, 1114]. Pleasures [145, 87, 146]. Plenary [1842]. Plot [682, 1622, 688]. Plots [1434, 1152]. Plotting [1324, 2]. Plus [1304, 487]. Podcast [1822, 1832]. Poems [1746]. Poetry [10, 131]. Point [570, 1590, 1181]. Points [1211, 1805]. Poisson [160, 612]. Pokémon [563]. Poker [1193, 1148, 1037, 358, 1222]. Police [1714]. Policy [628, 1324, 79, 708, 350, 284, 65, 1798]. Policy-making [284]. Political [680, 984, 747, 681]. Politics [337, 450, 1543, 1127, 984, 854, 1370]. Poll [184, 771, 1382]. Poll$ [1141]. Polling [1708]. Polls [1596, 303, 947, 182, 948]. Polyp [893]. Ponderings [601]. Pools [508, 374, 987]. Poor [1025, 256]. Popcorn [406]. Popular [58, 70]. Population [511, 1358, 246, 1767, 1820, 1115]. Populations [392, 345]. Portal [1743]. Portrait [745, 160]. Position [691]. Positive [1717]. Positives [990]. possibilities [997]. Possibility [1717]. Possible [1361, 60]. Possum [704]. Post [1450, 1144, 1758, 1382]. Post-2011 [1450]. Post-Doctoral [1758]. Post-Election [1144]. Posted [482]. Posterior [1332]. Postmenopausal [1151]. Potential [1762, 1360, 1857]. pots [578]. Pound [531]. Power [86, 1408, 60, 1599, 1409]. Powerful [1209]. Practical [1362]. Practice [1246, 1163, 1647, 1243, 1274]. Pragmatics [1613]. Praise [1257]. Pre [1464]. Pre-college [1464]. Preach [1246]. Precision [1647, 1820]. Precursors [486]. Predictability [266]. Predictable [856, 1382]. Predicted [426]. Predicting [167, 671, 1210, 1040, 12, 568, 1501, 1102].

Prediction [1283, 1708, 1141, 1481, 1292]. Predictions [971, 542, 1230]. Predictive [1440]. Preference [432]. Pregnant [170]. Premier [376]. Preparation [593, 594]. Prepare [1009]. Preparing [1454, 1468]. Preprocessing [900]. Presence [668]. Present [643, 1456, 1475]. Presentation [214]. Presentations [31]. Presenter [1842]. Presenting [1479]. Presents [1391, 1259]. Preservation [1681]. Preserving [1692]. President [1737, 1777, 692, 1410, 1370]. Presidential [1596, 947, 1109, 1146, 948, 85, 1760, 360, 797, 1505]. Presidents [1409]. Press [1249, 1250, 1247, 1248, 58]. Prevalence [1545, 1541, 1540]. Prevalences [1542]. Prevention [1546]. Prial [641]. Price [686, 834, 1777]. Price$ [1141]. pricing [556]. PRIM [1639]. Primary [970, 1488]. Prime [1662]. Principles [1321, 1282, 1709]. Prior [19, 1297]. prison [1594]. Privacy [1695, 14, 1548, 1293, 1433, 1612, 1689, 1790, 1471, 685, 1191, 1213, 1226, 1243, 1311, 1351, 1447, 1690, 98, 1681, 1692, 1214, 1192]. Privacy-Preserving [1692]. Private [1694, 1594]. Prize [1515, 1111, 96]. Prizes [546]. Pro [330, 983]. Probabilistic [137, 1371, 1393, 1837]. Probabilities [1183, 439, 443]. Probability [928, 446, 476, 455, 16, 119, 1289, 1343, 172, 1183, 1147, 1024, 1575, 1630, 1684, 1810, 1858, 1344, 1752, 1378, 1117, 1703]. Probability-Seventeen [1630]. Probable [149]. Probit [1752]. Problem [41, 564, 1036, 1270, 1856, 757]. Problems [125, 121, 1289, 1762, 1330]. Process [39, 1565, 865]. Processes [1346]. Producing [391]. Product [1160]. Productivity [1488]. Professional [167, 108, 1459, 1210, 280, 68]. Professor [1847, 1518]. Profile [64]. Profiling [628]. profits [556]. Program [1110, 654, 655, 528, 1150]. Programming [1264, 533, 630, 1864]. Programs [49, 228].

Progression [1035]. Project

[1702, 1419, 1223]. Projecting [1558, 1509]. Projection [112]. Projects [1700]. Prolific [1667]. Prologue [5]. Prominent [1670]. Promising [1545]. Promoting [1841, 1517]. Prone [1337]. Proof [758]. Propensity [302]. Proper [546]. Property [668]. Proportionality [206, 207, 208]. Proposed [258, 584]. Prospects [1509]. Prostate [375]. Protect [1565]. Protecting [912, 1293]. Protection [1311, 782]. Protective [1723]. Proteomics [739, 744]. Protests [667]. Prove [92, 1435, 247]. Proven [990]. Provide [1341]. Proxies [1552]. Proximity [170]. Psephologists [971]. Pseudo [215, 303]. Pseudo-Opinion [303]. Pseudo-Random [215]. Psychological [394]. Psychology [318]. Psychotherapy [118]. Public [1725, 297, 886, 400, 1702, 872, 1213, 1857, 1726, 43, 1820]. Publication [448]. Publicly [1844, 1517]. Publish [1342]. Published [857, 1605, 1621, 116]. Publisher [1249, 1840, 1250, 1247, 1248]. Pucksberry [1434]. Pull [1593, 1655]. Pulls [612]. Purloined [32]. Pursuit [1329]. Put [228]. Putt [534]. Putting [1697, 984, 268, 1441]. Puzzle [454, 1380, 119, 861, 813, 827, 842, 877, 894, 908, 922, 938, 953, 966, 992, 1011, 1664, 1002, 53, 1095, 1128, 503, 1010]. Puzzles [877, 1746]. Pythagorean [934]. Python [1703].

Q&A [1790]. Qualifying [1399]. Qualitative [1479]. Quality [322, 129, 138, 49, 952, 728, 1132]. Quantification [1400]. Quantified [1359]. Quantify [1080, 256]. Quantifying [1555]. Quantitative [1479, 959, 899, 1799]. Québec [372]. Quelques [736]. Question [379]. Questions [1241, 1805, 192, 99, 232]. Quetelet [10, 1224]. Quetelet/Nightingale [1224]. Queuing [916]. Quick [1744]. Quinta [1492]. Quota [1620].

R [96, 1264, 1855, 1863, 1775, 1864, 1436, 1571, 1809, 1497, 1861, 1250, 1365, 1616, 1769]. R. [499]. R.I.P. [835]. Race [353, 427, 317, 371, 708, 1631, 1787]. Race-Based [708]. Racetrack [263]. Racial [628, 609]. Racing [1225]. Raise [440, 1816]. Ramadan [943]. Random [472, 476, 215, 394, 1783, 1575, 1615, 1752, 338, 1166]. Randomization [1646]. Randomize [921]. Randomized [589]. Randomness [776]. Range [928]. Ranges [1489]. Ranking [472, 281, 1207]. Rankings [722]. Rao [1855]. Rapid [1723]. Rare [405]. Rate [1430]. Rated} [952]. Rates [1182]. Rather [1325]. Rating [1679]. Ratings [667]. Rational [350]. Raton [1250]. Ray [1831]. Re [1805, 312]. Re-Reading [1805]. Reactors [240]. Read [1510, 687, 289]. Readability [1762]. Readable [199]. Reader [536, 522]. Readiness [1813]. Reading [430, 445, 1805, 1373]. Reads [1401, 506, 521, 534, 550, 563, 575, 587, 602, 617, 631, 646, 660, 675, 687, 699, 712, 722, 734, 762, 777, 789, 798, 812, 841, 876, 891, 920, 1450, 950, 964, 1025, 1513, 1227, 1244, 1261, 1280, 1313, 1338, 1482, 382, 395, 408, 426, 444, 463, 468, 482, 1434, 1631]. Readying [284]. Real [1316, 1674, 648, 1373, 601, 1095, 1128]. Reality [893, 601, 62]. Really [93, 926, 567, 229, 376]. Reaping [1582]. Reason [648]. Reasoning [139]. Reasons [1404, 1603, 805]. Reauthorization [65]. Reborn [1184]. Rebounder [378]. Receiver [1509]. Recession [1221]. Recognition [757]. Recognizing [427, 1860]. Recollections [122]. Recommend [1852]. Recommendations [427, 1004]. Reconciling [1222]. Reconsidering [810, 1642]. Record [308, 1588, 756, 1544]. Recreating [1800]. Recruiting [1581]. Redesigned [246]. Redesigning [663]. Redistricting [984]. Reduce [1206, 248]. Reducing [1717, 1651, 984]. Reduction [19]. Reestimating [820]. Referendum [372]. Referrals [97]. Refining [570]. Reflect [1568]. Reflections [1318, 1528, 502, 438, 1797, 423]. Reflects [1796, 64]. Reform [337, 419]. Refracted [1824]. Registry [1026]. Regress [431]. Regression [1445, 1592, 417, 211, 1334, 1053, 1130]. Regressors [1603]. Regular [831]. Regulatory [904, 1288, 1517]. Reinventing [365]. Rejection [1185]. Rejoinder [608, 949, 543]. Related [1480, 959, 409]. Relations [310]. Relationships [317, 1453]. Relative [646, 1862, 546, 1238, 109, 1504, 1631]. Release [779]. Relevant [43]. Reliability [1160]. Reliably [875]. Relief [1730]. Remarkable [1230]. Remarks [90]. Remembering [1519, 1855, 1853, 1806, 1846]. Remembrance [760]. Reminiscences [123]. Renewal [65]. Rent [1565]. Repeated [835]. Replacement [1151]. Replicability [1737, 1644]. Replicating [1621]. Reply [804, 818, 931, 640, 656, 142, 726]. Report [1260, 74, 1132, 1583]. Reported [1720]. Reporting [1543, 1472, 676]. Reports [663, 952]. Representation [1209, 680, 681]. Representing [173]. Reproducibility [1271, 1644, 1270]. Reproducible [1404, 1548, 1472, 721]. Republication [1342]. Requirements [1740]. Resampling [141, 114]. Rescoring [936]. Rescued [1084]. Research [1110, 1271, 406, 1050, 1840, 353, 1550, 107, 1740, 1823, 1827, 857, 1605, 1559, 300, 1548, 1112, 611, 1424, 685, 721, 1243, 1272, 1534, 1649, 1226]. Researcher [1575]. Researchers [1670, 1369]. Reserve [831]. Residents [770]. Resizing [242]. Resolution [1636]. Resonance [788, 1294]. Resources [1743, 1163, 1462, 1465]. Respective [73]. Respondent [1829]. Respondent-Driven [1829]. Responds [869]. Response [917, 822, 1127, 623, 144, 1580, 1117]. Responses [1252, 785]. Responsibility [1599]. Responsible [1752]. Responsive [1646]. Restricted [1226]. Restricted-Use [1226]. Result [185]. Results [353, 541, 1717, 266, 373, 1128, 676, 874]. Retailer [1368]. Rethinking [1524]. Retiring [1010]. Retrospective [75, 76]. Return [145, 146, 1197, 626]. Returns [1655]. Revealed [704]. Revelations [1324, 44, 67, 99, 1495, 148, 176, 232, 199, 214, 285, 249, 269, 328, 309, 336, 363, 352, 402, 428, 414, 381, 486, 449, 532, 503, 518, 548, 573, 585, 561, 600, 615, 629, 644, 688, 663, 676, 700, 736, 713, 723, 747, 797, 760, 775, 838, 825, 859, 874, 892, 951, 936, 958, 1027, 1007, 1042, 990, 1056, 1101, 1084, 1069, 1132, 1152, 1167, 1117, 1199, 1212, 1230, 1281, 1245, 1262, 1296, 1341, 1314, 1374, 1426, 1390, 1406, 1451, 1483, 1505, 1547, 1526, 1585, 1631]. Reverend [1449]. Reversal [833]. Review [206, 1249, 86, 401, 140, 398, 470, 487, 1264, 1250, 1247, 1365, 1504, 1503, 1616, 1614, 1615, 1613, 1709, 1710, 1704, 1354, 1248]. Reviewed [330]. Reviewers [390]. Reviewing [517, 1194]. Reviews [234, 47, 69, 102, 151, 179, 202, 217, 252, 1231, 267, 283, 307, 326, 341, 351, 362, 383, 397, 411, 429, 447, 458, 473, 488, 525, 496, 510, 583, 547, 559, 571, 598, 613, 627, 642, 672, 684, 696, 732, 710, 720, 748, 1232]. Revised [526]. Revising [1696]. Revisited [964, 1002, 1590]. Revolution [746, 997]. Reward [1700, 618]. Reward-Based [1700]. Rhetoric [1835]. Richness [1767]. Rid [1406]. Right [1325, 1610]. Rights [1564, 1565, 211, 1562, 1566, 1208]. Rigorous [1325]. Rings [296]. Riots [247].

Rising [98]. Risk

[1579, 1705, 1555, 1041, 1786, 223, 1144, 1693, 1197, 1221, 109, 1816, 1471, 1795, 784, 1143, 1481, 618, 1681, 111, 781]. Risk-Limiting [1143]. Risk-Return [1197]. Risk-Weighted [1693]. Risks [170, 60, 325, 1258]. Risky [823]. Road [8, 5, 9, 7, 6, 1146]. Roadway [1705]. Robust [1694]. Rock [1700]. Rocket [1113]. Rods [1739]. Roger [1065, 1021]. Role [173, 1694, 788, 387, 1550, 860, 1500, 1374, 111]. ROM [1250]. Romantic [1646, 1387]. Rookie [1038]. Room [1718]. Rose [309]. Rosemary [1847]. Roses [1167]. Rotatable [585]. Roulette [1257]. Round [1591, 873]. Rounding [449]. Roxborough [498]. Roxy [498]. RS [23]. RS/Explore [23]. Rubber [935]. Rudin [1666]. Rule [1341, 1865]. Rules [95, 796, 1166, 1098, 1788, 1799]. Run [466, 1396, 506, 257, 636, 183]. Rundel [1536]. Runner [226, 527]. Runners [691]. Running [540]. Russia [1568]. Russian [1865]. Rusty [79]. Ruth [1499].

s [570, 1590, 487]. S-PLUS [487]. Sabotaging [1669]. Safe [1039]. Safeguards [860]. Safety [81, 262, 1288]. Said [229]. Salaries [1313, 409]. Salary [876]. Same [610]. Sample [424, 1352]. Samples [976]. Sampling [387, 1179, 335, 1829, 1620, 668, 345, 1615, 1501, 1208, 398, 933]. SAMSI [1417]. Samuel [1806]. San [196]. SAT [655, 654, 821, 768, 820, 936, 1406]. Satellite [830]. Saved [46]. Savvy [66]. say [1125]. Says [134, 329]. Scale [1635, 1283, 140]. Scaling [363, 518, 1530]. Scandal [1368]. Scatterplot [574]. Scatterplotting [673]. Scenario [668]. Scene [180]. Scenes [902]. Schedule [328, 1314]. Schedules [1568]. Scheduling [927]. Schizophrenia [806]. schleps} [974]. Schmidt [466]. School [1493, 708, 1411, 1462, 1457, 989]. Schools [1702]. Schrödinger [1117]. Schwarz [986].

Schwarzenegger [1164]. Science [1515, 1722, 450, 1793, 1757, 1702, 1805, 58, 1397, 1375, 1610, 1448, 1564, 1416, 1734, 1843, 1480, 1785, 1816, 1718, 1479, 1474, 1839, 1626, 1661, 174, 1360, 1351, 1727, 1530, 1517, 1213, 78, 1843]. Sciences [1353, 1743, 1624, 1345, 1285, 1560, 1779, 1857, 1463]. Scientific [125, 1766, 421, 847, 311, 1841, 1328, 1710, 883, 232, 1056, 848, 850, 849, 851, 852]. Scientist [1572, 1483]. Scientists [1817, 1343, 1758, 1533]. Score [663]. Scores [167, 26, 97, 821, 376, 431, 820]. Scoring [691, 1238, 853]. Scourge [532]. Scrabble [1244, 658]. Scraping [1652]. Screening [893]. Scrutinity [155]. Scylla [561]. Sea [1485]. Search [1675, 54, 186, 1668, 1303]. Searching [1508]. Season [1080]. Seasonality [1198]. Second [826, 223, 1610, 1448, 1709, 1710, 1858, 224, 1298]. Second-Hand [223, 224]. Secondary [1467]. Secret [1373, 1713]. Secrets [410]. Section [881, 1773]. SED [347]. See [1832, 1593]. Seed [444]. Seeing [928]. Seeming [933]. Seems [603]. Selected [77]. Selecting [391]. Selection [407, 452, 1019]. Selections [729]. Self [1600, 978, 405, 982, 605, 606, 607, 152]. Self-administered [1600]. Self-Defense [405]. Self-Experimentation [978, 982, 605, 606, 607]. Self-Help [152]. Sell [1658]. Sellers [1278]. Selves [1824]. Senate [926]. Sense [1765, 176]. Sense-Lining [176]. Sensitivity [1621]. Sentences [206]. Sequel [127]. Sequential [1754]. Series [1640, 463, 195]. Serious [187]. Serjev [1314]. Services [1502, 1004, 1027, 1102, 1003, 365]. Sesquicentennial [59]. Session [80, 881, 1411]. Sessions [1369]. Set [279]. Sets [1259]. Setting [1005, 1448, 728]. Settings [1125, 1369]. Seven [1528, 1573, 463]. Seventeen [1630]. Sex [1528, 866, 540, 543, 629]. Sexual [1564, 1540]. Sexuality [1564]. Shackles [1483]. Shades [292]. Shakespeare [128, 1440]. Shall [291, 1584]. Shane [1482]. Shape [1719]. Shaped [1112]. Shared [1623]. Shared/Hybrid [1623]. Shares [886]. Sharon [1241, 1247]. Shawangunks [296]. Sheet [1556]. Sheets [935]. Shelved [1182]. Shere [17]. Shlepping [103]. Shoot [456]. Shoot-Out [456]. Shooter [127]. Shooting [408]. Shootout [293]. Short [44, 67, 1630, 723]. Shortcomings [1268]. Shots [1094]. Should [1404, 1766, 170, 1332, 195, 456, 669, 99, 558, 1191, 1690, 929, 1801]. Show [534, 596, 1440]. Showcase [442]. Showdown [442]. Shown [1832]. Siberia [103]. SIBS [1306]. Side [48, 70, 103, 152, 180, 236, 203, 218, 272, 253, 289, 312, 332, 343, 355, 366, 388, 436, 420, 465, 507]. Sides [201]. Sierra [1277]. Sieving [579]. Signals [740, 794]. Significance [928, 1737, 1779, 1867, 1248]. Signs [742]. Silence [1594]. Silhouettes [1635]. Silicone [772]. Similarities [1037]. Simmons [1443]. Simple [1750, 1086, 1799, 892, 1598]. Simpson [275, 688, 461]. Simulating [1301]. Simulation [1116, 501, 1582, 1233]. Simulation-Based [1233]. Simulations [1552, 1857]. Simultaneous [1754]. Sing [1257]. Single [112]. Singpurwalla [1846]. Sir [96, 1515, 121, 122, 124, 120]. Sisters [610]. Sites [170]. Situation [52]. Six [30]. Six-Dimensional [30]. Size [546, 1767, 1143]. Sizes [646, 424]. Skaters [281]. Skating [667]. Skewed [1757]. Skiers [1035]. Skill [1280, 1736, 1587, 1037]. Skills [1813, 1448]. Skylines [1079]. Slave [1497]. Slavery [1539, 1607, 1546, 1545, 1543, 1541, 1542, 1830]. Sleep [1050, 605, 1051]. Sleepers [568]. Sleuthing [806]. Slices [425]. Slicing [585]. Slmple [564]. Slobodan [692]. SLOP [303]. Slower [859]. Slugging [871]. Small [474, 1730, 262, 976, 700]. Small-Area [474]. Smaller [1183]. Smallpox [810]. Smoke [223, 554, 224, 1501]. Smoking [629]. Smoothing [693, 600]. Snapshots [1132]. SNPs [901]. Soccer [1211, 707, 930, 558, 1094]. Social [1353, 425, 1324, 1734, 1413, 228, 1285, 1731, 820]. Social-Class [820]. Societal [1712, 111]. Society [212, 1794]. Software [138, 1502, 301, 39, 1264, 645, 733, 1864, 1594]. Solomon [1713]. Solution [1675, 454, 1543, 1664, 706]. Solutions [1000]. Solve [1746]. Solved [503]. Solving [853, 1330]. Some [1822, 16, 20, 90, 1289, 123, 1756, 1281, 1742, 1019]. Something [698]. Songbook [59]. Songs [1242, 1606]. Sonification [645]. Sort [195]. Sound [1816]. Sources [1210, 776]. South [1125, 1194]. Southwestern [1728]. Sovereignty [372]. Soviet [240]. Spain [1385]. Spam [1765]. Spanish [1677, 1492]. Spatial [621, 733, 1557]. Spatio [1394]. Spatio-Temporal [1394]. Speak [1628, 897]. Speaker [44, 67]. Speaking [1208]. Speaks [536, 522]. Special [484, 1722, 207, 1633, 881, 1689, 1309]. Species [1490, 704, 1489, 1767, 1335]. Specific [733]. Spectra [742]. Spectral [743]. Spectrometry [740, 744]. Speed [808, 1387]. Speeds [808]. Spend [814]. Spies [1167]. Spiral [827, 908, 992]. SPLOMS [585]. Spores [668]. Sport [541]. Sports [1520, 1401, 506, 521, 534, 550, 563, 575, 587, 602, 617, 631, 646, 660, 675, 687, 699, 712, 722, 734, 762, 777, 789, 798, 812, 841, 876, 891, 920, 766, 167, 1450, 1511, 950, 964, 1025, 1513, 318, 1227, 1244, 1261, 1280, 1313, 1338, 1482, 339, 382, 395, 408, 426, 444, 463, 468, 482, 1434, 543]. Spotlight [1773]. SPSS [470]. Spurious [1834]. Spy [1384, 1608]. Square [832, 499]. Square-cell [832]. Squares [1118, 1134, 1130, 13, 499]. Squid [1670]. Stability [133]. Stage [1050, 1448]. Stakes [1100]. Stamps [1023, 361]. Standard [1288, 645, 1867]. Standardized [26]. Standardizing [1103].

Standards [427, 517, 315, 728, 561].

Standing [1758]. Stanford [216]. Stanley [440, 918]. Star [1149]. Stars [1638, 269]. State [182, 1132, 1655]. Statement [1737]. States [1774, 774, 1545, 1115, 443]. Statgraphics [45]. Stationary [1346]. Statistic [54, 747]. Statistical [145, 822, 1454, 87, 146, 1737, 1679, 40, 920, 91, 337, 962, 129, 105, 1581, 112, 474, 502, 1033, 697, 799, 1793, 1005, 1299, 107, 138, 1805, 1792, 1460, 1862, 731, 745, 758, 811, 1577, 912, 79, 1297, 31, 416, 1057, 1163, 301, 1777, 693, 1000, 244, 1347, 806, 1288, 1264, 813, 827, 842, 894, 908, 922, 938, 953, 992, 1011, 1369, 377, 1469, 1250, 544, 168, 311, 1480, 1004, 754, 1773, 919, 178, 630, 1038, 1308, 1798, 1425, 1405, 1392, 1504, 1524, 1573, 1685, 1603, 721, 114, 1051, 1243, 1704, 10, 152, 180, 272, 253, 420]. Statistical [1738, 354, 62, 46, 66, 100, 149, 243, 820, 1547, 65, 98, 782, 932, 1079, 903, 30, 1384, 1608, 1867, 1082, 578, 1021, 1445, 1208, 1003, 1863, 1864, 1126, 303]. Statistically [512, 39]. Statistician [1401, 506, 521, 534, 550, 563, 575, 587, 602, 617, 631, 646, 660, 675, 687, 699, 712, 722, 734, 762, 777, 789, 798, 812, 841, 876, 891, 920, 1450, 1531, 1821, 1514, 950, 964, 1025, 1513, 1764, 1227, 1244, 1261, 1280, 1313, 1338, 1482, 382, 395, 408, 426, 444, 463, 468, 482, 355, 366, 1434, 1631, 111, 1845, 974]. Statisticians [295, 294, 997, 327, 77, 407, 971, 1397, 1246, 1369, 719, 1689, 961, 1756, 1474, 1500, 1442, 1684, 1533, 1690, 103, 256, 231, 131, 1231]. Statistico [404, 446, 485, 476, 462, 520, 505]. Statistics [516, 1467, 1510, 999, 1722, 1694, 1695, 1420, 1422, 1423, 1421, 1539, 1607, 692, 1567, 652, 1819, 1454, 446, 476, 275, 563, 1743, 448, 1275, 1624, 746, 759, 788, 340, 419, 400, 560, 614, 1389, 749, 799, 1550, 1546, 1702, 8, 1545, 1637, 156, 1410, 1514, 1638, 25, 26, 1578, 177, 200, 215, 233, 250, 270, 286, 310, 1228, 1263, 1297, 1246, 1278, 1361, 1325, 1342, 1315, 1375, 1391, 1437, 1445, 1522, 1670, 1597, 937, 965, 1493, 1569, 132, 872, 704, 823, 1556, 1276, 1415, 1464, 1456, 1649, 1408, 5, 1466, 1448, 1623, 1343, 192, 1564, 144, 1559]. Statistics [9, 881, 887, 637, 1345, 528, 753, 529, 1465, 172, 110, 1459, 1755, 1161, 1797, 1412, 1834, 50, 538, 1147, 118, 7, 1562, 1773, 59, 1785, 1776, 1599, 1436, 504, 519, 533, 549, 562, 574, 586, 601, 616, 630, 645, 530, 1024, 1758, 1659, 6, 368, 329, 347, 359, 386, 399, 418, 433, 451, 1866, 284, 1282, 1265, 1284, 1379, 1377, 1394, 1424, 1708, 1368, 662, 673, 685, 698, 711, 721, 733, 761, 776, 245, 342, 705, 1344, 1560, 391, 33, 1462, 1351, 431, 1498, 972, 1455, 576, 618, 36, 70, 332, 929, 1352, 1857, 1726, 1453, 1727]. Statistics [1233, 2, 561, 1084, 1245, 277, 43, 1739, 143, 475, 1583, 1440, 1272, 1835, 1229, 1318, 639, 1515, 441, 1611, 1503, 1710]. StatMost [311]. Stats [1652, 1208]. STATS4STEM [1462]. Statscan [193]. Status [907]. Stay [1460, 1102, 558]. Stay-tuned [1460]. Stays [1411]. Steering [609]. Stein [964]. Stellar [1637]. Step [1199]. Stephen [1353, 1354, 975, 1357]. Stern [1784]. Steroid [53]. Steroids [1066]. Steve [1351, 1356, 1519, 1355, 1352]. Stigler [1314]. Still [347]. Stochastic [462, 1298, 1347, 1346, 1023]. Stock [766, 569, 556]. stomach [1125]. Stonehenge [233]. Stop [1284]. Stopping [796, 1098]. Stories [661, 1630]. Story [1081, 1507, 1151, 1443, 1149, 730]. straight [983]. Straightforward [1200, 1424]. Strange [112, 1856]. Strategies [1827, 987, 1525, 255, 1540]. Strategy [468]. Streak [137, 127, 37, 1337, 1371, 319, 1067]. Streakiness [786, 1396]. Streaks [37, 1337, 1166]. Stream [347]. Street [556]. Strength [1437]. Strike [646]. Stripping [1377]. Strong [1006]. Structure [30]. Structuring [836]. Struggle [1764]. Student [1256, 122, 678, 729, 315, 1620, 1583, 985, 1845]. Student-Athletes [315]. Students [1852, 1404, 1005, 1813, 1559, 1465, 1468, 1785, 1462, 663]. Studies [860, 1559, 1717, 1336, 718, 1517]. Study [937, 1332, 1238, 405, 1288, 1559, 637, 1147, 594, 865, 972, 1240, 638, 1020, 1354]. Studying [774, 541]. Stuffy [55]. Stumbling [825]. Style [1254]. Stylometric [1667]. Stylometry [717, 715]. Subjects [1453]. submarines [1865]. Submerged [1491]. Submissions [1766]. Subprime [1375]. Subsequent [1621]. Success [455, 1507, 1416]. Successes [1754]. Successful [1718]. Succotash [343]. Sufferin [343]. Suffrage [1834]. Sugar [1042]. Summer [505, 137, 1306]. Sunbeams [1778]. Super [807]. Superbugs [1678]. Supercomputers [11, 662]. Superintelligence [1525]. superstars [1037]. Supervisor [1740]. Supported [866]. Supposed [945]. Supreme [356, 1430]. Surfing [1458]. Surgeon [110]. Surgeon-General [110]. Surprises [1630, 605]. Surprisingly [1179]. Surveillance [1578, 711, 755]. Survey [1600, 365, 584, 724, 814, 424, 371, 405, 1369, 1734, 1358, 246, 784, 461, 1019, 1006, 330, 1312]. Surveys [14, 107, 1277, 1542, 48, 1352, 1726]. Survival [1515, 126]. Survive [480]. Survived [1041]. Surviving [1411]. Suspects [1662]. Sustainability [1797]. Sustainable [776]. Swag [1822]. Sweden [14]. Swim [183]. Swimmer [1851]. Swimming [540]. Swiss [927]. Sydenham [1733]. Sydney [767]. Symbolic [1339]. Symmetries [1167]. Synchrony [1557]. Syndromic [755]. Synthesizers [1693]. Systat [470]. System [927, 1679, 513, 474, 502, 697, 735, 49, 707, 31, 930, 1197, 597, 330, 1195, 853, 1481, 98]. Systems [1828, 133, 799, 1298, 1637, 1805, 385, 244, 1563, 793, 1580, 1544].

Table [573, 1101]. Tables [1033, 779, 1621, 449]. Tabular [214]. Tabulation [1143]. Tackles [1689]. Tackling [1181]. Take [614, 1402]. Takes [440, 882, 1655]. Taking [1404, 212, 1260, 1323, 1340, 28, 1448, 1623, 1559, 1469, 1459, 1373, 1387, 1419, 1603, 1492, 1360, 1240, 1295, 1312, 1536, 975, 1357]. Tale [337, 317, 704, 1494, 222]. Talk [436]. Tamales [581]. Tango [678]. Tanur [906]. Targeting [1094]. Tarzans [675]. Task [1737]. Tasks [1610, 1698]. Tasting [512, 3, 562]. Tax [1319, 865]. Taxes [475]. Taylor [1250]. Teach [1454, 1246, 9]. Teacher [678]. Teachers [1454, 1459, 1177]. Teaches [1755]. Teaching [1743, 1624, 1623, 1273, 1254, 1785, 1758, 1560, 1779, 1731, 1857, 932, 1672]. Team [786, 1593, 1444, 657, 1274, 1181]. Team-Based [1274]. Teams [484, 1207, 1414]. Technique [1179, 1491]. Techniques [107, 578]. Techno [245]. Techno-Thriller [245]. Technology [129, 746, 847, 848, 850, 849, 851, 852]. Tectonics [1663]. telecast [807]. Tell [999]. Temperature [527]. Temperatures [624]. Temples [1825]. Temporal [621, 1394, 1557]. Temptations [259]. Tendency [1657]. Tennis [1403, 318, 394, 853, 1520]. Tension [98]. Tenure [1513, 1212]. Termed [1337]. Terror [769, 990]. Terrorism [912]. Test [1364, 1028, 1044, 1058, 1071, 1168, 1185, 1200, 1215, 1235, 1251, 1268, 1286, 1304, 1316, 1331, 1348, 1366, 1380, 593, 799, 26, 875, 97, 186, 228, 594, 185, 431, 676, 1086, 1103, 1119, 1135, 1153]. Test-Bed [186]. testimony [1208]. Testing [1521, 406, 1577, 1430, 668, 767, 1763, 1580, 1283, 1571, 1708, 1779, 561, 1374, 1801, 933, 1083]. Testosterone [765]. Tests [168, 603, 1742]. Texas [1378]. Text [1765]. Textbook [1383]. Textbooks [342]. thaliana} [998]. Thank [737]. Their [73, 814, 1408, 1547].

Them [1289, 1584]. Thematic [1245]. Theorem [985]. Theoretic [930, 1274]. Theory [1241, 1127, 1647, 1247, 1684, 1117, 1865]. Therapy [1151]. There [1401, 759, 1197, 901, 874]. there’d [1028]. These [46]. They’d [1325]. they’re [1445]. Thief [483]. Thin [425]. Thing [841, 1856, 1812, 1835]. Things [1719, 760]. Think [1572, 1483]. Thinking [1299, 1460, 1469, 1038, 1405, 663, 932, 1126]. Thirty [480]. Thomas [1010]. Thoroughbred [1225]. Those [567]. Thou [1548, 1293, 1471, 1191, 1213, 1226, 1243, 1311, 1447, 1214, 1192]. thought [1037]. Thoughts [433, 1742]. Threat [1320]. Three [1723, 1211, 1654, 1673, 1814, 610]. Thriller [245]. Throughput [1070]. Throwing [1147]. Thyself [355]. Ticket [1658]. Tidbits [1499]. Ties [293, 1207]. Tiger [631]. Tight [724, 1509]. Tiles [1244]. Tillman [1370]. Time [1714, 73, 378, 584, 814, 1640, 1279, 6, 339, 1133, 203, 465, 540, 768]. Time-Line [1133]. Time-Use [584]. Timeless [1825]. Times [915, 1380, 1589, 1399]. Timor [1125]. Tips [1776]. Titles [280]. Todai [249]. Today [1330]. Together [1111]. Toll [1687]. Tom [414]. Tomahto [843]. Tomato [843]. Too [1342, 951, 1445]. Tool [884, 836, 114]. Tools [630, 1424, 733, 573, 585, 600]. Top [1603]. topcoding [781]. Topical [434]. Topological [1719]. Topology [1634]. Topsy [1080]. Topsy-Turvy [1080]. Torah [454]. Toronto [187, 259]. Toss [1261]. Touchdown [570, 1590]. Tour [1238, 1264, 1284, 983, 1864]. Tournament [1592, 195, 172, 1210, 987, 1587, 1382]. Toxic [170]. TQM [209]. Tracing [1681]. Tracking [1578, 1482, 830]. Trade [1497]. Trading [667]. traditional [1383]. Traffic [866, 1717, 943, 1391]. Trafficking [1828, 1544, 1827, 1829, 1540]. Train [1568, 1314]. Training [1271, 1851, 1306, 1799]. Trait [899]. Trajectories [12, 582]. Trajectory [1065]. Tranche [1375]. Trans [1497]. Trans-Atlantic [1497]. transcription [1020, 1064]. Transformations [27]. Transforming [1724, 1274, 1778]. Transitioning [1050]. Transmission [190]. Transparency [1260, 1522]. Treatment [1645]. Treatments [1660]. Tree [296, 1255]. Treeline [1486]. trees [1102]. Trending [1446]. Trends [541, 540]. Trial [169, 1597, 1098, 625, 1804, 692, 1208, 279]. Trials [796, 978, 1263, 421, 860, 875, 1085, 1647, 422, 1646, 1565, 1648, 982, 1779, 589, 1680]. Triangulated [1543]. Triathlons [242]. Tribulations [1779]. Tribute [412]. Tricks [1326]. Trilinear [682]. Triple [575]. Triumph [1385]. triumphant [1865]. Trouble [17]. troubles [948]. Trout [428]. True [145, 935, 146, 661, 1766]. Truly [130]. Trust [1725, 542, 312, 785, 177]. Truth [1683, 1303, 1572, 1745, 1535, 149, 859]. Truthiness [1572, 1483]. Trygve [78]. Trying [563]. Tsunami [822]. tuned [1460]. Turkey [436, 943]. Turkish [393]. Turn [660]. Turned [1129]. Turnover [724]. Turtles [1485]. Turvy [1080]. Twenty [73, 771, 1523]. Twenty-First [73]. Twenty-One [1523]. Twin [1589]. Twins [965]. Twitter [1446]. Two [731, 1415, 877, 966, 1673, 369, 1494, 1184, 741, 486, 573, 713, 1181, 1865, 445]. Two-Child [1184]. Two-Dimensional [741]. Two-Point [1181]. Two-Way [573]. Two-Word [966]. type [1544].

U.S [4]. U.S. [308, 340, 387, 400, 79, 1728, 1003, 528, 15, 1146, 858]. U.S.A. [439]. Ultimate [54]. Umpire [646]. Uncertain [1558]. Uncertainty [1331, 998, 950, 314, 368, 1321, 1709, 1613]. Uncounted [695, 1240]. Underachievement [1588].

Underclassmen [766]. Undercount

[230, 248]. Underdogs [1385].

Undergraduate [1623]. Undergraduates [972]. Understand [263, 1127]. Understanding [1677, 836, 961, 1265, 1424, 1708, 985, 1517]. Undoing [171]. Unfairness [853]. Unfortunate [874]. unhelpful [638]. Unified [1818]. Uniform [1180]. United [1774, 774, 1545, 1115, 376, 443, 838]. Units [1308]. Universal [1418]. Universe [1635, 1300]. Universities [1003]. University [1247, 1248, 1005]. UNIX [86]. Unknown [1329, 1501]. Unlocking [1539, 1607]. Unmasked [762]. Unofficial [180]. Unpleasant [670]. Unpredictable [671]. Unrelated [1818]. Unsound [472]. Unsupervised [905]. Unsystematic [1704]. Until [990]. Untypical [866]. Unusual [1065, 668]. Update [401, 1288]. Updated [1583]. Uplifting [1860]. Upper [1623]. Upper-level [1623]. Upsets [444]. Urban [1079]. Urns [1428]. Use [91, 584, 814, 1297, 1277, 823, 1844, 538, 1226, 1206, 1594]. Used [97, 1003]. Useful [31, 303]. User [560, 286]. Users [86, 1430]. Uses [405]. USGA [597]. Using [682, 1162, 425, 927, 1822, 935, 621, 1491, 1862, 1127, 1636, 1578, 398, 1116, 1288, 192, 1559, 335, 172, 1844, 1542, 594, 389, 1267, 1405, 1504, 1336, 711, 1560, 972, 1274, 1335, 1497, 561, 892, 932, 692, 1020, 1064, 1102, 1517, 1639]. Utah [771, 730]. utility [781].

v [731, 738, 641]. Vaccine [1110, 579]. Validating [1160, 1070]. Validation [258]. Value [1813, 1513, 1337, 1658, 1038, 1177]. Value-Added [1177]. Values [1823, 1435]. VAMs [1495]. Variables [1575]. Variation [1210, 1166]. Variations [259]. Vasectomy [375]. Veerpalu [1397]. Veggies [414]. Vehicle [870, 1288]. Verify [1450]. Version [287, 330]. Versus [1622, 1600]. veto [1164]. Vetoed [1164]. via [1672]. Vibrant [1805].

Victims [1544]. Video [1193]. Videotape [1517]. Vie [1523]. Vietnam [532]. View [207, 1840, 674, 286, 1728, 1431, 629, 198, 208]. Vintage [322, 641]. Violence [511]. Virtual [459, 893, 1116, 601, 366]. Virtues [1478]. visibility [1517]. Visible [663]. Visiphilia [1403, 1389, 1418]. Visual [1324, 44, 67, 99, 1495, 148, 176, 232, 199, 214, 285, 249, 269, 328, 309, 336, 363, 352, 402, 428, 414, 381, 486, 449, 532, 503, 518, 548, 573, 585, 561, 600, 615, 629, 644, 688, 663, 676, 700, 736, 713, 723, 747, 797, 760, 775, 838, 825, 859, 874, 892, 951, 936, 958, 1027, 1007, 1042, 990, 1056, 1101, 1084, 1069, 1132, 1152, 1167, 1117, 1199, 1212, 1230, 1281, 1245, 1262, 1296, 1341, 1314, 1374, 1426, 1390, 1406, 1451, 1483, 1505, 1547, 1526, 1585, 1631]. Visualization [1654, 1832, 1771, 1405, 1056, 1748, 1672]. Visualizations [1254, 1800, 1585]. Visualizing [1176, 1497, 30]. Visually [417, 1139]. Voice [736]. Voices [1860]. Volleyball [1096]. Volume [288, 331, 435, 491, 535, 588, 647, 701, 750, 800]. Volumes [153, 235, 434]. Volunteer [1529, 1734]. Voodoo [26]. Vote [1512, 418, 916, 1143]. Vote-Tabulation [1143]. Voter [915, 1001]. Votes [1760, 947, 695, 1240, 603, 948]. Voting [595, 372, 680, 1142, 211, 695, 1382, 1115, 681]. vs [781, 1025, 1683, 1647, 1144, 66]. Vulnerability [1541].

W [1481]. W. [1800, 1547]. Wainer [1429]. Wait [379]. Waiting [915, 1296]. Walk [476, 1821]. Wall [709, 556]. Waller [1831]. Wallman [1853]. Walter [1]. Want [1332, 248]. Wants [818, 913, 819, 1775, 817, 914, 963]. War [1033, 1315, 990, 692, 1401, 717, 480]. Warfighter [1577]. Warm [193]. Warming [1551, 1609]. Was [128, 715, 1667, 945, 381, 820]. Washington [628, 327, 308, 353, 365, 340, 387, 419, 427, 400, 464, 474, 450, 502, 531, 517, 560, 584, 614, 599, 643, 674, 686, 697, 661, 749, 724, 735, 763, 790, 799, 814, 843, 268, 284, 180, 974, 277, 65, 43, 98, 147, 175, 213, 231, 198, 248]. Waste [170]. Water [1491, 160, 728]. Wave [1093, 1442]. Way [1081, 22, 779, 1625, 414, 573, 1639]. Ways [1765, 354, 1281]. weak [1125]. Wealth [1302, 1752, 1341]. Weather [322, 644]. Web [1600, 978, 549, 982, 342]. Webinar [1511]. Weight [978, 605, 1341]. Weighted [1693]. Welcome [346, 737, 1814, 441]. Weldon [1092]. Welfare [419, 49]. Well [1822, 1733, 1094]. Well-being [1733]. Were [657]. Western [240]. Where [1192, 1263, 1449, 1548, 1293, 1562, 197, 569, 1471, 1191, 1213, 1226, 1243, 1311, 1447, 1214]. Which [734, 1337, 1182]. Whigmaleeries [46, 66, 100, 149]. While [1206, 889]. WHIP [1065]. Whistle [1284]. Whistle-Stop [1284]. White [292]. Who [818, 999, 716, 378, 1129, 128, 310, 160, 1740, 1512, 715, 1755, 817, 914, 963, 1716, 360, 581, 313, 395, 498, 381, 1101, 229, 1594, 819, 913]. Whom [1584, 1369]. Wide [1509, 342]. Widths [363]. Wild [358, 1667, 1222]. Wild-Card [358]. Wildlife [345]. Wilkinson [1770]. Will [112, 834, 1102, 1197, 315, 569, 596, 444, 65, 1291, 1056]. Willard [1825]. William [380, 486, 975, 1357]. Williams [137]. Wilson [1852]. Win [891, 1211, 1593, 926]. WinBUGS [1267]. Window [628, 327, 308, 353, 365, 340, 387, 419, 427, 400, 464, 474, 450, 502, 531, 517, 572, 560, 584, 614, 599, 643, 674, 686, 697, 661, 749, 724, 735, 763, 790, 799, 814, 843, 268, 284, 65, 43, 98, 147, 175, 213, 231, 198, 248, 475, 441]. Windows [330]. Winds [340, 644]. Wine [322, 512]. Winner [1822, 1706, 1040]. Winners [108, 568, 854, 1068]. Winning [508, 841, 876, 1371, 439, 443, 1703, 555]. Wins [96, 1148, 1313, 1385]. Winter [667]. Wisdom [1573, 1199]. Wise [531]. Wit [1028, 1044, 1058, 1071, 1168, 1185, 1200, 1215, 1235, 1251, 1268, 1286, 1304, 1316, 1331, 1348, 1366, 1380, 1086, 1103, 1119, 1135, 1153]. Without [962, 1361, 1412, 1726, 1742, 783]. Witnesses [25]. Wittes [1410]. Wizard [410]. Woefully [1503, 1866]. Women [112, 1589, 1025, 1127, 170, 1151, 1415, 1408, 1416, 1207, 1533]. Won [999, 934]. Word [1675, 648, 564, 877, 966, 1668]. Words [1300]. Work [971, 1472, 984, 64, 114]. worker [1733]. Workforce [1813]. Working [1562, 1414]. Workplace [1430]. Works [1667, 654, 669, 655]. World [187, 1674, 572, 1194, 1025, 1468, 1385, 1329, 342, 1351, 2, 1742, 195, 1330, 1335]. Worth [1212, 755]. Worthless [380]. Would [1241, 669, 1247, 1865]. Wranglers [1318]. Wringing [317]. Writing [919, 709]. Wrong [1503, 116, 33, 1866]. Wrote [716].

xiv [1247]. xlv [1250]. xxiv [1248].

Yahoo [569]. Yahtzee [706]. Yale [1247]. Yankees [657]. Year [1591, 1238, 1839, 480, 147, 1544, 1391, 175, 248]. Years [1271, 502, 155, 1006, 771, 1276, 1649, 50, 569, 973]. Yeast [903]. Yellow [292]. Yonder [22]. York [1081]. Young [1845, 1035]. Yourself [1366]. Ysidro [471].

Z [1419]. Zealand [461]. Zero [1300, 936, 1262]. Zika [1498]. Zipf [1761]. Zone [506]. Zones [646]. Zulu [933].

References

Editors:1988:MWK

[1] Editors. In memoriam: Walter Kaufmann Buhler. Chance, 1(1):8, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1988:PMW

[2] Howard Wainer and David Thissen. Plotting in the modern world: Statistics packages and good graphics. Chance, 1(1):10–20, Winter 1988. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Madansky:1988:GTP

[3] Albert Madansky. The gentleman tasting pastrami. Chance, 1(1):22–26, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor,” Vol. 1, No. 3, 7.

Bounpane:1988:AUC

[4] Peter A. Bounpane and Thomas A. Jones. Automation of the 1990 U.S. Census. Chance, 1(1):28–35, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hill:1988:SRF

[5] William J. Hill and William G. Hunter. Statistics: A road to the future — prologue. Chance, 1(1):38–39, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Pfeifer:1988:SRF

[6] Charles G. Pfeifer, Donald W. Marquardt, and Ronald D. Snee. Statistics: A road to the future — a time for change. Chance, 1(1):39–41, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McDonald:1988:SRF

[7] Gary C. McDonald. Statistics: A road to the future — communicating with management. Chance, 1 (1):42–44, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Duncan:1988:SRF

[8] Joseph W. Duncan. Statistics: A road to the future — government economic statistics. Chance, 1(1):45, 52, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Joiner:1988:SRF

[9] Brian L. Joiner. Statistics: A road to the future — let’s change how we teach statistics. Chance, 1(1):53–54, Winter 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stone:1988:QPS

[10] Mervyn Stone. Quetelet and the poetry of statistical conjecture. Chance, 1(2):10–17, Spring 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Grier:1988:SMC

[11] David Alan Grier. Supercomputers and Monte Carlo experiments. Chance, 1(2):19–28, Spring 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Moore:1988:PIT

[12] Marc Moore. Predicting iceberg trajectories. Chance, 1(2):30–36, Spring 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Proschan:1988:ILS

[13] Frank Proschan. Investigation of latin squares. Chance, 1(2):37–41, Spring 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dalenius:1988:DPS

[14] Tore Dalenius. The debate on privacy and surveys in Sweden. Chance, 1(2):43–47, Spring 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mann:1988:WIH

[15] Nancy R. Mann. Why it happened in Japan and not in the U.S. Chance, 1(3):8–15, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Duncan:1988:SAA

[16] Otis Dudley Duncan. Some ancient anticipations of probability. Chance, 1(3):16–24, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Streitfeld:1988:SHT

[17] David Streitfeld. Shere Hite and the trouble with numbers. Chance, 1 (3):26–31, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor” Vol. 2, No. 2, 8.

Cliff:1988:MBA

[18] Norman Cliff. A model-based analysis. Chance, 1(3):32–36, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 2, No. 1, 8.

Ehrenberg:1988:DRP

[19] A. S. C. Ehrenberg. Data reduction and prior knowledge. Chance, 1(3):37–42, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Emerson:1988:ESA

[20] John D. Emerson, David C. Hoaglin, John W. Tukey, and George Y. Wong. Exploring some adverb–adjective data. Chance, 1(3):42–48, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See comment [21].

Tanur:1988:CAA

[21] Judith M. Tanur. Comment on adverb–adjective analyses. Chance, 1 (3):48–50, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [20].

Belizaire:1988:WYN

[22] Belizaire. Way down yonder in New Orleans. Chance, 1(3):52–54, Summer 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

DuMouchel:1988:REE

[23] William DuMouchel. Rs/explore: The electronic data analysis advisor. Chance, 1(4):12–22, Fall 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Greenhouse:1988:AD

[24] Joel B. Greenhouse and Samuel W. Greenhouse. An aspirin a day …? Chance, 1(4):24–31, Fall 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1988:SEW

[25] Stephen E. Fienberg, Samuel Krislov, and Miron L. Straf. Statistics, expert witnesses, and the courts. Chance, 1(4):32–37, Fall 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fiske:1988:STS

[26] Edward B. Fiske. Standardized test scores: Voodoo statistics for the nation. Chance, 1(4):38–39, Fall 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoaglin:1988:TEE

[27] David C. Hoaglin. Transformations in everyday experience. Chance, 1(4):40– 45, Fall 1988. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 2, No. 2, 8.

Frosch:1989:TCI

[28] Robert A. Frosch. Taking chances in industry and in government. Chance, 2(1):11–15, Winter 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tversky:1989:CFA

[29] Amos Tversky and Gilovich Thomas. The cold facts about the ‘hot hand’ in basketball. Chance, 2(1):16–21, Winter 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Young:1989:VSD

[30] Forrest W. Young. Visualizing sixdimensional structure with dynamic statistical graphics. Chance, 2(1):22– 30, Winter 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1989:UCS

[31] Jane F. Gentleman. A useful classification system for statistical presentations and conference papers. Chance, 2(1):31–32, Winter 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 2, No. 3, 7.

Finkelstein:1989:CPN

[32] Michael O. Finkelstein and Bruce Levin. The case of the purloined notices & answers. Chance, 2(1):33, 54, Winter 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Shanker:1989:WWP

[33] Albert Shanker. What’s wrong with this picture? Lies, damned lies, and statistics. Chance, 2(1):34–35, Winter 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1989:BMa

[34] Anonymous. Chance musings. Chance, 2(2):4–7, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1989:LEa

[35] Anonymous. Letters to the Editors. Chance, 2(2):8, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stoto:1989:OSA

[36] Michael A. Stoto. Opinion: Statistics for the AIDS epidemic. Chance, 2(2):9, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gould:1989:SS

[37] Stephen Jay Gould. The streak of streaks. Chance, 2(2):10–16, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 3, No. 3, 6.

Berger:1989:IGO

[38] Marc A. Berger. Images generated by orbits of 2-D Markov chains. Chance, 2(2):18–28, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Humphrey:1989:SMS

[39] Watts S. Humphrey. Statistically managing the software process. Chance, 2 (2):30–39, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:1989:SA

[40] Jonathan Berkowitz. Statistical acrostic. Chance, 2(2):40, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Abelson:1989:BP

[41] Robert P. Abelson. A Bayesball problem. Chance, 2(2):41–43, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1989:CPD

[42] Belizaire. A capital place for dining. Chance, 2(2):44–46, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1989:WWR

[43] Katherine K. Wallman and Vincent P. Barabba. Window on Washington: Relevant Federal statistics and the public interest. Chance, 2(2):47–49, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Paller:1989:VRSa

[44] Alan Paller. Visual revelations: A short cut to becoming a competent speaker: Part 1. Chance, 2(2):50–51, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schervish:1989:PCS

[45] Mark J. Schervish. Personal crunching: Statgraphics. Chance, 2(2):53–58, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vaisrub:1989:SWC

[46] Naomi Vaisrub. Statistical whigmaleeries: Can these data be saved? Chance, 2(2):59, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1989:BRa

[47] Michael M. Meyer. Book reviews. Chance, 2(2):60–62, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1989:LSO

[48] Miron L. Straf. On the lighter side: Opinion surveys. Chance, 2(2):64, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fairley:1989:WQC

[49] William B. Fairley and Alan J. Izenman. Welfare quality control programs: How much is misspent by the welfare system? Chance, 2(3):10–17, Summer 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mason:1989:CYS

[50] Robert L. Mason. Celebrating 150 years of statistics in America. Chance, 2(3):18–22, Summer 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1989:AC

[51] Stephen E. Fienberg. An adjusted census in 1990? Chance, 2(3):23–25, Summer 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Norwood:1989:ESH

[52] Janet L. Norwood. The employment situation: How do we know about it? Chance, 2(3):26–29, Summer 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Smykla:1989:SP

[53] Margaret L. Smykla. The steroid puzzle. Chance, 2(3):30–35, Summer 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoffman:1989:SUB

[54] Tony Hoffman. The search for the ultimate baseball statistic. Chance, 2(3):37–39, Summer 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Solow:1989:IGS

[55] Andrew R. Solow. Is it getting stuffy in here, or is it just my imagination? Chance, 2(3):40–46, Summer 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1989:BMb

[56] Anonymous. Chance musings. Chance, 2(4):3–6, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1989:LEb

[57] Anonymous. Letters to the Editors. Chance, 2(4):7, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1989:OSP

[58] Stephen E. Fienberg. Opinion: Science and the popular Press: From 60 Hertz electrical and magnetic fields to the cholesterol myth. Chance, 2(4):8, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Minton:1989:MDS

[59] Paul Minton. Music by design: A sesquicentennial statistics songbook. Chance, 2(4):9–11, Fall 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morgan:1989:EMF

[60] M. Granger Morgan. Electric and magnetic fields from 60 Hz electric power: Possible health risks? Chance, 2 (4):12–20, 37, Fall 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larkey:1989:OBH

[61] Patrick D. Larkey, Richard A. Smith, and Joseph B. Kadane. It’s okay to believe in the “hot hand”. Chance, 2(4): 22–30, Fall 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See comment [62].

Tversky:1989:HHS

[62] Amos Tversky and Thomas Gilovich. The “hot hand”: Statistical reality or cognitive illusion? Chance, 2(4):31–34, Fall 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [61].

Hooke:1989:BBN

[63] Robert Hooke. Basketball, baseball, and the null hypothesis. Chance, 2(4): 35–37, Fall 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Porter:1989:PDA

[64] Henry Porter. Profile: A decision analyst reflects on his work. Chance, 2(4): 38–40, Fall 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1989:WWF

[65] Katherine K. Wallman. Window on Washington: Federal statistical policy: Will reauthorization bring renewal? Chance, 2(4):41–47, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vaisrub:1989:SWS

[66] Naomi Vaisrub. Statistical whigmaleeries: Statistical savvy vs. marital harmony. Chance, 2(4):45–47, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Paller:1989:VRSb

[67] Alan Paller. Visual revelations: A short cut to becoming a competent speaker: Part 2. Chance, 2(4):46–47, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schervish:1989:PCD

[68] Mark J. Schervish. Personal crunching: Data Desk Professional. Chance, 2(4): 48–52, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1989:BRb

[69] Michael M. Meyer. Book reviews. Chance, 2(4):53–54, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1989:LSP

[70] Miron L. Straf. On the lighter side: Popular statistics. Chance, 2(4):56, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1989:EBP

[71] Anonymous. Editorial Board page. Chance, 2(4):ebi, 1989. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Editors:1990:MMH

[72] Editors. A memoriam: Morris H. De-Groot: 1931–1989. Chance, 3(1):11, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anderson:1990:ATR

[73] Margo Anderson. “According to their respective numbers… ” for the twentyfirst time. Chance, 3(1):12–18, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1990:ACI

[74] Stephen E. Fienberg. An adjusted census in 1990? An interim report. Chance, 3(1):19–21, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barnard:1990:FR

[75] George Barnard. Fisher: A retrospective. Chance, 3(1):22–28, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lindley:1990:FR

[76] D. V. Lindley. Fisher: a retrospective. Chance, 3(1):31–32, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Editors:1990:CSS

[77] Editors. Comments from selected statisticians on Fisher. Chance, 3 (1):29–32, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zellner:1990:THN

[78] Arnold Zellner. Trygve Haalvelmo, 1989 Nobel Laureate in Economic Science. Chance, 3(1):33–35, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fuerbringer:1990:RSC

[79] Jonathan Fuerbringer. A rusty statistical compass for U.S. policy mappers. Chance, 3(1):36–38, 56, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1990:SIG

[80] Belizaire. 47th session of the ISI: Gastronomie et Paris. Chance, 3 (1):39–42, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barnett:1990:ASE

[81] Arnold Barnett. Air safety: End of the Golden Age? Chance, 3(2):8–12, Spring 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Goldfield:1990:WWE

[82] Edwin D. Goldfield. What we expect to learn from the 1990 Decennial Census. Chance, 3(2):13–19, Spring 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoffman:1990:CEW

[83] Tony Hoffman. Chatting with extraterrestrials — what are the odds? Chance, 3(2):20–31, Spring 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1990:ACBa

[84] Stephen E. Fienberg. An adjusted census in 1990? Back to court again. Chance, 3(2):32–35, Spring 1990. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Spisak:1990:MPE

[85] Andrew W. Spisak. Modeling the 1988 Presidential Election. Chance, 3(2):36–41, Spring 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eddy:1990:HRD

[86] William F. Eddy. Hardware review: The DECstation 3100 — UNIX for power users. Chance, 3(2):42–47, Spring 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1990:SPC

[87] Belizaire. The statistical pleasures of California Cabernets. Chance, 3 (2):48–50, Spring 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1990:BM

[88] Anonymous. Chance musings. Chance, 3(3):3–5, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1990:LE

[89] Anonymous. Letters to the Editors. Chance, 3(3):6–7, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Geisser:1990:OSR

[90] Seymour Geisser. Opinion: Some remarks on DNA fingerprinting. Chance, 3(3):8–9, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Boland:1990:USE

[91] Philip J. Boland and Michael Proschan. The use of statistical evidence in allegations of exam cheating. Chance, 3(3):10–14, Summer 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:1990:DFW

[92] Donald A. Berry. DNA fingerprinting: What does it prove? Chance, 3 (3):15–25, 36, Summer 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cohen:1990:DFW

[93] Joel E. Cohen. DNA fingerprinting: What (really) are the odds? Chance, 3(3):26–32, Summer 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1990:ACBb

[94] Stephen E. Fienberg. An adjusted census in 1990? Back to court again. Chance, 3(3):32–35, Spring 1990. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1990:ACJ

[95] Stephen E. Fienberg. An adjusted census in 1990? The judge rules and the PES begins. Chance, 3 (3):33–36, Summer 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1990:SDR

[96] Anonymous. Sir David R. Cox wins Kettering Prize. Chance, 3(3):37, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hartigan:1990:AMG

[97] J. A. Hartigan. Adjustments of minority group test scores used in employment referrals. Chance, 3(3):38– 44, Summer 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1990:WWF

[98] Katherine K. Wallman, George T. Duncan, and Virginia A. de Wolf. Window on Washington: The Federal statistical system amidst the rising tension between privacy and data access. Chance, 3(3):45–48, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Paller:1990:VRQ

[99] Alan Paller. Visual revelations: Questions every EIS implementor should be able to answer. Chance, 3(3):49–56, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vaisrub:1990:SWA

[100] Naomi Vaisrub. Statistical whigmaleeries: Amateur hour. Chance, 3 (3):52–56, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schervish:1990:PC

[101] Mark J. Schervish. Personal crunching. Chance, 3(3):53–56, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1990:BRa

[102] Michael M. Meyer. Book reviews. Chance, 3(3):58–61, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1990:LSSa

[103] Miron L. Straf. On the lighter side: Shlepping statisticians to Siberia. Chance, 3(3):62–63, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1990:EBPa

[104] Anonymous. Editorial Board page. Chance, 3(3):ebi, 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Brookmeyer:1990:SHA

[105] Ron Brookmeyer and Mitchell H. Gail. A statistical history of the AIDS epidemic. Chance, 3(4):9–14, Fall 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1990:BBH

[106] David Bellhouse. A Bayesian busman’s holiday. Chance, 3(4):15–17, 47, Fall 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dutka:1990:MAS

[107] Solomon Dutka and Lester R. Frankel. Misuses and abuses of statistical techniques in market research surveys. Chance, 3(4):18–23, Fall 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larkey:1990:FBW

[108] Patrick D. Larkey. Fair bets on winners in professional golf. Chance, 3(4):24– 26, Fall 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meier:1990:EER

[109] Paul Meier. The experimental evaluation of relative risk. Chance, 3(4):28– 32, Fall 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lave:1990:DSG

[110] Lester B. Lave. Does the Surgeon-General need a statistics advisor? Chance, 3(4):33–40, Fall 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ware:1990:REA

[111] James H. Ware. The role of epidemiology in the assessment of societal risk: A statistician’s perspective. Chance, 3(4):41–46, Fall 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cherlin:1991:WSW

[112] Andrew Cherlin. Will single women ever marry? The strange career of a statistical projection. Chance, 4 (1):11–15, Winter 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Janssen:1991:GG

[113] Christian T. L. Janssen and Robert W. Schutz. The Great Gretzky. Chance, 4(1):16–21, Winter 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Simon:1991:RTE

[114] Julian L. Simon and Peter C. Bruce. Resampling: A tool for everyday statistical work. Chance, 4(1):22–32, Winter 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also, “Letters to the Editor”, Vol. 4, No. 3, 7, 29.

Iyengar:1991:MAA

[115] Satish Iyengar. Much ado about metaanalysis. Chance, 4(1):33–40, Winter 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rousseeuw:1991:WWP

[116] Peter J. Rousseeuw. Why the wrong papers get published. Chance, 4 (1):41–43, Winter 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

DeGroot:1991:BN

[117] M. H. DeGroot and B. Natvig. Bayesians in the night. Chance, 4(1): 45, Winter 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McCown:1991:LCS

[118] William G. McCown, Judith Johnson, and Z. Harry Galina. Last chance: Statistics and cognitive psychotherapy. Chance, 4(1):46–47, Winter 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Engel:1991:MHP

[119] Eduardo Engel and Achilles Venetoulias. Monty Hall’s probability puzzle. Chance, 4(2):6–9, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lindley:1991:SHJ

[120] D. V. Lindley. Sir Harold Jeffreys. Chance, 4(2):10–14, 21, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bolt:1991:SHJ

[121] Bruce A. Bolt. Sir Harold Jeffreys and geophysical inverse problems. Chance, 4(2):15–17, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Huzurbazar:1991:SHJ

[122] Vasant S. Huzurbazar. Sir Harold Jeffreys: Recollections of a student. Chance, 4(2):18–21, Spring 1991. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jeffreys:1991:HJS

[123] Lady Jeffreys. Harold Jeffreys: Some reminiscences. Chance, 4(2):22–23, 26, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Knopoff:1991:SHJ

[124] Leon Knopoff. Sir Harold Jeffreys: The Earth: Its origin, history, and physical constitution. Chance, 4(2):24–26, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bailar:1991:SIE

[125] John C. Bailar III. Scientific inferences and environmental health problems. Chance, 4(2):27–38, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Grove:1991:SBL

[126] M. A. Grove. Survival on the bestseller list. Chance, 4(2):39–45, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Forthofer:1991:SSS

[127] Ronald Forthofer. Streak shooter — the sequel. Chance, 4(2):46–48, Spring 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Elliott:1991:WWS

[128] Ward Elliott and Robert Valenza. Who was Shakespeare? Chance, 4(3):8–14, Summer 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bowman:1991:SQC

[129] K. O. Bowman, T. H. Hopp, R. N. Kacker, and R. J. Lundegard. Statistical quality control technology in Japan. Chance, 4(3):15–21, Summer 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kadell:1991:LPG

[130] Dan Kadell and Donald Ylvisaker. Lotto play: The good, the fair, and the truly awful. Chance, 4(3):22– 25, 57, Summer 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Young:1991:PS

[131] David A. Young, R. W. Hoyer, Colin Mallows, and Miriam D. Rosenthal. Poetry for statisticians. Chance, 4 (3):26–28, Summer 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Grier:1991:SID

[132] David Alan Grier. Statistics and the introduction of digital computers. Chance, 4(3):30–36, Summer 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Chatterjee:1991:PBS

[133] Sangit Chatterjee and Mustafa Yilmaz. Parity in baseball: Stability of evolving systems? Chance, 4(3):37–42, Summer 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1991:ACC

[134] Stephen E. Fienberg. An adjusted census in 1990? Commerce says ‘no’. Chance, 4(3):44–52, Summer 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1991:CM

[135] Anonymous. Chance musings. Chance, 4(4):3–6, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eddy:1991:LE

[136] William F. Eddy and Stephen E. Fienberg. Letter from the Editors. Chance, 4(4):7, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:1991:SPA

[137] Scott Berry. The summer of ’41: A probabilistic analysis of DiMaggio’s ‘streak’ and Williams’s average of .406. Chance, 4(4):8–11, Fall 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eddy:1991:QSS

[138] William F. Eddy and Lawrence H. Cox. The quality of statistical software: Controversy in the statistical software industry. Chance, 4(4):12–18, Fall 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evett:1991:FRC

[139] Ian W. Evett and Bruce S. Weir. Flawed reasoning in court. Chance, 4(4):19–21, Fall 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1991:ACF

[140] Stephen E. Fienberg. An adjusted census in 1990? A full-scale judicial review approaches. Chance, 4 (4):22–24, 29, Fall 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Boomsma:1991:RMC

[141] Anne Boomsma and Ivo W. Molenaar. Resampling with more care. Chance, 4(4):25–29, Fall 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See reply [142]. See also “Letters to the Editor”, Vol. 5, No. 3–4, 4, 7.

Simon:1991:RBM

[142] Julian L. Simon and Peter C. Bruce. Reply to Boomsma and Molenaar. Chance, 4(4):30–40, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [141].

Williams:1991:DS

[143] W. H. Williams. Dispensable statistics. Chance, 4(4):32–33, Fall 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Janssen:1991:RDS

[144] Christian T. L. Janssen and Robert W. Schutz. Response to “dispensable statistics”. Chance, 4(4):33–35, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1991:SRL

[145] Anonymous. A statistical return to the land of true gastronomic pleasures “Belizaire”. Chance, 4(4):36–40, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1991:SRL

[146] Belizaire. A statistical return to the land of true gastronomic pleasures. Chance, 4(4):36–40, Fall 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1991:WWM

[147] Katherine K. Wallman, Mary Anne Freedman, and Manning Feinleib. Window on Washington: Monitoring the Nation’s health objectives for the year 2000. Chance, 4(4):41–44, 1991. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1991:VRE

[148] Howard Wainer. Visual revelations: Elegance, grace, impact, and graphical displays. Chance, 4(4):45–47, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vaisrub:1991:SWP

[149] Naomi Vaisrub. Statistical whigmaleeries: The probable truth. Chance, 4(4):48–53, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schervish:1991:PCE

[150] Mark J. Schervish. Personal crunching: EasyPlot. Chance, 4(4):50–53, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1991:BRb

[151] Michael M. Meyer. Book reviews. Chance, 4(4):54–56, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1991:LSSb

[152] Miron L. Straf. On the lighter side: Statistical self-help. Chance, 4(4):60– 56, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1991:BCI

[153] Anonymous. Chance cumulative index: Volumes 1–4. Chance, 4(4):61–64, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1991:EBPb

[154] Anonymous. Editorial Board page. Chance, 4(4):ebi, 1991. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Editors:1992:FHY

[155] Editors. Five hundred years of scrutinity. Chance, 5(1–2):36–37, 39, Winter– Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Editors:1992:PDM

[156] Editors. Pico della Mirandola: A. philosophical forefather of applied statistics? Chance, 5(1–2):40– 42, Winter–Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Safran:1992:FN

[157] Charles Safran. The Fibonacci numbers. Chance, 5(1–2):43–46, Winter– Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Goldberger:1992:COH

[158] Ary L. Goldberger and Bruce J. West. Chaos and order in the human body. Chance, 5(1–2):47–55, Winter–Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Editors:1992:CBD

[159] Editors. Charles Babbage describes his calculating engines. Chance, 5(1–2): 56–62, Winter–Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Geller:1992:PPF

[160] B. Geller and Y. Bruk. A portrait of Poisson: A fish out of water who found his calling. Chance, 5(1–2): 63–67, Winter–Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vavilov:1992:CP

[161] Valery Vavilov. Calculating pi. Chance, 5(1–2):73–77, Winter–Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dubrovsky:1992:WP

[162] Vladimir Dubrovsky. What is pi? Chance, 5(1–2):78–79, Winter–Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1992:LE

[163] Anonymous. Letter from the Editor. Chance, 5(3–4):3, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Haider:1992:LE

[164] Trond Haider and T. H. Wonnacott. Letters to the Editor. Chance, 5(3– 4):4–7, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1992:BM

[165] Anonymous. Chance musings. Chance, 5(3–4):8–11, Summer–Fall 1992. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Snell:1992:CCC

[166] J. Laurie Snell and John Finn. A course called “Chance”. Chance, 5(3– 4):12–16, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cooper:1992:PPS

[167] Harris Cooper, Kristina M. Deneve, and Frederick Mosteller. Predicting professional sports game outcomes from intermediate game scores. Chance, 5(3–4):18–22, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Klein:1992:SEC

[168] Stephen P. Klein. Statistical evidence of cheating on multiple-choice tests. Chance, 5(3–4):23–27, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1992:ACT

[169] Stephen E. Fienberg. An adjusted census in 1990? The trial. Chance, 5(3– 4):28–38, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Geschwind:1992:SPW

[170] Sandra A. Geschwind. Should pregnant women move? Linking risks for birth defects with proximity to toxic waste sites. Chance, 5(3–4):40– 45, 86, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fortney:1992:UMG

[171] William G. Fortney, Steven Hotovy, Fritz Klein, and Fritz Scholz. The undoing of Maurice G. Kendall. Chance, 5(3–4):46–48, Summer–Fall 1992. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also, “Letters to the Editor”, Vol. 6, No. 1, 4.

Ladwig:1992:UPS

[172] J. A. Ladwig and N. C. Schwertman. Using probability and statistics to analyze tournament competitions. Chance, 5(3–4):49–53, Summer–Fall 1992. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also, “Letters to the Editor”, Vol. 6, No. 2, 3.

Allen:1992:ERB

[173] Terry Allen and Glen Buckner. Expanding the role of the bar chart in representing crime data. Chance, 5(3– 4):54–59, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Roberts:1992:SCC

[174] Leslie Roberts. Science in court: A culture clash. Chance, 5(3–4): 60–66, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1992:WWT

[175] Katherine K. Wallman. Window on Washington: Toward the Year 2000 Census. Chance, 5(3–4):67–70, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1992:VRS

[176] Howard Wainer. Visual revelations: Sense-lining. Chance, 5(3–4):71–72, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1992:ESC

[177] Lionel Galway. On the edge: Statistics and computing: Nuts and bolts: How much can you trust your hardware? Chance, 5(3–4):73–75, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morton:1992:PCN

[178] Sall C. Morton. Personal crunching: New advances in statistical dendrology. Chance, 5(3–4):76–79, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1992:BR

[179] Michael M. Meyer. Book reviews. Chance, 5(3–4):80–83, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1992:LSW

[180] Miron L. Straf. On the lighter side: The Washington statistical scene: The unofficial guide. Chance, 5(3–4): 88–70, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1992:EBP

[181] Anonymous. Editorial Board page. Chance, 5(3–4):ebi, Summer–Fall 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mitofsky:1993:SSE

[182] Warren J. Mitofsky. The state of state election polls. Chance, 6(1):9–16, Winter 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1993:HMM

[183] Howard Wainer. How much more efficiently can humans run than swim? Chance, 6(1):17–21, Winter 1993. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ansolabehere:1993:PFQ

[184] Stephen Ansolabehere and Thomas R. Belin. Poll faulting. Chance, 6 (1):22–27, Winter 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also, “Letters to the Editor”, Vol. 6, No. 3, p. 3 and Vol. 7, No. 1, p. 3.

Royko:1993:BTR

[185] Mike Royko. Blood test ‘results’ already legendary. Chance, 6(1):28–29, Winter 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

OGeran:1993:MTB

[186] J. H. O’Geran, H. P. Wynn, and A. A. Zhiglyavsky. Mastermind as a testbed for search algorithms. Chance, 6(1):31–37, Winter 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1993:TWS

[187] Belizaire. Toronto: A world serious city. Chance, 6(1):38–40, Winter 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jackson:1993:LE

[188] David A. Jackson, Neil C. Schwertman, and Jack A. Ladwig. Letter to the Editor. Chance, 6(2):3, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1993:CMa

[189] Anonymous. Chance musings. Chance, 6(2):5–8, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaplan:1993:WHH

[190] Edward H. Kaplan and Robert Heimer. What happened to HIV transmission among drug injectors in new haven? Chance, 6(2):9–14, Spring 1993. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ratzan:1993:LMC

[191] Lee Ratzan. L’homme moyen (the canonical man). Chance, 6(2):15–16, Spring 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

James:1993:AQA

[192] Bill James, Jim Albert, and Hal S. Stern. Answering questions about baseball using statistics. Chance, 6(2): 17–22, 30, Spring 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Strauss:1993:SDN

[193] Stephen Strauss. Statscan devises a new math for counting warm bodies. Chance, 6(2):23, 49, Spring 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morrison:1993:MTM

[194] Peter A. Morrison. More than meets the eye. Chance, 6(2):24–30, Spring 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hurley:1993:WST

[195] William Hurley. What sort of tournament should the World Series be? Chance, 6(2):31–33, Spring 1993. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1993:SFF

[196] Belizaire. A San Francisco feast. Chance, 6(2):34–36, Spring 1993. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mosteller:1993:WEF

[197] Frederick Mosteller and Cleo Youtz. Where eagles fly. Chance, 6(2):37–42, Spring 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1993:WWV

[198] Katherine K. Wallman. Window on Washington: The view beyond. Chance, 6(2):43–45, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1993:VRM

[199] Howard Wainer. Visual revelations: Making readable overhead displays. Chance, 6(2):46–49, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1993:ESCa

[200] Lionel Galway. On the edge: Statistics and computing: What good is a benchmark? Chance, 6(2):50–53, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Adams:1993:PCL

[201] John Adams. Personal crunching: I have looked at clouds from both sides now. Chance, 6(2):54–58, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1993:BRa

[202] Michael M. Meyer. Book reviews. Chance, 6(2):59–61, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1993:LSH

[203] Miron L. Straf. On the lighter side: Having a difficult time adjusting. Chance, 6(2):64–58, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1993:LE

[204] Anonymous. Letter to the Editors. Chance, 6(3):3, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1993:CMb

[205] Anonymous. Chance musings. Chance, 6(3):4–7, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1993:PRN

[206] Anonymous. The proportionality review of new Jersey death sentences. Chance, 6(3):8, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Baldus:1993:PVS

[207] David Baldus and George Woodworth. Proportionality: The view of the special master. Chance, 6(3):9–17, Summer 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Weisberg:1993:PAV

[208] Herbert I. Weisberg. Proportionality: An alternative view. Chance, 6(3):18–24, Summer 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Roberts:1993:CTI

[209] Harry V. Roberts. Can TQM improve athletic performance? Chance, 6(3):25–29, Summer 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morrison:1993:BNF

[210] Donald G. Morrison and Manohar U. Kalwani. The best NFL field goal kickers: Are they lucky or good? Chance, 6(3):30–37, Summer 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Klein:1993:ERV

[211] Stephen P. Klein and David A. Freedman. Ecological regression in voting rights cases. Chance, 6(3):38–43, Summer 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bryant:1993:GCC

[212] Barbara Everitt Bryant. Guest commentary: Census-taking for a litigious, data-driven society. Chance, 6(3):44– 49, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1993:WWE

[213] Katherine K. Wallman and Stanley Presser. Window on Washington: To enhance and enrich. Chance, 6(3):50– 51, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1993:VRT

[214] Howard Wainer. Visual revelations: Tabular presentation. Chance, 6(3):52– 56, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1993:ESCb

[215] Lionel Galway. On the edge: Statistics and computing: John von Neumann and the origin of pseudo-random number generators. Chance, 6(3):57–59, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Marsh:1993:PCS

[216] Paul Marsh. Personal crunching: Stanford Graphics. Chance, 6(3):60–65, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1993:BRb

[217] Michael M. Meyer. Book reviews. Chance, 6(3):66–69, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1993:LSM

[218] Miron L. Straf. On the lighter side: More journals! Chance, 6(3):72–69, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rolph:1993:LE

[219] John E. Rolph. Letter from the Editor. Chance, 6(4):3, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bienen:1993:LE

[220] Leigh B. Bienen and Douglas L. Mills. Letters to the Editor. Chance, 6(4):4–9, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1993:CMc

[221] Anonymous. Chance musings. Chance, 6(4):5–7, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vaisrub:1993:GT

[222] Naomi Vaisrub. The gambler’s tale. Chance, 6(4):10, 37, Fall 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gross:1993:DES

[223] Alan J. Gross. Does exposure to second-hand smoke increase lung cancer risk? Chance, 6(4):11–14, Fall 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 7, No. 1, 4; No. 2, 4–5; No. 3, 3.

Rockette:1993:WEN

[224] Howard E. Rockette. What evidence is needed to link lung cancer and secondhand smoke? Chance, 6(4):15–18, Fall 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 7, No. 1, 4; No. 2, 4–5; No. 3, 3.

Reidpath:1993:CDC

[225] Daniel D. Reidpath, Mark R. Diamond, and K. Vijayan. Catching drug couriers. Chance, 6(4):19–23, 30, Fall 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Buoncristiani:1993:FAR

[226] John F. Buoncristiani and David E. Martin. Factors affecting runners’ Marathon performance. Chance, 6(4): 24–30, Fall 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

MacCoun:1993:BF

[227] Robert Maccoun. Blaming others to a fault? Chance, 6(4):31–33, Fall 1993.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Passell:1993:LND

[228] Peter Passell. Like a new drug, social programs are put to the test. Chance, 6(4):34–36, 57, Fall 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Woemer:1993:WRS

[229] Denise Woemer. Who really said it? Chance, 6(4):37, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ericksen:1993:GCB

[230] Eugene P. Ericksen and Teresa K. Defonso. Guest commentary: Beyond the net undercount: How to measure census error. Chance, 6(4):38–??, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1993:WWF

[231] Katherine K. Wallman, Stephen E. Fienberg, Margaret E. Martin, and Miron L. Straf. Window on Washington: Fostering interactions between statisticians in academia and government. Chance, 6(4):44–47, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1993:VRG

[232] Howard Wainer. Visual revelations: Graphical answers to scientific questions. Chance, 6(4):48–50, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1993:ESCc

[233] Lionel Galway. On the edge: Statistics and computing: Statistics at Stonehenge. Chance, 6(4):51–53, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Greenhouse:1993:BR

[234] Joel B. Greenhouse and Michael M. Meyer. Book reviews. Chance, 6(4): 54–57, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1993:BCI

[235] Anonymous. Chance cumulative index: Volumes 5–6. Chance, 6(4):61–63, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1993:LSA

[236] Miron L. Straf. On the lighter side: The academiks are coming! The academiks are coming! Chance, 6(4):64–63, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1993:EBP

[237] Anonymous. Editorial Board page. Chance, 6(4):ebi, 1993. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lachenbruch:1994:LE

[238] Peter A. Lachenbruch, Thomas R. Belin, Stephen Ansolabehere, Dorothy P. Rice, and Stanton A. Glantz. Letters to the Editor. Chance, 7(1):3–4, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1994:CMa

[239] Anonymous. Chance musings. Chance, 7(1):5–8, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sturm:1994:DSN

[240] Roland Sturm. Do Soviet nuclear reactors outperform Western reactors? Chance, 7(1):9–16, 27, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Greiner:1994:BBG

[241] Keith Greiner. The Baby Boom Generation and how they grow. Chance, 7(1):17–19, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1994:RTF

[242] Howard Wainer and Richard D. De Veaux. Resizing triathlons for fairness. Chance, 7(1):20–25, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vaisrub:1994:GSC

[243] Naomi Vaisrub. Guidelines for statistical consulting. Chance, 7(1):26–27, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hand:1994:SES

[244] David J. Hand. Statistical expert systems. Chance, 7(1):28–31, 34, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossman:1994:TTS

[245] Allan J. Rossman. “techno-thriller” statistics: Chance in the fiction of Michael crichton. Chance, 7(1):32–34, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Plewes:1994:FAI

[246] Thomas J. Plewes. Federal agencies introduce redesigned current population survey. Chance, 7(1):35–41, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zorn:1994:MRP

[247] Eric Zorn. Math riots prove fun incalculable. Chance, 7(1):42, 61, Winter 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:1994:WWD

[248] Katherine K. Wallman and Richard C. Rockwell. Window on Washington: Do you want to reduce the differential undercount and control costs in the Year 2000 Census? Chance, 7(1):43–45, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1994:VRT

[249] Howard Wainer. Visual revelations: Todai Moto Kurashi. Chance, 7(1):46– 49, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1994:ESCa

[250] Lionel Galway. On the edge: Statistics and computing: Computational experiments and exploratory modeling. Chance, 7(1):50–57, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Marsh:1994:PCO

[251] Paul Marsh. Personal crunching: Origin. Chance, 7(1):52–57, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:1994:BR

[252] Michael M. Meyer. Book reviews. Chance, 7(1):58–60, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1994:LSI

[253] Miron L. Straf. On the lighter side: International statistical junkets. Chance, 7(1):64–61, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Klein:1994:CCL

[254] Stephen P. Klein. Coping with catastrophes (and lesser events). Chance, 7(2):10–13, Spring 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Taylor:1994:BSF

[255] Jeremy M. G. Taylor. Betting strategies in final jeopardy. Chance, 7 (2):14–19, 27, Spring 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tweedie:1994:GGC

[256] Richard L. Tweedie, Kerrie L. Mengerson, and John A. Eccleston. Garbage in, garbage out: Can statisticians quantify the effects of poor data?

Chance, 7(2):20–27, Spring 1994. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Edwards:1994:BER

[257] Christopher T. Edwards. Baseball’s earned run average: What kind of an average is it? Chance, 7(2):28–30, Spring 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1994:EEP

[258] Constance F. Citro and Eric A. Hanushek. Estimating the effects of proposed legislation: The case for model validation. Chance, 7(2):31–40, Spring 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1994:TTE

[259] Belizaire. Temptations in Toronto: Ethnic variations. Chance, 7(2):41–44, Spring 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Humbaugh:1994:LE

[260] Philip R. Humbaugh and Betty C. Jung. Letters to the Editor. Chance, 7(3):3, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1994:CMb

[261] Anonymous. Chance musings. Chance, 7(3):4–8, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:1994:SCB

[262] Leonard Evans. Small cars, big cars: What is the safety difference? Chance, 7(3):9–16, Summer 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Brown:1994:RBD

[263] Lawrence D. Brown, Rebecca D’Amato, and Randy Gertner. Racetrack betting: Do bettors understand the odds? Chance, 7(3):17–23, Summer 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Frohlich:1994:BPN

[264] Cliff Frohlich. Baseball: Pitching nohitters. Chance, 7(3):24–30, Summer 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fairley:1994:BBB

[265] William B. Fairley, Alan J. Izenman, and A. Rhett Whitlock. Bricks, buildings, and the Bronx: Estimating masonry deterioration. Chance, 7(3):32–37, Summer 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Phillips:1994:CPF

[266] Haynes Phillips and Denise Phillips Wilburn. Consistency and predictability of football results. Chance, 7 (3):38–41, Summer 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1994:BRa

[267] Stanley Wasserman. Book reviews. Chance, 7(3):42–46, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Riche:1994:WWP

[268] Martha Farnsworth Riche and Jeffrey Passel. Window on Washington: Putting immigration in perspective. Chance, 7(3):47–48, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1994:VRV

[269] Howard Wainer. Visual revelations: Visual aids when comparing an Apple to the stars. Chance, 7(3):50–53, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1994:ESCb

[270] Lionel Galway. On the edge: Statistics and computing: E-books and ejournals: Why not? Chance, 7(3):54– 56, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lachenbruch:1994:PCL

[271] Peter A. Lachenbruch. Personal crunching: Logical decisions. Chance, 7(3):57–63, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1994:LSE

[272] Miron L. Straf. On the lighter side: Enhancing statistical performance. Chance, 7(3):64–63, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cargal:1994:LE

[273] Jim Cargal, Julian L. Simon, Charles D. Jones, Eugene P. Ericksen, and Teresa K. Defonso. Letters to the Editor. Chance, 7(4):3–4, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1994:CMc

[274] Anonymous. Chance musings. Chance, 7(4):5–8, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:1994:DSS

[275] Donald A. Berry. DNA, statistics, and the Simpson case. Chance, 7(4):9–12, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Matthews:1994:IOJ

[276] Robert Matthews. Improving the odds on justice? Chance, 7(4):13–15, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallis:1994:SW

[277] W. Allen Wallis. Statistics in Washington, 1935–1945. Chance, 7(4):16–20, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Manton:1994:WFA

[278] K. G. Manton and Burton Singer. What’s the fuss about compression of mortality? Chance, 7(4):21–30, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:1994:ACT

[279] Stephen E. Fienberg. An adjusted census in 1990? Trial judgment set aside. Chance, 7(4):31–32, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sloan:1994:LGP

[280] Jeff A. Sloan. Layperson’s guide to professional titles. Chance, 7(4):33, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bring:1994:PRF

[281] Johan Bring and Kenneth Carling. A paradox in the ranking of figure skaters. Chance, 7(4):34–37, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Scanlan:1994:DD

[282] James P. Scanlan. Divining difference. Chance, 7(4):38–39, 48, Fall 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1994:BRb

[283] Stanley Wasserman. Book reviews. Chance, 7(4):40–54, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Riche:1994:WWR

[284] Martha Farnsworth Riche. Window on Washington: Readying statistics on families and children for policy-making. Chance, 7(4):42–43, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1994:VRI

[285] Howard Wainer. Visual revelations: Implicit graphs. Chance, 7(4):44–48, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1994:ESCc

[286] Lionel Galway. On the edge: Statistics and computing: The Internet — a user’s view. Chance, 7(4):49–54, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lachenbruch:1994:PCE

[287] Peter A. Lachenbruch and Jerry Dallal. Personal crunching: Epi Info version 5.01. Chance, 7(4):55–58, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1994:BCI

[288] Anonymous. Chance cumulative index: Volume 7. Chance, 7(4):62–63, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1994:LSR

[289] Miron L. Straf. On the lighter side: To read or not to read? That is, the instructions. Chance, 7(4):64–59, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1994:EBP

[290] Anonymous. Editorial Board page. Chance, 7(4):ebi, 1994. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evinger:1995:HSW

[291] Suzann Evinger. How shall we measure our nation’s diversity? Chance, 8(1):7–14, Winter 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anderson:1995:BWS

[292] Margo Anderson and Stephen E. Fienberg. Black, white, and shades of gray (and brown and yellow). Chance, 8(1):15–18, Winter 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hurley:1995:OSD

[293] William Hurley. Overtime shootout: Deciding ties in hockey. Chance, 8 (1):19–22, Winter 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bailar:1995:HSC

[294] John C. Bailar III. How statisticians can help the news media do a better job. Chance, 8(1):24, 26–27, Winter 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [295].

Bailar:1995:HNM

[295] John C. Bailar III. How news media can help statisticians do a better job. Chance, 8(1):25, 28–29, Winter 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [294].

Heiner:1995:TRN

[296] Karlyn S. Heiner and Karl W. Heiner. Tree rings in the Northern Shawangunks. Chance, 8(1):30–37, Winter 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bennett:1995:CEP

[297] Jay Bennett. Cost-effectiveness in public education. Chance, 8(1):38–41, Winter 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 8, No. 2, 3.

Hanushek:1995:LE

[298] Eric A. Hanushek. Letter to the Editor. Chance, 8(2):3, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1995:CMa

[299] Anonymous. Chance musings. Chance, 8(2):4–45, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jones:1995:WDY

[300] Albyn C. Jones. What do you get when you combine a Pentium and a research Grant? Chance, 8(2):7–8, Spring 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gould:1995:PBS

[301] William Gould. The Pentium bug and statistical software. Chance, 8 (2):9–13, 25, Spring 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Merz:1995:PAP

[302] Jon F. Merz and Jonathan P. Caulkins. Propensity to abuse — propensity to murder? Chance, 8(2):14, Spring 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Horvitz:1995:POP

[303] Dan Horvitz, Daniel Koshland, Donald Rubin, Albert Gollin, Tom Sawyer, and Judith M. Tanur. Pseudo-opinion polls: SLOP or useful data? Statistical cryptic. Chance, 8(2):16–25, Spring 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1995:O

[304] Belizaire. Off to Orlando! Chance, 8 (2):26–27, 32, Spring 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Watkins:1995:LA

[305] Ann E. Watkins. The law of averages. Chance, 8(2):28–32, Spring 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaigh:1995:FBG

[306] W. D. Kaigh. Forecasting baseball games. Chance, 8(2):33–37, Spring 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1995:BRa

[307] Stanley Wasserman. Book reviews. Chance, 8(2):38–42, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1995:WWU

[308] Constance Citro. Window on Washington: A U.S. administrative record census? Chance, 8(2):43–45, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1995:VRR

[309] Howard Wainer. Visual revelations: A Rose by another name. Chance, 8(2): 46–51, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Galway:1995:ESC

[310] Lionel Galway. On the edge: Statistics and computing: Relations — who needs ’em? Chance, 8(2):52–55, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lachenbruch:1995:PCS

[311] Peter A. Lachenbruch. Personal crunching: StatMost statistical analysis and scientific graphs. Chance, 8(2): 56–61, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1995:LSR

[312] Miron Straf. On the lighter side: In Re: We trust. Chance, 8(2):64–61, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1995:WNC

[313] Hal S. Stern. Who’s Number 1 in college football? … and how might we decide? Chance, 8(3):7–14, Summer 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hammitt:1995:CMI

[314] James K. Hammitt. Can more information increase uncertainty? Chance, 8(3):15–17, 36, Summer 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Klein:1995:HWN

[315] Stephen P. Klein and Robert M. Bell. How will the NCAA’s new standards affect minority student-athletes? Chance, 8(3):18–21, Summer 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kim:1995:APF

[316] Jinho Kim and Hyung Jun Kim. The advantage of playing first in Go. Chance, 8(3):22–26, Summer 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Devlin:1995:WBC

[317] Bernie Devlin, Stephen E. Fienberg, Daniel P. Resnick, and Kathryn Roeder. Wringing the bell curve: A cautionary tale about the relationships among race, genes, and IQ. Chance, 8(3):27–36, Summer 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jackson:1995:TLF

[318] David Jackson. Tennis in Lilliput: A fable on sports and psychology. Chance, 8(3):37–40, Summer 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Warrack:1995:GS

[319] Giles Warrack. The great streak. Chance, 8(3):41–43, 60, Summer 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rolph:1995:LE

[320] John E. Rolph. Letter from the Editor. Chance, 8(4):3, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1995:CMb

[321] Anonymous. Chance musings. Chance, 8(4):4–6, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ashenfelter:1995:BWV

[322] Orley Ashenfelter, David Ashmore, and Robert Lalonde. Bordeaux wine vintage quality and the weather. Chance, 8 (4):7–14, Fall 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 9, No. 2, 7.

Laviolette:1995:FLG

[323] Michael Laviolette. Fuzzy logic: Great hope or grating hype? Chance, 8(4): 15–19, Fall 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Moore:1995:HBO

[324] Kris Moore, Jonathan Trower, and Kent Borowick. How birth order influences individual characteristics. Chance, 8(4):20–24, 48, Fall 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 9, No. 2, 5–6, includes Author’s Reply.

Nash:1995:MR

[325] John C. Nash and Mary M. Nash. Managing risks. Chance, 8(4):25–31, Fall 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1995:BRb

[326] Stanley Wasserman. Book reviews. Chance, 8(4):32–39, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1995:WWF

[327] Constance F. Citro, John Adams, and Audrey Burnam. Window on Washington: Federal block grants: What do statisticians have to contribute? Chance, 8(4):40–42, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1995:VRP

[328] Howard Wainer. Visual revelations: Picturing an L.A. bus schedule. Chance, 8(4):44–48, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1995:ESC

[329] R. Allan Reese. On the edge: Statistics and computing: Means what He says, says what He means. Chance, 8(4):49– 52, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lachenbruch:1995:PCB

[330] Peter A. Lachenbruch and Pamela Getson. Personal crunching: BMDP new system for Windows, version 1.11, reviewed by Pamela Getson, and Survey Pro for Windows, version 1.8. Chance, 8(4):53–59, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1995:BCI

[331] Anonymous. Chance cumulative index: Volume 8. Chance, 8(4):62–63, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1995:LSS

[332] Miron Straf. On the lighter side: Statistics happens. Chance, 8(4):64–52, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1995:EBP

[333] Anonymous. Editorial Board page. Chance, 8(4):ebi, 1995. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Styan:1996:LE

[334] George P. H. Styan. Letter from the Editor. Chance, 9(1):3, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Johnson:1996:DCE

[335] Glen D. Johnson, Barry D. Nussbaum, Ganapati P. Patil, and N. Phillip Ross. Designing cost-effective environmental sampling using concomitant information. Chance, 9(1):4–11, Winter 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1996:VRN

[336] Howard Wainer and John W. Durso. Visual revelations: A Nobel graph. Chance, 9(1):12–16, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bonnen:1996:PSR

[337] James T. Bonnen. The politics of statistical reform: A cautionary tale, 1978–1980. Chance, 9(1):17–26, Winter 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Worsley:1996:GRI

[338] Keith J. Worsley. The geometry of random images. Chance, 9(1):27–40, Winter 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sackrowitz:1996:TMS

[339] Harold Sackrowitz and David Sackrowitz. Time management in sports: Ball control and other myths. Chance, 9(1):41–49, Winter 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1996:WWW

[340] Constance F. Citro and Janet L. Norwood. Window on Washington: The winds of change: Opportunity or disaster for U.S. Federal statistics? Chance, 9(1):50–52, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1996:BRa

[341] Stanley Wasserman. Book reviews. Chance, 9(1):53–55, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ruxton:1996:CWW

[342] Graeme D. Ruxton. Can the world wide Web make statistics textbooks more fun? Chance, 9(1):56–57, Winter 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1996:LSSa

[343] Miron L. Straf. On the lighter side: Sufferin’ succotash! Another ASA cartoon meeting! Chance, 9(1):64–63, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1996:LE

[344] Anonymous. Letters to the Editor. Chance, 9(2):5–7, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Manly:1996:SWP

[345] Bryan F. J. Manly and Lyman L. McDonald. Sampling wildlife populations. Chance, 9(2):9–19, Spring 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1996:PCW

[346] Robert Gentleman. Personal crunching: Welcome from the new “Personal Crunching” Column Editor. Chance, 9(2):20–22, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1996:ESCa

[347] R. Allan Reese. On the edge: Statistics and computing: SED Moveat — stream editing still goes. Chance, 9(2): 23–26, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1996:CCB

[348] Anonymous. Coming to Chicago: “Belizaire”. Chance, 9(2):27–33, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1996:CC

[349] Anonymous. Coming to Chicago. Chance, 9(2):27–33, Spring 1996. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Paltiel:1996:MHP

[350] A. David Paltiel and Aaron A. Stinnett. Making health policy decisions: Is human instinct rational? Is rational choice human? Chance, 9(2):34–39, Spring 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1996:BRb

[351] Stanley Wasserman. Book reviews. Chance, 9(2):40–42, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1996:VRW

[352] Howard Wainer. Visual revelations: Why Playfair? Chance, 9(2):43–52, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1996:WWR

[353] Constance F. Citro. Window on Washington: Results of CPS research on the measurement of race and ethnicity. Chance, 9(2):53–56, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Trickett:1996:CWS

[354] Jennifer E. Trickett and Paul M. Sommers. Congressional ways and statistical means. Chance, 9(2):57–59, Spring 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1996:LSSb

[355] Miron L. Straf. On the lighter side: Statistician: Measure thyself. Chance, 9(2):64–63, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anderson:1996:ACS

[356] Margo Anderson and Stephen E. Fienberg. An adjusted census in 1990: The Supreme Court decides. Chance, 9(3):4–9, Summer 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Haff:1996:FFM

[357] L. R. Haff. The first four moves in the game of Go. Chance, 9 (3):10–16, Summer 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Emert:1996:IWC

[358] John Emert and Dale Umbach. Inconsistencies of “wild-card” poker. Chance, 9(3):17–22, Summer 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1996:ESCb

[359] R. Allan Reese. On the edge: Statistics and computing: Graphical literacy. Chance, 9(3):23–28, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sommers:1996:PCW

[360] Paul M. Sommers. Presidential candidates who measure up. Chance, 9 (3):29–32, Summer 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sommers:1996:SA

[361] Paul M. Sommers. Stamps of approval. Chance, 9(3):33–37, Summer 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 10, No. 3, 48.

Wasserman:1996:BRc

[362] Stanley Wasserman. Book reviews. Chance, 9(3):38–42, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1996:VRS

[363] Howard Wainer. Visual revelations: Scaling the heights (and widths). Chance, 9(3):43–49, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1996:PCP

[364] Robert Gentleman. Personal crunching: A peek at the Internet. Chance, 9(3):50–53, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1996:WWS

[365] Constance F. Citro, Edward L. Hunter, and Ross Arnett III. Window on Washington: Survey “reinventing” at Health and Human Services. Chance, 9(3):54– 57, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1996:LSV

[366] Miron L. Straf. On the lighter side: The virtual statistician. Chance, 9(3): 64–63, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1996:EBP

[367] Anonymous. Editorial Board page. Chance, 9(3):ebi, 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rao:1996:USC

[368] C. Radhakrishna Rao. Uncertainty, statistics, and creation of new knowledge. Chance, 9(4):5–11, Fall 1996.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lackritz:1996:TBG

[369] James R. Lackritz. Two of baseball’s great marks: Can they ever be broken? Chance, 9(4):12–18, Fall 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fricker:1996:MCB

[370] Ronald D. Fricker, Jr. The mysterious case of the blue M&M’s. Chance, 9(4): 19–22, Fall 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gilberg:1996:SBW

[371] Chris Gilberg, Julie L. Cox, Hidetaro Kashima, and Kurt Eberle. Survey biases: When does the interviewer’s race matter? Chance, 9(4):23–25, Fall 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cawley:1996:VIR

[372] Jason J. Cawley and Paul M. Sommers. Voting irregularities in the 1995 referendum on Québec sovereignty. Chance, 9(4):26–27, Fall 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor,” Vol. 10, No. 4, p. ??.

Sommers:1996:MOR

[373] Paul M. Sommers. Meddling with the 1996 Olympic results. Chance, 9(4):28– 30, Fall 1996. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also “Letters to the Editor”, Vol. 10, No. 2, 64.

Breiter:1997:HPO

[374] David J. Breiter and Bradley P. Carlin. How to play Office pools if you must. Chance, 10(1):5–11, Winter 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

DerSimonian:1997:DVC

[375] Rebecca Dersimonian and John Clemens. Does vasectomy cause prostate cancer? Chance, 10(1):12–14, Winter 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lee:1997:MSP

[376] Alan J. Lee. Modeling scores in the Premier League: Is Manchester United Really the best? Chance, 10(1):15–19, Winter 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaplan:1997:SEB

[377] Jack Kaplan. A statistical error in The Bell Curve. Chance, 10(1):20–21, Winter 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Brooks:1997:WGR

[378] George Brooks, Sangit Chatterjee, and Mustafa R. Yilmaz. Who is the greatest rebounder of all time? Chance, 10(1):22–25, Winter 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See also, “Letters to the Editor”, Vol. 3, No. 3, 6–7.

vanDijk:1997:WWQ

[379] Nico M. van Dijk. To wait or not to wait: That is the question. Chance, 10(1):26–30, Winter 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Spence:1997:WPD

[380] Ian Spence and Howard Wainer. William Playfair: A daring worthless fellow. Chance, 10(1):31–34, Winter 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1997:VRW

[381] Howard Wainer. Visual revelations: Who was Playfair? Chance, 10(1):35– 37, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1997:SRSa

[382] Hal S. Stern. A statistician reads the sports pages: Baseball by the numbers. Chance, 10(1):38–41, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1997:BRa

[383] Stanley Wasserman. Book reviews. Chance, 10(1):42–45, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:Ca

[384] Anonymous. Correction. Chance, 10 (1):45, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1997:PCO

[385] Robert Gentleman. Personal crunching: Operating systems for PCs. Chance, 10(1):46–48, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1997:ESCa

[386] R. Allan Reese. On the edge: Statistics & computing: What is data? Chance, 10(1):49–52, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1997:WWA

[387] Constance F. Citro. Window on Washington: Always a bridesmaid, never a bride? The role of sampling in the U.S. Decennial Census. Chance, 10(1):53–54, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1997:LSLa

[388] Miron L. Straf. On the lighter side: Lost in cyberspace. Chance, 10(1):64– 48, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Redelmeier:1997:UCP

[389] Donald A. Redelmeier and Robert J. Tibshirani. Is using a car phone like driving drunk? Chance, 10(2):5–9, Spring 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Agresti:1997:EAD

[390] Alan Agresti and Larry Winner. Evaluating agreement and disagreement among movie reviewers. Chance, 10 (2):10–14, Spring 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Searle:1997:SHP

[391] Shayle R. Searle. Selecting highproducing farm animals — with help from statistics. Chance, 10(2):15–18, Spring 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Losinger:1997:LCA

[392] Willard C. Losinger. The Lorenz curve applied to livestock populations. Chance, 10(2):19–22, Spring 1997. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:AAT

[393] Anonymous. Anaheim appetite and Turkish delight “Belizaire”. Chance, 10(2):23–26, Spring 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jackson:1997:HDT

[394] David Jackson and Krzysztof Mosurski. Heavy defeats in tennis: Psychological momentum or random effect? Chance, 10(2):27–34, Spring 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1997:SRSb

[395] Hal S. Stern. A statistician reads the sports pages: Judging who’s hot and who’s not. Chance, 10(2):40–43, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:Cb

[396] Anonymous. Corrections. Chance, 10 (2):43, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1997:BRb

[397] Stanley Wasserman. Book reviews. Chance, 10(2):44–47, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1997:PCR

[398] Robert Gentleman. Personal crunching: A review of BUGS: Bayesian inference Using Gibbs Sampling. Chance, 10(2):48–51, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1997:ESCb

[399] R. Allan Reese. On the edge: Statistics & computing: What goes down also goes up. Chance, 10(2):52–53, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1997:WWW

[400] Constance F. Citro. Window on Washington: What the public needs to know about U.S. federal statistics. Chance, 10(2):54–56, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evinger:1997:UOR

[401] Suzann Evinger. Update on OMB’s review of Directive No. 15. Chance, 10(2): 57, Spring 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1997:VRG

[402] Howard Wainer. Visual revelations: Graphical birth announcements. Chance, 10(2):58–60, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sommers:1997:LE

[403] Paul M. Sommers. Letter to the Editor. Chance, 10(2):64–63, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1997:CSC

[404] David R. Bellhouse. Il campanile statistico: Campanologia et ars combinatoria. Chance, 10(3):3–5, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hemenway:1997:MMA

[405] David Hemenway. The myth of millions of annual self-defense gun uses or a case study of survey overestimates of rare events. Chance, 10(3):6–10, Summer 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bradlow:1997:FFP

[406] Eric T. Bradlow. Forbes, free popcorn, and experimental design or a note on multivariate testing in marketing research. Chance, 10(3):11–13, Summer 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Elliott:1997:NSS

[407] Marc N. Elliott. NonRANDom selection of statisticians? Chance, 10 (3):14–15, Summer 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1997:SRSc

[408] Hal S. Stern and Wade Wilcox. A statistician reads the sports pages: Shooting darts. Chance, 10(3):16–19, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Williams:1997:SNH

[409] Bill Williams and David Williams. Are salaries in the National Hockey League related to nationality? Chance, 10 (3):20–24, Summer 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cross:1997:SLW

[410] Edward M. Cross and Edward P. Markowski. Secrets of the lottery wizard. Chance, 10(3):25–32, Summer 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1997:BRc

[411] Stanley Wasserman. Book reviews. Chance, 10(3):33–34, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:BTJ

[412] Anonymous. From the bookshelf: A tribute to Johnson and Kotz. Chance, 10(3):35, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jovanovic:1997:LLD

[413] Borko D. Jovanovic, Paul S. Levy, and Jacob A. Brody. A long line of dead men. Chance, 10(3):36–39, Summer 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1997:VRT

[414] Howard Wainer. Visual revelations: Tom’s veggies and the American way. Chance, 10(3):40–42, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ludwig:1997:LED

[415] David A. Ludwig and Paul M. Sommers. Letter to the Editor: Disputing discernible height. Chance, 10(3):48, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1997:PCS

[416] Robert Gentleman. Personal crunching: Statistical computing environments. Chance, 10(3):49–52, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hung:1997:VFL

[417] Rudy Hung. Visually fitting a least-absolute-deviation regression line. Chance, 10(3):53–52, Summer 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1997:ESCc

[418] R. Allan Reese. On the edge: Statistics & computing: Vote for Apple pie. Chance, 10(3):54–56, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1997:WWM

[419] Constance F. Citro. Window on Washington: Maintaining national statistics in an era of devolution: The challenge of welfare reform. Chance, 10(3):57–59, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1997:LSS

[420] Miron L. Straf. On the lighter side: Statistical partnering. Chance, 10(3): 64–63, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

George:1997:PSM

[421] Stephen L. George. Perspectives on scientific misconduct and fraud in clinical trials. Chance, 10(4):3–5, Fall 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Harrington:1997:AFD

[422] David Harrington. The aftermath of falsified data in breast cancer trials. Chance, 10(4):6–8, Fall 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ofallon:1997:RNA

[423] Judith Rich O’fallon. Reflections on the NSABP affair. Chance, 10(4):9–11, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dawes:1997:LES

[424] Robyn M. Dawes. Letter to the Editor: Survey biases and sample sizes. Chance, 10(4):11, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ambady:1997:JSB

[425] Nalini Ambady and Robert Rosenthal. Judging social behavior using “thin slices”. Chance, 10(4):12–51, Fall 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1997:SRSd

[426] Hal S. Stern. A statistician reads the sports pages: How accurately can sports outcomes be predicted? Chance, 10(4):19–23, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1997:WWR

[427] Constance F. Citro. Window on Washington: Recognizing diversity:

Recommendations to OMB on standards for data on race and ethnicity. Chance, 10(4):26–31, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1997:VRL

[428] Howard Wainer. Visual revelations: Like a trout in the milk. Chance, 10(4): 34–36, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1997:BRd

[429] Stanley Wasserman. Book reviews. Chance, 10(4):37–40, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:BBH

[430] Anonymous. From the bookshelf: A Chance holiday reading list. Chance, 10 (4):40, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Smith:1997:DST

[431] Gary Smith. Do statistics test scores regress toward the mean? Chance, 10(4):42–45, Fall 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Salsburg:1997:DPB

[432] David Salsburg. Digit preference in the Bible. Chance, 10(4):46–48, Fall 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1997:ESCd

[433] R. Allan Reese. On the edge: Statistics & computing: IQ — abnormal thoughts. Chance, 10(4):49–51, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:BTI

[434] Anonymous. Chance topical index: Volumes 1–10 (1988–1997). Chance, 10 (4):54–58, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:BCI

[435] Anonymous. Chance cumulative index: Volume 10 (1997). Chance, 10(4):59– 63, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1997:LSLb

[436] Miron L. Straf. On the lighter side: Let’s talk turkey. Chance, 10(4):64–63, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1997:EBP

[437] Anonymous. Editorial Board page. Chance, 10(4):ebi, 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fallon:1997:RNA

[438] Judith Rich Fallon. Reflections on the NSABP affair. Chance, 10(4):??, Faill 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Searle:1997:WPL

[439] Shayle R. Searle. Winning probabilities of lotto in the U.S.A. Chance, 10(4):??, Fall 1997. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morrison:1998:ITH

[440] Donald G. Morrison and David C. Schmittlein. It takes a hot goalie to raise the Stanley Cup. Chance, 11(1):3– 7, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wolfson:1998:WOW

[441] Michael C. Wolfson. Window on Ottawa: Welcome to Statistics Canada. Chance, 11(1):8–13, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Grosjean:1998:BSS

[442] James H. Grosjean. Beating the showcase showdown. Chance, 11(1):14–19, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Searle:1998:WPL

[443] Shayle R. Searle. Winning probabilities of lotto in the United States. Chance, 11(1):20–25, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1998:SRSa

[444] Hal S. Stern and Barbara Mock. A statistician reads the sports pages: College basketball upsets: Will a 16-seed ever beat a 1-seed? Chance, 11(1):26– 31, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1998:BBH

[445] Anonymous. From the bookshelf: A Chance holiday reading list — Part Two. Chance, 11(1):32–57, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1998:CSLa

[446] David R. Bellhouse. Il campanile statistico: Literary probability and statistics. Chance, 11(1):34–37, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1998:BRa

[447] Stanley Wasserman. Book reviews. Chance, 11(1):38–41, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bradlow:1998:PDS

[448] Eric T. Bradlow and Howard Wainer. Publication delays in statistics journals. Chance, 11(1):42–45, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1998:VRR

[449] Howard Wainer. Visual revelations: Rounding tables. Chance, 11(1):46–50, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1998:WWS

[450] Constance F. Citro and Janet L. Norwood. Window on Washington: Science and politics in numbers. Chance, 11(1):51–53, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reese:1998:ESC

[451] R. Allan Reese. On the edge: Statistics & computing: What’s normal? Chance, 11(1):54–57, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1998:SSA

[452] Howard Wainer, Samuel Palmer, and Eric T. Bradlow. A selection of selection anomalies. Chance, 11(2):3–7, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morton:1998:HBE

[453] R. Hugh Morton. The human body as an engine. Chance, 11(2):8–12, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bar-Hillel:1998:TCP

[454] Maya Bar-Hillel, Dror Bar-Natan, and Brendan McKay. The torah codes: Puzzle and solution. Chance, 11(2):13– 19, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bilder:1998:GAP

[455] Christopher R. Bilder and Thomas M. Loughin. “it’s good!” an analysis of the probability of success for placekicks. Chance, 11(2):20–30, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hurley:1998:HSY

[456] W. J. Hurley. How should you order players in a shoot-out? Chance, 11(2): 25–26, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hurley:1998:LE

[457] W. J. Hurley. Letter to the Editor. Chance, 11(2):26, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1998:BRb

[458] Stanley Wasserman. Book reviews. Chance, 11(2):27–30, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Belizaire:1998:VDD

[459] Belizaire. Virtual Dallas dining. Chance, 11(2):31–33, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1998:PCD

[460] Robert Gentleman and Murray Jorgensen. Personal crunching: Data mining. Chance, 11(2):34–42, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Westbrooke:1998:SPE

[461] Ian Westbrooke. Simpson’s paradox: An example in a New Zealand survey of jury composition. Chance, 11(2):40–42, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1998:CSS

[462] David R. Bellhouse. Il campanile statistico: Stochastic music. Chance, 11(2): 43–45, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1998:SRSb

[463] Hal S. Stern. A statistician reads the sports pages: Best of seven playoff series. Chance, 11(2):46–49, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1998:WWC

[464] Constance F. Citro and David B. McMillen. Window on Washington: The Census under attack. Chance, 11 (2):50–55, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1998:LST

[465] Miron L. Straf. On the lighter side: On time. Chance, 11(2):64–63, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Albert:1998:HRH

[466] Jim Albert. The home-run hitting of Mike Schmidt. Chance, 11(3):3–11, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jarvis:1998:CIB

[467] John F. Jarvis. Chance and intent: A baseball paradox. Chance, 11(3):12–19, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1998:SRSc

[468] Hal S. Stern. A statistician reads the sports pages: Football strategy: Go for it! Chance, 11(3):20–24, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Terrin:1998:CCL

[469] Norma Terrin and Joseph B. Kadane. Counting cars in a legal case involving differential enforcement. Chance, 11(3):25–27, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1998:PCRa

[470] Robert Gentleman. Personal crunching: A review of Systat 7.0 by SPSS. Chance, 11(3):28–30, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Farebrother:1998:GCD

[471] Richard William Farebrother. A genealogy of Charles Darwin, Francis Galton, and Francis Ysidro Edgeworth. Chance, 11(3):31–33, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Andersson:1998:RRH

[472] Jonas Andersson, Kenneth Carling, and Stefan Mattson. Random ranking of hospitals is unsound. Chance, 11(3): 34–39, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1998:BRc

[473] Stanley Wasserman. Book reviews. Chance, 11(3):38–39, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1998:WWM

[474] Constance F. Citro. Window on Washington: Model-based small-area estimates: The next major advance for the Federal statistical system for the 21st Century. Chance, 11(3):40–50, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wolfson:1998:WOS

[475] Michael C. Wolfson. Window on Ottawa: Statistics, taxes, and PIPES. Chance, 11(3):42–46, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1998:CSR

[476] David R. Bellhouse. Il campanile statistico: A random walk into Marxist– Leninist probability and statistics. Chance, 11(3):47–50, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1998:LEa

[477] Anonymous. Letters to the Editor. Chance, 11(3):51–58, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1998:E

[478] Anonymous. Errata. Chance, 11(3):61, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Styan:1998:LE

[479] George P. H. Styan. Letter from the Editor. Chance, 11(4):3, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Theus:1998:HST

[480] Martin Theus and Robert Erber. How to survive the Thirty Years’ War: Analyzing historical data of the City of Augsburg. Chance, 11(4):4–10, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Styan:1998:CAP

[481] Evelyn Matheson Styan. Chance in Aristotle’s physics. Chance, 11(4):11– 16, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1998:SRSd

[482] Hal S. Stern. A statistician reads the sports pages: How accurate are the posted odds? Chance, 11(4):17–21, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Manly:1998:HCT

[483] Bryan F. J. Manly and Richard Thomson. How to catch a thief. Chance, 11(4):22–25, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anderson-Cook:1998:MHS

[484] Christine M. Anderson-Cook and Tim Thornton. Measuring hockey’s special teams efficiency. Chance, 11(4):26–38, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1998:CSLb

[485] David R. Bellhouse. Il campanile statistico: Lotto mania. Chance, 11(4):35– 38, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1998:VRG

[486] Howard Wainer. Visual revelations: The graphical inventions of Dubourg and Ferguson: Two precursors to William Playfair. Chance, 11(4):39–41, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentleman:1998:PCRb

[487] Robert Gentleman. Personal crunching: Review of S-PLUS 4.5. Chance, 11(4):42–45, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1998:BRd

[488] Stanley Wasserman. Book reviews. Chance, 11(4):44–45, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1998:B

[489] Anonymous. From the bookshelf. Chance, 11(4):46–53, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1998:LEb

[490] Anonymous. Letters to the Editor. Chance, 11(4):47–53, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1998:BCI

[491] Anonymous. Chance cumulative index: Volume 11 (1998). Chance, 11(4):61– 64, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1998:EBP

[492] Anonymous. Editorial Board page. Chance, 11(4):ebi, 1998. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:LEE

[493] Anonymous. Letter from the executive Editor. Chance, 12(1):3, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:LEa

[494] Anonymous. Letters to the Editor. Chance, 12(1):4, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:CMa

[495] Anonymous. Chance musings. Chance, 12(1):5–6, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1999:BRa

[496] Stanley Wasserman. Book reviews. Chance, 12(1):7–9, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Warner:1999:CCA

[497] Brad Warner and Jim Rutledge. Checking the chips ahoy! Guarantee. Chance, 12(1):10–14, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1999:MWM

[498] Hal Stern. The man who makes the odds: An interview with “Roxy” Roxborough. Chance, 12(1):15–21, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ullrich:1999:ESB

[499] Peter Ullrich. An Eulerian square before Euler and an experimental design before R. A. Fisher: On the early history of latin squares. Chance, 12(1): 22–27, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Farebrother:1999:BHB

[500] Richard William Farebrother. A brief history of the British National Lottery 1567–1826. Chance, 12(1):27–31, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hastings:1999:BBS

[501] Kevin J. Hastings. Building a baseball simulation game. Chance, 12(1):32–37, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1999:WWF

[502] Constance F. Citro. Window on Washington: Five years later — reflections from inside the Federal statistical system. Chance, 12(1):38–41, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1999:VRP

[503] Howard Wainer. Visual revelations: A puzzle solved? Chance, 12(1):42–43, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:1999:ESCa

[504] Duncan J. Murdoch. On the edge: Statistics & computing: What’s Java? Chance, 12(1):44–47, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1999:CSW

[505] David R. Bellhouse. Il campanile statistico: What I did on my summer holidays. Chance, 12(1):48–50, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:1999:SRSa

[506] Scott M. Berry. A statistician reads the sports pages: Does “The Zone” exist for home-run hitters? Chance, 12(1): 51–56, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:1999:LSG

[507] Miron L. Straf. On the lighter side: Garbology. Chance, 12(1):64–63, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:LEW

[508] Anonymous. Letter to the Editor: On winning college basketball pools. Chance, 12(3):3, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:CMb

[509] Anonymous. Chance musings. Chance, 12(3):4–6, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1999:BRb

[510] Stanley Wasserman. Book reviews. Chance, 12(3):7–11, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Blumstein:1999:CVN

[511] Alfred Blumstein and Jeff Benedict. Criminal violence of NFL players compared to the general population. Chance, 12(3):12–15, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ashenfelter:1999:AWT

[512] Orley Ashenfelter and Richard Quandt. Analyzing a wine tasting statistically. Chance, 12(3):16–20, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Carlin:1999:DCF

[513] Bradley P. Carlin and Hal S. Stern. Designing a college football playoff system. Chance, 12(3):21–26, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hill:1999:DFD

[514] Theodore P. Hill. The difficulty of faking data. Chance, 12(3):27–31, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kadane:1999:AEC

[515] Joseph B. Kadane. An allegation of examination copying. Chance, 12(3):32– 36, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Akdeniz:1999:HSO

[516] Fikri Akdeniz and Doggan Dönmez. The history of statistics in the Ottoman Empire. Chance, 12(3):37–39, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1999:WWR

[517] Constance F. Citro and James D. Fitzsimmons. Window on Washington: Reviewing the metropolitan area standards. Chance, 12(3):40–42, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1999:VRS

[518] Howard Wainer. Visual revelations: Scaling the market. Chance, 12(3):43– 44, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:1999:ESCb

[519] Duncan J. Murdoch and Angela L. Jonkhans. On the edge: Statistics & computing: ActivStats 2.0. Chance, 12(3):45–47, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bellhouse:1999:CSB

[520] David R. Bellhouse. Il campanile statistico 2001: A Bayes odyssey. Chance, 12(3):48–50, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:1999:SRSb

[521] Scott M. Berry. A statistician reads the sports pages: A geometry model for NFL field goal kickers. Chance, 12(3): 51–56, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:1999:RSH

[522] Hal Stern. The reader speaks out an honest comparison. Chance, 12(3):64, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:LEb

[523] Anonymous. Letters to the Editor. Chance, 12(4):3–5, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:CMc

[524] Anonymous. Chance musings. Chance, 12(4):6–8, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:1999:BRBa

[525] Stanley Wasserman. Book reviews: From the bookshelf. Chance, 12(4):9– 10, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wright:1999:CER

[526] Tommy Wright and Howard Hogan. Census 2000: Evolution of the revised plan. Chance, 12(4):11–19, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Martin:1999:ETM

[527] David E. Martin and John F. Buoncristiani. The effects of temperature on Marathon runners’ performance. Chance, 12(4):20–24, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ker:1999:SUC

[528] Alan P. Ker and Barry K. Goodwin. Statistics and the U.S. Crop Insurance Program. Chance, 12(4):25–30, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kotz:1999:LLP

[529] Samuel Kotz and Norman L. Johnson. Lifelengths of “leading personalities” in statistics. Chance, 12(4):31–35, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murrell:1999:SM

[530] Paul R. Murrell. The statistics of monopoly. Chance, 12(4):36–40, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:1999:WWP

[531] Constance F. Citro. Window on Washington: Penny wise, pound foolish: The case of environmental accounting. Chance, 12(4):41–43, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:1999:VRG

[532] Howard Wainer. Visual revelations: A graphical investigation of the scourge of Vietnam. Chance, 12(4):44–46, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:1999:ESCc

[533] Duncan J. Murdoch. On the edge: Statistics & computing: Programming with components. Chance, 12(4):47–49, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:1999:SRSc

[534] Scott M. Berry. A statistician reads the sports pages: Drive for show and putt for dough. Chance, 12(4):50–55, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:BCI

[535] Anonymous. Chance cumulative index: Volume 12 (1999). Chance, 12(4):62– 63, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sawyer:1999:RST

[536] Congressman Tom Sawyer. The reader speaks out toward a more accurate census. Chance, 12(4):64, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:1999:EBP

[537] Anonymous. Editorial Board page. Chance, 12(4):ebi, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mauser:2000:LED

[538] Gary A. Mauser. Letter to the Editor: On defensive gun use statistics. Chance, 13(1):3–6, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2000:CMa

[539] Anonymous. Chance musings. Chance, 13(1):7–9, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2000:ATT

[540] Howard Wainer, Catherine Njue, and Samuel Palmer. Assessing time trends in sex differences in swimming & running. Chance, 13(1):10–15, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Martin:2000:CST

[541] David E. Martin. Comment: Studying trends in sport participation by modeling results of elite-level athletic performance. Chance, 13(1):16–17, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Price:2000:CCW

[542] Phillip N. Price. Comment: Can we trust a method just because we like its predictions? Chance, 13(1):18–20, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2000:SSR

[543] Howard Wainer, Samuel Palmer, and Catherine Njue. Sex and sports: A rejoinder. Chance, 13(1):21, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kerr:2000:PSI

[544] M. Kathleen Kerr. Patents and statistical inventions. Chance, 13(1):22–25, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hankins:2000:BDN

[545] Thomas L. Hankins. Blood, dirt, and nomograms. Chance, 13(1):26–37, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gerchak:2000:PRS

[546] Yigal Gerchak. On the “proper” relative size of prizes in competitions. Chance, 13(1):38–42, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2000:BRa

[547] Stanley Wasserman. Book reviews. Chance, 13(1):43–46, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2000:VRK

[548] Howard Wainer. Visual revelations: Kelley’s paradox. Chance, 13(1):47–48, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2000:ESCa

[549] Duncan J. Murdoch. On the edge: Statistics & computing: Statistics on the Web. Chance, 13(1):49–51, 2000.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2000:SRSa

[550] Scott M. Berry. A statistician reads the sports pages: Modeling acceptance to the Major League Baseball Hall of Fame. Chance, 13(1):52–57, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2000:OLL

[551] Hal S. Stern. Opinion: A lesson learned from Census 2000. Chance, 13(1):64, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2000:LE

[552] Anonymous. Letters to the Editor. Chance, 13(2):3–5, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2000:CMb

[553] Anonymous. Chance musings. Chance, 13(2):6–9, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kovar:2000:FMA

[554] Mary Grace Kovar. Four million adolescents smoke: Or do they? Chance, 13(2):10–14, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stefani:2000:OWP

[555] Raymond T. Stefani. Olympic winning performances: no longer “Citius, altius, fortius”. Chance, 13(2):15–19, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Strong:2000:WSP

[556] Robert A. Strong and John F. Buoncristiani. Wall Street profits, arbitrage, and the pricing of stock options. Chance, 13(2):20–24, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lu:2000:FPA

[557] Shiyong Lu and Steven Skiena. Filling a Penny album. Chance, 13(2):25–28, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Poulsen:2000:SSS

[558] Rolf Poulsen. Should He stay or should He go? Estimating the effect of firing the manager in soccer. Chance, 13(2):29–32, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2000:BRb

[559] Stanley Wasserman. Book reviews. Chance, 13(2):33–38, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:2000:WWF

[560] Constance F. Citro. Window on Washington: Federal statistics in the new Millennium — a user perspective. Chance, 13(2):39–41, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2000:VRT

[561] Howard Wainer. Visual revelations: Testing the disabled: Using statistics to navigate between the Scylla of standards and the Charybdis of court decisions. Chance, 13(2):42–44, 2000.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2000:ESCb

[562] Duncan J. Murdoch. On the edge: Statistics & computing: Omega tasting. Chance, 13(2):45–50, 2000. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2000:SRSb

[563] Scott M. Berry. A statistician reads the sports pages: Statistics and the Pokémon craze: Trying to impress my kids. Chance, 13(2):51–56, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Farebrother:2000:LWS

[564] Richard William Farebrother. The last word a slmple combinatorial problem. Chance, 13(2):64, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zaslavsky:2000:LE

[565] Alan Zaslavsky. Letters to the Editor. Chance, 13(3):3–5, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2000:CMc

[566] Anonymous. Chance musings. Chance, 13(3):6–7, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Redelmeier:2000:ODR

[567] Donald A. Redelmeier and Robert J. Tibshirani. Are those other drivers really going faster? Chance, 13(3):8–14, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Simonoff:2000:PMG

[568] Jeffrey S. Simonoff and Ilana R. Sparrow. Predicting movie grosses: Winners and losers, blockbusters and sleepers. Chance, 13(3):15–24, 2000. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Polson:2000:WWY

[569] Nicholas G. Polson and Jeffrey Yasumoto. Where will Yahoo! stock be in five years? Chance, 13(3):25–28, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sackrowitz:2000:RPA

[570] Harold Sackrowitz. Refining the point(s)-after-touchdown decision. Chance, 13(3):29–34, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2000:BRc

[571] Stanley Wasserman. Book reviews. Chance, 13(3):35–37, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:2000:WW

[572] Constance F. Citro. Window on the world. Chance, 13(3):38–42, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2000:VRNa

[573] Howard Wainer. Visual revelations: New tools for exploratory data analysis: I. automating two-way table decomposition. Chance, 13(3):43–47, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2000:ESCc

[574] Duncan J. Murdoch. On the edge: Statistics & computing: Drawing a scatterplot. Chance, 13(3):53–55, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2000:SRSc

[575] Scott M. Berry. A statistician reads the sports pages: My Triple Crown. Chance, 13(3):56–61, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2000:ONA

[576] Hal Stern. Opinion: A new age of statistics? Chance, 13(3):64, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2000:CMd

[577] Anonymous. Chance musings. Chance, 13(4):3–7, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berg:2000:PPA

[578] Ina Berg and Selwyn Blieden. The pots of Phylakopi: applying statistical techniques to archaeology. Chance, 13(4): 8–15, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gilbert:2000:DAV

[579] Peter Gilbert. Developing an AIDS vaccine by sieving. Chance, 13(4):16– 21, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anderson:2000:CC

[580] Margo Anderson and Stephen E. Fienberg. Census 2000 controversies. Chance, 13(4):22–30, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Soto-Johnson:2000:WLA

[581] Hortensia Soto-Johnson and Roger W. Johnson. Who lied about the Hot Tamales? Chance, 13(4):31–34, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schall:2000:CTB

[582] Teddy Schall and Gary Smith. Career trajectories in baseball. Chance, 13(4): 35–38, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2000:BRBa

[583] Stanley Wasserman. Book reviews: From the bookshelf. Chance, 13(4):39– 40, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:2000:WWP

[584] Constance F. Citro and Diane Hen. Window on Washington: The proposed American time-use survey. Chance, 13 (4):41–44, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2000:VRNb

[585] Howard Wainer. Visual revelations: New tools for exploratory data analysis: II. rotatable SPLOMS and a slicing engine. Chance, 13(4):45–47, 2000.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2000:ESCd

[586] Duncan J. Murdoch. On the edge: Statistics & computing: Markov chain Monte Carlo. Chance, 13(4):48–51, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2000:SRSd

[587] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: Modeling offensive ability in baseball. Chance, 13(4):56–59, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2000:IV

[588] Anonymous. Index to volume 13 (2000). Chance, 13(4):62–63, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2000:ORT

[589] Hal Stern. Opinion: Randomized trials in medicine. Chance, 13(4):64, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2000:EBP

[590] Anonymous. Editorial Board page. Chance, 13(4):ebi, 2000. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Nettler:2001:LE

[591] Gwynne Nettler, James A. Wollack, Joel R. Levin, Joseph B. Kadane, Stephen Senn, Robert A. Strong, and John F. Buoncristiani. Letters to the Editor. Chance, 14(1):3–5, 2001.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:CMa

[592] Anonymous. Chance musings. Chance, 14(1):6–9, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Briggs:2001:EAT

[593] Derek C. Briggs. The effect of admissions test preparation: Evidence from NELS:88. Chance, 14(1):10–18, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Powers:2001:CUN

[594] Donald E. Powers. Comment: Using National Education Longitudinal Study (NELS) data to evaluate the effects of commercial test preparation. Chance, 14(1):19–21, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Adams:2001:VIP

[595] Greg Adams. Voting irregularities in Palm beach, Florida. Chance, 14(1): 22–24, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Riche:2001:WCW

[596] Martha Farnsworth Riche. What Census 2000 will show. Chance, 14(1):25– 29, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kupper:2001:UGH

[597] Lawrence L. Kupper, Leonard B. Hearne, Sandra L. Martin, and Jeffrey M. Griffin. Is the USGA golf handicap system equitable? Chance, 14(1): 30–35, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2001:BRa

[598] Stanley Wasserman. Book reviews. Chance, 14(1):36–37, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:2001:WWW

[599] Constance F. Citro, Peggy G. Carr, and Janet Johnson. Window on Washington: What’s in NAEP’s black box. Chance, 14(1):38–42, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2001:VRN

[600] Howard Wainer. Visual revelations: New tools for exploratory data analysis: III. smoothing and nearness engines. Chance, 14(1):43–46, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2001:ESCa

[601] Duncan J. Murdoch and Dianne Cook. On the edge: Statistics & computing: Virtual reality: Real ponderings. Chance, 14(1):47–51, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2001:SRSa

[602] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: Luck in sports. Chance, 14(1):52–57, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2001:OGT

[603] Hal Stern. Opinion: Grading tests and counting votes — not as easy as it seems. Chance, 14(1):64, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:CMb

[604] Anonymous. Chance musings. Chance, 14(2):3–6, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Roberts:2001:SSE

[605] Seth Roberts. Surprises from selfexperimentation: Sleep, mood, and weight. Chance, 14(2):7–14, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rosenthal:2001:CLS

[606] Robert Rosenthal. Comment: Lessons from self-experimentation: Counting for content, inspecting for insight. Chance, 14(2):15, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rubin:2001:CSE

[607] Donald B. Rubin. Comment: Selfexperimentation for causal effects. Chance, 14(2):16–17, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Roberts:2001:R

[608] Seth Roberts. Rejoinder. Chance, 14 (2):18, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Connor:2001:ARS

[609] Jason T. Connor and Joseph B. Kadane. Alleged racial steering in an apartment complex. Chance, 14(2):19– 22, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wetzel:2001:TSG

[610] Nathan Wetzel. Three sisters give birth on the same day. Chance, 14(2):23–25, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mason:2001:MRP

[611] Randy Mason. A market research paradox of means. Chance, 14(2):26–30, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zaman:2001:CMP

[612] Zia Zaman. Coach Markov pulls goalie Poisson. Chance, 14(2):31–35, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2001:BRb

[613] Stanley Wasserman. Book reviews. Chance, 14(2):36–38, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Citro:2001:WWC

[614] Constance F. Citro. Window on Washington: The credibility of Federal statistics: Don’t take it for granted. Chance, 14(2):39–42, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2001:VRO

[615] Howard Wainer. Visual revelations: Order in the court. Chance, 14(2):43– 46, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2001:ESCb

[616] Duncan J. Murdoch, Radu V. Craiu, and Xiao-Li Meng. On the edge: Statistics & computing: Chance and fractals. Chance, 14(2):47–52, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2001:SRSb

[617] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: Do you feel a draft in here? Chance, 14(2):53–57, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2001:ORR

[618] Hal Stern. Opinion: Risk, reward, and the college statistics course. Chance, 14 (2):64, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:LEa

[619] Anonymous. Letter to the Editor. Chance, 14(3):3, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:CMc

[620] Anonymous. Chance musings. Chance, 14(3):4–6, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bernstein:2001:MBM

[621] Elliot Bernstein, Kevin Cottrell, Naomi Altman, André Dhondt, Wesley Hochachka, Roger Slothower, and Frances C. James. Measuring bird migration using spatial and temporal counts. Chance, 14(3):7–14, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

James:2001:C

[622] Frances C. James. Comment. Chance, 14(3):13, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hochachka:2001:R

[623] Wesley Hochachka, André Dhondt, and Naomi Altman. Response. Chance, 14 (3):14, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Deshotels:2001:DET

[624] Bob Deshotels and Mark Fitzgerald. Designing for extreme temperatures. Chance, 14(3):15–22, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Thall:2001:BCT

[625] Peter F. Thall. Bayesian clinical trial design in a cancer center. Chance, 14 (3):23–28, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Willoughby:2001:RMF

[626] Keith A. Willoughby. The return of a missed field goal in Canadian football. Chance, 14(3):29–33, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2001:BRc

[627] Stanley Wasserman. Book reviews. Chance, 14(3):34–39, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ahmed:2001:WWR

[628] Wendy Ahmed, Evi Rezmovic, and Terri Ann Lowenthal. Window on Washington: Racial profiling: A policy issue in need of better answers. Chance, 14(3):40–41, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2001:VRS

[629] Howard Wainer. Visual revelations: Sex, smoking, and life insurance: A graphic view. Chance, 14(3):42–45, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2001:ESCc

[630] Duncan J. Murdoch and Vincent J. Carey. On the edge: Statistics & computing: Literate statistical programming: Concepts and tools. Chance, 14 (3):46–50, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2001:SRSc

[631] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: How ferocious is Tiger? Chance, 14(3): 51–56, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2001:OBA

[632] Hal Stern. Opinion: The buzz about Bayes. Chance, 14(3):64, 2001. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2001:LEE

[633] Hal S. Stern. Letter from the executive Editor. Chance, 14(4):3, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:LEb

[634] Anonymous. Letters to the Editor. Chance, 14(4):4, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:CMd

[635] Anonymous. Chance musings. Chance, 14(4):5–7, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rodgers:2001:DHB

[636] Joseph Lee Rodgers and Debby Doughty. Does having Boys or girls run in the family? Chance, 14(4):8–13, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaplan:2001:MSS

[637] Jack Kaplan, Conor V. Dolan, and Arthur R. Jensen. Misuses of statistics in the study of intelligence: The case of Arthur Jensen. Chance, 14(4):14–26, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [638, 639] and reply [640].

Dolan:2001:CUC

[638] Conor V. Dolan. Comment: unhelpful criticism in the study of intelligence. Chance, 14(4):20–21, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [637].

Jensen:2001:CMC

[639] Arthur R. Jensen. Comment: misleading caricatures of Jensen’s statistics. Chance, 14(4):22–24, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [637].

Kaplan:2001:R

[640] Jack Kaplan. Reply. Chance, 14(4): 25–26, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [637].

Weil:2001:PVP

[641] Roman L. Weil. Parker v. Prial: the death of the vintage chart. Chance, 14(4):27–31, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2001:BRd

[642] Stanley Wasserman. Book reviews. Chance, 14(4):33–36, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2001:WWC

[643] Daniel Cork and Michael Cohen. Window on Washington: Census Directors past and adjustment decisions present: The 2000 Census Adjustment debate. Chance, 14(4):37–43, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2001:VRW

[644] Howard Wainer. Visual revelations: Winds across Europe: Francis Galton and the graphic discovery of weather patterns. Chance, 14(4):44–47, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Murdoch:2001:ESCd

[645] Duncan J. Murdoch and Erich Neuwirth. On the edge: Statistics & computing: Listen to your data — sonification with standard statistics software. Chance, 14(4):48–50, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2001:SRSd

[646] Scott M. Berry. A statistician reads the sports pages: The relative sizes of MLB umpires’ strike zones. Chance, 14(4): 51–54, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:IV

[647] Anonymous. Index to volume 14 (2001). Chance, 14(4):62–63, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Chameleon:2001:LWR

[648] T. Chameleon. The last word the real reason. Chance, 14(4):64, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2001:EBP

[649] Anonymous. Editorial Board page. Chance, 14(4):ebi, 2001. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2002:LEE

[650] Dalene Stangl. Letter from the executive Editor. Chance, 15(1):3, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:CMa

[651] Anonymous. Chance musings. Chance, 15(1):4–7, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Banks:2002:SHD

[652] David L. Banks. Statistics for homeland defense. Chance, 15(1):8–10, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:LEa

[653] Anonymous. Letters to the Editor. Chance, 15(1):11, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaplan:2002:SCP

[654] Jack Kaplan. An SAT coaching program that works. Chance, 15(1):12–17, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See commend [655].

Briggs:2002:JKS

[655] Derek C. Briggs. Jack Kaplan’s “An SAT coaching program with works”. Chance, 15(1):18–19, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [654].

Kaplan:2002:R

[656] Jack Kaplan. Reply. Chance, 15(1): 20–22, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Simon:2002:WYB

[657] Gary A. Simon and Jeffrey S. Simonoff. Were the 1996–2000 Yankees the best baseball team ever? Chance, 15(1):23–29, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robinove:2002:LFB

[658] Charles J. Robinove. Letter-frequency bias in an electronic Scrabble game. Chance, 15(1):30–31, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reed:2002:AGC

[659] Matt Reed and Vyvyan Howard. Ancient geometry comes to the aid of modern medics. Chance, 15(1):32–40, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2002:SRSa

[660] Scott M. Berry. A statistician reads the sports pages: Turn! Turn! Turn! Chance, 15(1):41–46, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2002:WWT

[661] Daniel Cork, Michael Cohen, Michael R. Rand, and Callie M. Rennison. Window on Washington: True crime stories? Accounting for differences in our national crime indicators. Chance, 15 (1):47–51, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2002:ESCa

[662] A. J. Rossini and Thomas Lumley. On the edge: Statistics & computing: Beowulfs meet statistics: An introduction to cluster supercomputers. Chance, 15(1):52–55, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2002:VRC

[663] Howard Wainer. Visual revelations: Clear thinking made visible: Redesigning score reports for students. Chance, 15(1):56–58, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:CMb

[664] Anonymous. Chance musings. Chance, 15(2):3–7, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:LEb

[665] Anonymous. Letters to the Editor. Chance, 15(2):8–9, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kadane:2002:AJC

[666] Joseph B. Kadane. Anatomy of a jury challenge. Chance, 15(2):10–13, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2002:GMP

[667] Scott M. Berry. Gold medals, protests, and trading ratings: The 2002 Winter Olympics figure skating. Chance, 15(2): 14–18, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Levy:2002:SMP

[668] Paul S. Levy, Jason Hsia, Borko Jovanovic, and Douglas Passaro. Sampling mailrooms for presence of anthrax spores: A curious property of the hypergeometric distribution under an unusual hypothesis testing scenario. Chance, 15(2):19–21, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Laska:2002:IPW

[669] Eugene M. Laska and Paul Leber. If placebo works, could, should or would it be approved for marketing? Chance, 15(2):22–26, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Senn:2002:UP

[670] Stephen Senn. The unpleasant placebo? Chance, 15(2):27–29, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Grier:2002:PUE

[671] David Alan Grier. Predicting the unpredictable election. Chance, 15(2):30– 35, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2002:BRB

[672] Stanley Wasserman. Book reviews: From the bookshelf. Chance, 15(2):36– 38, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2002:ESCb

[673] A. J. Rossini and Thomas Lumley. On the edge: Statistics & computing:

Scatterplotting. Chance, 15(2):39–42, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2002:WWC

[674] Daniel Cork and Michael Cohen. Window on Washington: Census adjustment: The view from Capitol Hill. Chance, 15(2):43–48, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2002:SRSb

[675] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: One modern man or 15 Tarzans? Chance, 15(2):49–53, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2002:VRR

[676] Howard Wainer. Visual revelations: Reporting test results to institutions and nations. Chance, 15(2):58–61, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:CMc

[677] Anonymous. Chance musings. Chance, 15(3):3–8, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Johnson:2002:TCE

[678] Valen E. Johnson. Teacher course evaluations and student grades: An academic tango. Chance, 15(3):9–16, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kleinman:2002:HGI

[679] Kim Kleinman. How graphical innovations assisted Edgar Anderson’s discoveries in evolutionary biology. Chance, 15(3):17–21, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2002:VFP

[680] Andrew Gelman. Voting, fairness, and political representation. Chance, 15(3): 22–26, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ansolabehere:2002:CVF

[681] Stephen Ansolabehere. Comment on “Voting, fairness, and political representation”. Chance, 15(3):27–28, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Allen:2002:UIT

[682] Terry Allen. Using and interpreting the trilinear plot. Chance, 15(3):29–35, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Boll:2002:B

[683] Heinrich Böll. At the bridge. Chance, 15(3):36, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2002:BRa

[684] Stanley Wasserman. Book reviews. Chance, 15(3):37–39, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2002:ESCc

[685] A. J. Rossini and Thomas Lumley. On the edge: Statistics & computing: Database linkage, privacy, and medical research. Chance, 15(3):40–42, 2002.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2002:WWD

[686] Daniel Cork and Michael Cohen. Window on Washington: Determining data needs for the consumer price index. Chance, 15(3):43–47, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2002:SRSc

[687] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: Read during your leisure lime. Chance, 15(3):48–55, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2002:VRB

[688] Howard Wainer. Visual revelations: The BK-Plot: Making Simpson’s paradox clear to the masses. Chance, 15(3): 60–62, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lewis:2002:LE

[689] Hal Lewis and Santa Barbara. Letters to the Editor. Chance, 15(4):3–4, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:CMd

[690] Anonymous. Chance musings. Chance, 15(4):4–7, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Albert:2002:HRS

[691] Jim Albert. Hitting with runners in scoring position. Chance, 15(4):8–16, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ball:2002:SSU

[692] Patrick Ball and Jana Asher. Statistics and Slobodan: using data analysis and statistics in the war crimes trial of former president Milosevic. Chance, 15(4):17–24, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hall:2002:SSI

[693] Peter Hall. Statistical smoothing and the investigation of Flight 587. Chance, 15(4):25–26, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kadane:2002:CLP

[694] Joseph B. Kadane. Crossing lines in a patent case. Chance, 15(4):27–32, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:2002:UVD

[695] Mary C. Meyer. Uncounted votes: Does voting equipment matter? Chance, 15(4):33–38, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2002:BRc

[696] Stanley Wasserman. Book reviews. Chance, 15(4):39–40, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2002:WWI

[697] Daniel Cork and Michael Cohen. Window on Washington: Innovations in the Federal statistical system. Chance, 15 (4):41–45, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2002:ESCd

[698] A. J. Rossini and Thomas Lumley. On the edge: Statistics & computing: Something in the air? Chance, 15(4): 46–49, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2002:SRSd

[699] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: A juiced analysis. Chance, 15(4):50–53, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2002:VRS

[700] Howard Wainer. Visual revelations: A small hurrah for the black death. Chance, 15(4):58–60, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:IV

[701] Anonymous. Index to volume 15 (2002). Chance, 15(4):63–64, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2002:EBP

[702] Anonymous. Editorial Board page. Chance, 15(4):ebi, 2002. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2003:CMa

[703] Anonymous. Chance musings. Chance, 16(1):3–7, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hall:2003:PTH

[704] Peter Hall. A possum’s tale — how statistics revealed a new mammal species. Chance, 16(1):8–13, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ryan:2003:BSM

[705] Thomas P. Ryan. Basketball statistics: A matter of adjustment. Chance, 16(1): 14–17, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Woodward:2003:YS

[706] Phil Woodward. Yahtzee^{(}R): The solution. Chance, 16(1):18–22, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fernandez-Cantelli:2003:IAS

[707] Eduardo Fernandez-Cantelli and Glen Meeden. An improved award system for soccer. Chance, 16(1):23–29, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morrison:2003:CRB

[708] Peter A. Morrison. Confronting a race-based school admissions policy. Chance, 16(1):30–34, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sommers:2003:WW

[709] Paul M. Sommers. The writing on the wall. Chance, 16(1):35–38, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2003:BRa

[710] Stanley Wasserman. Book reviews. Chance, 16(1):39–43, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2003:ESCa

[711] A. J. Rossini, Thomas Lumley, and W. B. Lober. On the edge: Statistics & computing: Bioterrorism surveillance using computerized clinical data. Chance, 16(1):44–48, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2003:SRSa

[712] Scott M. Berry. A statistician reads the sports pages: The great one. Chance, 16(1):49–54, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2003:VRG

[713] Howard Wainer. Visual revelations: A graphical legacy of Charles Joseph Minard: Two jewels from the past. Chance, 16(1):58–62, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2003:CMb

[714] Anonymous. Chance musings. Chance, 16(2):3–4, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Holmes:2003:WWA

[715] David I. Holmes and Judit Kardos. Who was the author? An introduction to stylometry. Chance, 16(2):5–8, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Binongo:2003:WWB

[716] José Nilo G. Binongo. Who wrote the 15th Book of Oz? An application of multivariate analysis to authorship attribution. Chance, 16(2):9–17, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Holmes:2003:SCW

[717] David I. Holmes. Stylometry and the Civil War: The case of the Pickett Letters. Chance, 16(2):18–25, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rudman:2003:CPN

[718] Joseph Rudman. Cherry picking in nontraditional authorship attribution studies. Chance, 16(2):26–32, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2003:MCO

[719] Nicole A. Lazar and David S. Sidore. Math chaps and other oddities: Statisticians in literature. Chance, 16(2):33– 37, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2003:BRb

[720] Stanley Wasserman. Book reviews. Chance, 16(2):38–40, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2003:ESCb

[721] A. J. Rossini, Thomas Lumley, and Friedrich Leisch. On the edge: Statistics & computing: Reproducible statistical research. Chance, 16(2):41–45, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2003:SRSb

[722] Scott M. Berry. A statistician reads the sports pages: College football rankings: The BCS and the CLT. Chance, 16(2):46–49, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2003:VRH

[723] Howard Wainer. Visual revelations: How long is short? Chance, 16(2):55– 57, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2003:WWD

[724] Daniel Cork, Michael Cohen, and Katharine G. Abraham. Window on Washington: Data for managing tight labor markets: The new job openings and labor turnover survey. Chance, 16(2):58–62, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2003:LE

[725] Anonymous. Letters to the Editor. Chance, 16(3):3, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ryan:2003:AR

[726] Tom Ryan. Author’s reply. Chance, 16 (3):3, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2003:CMc

[727] Anonymous. Chance musings. Chance, 16(3):4–9, 2003. CODEN CNDCE4.

ISSN 0933-2480 (print), 1867-2280 (electronic).

Qian:2003:SSW

[728] Song S. Qian and Michael Lavine. Setting standards for water quality in the Everglades. Chance, 16(3):10–16, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Johnson:2003:CCS

[729] Valen E. Johnson. Consumerism in the classroom: Student grades and student course selections. Chance, 16(3):17–20, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wolfson:2003:UCB

[730] Lara J. Wolfson and Thomas R. Lee. Utah’s Census battles: The inside story. Chance, 16(3):21–29, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2003:BBB

[731] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: Bush v. Gore: Two neglected lessons from a statistical perspective. Chance, 16(3): 30–36, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2003:BRBa

[732] Stanley Wasserman. Book reviews: From the bookshelf. Chance, 16(3):37– 38, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2003:ESCc

[733] A. J. Rossini, Thomas Lumley, Lance A. Waller, and Andrew B. Barclay. On the edge: Statistics & computing: “agile” GIS: Building applicationspecific spatial analytic software from freely available software tools. Chance, 16(3):39–44, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2003:SRSc

[734] Scott M. Berry. A statistician reads the sports pages: “A game of which I am not familiar”. Chance, 16(3):45–48, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2003:WWF

[735] Daniel Cork, Michael Cohen, Thomas B. Jabine, and Thomas A. Louis. Window on Washington: Funding formulas in a Federal system. Chance, 16(3):49–53, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2003:VRD

[736] Howard Wainer. Visual revelations: La diffusion de quelques idées: a master’s voice. Chance, 16(3):58–61, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2003:EWT

[737] Dalene Stangl. From the Editor welcome and thank you. Chance, 16(4):3, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Smith:2003:LED

[738] Richard L. Smith. Letter to the Editor: Discussion of “Bush v. Gore… ”. Chance, 16(4):4–5, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Banks:2003:PFB

[739] David Banks. Proteomics: A frontier between genomics and metabolomics. Chance, 16(4):6–7, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Banks:2003:FCS

[740] David Banks and Emmanuel Petricoin. Finding cancer signals in mass spectrometry data. Chance, 16(4):8–12, 57, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Roy:2003:ATD

[741] Anindya Roy, Françoise Seillier-Moiseiwitsch, Kwan R. Lee, Yaming Hang, Mark Marten, and Babu Raman. Analyzing two-dimensional gel images. Chance, 16(4):13–18, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hawkins:2003:EBS

[742] Douglas M. Hawkins, Russell D. Wolfinger, Li Liu, and S. Stanley Young. Exploring blood spectra for signs of ovarian cancer. Chance, 16(4): 19–23, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Liggett:2003:LMS

[743] Walter Liggett, Lisa Cazares, and O. John Semmes. A look at mass spectral measurement. Chance, 16(4):24– 28, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Thiele:2003:MSB

[744] Herbert Thiele. Mass spectrometry and bioinformatics in proteomics. Chance, 16(4):29–36, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2003:BBS

[745] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: A statistical portrait of the handgun industry. Chance, 16(4):37–40, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Christiansen:2003:HYH

[746] Cindy Christiansen and Mark Glickman. Here’s to your health: Statistics: Keeping pace with the medical technology revolution. Chance, 16(4):41–44, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2003:VRP

[747] Howard Wainer. Visual revelations: A political statistic. Chance, 16(4):45–47, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wasserman:2003:BRc

[748] Stanley Wasserman. Book reviews. Chance, 16(4):48–49, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2003:WWC

[749] Daniel Cork and Michael Cohen. Window on Washington: The Congress and statistics. Chance, 16(4):54–57, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2003:IV

[750] Anonymous. Index to volume 16. Chance, 16(4):62–63, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2003:EBP

[751] Anonymous. Editorial Board page. Chance, 16(4):ebi, 2003. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2004:CMa

[752] Anonymous. Chance musings. Chance, 17(1):3–5, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kettenring:2004:SC

[753] Jon Kettenring. Statistics and counterterrorism. Chance, 17(1):6–7, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Marchette:2004:SAN

[754] David J. Marchette and Edward J. Wegman. Statistical analysis of network data for cybersecurity. Chance, 17(1):8–18, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stoto:2004:SSI

[755] Michael A. Stoto, Matthias Schonlau, and Louis T. Mariano. Syndromic surveillance: Is it worth the effort? Chance, 17(1):19–24, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gomatam:2004:RLC

[756] Shanti Gomatam and Michael D. Larsen. Record linkage and counterterrorism. Chance, 17(1):25–29, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rukhin:2004:RPB

[757] Andrew L. Rukhin. The recognition problem of biometrics. Chance, 17(1): 30–34, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2004:BBSa

[758] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: Statistical proof of discrimination in peremptory challenges. Chance, 17(1):35–38, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Christiansen:2004:HYHa

[759] Cindy Christiansen and Mark Glickman. Here’s to your health: Is there one best model of care for children orphaned by AIDS?: A multinational perspective and how statistics can help. Chance, 17(1):39–43, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2004:VRNa

[760] Howard Wainer and Henry Braun. Visual revelations: Numbers and the remembrance of things past. Chance, 17 (1):44–48, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2004:ESCa

[761] A. J. Rossini and Thomas Lumley. On the edge: Statistics & computing: Building a better mailtrap. Chance, 17(1):49–52, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2004:SRSa

[762] Scott M. Berry. A statistician reads the sports pages: The black cat unmasked. Chance, 17(1):53–56, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2004:WW

[763] Daniel Cork and Michael Cohen. Window on Washington. Chance, 17(1):59– 62, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2004:CMb

[764] Anonymous. Chance musings. Chance, 17(2):3–4, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2004:IAT

[765] Donald A. Berry and LeeAnn Chastain. Inferences about testosterone abuse among athletes. Chance, 17(2):5–8, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bishop:2004:DCS

[766] Todd Bishop and Byron J. Gajewski. Drafting a career in sports: Determining underclassmen college players’ stock in the NBA draft. Chance, 17(2): 9–12, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mitchell:2004:DDT

[767] Heather Mitchell and Mark F. Stewart. Does drug-testing deter participation in athletic events? EPO and the Sydney Olympics. Chance, 17(2):13–18, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2004:HMD

[768] Howard Wainer, Brent Bridgeman, Michelle Najarian, and Catherine Trapani. How much does extra time on the SAT help? Chance, 17(2):19–24, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Martonosi:2004:TA

[769] Susan Martonosi and Arnold Barnett. Terror is in the air. Chance, 17(2):25– 27, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2004:RDN

[770] Andrew Gelman. 55,000 residents desperately need your help! Chance, 17 (2):28–31, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Grimshaw:2004:TYU

[771] Scott D. Grimshaw, Howard B. Christensen, David B. Magleby, and Kelly D. Patterson. Twenty years of the Utah colleges exit poll: Learning by doing. Chance, 17(2):32–38, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2004:BBE

[772] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: Epidemiologic evidence in the silicone breast implant cases. Chance, 17(2):39–43, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2004:C

[773] Anonymous. Correction. Chance, 17 (2):43, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Christiansen:2004:HYHb

[774] Cindy Christiansen, Mark Glickman, and Bei-Hung Chang. Here’s to your health: Studying complementary and alternative medicine in the United States. Chance, 17(2):44–47, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2004:VRNb

[775] Howard Wainer and Michael Friendly. Visual revelations: Nobody’s perfect. Chance, 17(2):48–51, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rossini:2004:ESCb

[776] A. J. Rossini and Thomas Lumley. On the edge: Statistics & computing: Sustainable sources of randomness. Chance, 17(2):52–54, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2004:SRSb

[777] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: Hired to be fired. Chance, 17(2):55–59, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2004:CMc

[778] Anonymous. Chance musings. Chance, 17(3):3–4, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:2004:MRC

[779] Stephen E. Fienberg and Aleksandra B. Slavkovic. Making the release of confidential data from multi-way tables count. Chance, 17(3):5–10, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reiter:2004:NAD

[780] Jerome P. Reiter. New approaches to data dissemination: A glimpse into the future (?). Chance, 17(3):11–15, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Duncan:2004:DRV

[781] George T. Duncan and S. Lynne Stokes. Disclosure risk vs. data utility: the R–U confidentiality map as applied to topcoding. Chance, 17(3):16–20, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wallman:2004:ICI

[782] Katherine K. Wallman and Brian A. Harris-Kojetin. Implementing the Confidential Information Protection and Statistical Efficiency Act of 2002. Chance, 17(3):21–25, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Karr:2004:AID

[783] Alan F. Karr, Xiaodong Lin, Ashish P. Sanil, and Jerome P. Reiter. Analysis of integrated data without data integration. Chance, 17(3):26–29, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Singer:2004:CRP

[784] Eleanor Singer. Confidentiality, risk perception, and survey participation. Chance, 17(3):30–34, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:2004:FTM

[785] Miron L. Straf. Fostering trust to motivate responses: It’s not just a matter of confidentiality. Chance, 17(3):35–36, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Albert:2004:STP

[786] Jim Albert. Streakiness in team performance. Chance, 17(3):37–43, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2004:BBB

[787] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: Bullet lead as forensic evidence. Chance, 17(3):44– 47, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Christiansen:2004:HYHc

[788] Cindy Christiansen, Mark Glickman, and Nicole A. Lazar. Here’s to your health: Numbers on the brain: The role of statistics in the analysis of functional magnetic resonance images.

Chance, 17(3):48–51, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2004:SRSc

[789] Scott M. Berry and Berry Consultants. A statistician reads the sports pages: Order, by “The Book”. Chance, 17(3): 52–55, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2004:WWB

[790] Daniel Cork, Michael Cohen, David M. Talan, James R. Spletzer, and Richard L. Clayton. Window on Washington: Business employment dynamics. Chance, 17(3):58–61, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2004:CMd

[791] Anonymous. Chance musings. Chance, 17(4):3–6, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Beran:2004:MCF

[792] Jan Beran. Music — chaos, fractals, and information. Chance, 17(4):7–16, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Raphael:2004:MAS

[793] Christopher Raphael. Musical accompaniment systems. Chance, 17(4):17– 22, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Godsill:2004:CMM

[794] Simon Godsill. Computational modeling of musical signals. Chance, 17(4):

23–29, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Irizarry:2004:PMI

[795] Rafael A. Irizarry. Parameters with musical interpretations. Chance, 17(4): 30–38, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2004:BBSb

[796] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: Stopping rules in clinical trials. Chance, 17(4): 39–42, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2004:VRC

[797] Howard Wainer. Visual revelations: Curbstoning IQ and the 2000 Presidential Election. Chance, 17(4):43–46, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2004:SRSd

[798] Scott M. Berry, Berry Consultants, and Craig Wood. A statistician reads the sports pages: The cold-foot effect. Chance, 17(4):47–51, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2004:WWS

[799] Daniel Cork, Michael Cohen, and Arthur Fries. Window on Washington: Statistics counts! Examples of the impact of statistical analyses on the operational test and evaluation of major defense systems. Chance, 17(4):55–60, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2004:IV

[800] Anonymous. Index to volume 17. Chance, 17(4):61–62, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2004:EBP

[801] Anonymous. Editorial Board page. Chance, 17(4):ebi, 2004. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:EBa

[802] Anonymous. Editorial Board. Chance, 18(1):2, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Potthoff:2005:LEN

[803] Richard F. Potthoff. Letter to the Editor: A neglected issue in Florida in 2000? Chance, 18(1):3–4, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See reply [804].

Anonymous:2005:ER

[804] Anonymous. The Editors’ reply. Chance, 18(1):4, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [803].

Rosenbaum:2005:RE

[805] Paul R. Rosenbaum. Reasons for effects. Chance, 18(1):5–10, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Horton:2005:SSD

[806] Nicholas J. Horton and Emily C. Shapiro. Statistical sleuthing during epidemics: Maternal influenza and schizophrenia. Chance, 18(1):11–19, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Redelmeier:2005:DFC

[807] Donald A. Redelmeier and Craig L. Stewart. Do fatal crashes increase following a Super Bowl telecast? Chance, 18(1):19–24, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Falk:2005:ASB

[808] Ruma Falk, Avital Lann, and Shmuel Zamir. Average speed bumps: Four perspectives on averaging speeds. Chance, 18(1):25–32, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gideon:2005:DBH

[809] Rudy A. Gideon. The distribution of bridge hands in newspapers. Chance, 18(1):33–37, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Christiansen:2005:HYH

[810] Cindy Christiansen, Mark. Glickman, Sarah Michalak, and Greg Wilson. Here’s to your health: Reconsidering smallpox. Chance, 18(1):38–43, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2005:BBS

[811] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: Statistical evidence in an education finance case. Chance, 18(1):44–47, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2005:SRSa

[812] Scott M. Berry. A statistician reads the sports pages: Nature, nurture, or culture. Chance, 18(1):48–53, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2005:PCSa

[813] Thomas B. Jabine. Puzzle corner statistical cryptic puzzle no. 17. Chance, 18(1):54–55, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2005:WWH

[814] Daniel Cork, Michael Cohen, and Mary Dorinda Allard. Window on Washington: How Americans spend their time: A first look at the new American time use survey. Chance, 18 (1):56–62, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:AAa

[815] Anonymous. About the authors. Chance, 18(1):63–64, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:EBb

[816] Anonymous. Editorial Board. Chance, 18(2):2, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:2005:WWA

[817] Mary C. Meyer and Tremika Finney. Who wants airbags? Chance, 18(2): 3–16, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See discussion [818, 913, 914, 963].

Anonymous:2005:RDW

[818] Anonymous. Reply to discussion of “who wants airbags?”. Chance, 18(2): 16, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [817].

Kahane:2005:DWW

[819] Charles J. Kahane. Discussion: “Who wants airbags?”. Chance, 18(2):16, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2005:WIE

[820] Howard Wainer and William P. Skorupski. Was it ethnic and social-class bias or statistical artifact? Logical and empirical evidence against Freedle’s method for reestimating SAT scores. Chance, 18(2):17–24, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaplan:2005:ESC

[821] Jack Kaplan. The effectiveness of SAT coaching on math SAT scores. Chance, 18(2):25–34, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Banks:2005:SAT

[822] David Banks, Maria J. Sirois, and Raphel O. Thattil. Statistical aspects of tsunami response. Chance, 18(2):35– 40, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hannum:2005:RBU

[823] Robert C. Hannum. Risky business: The use and misuse of statistics in casino gaming. Chance, 18(2):41–47, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hombas:2005:GGA

[824] Vassilios C. Hombas and Christos P. Baloglou. Gambling in Greek antiquity. Chance, 18(2):49–50, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2005:VRS

[825] Howard Wainer. Visual revelations: Stumbling on the path toward the visual communication of complexity. Chance, 18(2):53–54, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2005:SPF

[826] Scott M. Berry. Is second place the first loser? Chance, 18(2):55–59, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2005:PCSb

[827] Thomas B. Jabine. Puzzle corner statistical spiral cryptic no. 7. Chance, 18(2):60–61, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:AAb

[828] Anonymous. About the authors. Chance, 18(2):63–64, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:EBc

[829] Anonymous. Editorial Board. Chance, 18(3):2, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wilburn:2005:MDT

[830] John Bart Wilburn. Missile detection and tracking in satellite imagery. Chance, 18(3):3–7, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Weil:2005:ARR

[831] Roman L. Weil. Analysis of reserve and regular bottlings: Why pay for a difference only the critics claim to notice? Chance, 18(3):9–15, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Durdan:2005:ESC

[832] Simon F. Durdan, Daniela Ferrero, and Frank Harary. On the evolution of square-cell animals. Chance, 18(3):16– 19, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Helman:2005:RF

[833] Danny Helman. Reversal of fortunes. Chance, 18(3):20–22, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Erosheva:2005:WMB

[834] Elena A. Erosheva. Will the market bear your asking price? Chance, 18(3): 23–28, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McCulloch:2005:RMA

[835] Charles E. McCulloch. Repeated measures ANOVA, R.I.P.? Chance, 18(3): 29–33, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2005:CAT

[836] Andrew Gelman. Comment: Anova as a tool for structuring and understanding hierarchical models. Chance, 18(3): 33, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Finkelstein:2005:BBM

[837] Michael O. Finkelstein and Bruce Levin. Chance at the Bar: The machinery of death. Chance, 18(3):34–37, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2005:VRO

[838] Howard Wainer and Danielle Vasilescu. Visual revelations: Old Mother Hubbard and the United Nations: an adventure in exploratory data analysis. Chance, 18(3):38–45, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Weisberg:2005:C

[839] Sanford Weisberg. Comment. Chance, 18(3):46–48, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2005:HYHa

[840] Mark Glickman, Jeffrey D. Dawson, Patricia H. Davis, and Cindy Christiansen. Here’s to your health: What’s the big deal about BMI? Chance, 18(3):

49–52, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2005:SRSb

[841] Scott M. Berry. A statistician reads the sports pages: Winning isn’t everything; it’s the only thing. Chance, 18(3):53– 56, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2005:PCSc

[842] Thomas B. Jabine. Puzzle corner statistical cryptic puzzle no. 18. Chance, 18(3):57–58, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cork:2005:WWT

[843] Daniel Cork and Michael Cohen. Window on Washington: Tomato–tomahto: What to do about competing estimates? Chance, 18(3):59–62, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:AAc

[844] Anonymous. About the authors. Chance, 18(3):63–64, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:EBd

[845] Anonymous. Editorial Board. Chance, 18(4):2, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lavine:2005:EC

[846] Michael Lavine. Editor’s column. Chance, 18(4):3, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Keller-McNulty:2005:ITS

[847] S. Keller-McNulty, A. G. Wilson, and G. Wilson. The impact of technology on the scientific method. Chance, 18 (4):4–8, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See comments [848, 849, 850, 851, 852].

Ghosh-Dastidar:2005:CIT

[848] Bonnie Ghosh-Dastidar and Susan Paddock. Comment: “The impact of technology on the scientific method”. Chance, 18(4):9–12, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [847].

Landwehr:2005:CIT

[849] James M. Landwehr, Colin L. Mallows, and Patrick Tendick. Comment: “The impact of technology on the scientific method”. Chance, 18(4):12–13, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [847].

Karr:2005:CIT

[850] Alan F. Karr. Comment: “The impact of technology on the scientific method”. Chance, 18(4):13–15, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [847].

Nair:2005:CIT

[851] Vijay Nair. Comment: “The impact of technology on the scientific method”. Chance, 18(4):15–16, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [847].

Stufken:2005:CIT

[852] John Stufken. Comment: “The impact of technology on the scientific method”. Chance, 18(4):16, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [847].

Pollard:2005:SAU

[853] Graham Pollard. On solving an aspect of unfairness in the tennis doubles scoring system. Chance, 18(4):17–19, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stefani:2005:PDO

[854] Raymond T. Stefani. Politics, drugs, and the Olympics: And the winners are… politics and drugs. Chance, 18 (4):20–28, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Singer:2005:ADE

[855] Judith D. Singer. Afraid to discuss evolution? Chance, 18(4):29–31, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Pardoe:2005:JHP

[856] Iain Pardoe. Just how predictable are the Oscars? Chance, 18(4):32–39, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ioannidis:2005:WMP

[857] John P. A. Ioannidis. Why most published research findings are false. Chance, 18(4):40–47, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2005:BCF

[858] Howard Wainer. Chance conversation: Former Director of the U.S. Census Bureau gets personal. Chance, 18(4):48– 51, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2005:VRT

[859] Howard Wainer. Visual revelations: Truth is slower than fiction: Francis Galton as an illustration. Chance, 18 (4):52–54, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2005:HYHb

[860] Mark Glickman and James D. Neaton. Here’s to your health: Safeguards in studies: the role of data monitoring committees in clinical trials. Chance, 18(4):55–58, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2005:PCC

[861] Thomas B. Jabine. Puzzle corner cryptic acrostic no. 6. Chance, 18(4):59–61, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2005:AAd

[862] Anonymous. About the authors. Chance, 18(4):63–64, 2005. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:EBa

[863] Anonymous. Editorial Board. Chance, 19(1):2, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lavine:2006:EC

[864] Michael Lavine. Editor’s column. Chance, 19(1):3, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Singmaster:2006:IFP

[865] Rob Singmaster and Lissa Redmiles. The impact of the follow-up process on the 2002 Foreign Tax Credit Study Data. Chance, 19(1):4–9, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2006:ISD

[866] Leonard Evans. Innate sex differences supported by untypical traffic fatalities. Chance, 19(1):10–15, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See comments [867, 868] and reply [869].

Anonymous:2006:CDC

[867] Anonymous. Comment: David C. Geary. Chance, 19(1):16–17, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [866].

Anonymous:2006:CMM

[868] Anonymous. Comment: Mary Meyer. Chance, 19(1):18–19, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [866].

Anonymous:2006:ER

[869] Anonymous. Evans responds. Chance, 19(1):20–21, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Giesbrecht:2006:MEV

[870] Lee Giesbrecht, Jonaki Bose, and Michael P. Cohen. Measuring exposure to vehicle crashes. Chance, 19(1):22– 27, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schaffer:2006:EES

[871] Jay Schaffer and Erik L. Heiny. The effects of elevation on slugging percentage in major league baseball. Chance, 19(1):28–34, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Halberstadt:2006:GPH

[872] Steffanie Halberstadt and Stacey Wood. Global public health: The new frontier of statistics. Chance, 19(1):35– 40, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Noble:2006:FCF

[873] Robert B. Noble, A. John Bailer, and Lary L. Faris. Finding consolation in first-round losses. Chance, 19(1):41–48, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2006:VRF

[874] Howard Wainer, Marc Gessaroli, and Monica Verdi. Visual revelations: Finding what is not there through the unfortunate binning of results: The Mendel effect. Chance, 19(1):49–52, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2006:HYHa

[875] Mark Glickman and Michael Proschan. Here’s to your health: Clinical trials reliably test dietary claims. Chance, 19 (1):53–56, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2006:SRSa

[876] Scott M. Berry. A statistician reads the sports pages: Budgets and baseball concave or convex: Winning the salary game. Chance, 19(1):57–60, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2006:PCB

[877] Thomas B. Jabine. Puzzle corner from the beginning to the end: Two word puzzles. Chance, 19(1):61–62, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:AAa

[878] Anonymous. About the authors. Chance, 19(1):63–64, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:EBb

[879] Anonymous. Editorial Board. Chance, 19(2):2, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rodgers:2006:LE

[880] Joe Rodgers. Letter to the Editor. Chance, 19(2):3–4, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jones:2006:SSF

[881] Albyn Jones. Special section on forensic statistics grows from JSM session. Chance, 19(2):5–6, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson:2006:ATF

[882] David W. Peterson. Analyst takes fresh look at pay discrimination. Chance, 19(2):7–16, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Spiegelman:2006:DIS

[883] Clifford H. Spiegelman and Karen Kafadar. Data integrity and the scientific method: the case of bullet lead data as forensic evidence. Chance, 19(2):17–25, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Carriquiry:2006:CFA

[884] Alicia Carriquiry. Comment: Further arguments against CABL as a forensic tool. Chance, 19(2):25–26, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Egeland:2006:EMI

[885] T. Egeland, B. Kulle, and R. Andreassen. Esséen–Möller and identification based on DNA. Chance, 19(2): 27–31, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bura:2006:AFF

[886] Efstathia Bura and Joseph L. Gastwirth. Assessing the fairness of a firm’s allocation of shares in initial public offerings: Adapting a measure from biostatistics. Chance, 19(2):32–37, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kadane:2006:MBS

[887] Joseph B. Kadane. Misuse of Bayesian statistics in court. Chance, 19(2):38– 40, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See comment [888].

Becker:2006:C

[888] Elizabeth Becker. Comment. Chance, 19(2):41–42, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [887].

Khadjavi:2006:DWB

[889] Lily S. Khadjavi. Driving while black in the City of Angels. Chance, 19(2): 43–46, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Scanlan:2006:GEC

[890] James P. Scanlan. Guest editorial: Can we actually measure health disparities? Chance, 19(2):47–51, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2006:SRSb

[891] Scott M. Berry. A statistician reads the sports pages: Does defense win championships? Chance, 19(2):52–54, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2006:VRU

[892] Howard Wainer. Visual revelations: Using graphs to make the complex simple: the Medicare drug plan as an example. Chance, 19(2):55–56, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Christiansen:2006:HYH

[893] Cindy Christiansen and Valerie L. Durkalski. Here’s to your health: Virtual colonoscopy: the reality of colon polyp and colorectal cancer screening. Chance, 19(2):57–60, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2006:PCSa

[894] Thomas B. Jabine. Puzzle corner statistical cryptic puzzle no. 19. Chance, 19(2):61–62, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:AAb

[895] Anonymous. About the authors. Chance, 19(2):63–64, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:EBc

[896] Anonymous. Editorial Board. Chance, 19(3):2, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schwender:2006:DYS

[897] Holger Schwender, Sylvia Rabstein, and Katja Ickstadt. Do you speak genomish? Chance, 19(3):3–8, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See glossary [898].

Anonymous:2006:G

[898] Anonymous. Glossary. Chance, 19(3): 9–10, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [897].

Simmons:2006:BHM

[899] Susan J. Simmons and Ann E. Stapleton. Bayesian hierarchical models to detect quantitative trait loci. Chance, 19(3):11–14, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schwender:2006:CPM

[900] Holger Schwender and Anton Belousov. Comparison of preprocessing methods for Affymetrix microarrays. Chance, 19(3):15–20, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ickstadt:2006:AST

[901] Katja Ickstadt, Tina Mueller, and Holger Schwender. Analyzing SNPs: Are there needles in the haystack? Chance, 19(3):21–26, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Altman:2006:BSP

[902] Naomi Altman, Jim Leebens-Mack, Laura Zahn, André Chanderbali, Donglan Tian, Lillian Werner, Hong Ma, and Claude dePamphilis. Behind the scenes: Planning a multispecies microarray experiment. Chance, 19(3): 27–38, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wise:2006:YTA

[903] A. Wise, J. Hardin, and L. Hoopes. Yeast through the ages: a statistical analysis of genetic changes in aging yeast. Chance, 19(3):39–44, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Conlon:2006:CGA

[904] Erin M. Conlon, Gulhan Alpargu, and Jeffrey L. Blanchard. Comparative genomics approaches to identifying genetic regulatory networks. Chance, 19(3):45–48, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Huang:2006:UCA

[905] Haiyan Huang and Kyungpil Kim. Unsupervised clustering analysis of gene expression. Chance, 19(3):49–51, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2006:BCJ

[906] Howard Wainer. Chance conversation: Judith Tanur. Chance, 19(3):52–57, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2006:HYHb

[907] Mark Glickman, Ron Neath, and Brad Carlin. Here’s to your health: Does insurance status affect access to ER follow-up care? Chance, 19(3):58–60, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2006:PCSb

[908] Thomas B. Jabine. Puzzle corner statistical spiral cryptic no. 8. Chance, 19(3):61–62, 2006. CODEN CNDCE4.

ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:AAc

[909] Anonymous. About the authors. Chance, 19(3):63–64, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:EBd

[910] Anonymous. Editorial Board. Chance, 19(4):2, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:EL

[911] Anonymous. Editor’s letter. Chance, 19 (4):3, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fricker:2006:PAB

[912] Ronald D. Fricker, Jr. and Henry Rolka. Protecting against biological terrorism: Statistical issues in electronic biosurveillance. Chance, 19(4): 4–14, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Farmer:2006:ALM

[913] Charles M. Farmer. Another look at Meyer and Finney’s ‘Who Wants Airbags?’. Chance, 19(4):15–22, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [817].

Meyer:2006:CAL

[914] Mary Meyer. Commentary on “Another Look at Meyer and Finney’s ‘Who Wants Airbags?”’. Chance, 19 (4):23–24, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [817].

Allen:2006:MVW

[915] Theodore Allen and Mikhail Bernshteyn. Mitigating voter waiting times. Chance, 19(4):25–34, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sekhon:2006:CQV

[916] Jasjeet S. Sekhon. Comment: Queuing to vote in Franklin County, Ohio, in 2004. Chance, 19(4):35–36, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:AR

[917] Anonymous. Authors’ response. Chance, 19(4):36, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rump:2006:AMS

[918] Christopher M. Rump. Andrei Markov in the Stanley Cup playoffs. Chance, 19(4):37–42, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Miller:2006:HCS

[919] Jane E. Miller. How to communicate statistical findings: an expository writing approach. Chance, 19(4):43–49, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2006:SRSc

[920] Scott M. Berry. A statistician reads the sports pages: Statistical fallacies in sports. Chance, 19(4):50–56, 2006.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2006:HYHc

[921] Mark Glickman, Stuart G. Baker, Karen S. Lindeman, and Barnett S. Kramer. Here’s to your health: The paired availability design: If you can’t randomize, perhaps this applies. Chance, 19(4):57–60, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2006:PCSc

[922] Thomas B. Jabine. Puzzle corner statistical cryptic puzzle no. 20. Chance, 19(4):61–62, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2006:AAd

[923] Anonymous. About the authors. Chance, 19(4):63–64, 2006. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:EBa

[924] Anonymous. Editorial Board. Chance, 20(1):2, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:AAa

[925] Anonymous. About the authors. Chance, 20(1):3–4, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Marker:2007:LED

[926] David Marker. Letter to the Editor: Did the Democrats really win control of the Senate? Chance, 20(1):5, 2007.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Annis:2007:ICF

[927] David H. Annis and Samuel S. Wu. Improved college football scheduling using a modified Swiss system. Chance, 20(1):6–10, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ayres:2007:SSP

[928] Ian Ayres, Antonia R. Ayres-Brown, and Henry J. Ayres-Brown. Seeing significance: Is the 95% probability range easier to perceive? Chance, 20(1):11– 16, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Switzer:2007:WYD

[929] Suzanne S. Switzer and Nicholas J. Horton. What your doctor should know about statistics (but perhaps doesn’t…). Chance, 20(1):17–21, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Haugen:2007:IAS

[930] Kjetil K. Haugen. An improved award system for soccer: a (game-theoretic) comment. Chance, 20(1):22–24, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fernandez-Cantelli:2007:R

[931] Eduardo Fernandez-Cantelli and Glen Meeden. Reply. Chance, 20(1):24, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wang:2007:TST

[932] Steve C. Wang. Teaching statistical thinking using the baseball Hall of fame. Chance, 20(1):25–31, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Yang:2007:OZS

[933] Andrew Yang. Is Oprah Zulu? Sampling and seeming certainty in DNA ancestry testing. Chance, 20(1):32–39, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Miller:2007:DPW

[934] Steven J. Miller. A derivation of the Pythagorean won–loss formula in baseball. Chance, 20(1):40–48, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barber:2007:URS

[935] Jarrett J. Barber. Using rubber sheets to infer true map location. Chance, 20(1):49–54, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2007:VRIa

[936] Howard Wainer. Visual revelations: Insignificant is not zero: Rescoring the SAT as an example. Chance, 20(1):55– 58, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2007:HYHa

[937] Mark Glickman and R. Todd Ogden. Here’s to your health: Statistics and brain imaging in the study of mental illness. Chance, 20(1):59–62, 2007.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2007:PCSa

[938] Thomas B. Jabine. Puzzle corner statistical cryptic acrostic no. 7. Chance, 20(1):63–64, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:EBb

[939] Anonymous. Editorial Board. Chance, 20(2):2, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:AAb

[940] Anonymous. About the authors. Chance, 20(2):3–4, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lavine:2007:ELa

[941] Michael Lavine. Editor’s letter. Chance, 20(2):5, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barnett:2007:CEC

[942] Arnold Barnett and Edward H. Kaplan. A cure for the Electoral College? Chance, 20(2):6–9, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tolon:2007:EFD

[943] Mahmut Tolon and Herman Chernoff. The effect of fasting during Ramadan on traffic accidents in Turkey. Chance, 20(2):10–18, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Emerson:2007:CI

[944] John W. Emerson. Chance, on and off the ice. Chance, 20(2):19–21, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rathgeber:2007:WGW

[945] Andreas Rathgeber and Harriet Rathgeber. Why Germany was supposed to be drawn in the group of death and why it escaped. Chance, 20(2):22– 24, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hofmann:2007:ICC

[946] Heike Hofmann. Interview with a centennial chart. Chance, 20(2):26–35, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kyle:2007:EDB

[947] Susan Kyle, Douglas A. Samuelson, Fritz Scheuren, Nicole Vicinanza, Scott Dingman, and Warren Mitofsky. Explaining discrepancies between official votes and exit polls in the 2004 Presidential Election. Chance, 20(2):36–43, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See discussion [948, 949].

Sekhon:2007:CDT

[948] Jasjeet S. Sekhon. Comment: Data troubles: Explaining discrepancies between official votes and exit polls in the 2004 Presidential Election [MR 2370724]. Chance, 20(2):43–45, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [947].

Scheuren:2007:RM

[949] Fritz Scheuren. Rejoinder [MR2370724]. Chance, 20(2):45–47, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [947].

Everson:2007:SRSa

[950] Phil Everson. A statistician reads the sports pages: Describing uncertainty in games already played. Chance, 20(2): 49–56, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2007:VRG

[951] Howard Wainer. Visual revelations: Galton’s normal is too platykurtic. Chance, 20(2):57–58, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2007:HYHb

[952] Mark Glickman, A. John Bailer, Jane K. Straker, Robert B. Noble, Michael R. Hughes, Kyoungah See, and Cindy Christiansen. Here’s to your health: Considering nursing home quality: Is Places Rated or Consumer Reports a better model? Chance, 20(2): 59–62, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2007:PCSb

[953] Thomas B. Jabine. Puzzle corner statistical cryptic puzzle no. 21. Chance, 20(2):63–64, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:EBc

[954] Anonymous. Editorial Board. Chance, 20(3):2, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:AAc

[955] Anonymous. About the authors. Chance, 20(3):3–4, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lavine:2007:ELb

[956] Michael Lavine. Editor’s letter. Chance, 20(3):5, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lewis:2007:LE

[957] Hal Lewis. Letters to the Editor. Chance, 20(3):6–7, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2007:VRIb

[958] Howard Wainer. Visual revelations: Improving data displays: Ours and the media’s. Chance, 20(3):8–15, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Said:2007:QAA

[959] Yasmin H. Said and Edward J. Wegman. Quantitative assessments of alcohol-related outcomes. Chance, 20 (3):17–25, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fan:2007:AAD

[960] Juanjuan Fan and Richard A. Levine. To amnio or not to amnio: That is the decision for Bayes. Chance, 20(3):26– 32, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McDonald:2007:UPW

[961] John G. McDonald and Joe Lesko. Understanding phishing and why statisticians are needed. Chance, 20(3):33–37, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bordley:2007:SDM

[962] Robert Bordley. Statistical decisionmaking without math. Chance, 20(3): 39–44, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Meyer:2007:CAL

[963] Mary Meyer. Commentary on “Another look at Meyer and Finney’s ‘Who wants airbags?”’ [MR2149936]. Chance, 20(3):45–47, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). See [817].

Everson:2007:SRSb

[964] Phil Everson. A statistician reads the sports pages: Stein’s paradox revisited. Chance, 20(3):49–56, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2007:HYHc

[965] Mark Glickman, Stuart G. Baker, Jaakko Kaprio, and Cindy Christiansen. Here’s to your health: Genes, cancer, and twins: The statistics behind the headlines. Chance, 20(3):59– 62, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jabine:2007:PCT

[966] Thomas B. Jabine. Puzzle corner twoword loops: A word puzzle. Chance, 20 (3):64, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:EBd

[967] Anonymous. Editorial Board. Chance, 20(4):2, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2007:AAd

[968] Anonymous. About the authors. Chance, 20(4):3, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lavine:2007:ELc

[969] Michael Lavine. Editor’s letter. Chance, 20(4):5, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2007:ECA

[970] Howard Wainer and Andrew Gelman. Editorial: A catch-22 in assigning primary delegates. Chance, 20(4):6–7, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fienberg:2007:MEN

[971] Stephen E. Fienberg. Memories of Election Night predictions past: Psephologists and statisticians at work. Chance, 20(4):8–17, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2007:UBE

[972] Dalene Stangl. Using Chance to engage undergraduates in the study of statistics. Chance, 20(4):18–25, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stigler:2007:CYO

[973] Stephen M. Stigler. Chance is 350 years old. Chance, 20(4):26–30, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Straf:2007:MSS

[974] Miron L. Straf. Mr. Statistician schleps to Washington. Chance, 20(4):31–32, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2007:TCI

[975] Howard Wainer. Taking a CHANCE: An interview with William F. Eddy and Stephen E. Fienberg. Chance, 20(4): 33–39, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stern:2007:BDS

[976] Hal Stern and Adam Sugano. Baseball decisions and small samples. Chance, 20(4):40–47, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eddy:2007:ILB

[977] William F. Eddy, Rebecca McNamee, and Kary L. Myers. Imaging the living brain. Chance, 20(4):48–58, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2007:WLS

[978] Andrew Gelman and Seth Roberts. Weight loss, self-experimentation, and Web trials: A conversation. Chance, 20(4):59–63, 2007. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:EBa

[979] Anonymous. Editorial Board. Chance, 21(1):2, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:AAa

[980] Anonymous. About the authors. Chance, 21(1):3–4, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2008:ELa

[981] Mike Larsen. Editor’s letter. Chance, 21(1):5, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-008-0046-y.pdf.

Proschan:2008:SEW

[982] Michael Proschan. Self-experimentation and Web trials. Chance, 21(1):7–9, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0001-y.

Heiny:2008:PTP

[983] Erik L. Heiny. PGA tour pro: Long but not so straight. Chance, 21(1):11–21, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0005-7.

Peterson:2008:PCW

[984] David W. Peterson. Putting chance to work: Reducing the politics in political redistricting. Chance, 21(1):22–26, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0006-6.

Lunsford:2008:ASU

[985] M. Leigh Lunsford, Ginger Holmes Rowell, and Tracy Goodson-Espy. Assessment of student understanding of the Central Limit Theorem. Chance, 21(1):27–32, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0007-5.

Kadane:2008:LFC

[986] Joseph B. Kadane. Load factors, crowdedness, and the Cauchy–Schwarz inequality. Chance, 21(1):33–34, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00144-008-0008-4.

Niemi:2008:CSN

[987] Jarad B. Niemi, Bradley P. Carlin, and Jonathan M. Alexander. Contrarian strategies for NCAA tournament pools: A cure for March madness? Chance, 21(1):35–42, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0009-3.

Moore:2008:GDE

[988] Kris Moore, Dawn Carlson, Dwayne Whitten, and Aimee Clement. Gender differences at the executive level: Perceptions and experiences. Chance, 21(1):43–48, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0010-x.

Sink:2008:FEG

[989] Christopher Sink, Matthew Sink, Jonathan Stob, and Kevin Taniguchi. Further evidence of gender differences in high school–level computer literacy. Chance, 21(1):49–53, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0011-9.

Wainer:2008:VRU

[990] Howard Wainer and Sam Savage. Visual revelations: Until proven guilty: False positives and the war on terror. Chance, 21(1):55–58, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0012-8.

Freedman:2008:EOM

[991] David A. Freedman. Editorial: Oasis or mirage? Chance, 21(1):59–61, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00144-008-
0047-x.

Jabine:2008:PCSa

[992] Thomas B. Jabine. Puzzle corner: Statistical spiral cryptic no. 9. Chance, 21(1):62–64, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0013-7.

Anonymous:2008:HCa

[993] Anonymous. Help & contacts. Chance, 21(1):??, March 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:EBb

[994] Anonymous. Editorial Board. Chance, 21(2):2, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:AAb

[995] Anonymous. About the authors. Chance, 21(2):3, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2008:ELb

[996] Mike Larsen. Editor’s letter. Chance, 21(2):4, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-008-0048-9.pdf.

Banks:2008:SMC

[997] David Banks. Statisticians and metabolomics: collaborative possibilities for the next *omics revolution? Chance, 21(2):5–11, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0002-x.

Dixon:2008:MDU

[998] Philip M. Dixon and Geng Ding. A metabolomic data uncertainty budget for the plant Arabidopsis thaliana. Chance, 21(2):12–17, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://link.springer.com/
article/10.1007/s00144-008-0014-
6.

Ash:2008:FCS

[999] Arlene Ash and John Lamperti. Florida 2006: Can statistics tell us who won Congressional District-13? Chance, 21(2):18–24, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0015-5.

Hall:2008:SSE

[1000] Joseph Lorenzo Hall. Statistical solutions to election mysteries. Chance, 21(2):24–25, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-008-0016-
4.

Mebane:2008:CFV

[1001] Walter R. Mebane, Jr. Counting frustrated voter intentions. Chance, 21(2):26–27, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/

article/10.1007/s00144-008-0017-
3.

Pinkham:2008:HPM

[1002] Roger Pinkham. A historical puzzle:

de Mere’s paradox revisited. Chance, 21(2):28–30, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0018-2.

Johnson:2008:CSC

[1003] H. Dean Johnson, Sarah Boslaugh, Murray Clayton, Chris Holloman, and Linda J. Young. Compensation for statistical consulting services: Approaches used at four U.S. universities. Chance, 21(2):31–35, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0019-1.

Louis:2008:CSC

[1004] Thomas A. Louis. Compensation for statistical consulting services: Observations and recommendations. Chance, 21(2):36–37, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-008-0020-
8.

Derr:2008:SCU

[1005] Janice Derr. Statistical consulting in a University setting: Don’t forget the students. Chance, 21(2):38–39, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00144-008-0021-7.

Gentleman:2008:NHI

[1006] Jane F. Gentleman. The National Health Interview Survey: 50 years and going strong. Chance, 21(2):40–45, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00144-008-
                                        0022-6.

Wainer:2008:VRI

[1007] Howard Wainer. Visual revelations: Improving graphic displays by controlling creativity. Chance, 21(2):46–52, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00144-008-

0023-5. See commentary [1008].

Pickle:2008:CIG

[1008] Linda Williams Pickle. Commentary on ‘Improving graphic displays by controlling creativity’. Chance, 21(2):53, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00144-
                                        008-0049-8.pdf. See [1007].

Glickman:2008:HYHa

[1009] Mark Glickman, Gerardo Chowell, N. W. Hengartner, Catherine Ammon, and Mac Hyman. Here’s to your health: Learning from the past to prepare for the future: Modeling the impact of hypothetical interventions during the Great Influenza Pandemic of 1918. Chance, 21(2):55–60, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00144-008-
0044-0.

Larsen:2008:ITJ

[1010] Michael Larsen. An interview with Thomas Jabine, retiring Puzzle Corner Editor. Chance, 21(2):61–62, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00144-008-0045-z.

Jabine:2008:PCSb

[1011] Thomas B. Jabine. Puzzle corner: Statistical cryptic puzzle no. 22. Chance, 21(2):63–64, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-008-0050-
2.

Anonymous:2008:HCb

[1012] Anonymous. Help & contacts.

Chance, 21(2):??, June 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:EBc

[1013] Anonymous. Editorial Board. Chance, 21(3):2, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:AAc

[1014] Anonymous. About the authors.

Chance, 21(3):3–4, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2008:ELc

[1015] Mike Larsen. Editor’s letter. Chance, 21(3):5, September 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s144-008-0051-1.pdf.

Anonymous:2008:C

[1016] Anonymous. Correction. Chance, 21(3):6, September 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s144-008-0054-y.pdf.

Anonymous:2008:LEB

[1017] Anonymous. Letter to the Editor: The best graph may be no graph. Chance, 21(3):6, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berry:2008:BGM

[1018] Donald Berry and Frank T. Mc-Graw. The best graph may be no graph. Chance, 21(3):6, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s144-008-0052-0.pdf.

Krenzke:2008:FBS

[1019] Tom Krenzke and David Judkins. Filling in the blanks: some guesses are better than others. Illustrating the impact of covariate selection when imputing complex survey items. Chance, 21(3):7–13, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-008-0003-9.

Elliott:2008:HLA

[1020] Michael R. Elliott. Healthy for life:

accounting for transcription errors using multiple imputation. Application to a study of childhood obesity. Chance, 21(3):14–23, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-008-0024-4.

Bradlow:2008:SLR

[1021] Eric T. Bradlow, Shane T. Jensen, Justin Wolfers, and Abraham J. Wyner. A statistical look at Roger Clemens’ pitching career. Chance, 21(3):24–30, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-008-0025-3.

Freeman:2008:AFF

[1022] Peter Freeman, Joseph Richards, Chad Schafer, and Ann Lee. Astrostatistics: The final frontier. Chance, 21(3):31–35, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://link.springer.com/
article/10.1007/s144-008-0026-2.

Puntanen:2008:SSP

[1023] Simo Puntanen and George P. H.

Styan. Stochastic stamps: A philatelic introduction to chance. Chance, 21(3):36–41, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-008-0027-1.

Olofsson:2008:PSE

[1024] Peter Olofsson. Probability, statistics, evolution, and intelligent design. Chance, 21(3):42–45, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s144-008-0028-0.

Everson:2008:SRS

[1025] Phil Everson. A statistician reads the sports pages: Good offense vs. poor defense: The 2007 Women’s World Cup. Chance, 21(3):46–54, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s144-008-0029-z.

Glickman:2008:HYHb

[1026] Mark Glickman, Yulei He, Recai Yucel, and Alan M. Zaslavsky. Here’s to your health: Misreporting, missing data, and multiple imputation: Improving accuracy of cancer registry databases. Chance, 21(3):55–58, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s144-008-0030-6.

Wainer:2008:VRG

[1027] Howard Wainer, Grace Lee, and Paul Velleman. Visual revelations: Giving the finger to dating services. Chance, 21(3):59–61, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s144-008-0031-5.

Berkowitz:2008:GWTa

[1028] Jonathan Berkowitz. Goodness of wit test: “If nothing ever changed, there’d be no butterflies”. Chance, 21(3):62–64, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-008-0032-4.

Anonymous:2008:HCc

[1029] Anonymous. Help & contacts. Chance, 21(3):??, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:EBd

[1030] Anonymous. Editorial Board. Chance, 21(4):2, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2008:AAd

[1031] Anonymous. About the authors.

Chance, 21(4):3–4, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2008:ELd

[1032] Mike Larsen. Editor’s letter. Chance, 21(4):5, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00144-
008-0053-z.pdf.

Clauser:2008:WES

[1033] Brian E. Clauser. War, enmity, and statistical tables. Chance, 21(4):6– 11, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00144-008-
0004-8.

Stigler:2008:FL

[1034] Stephen Stigler. Fisher and the 5% level. Chance, 21(4):12, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-008-0033-3.pdf.

Alam:2008:HDP

[1035] Moudud Alam, Kenneth Carling, Rui Chen, and Yuli Liang. How to determine the progression of Young skiers? Chance, 21(4):13–19, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0034-2.

Hurley:2008:BMP

[1036] W. J. Hurley. The birthday matching problem when the distribution of birthdays is nonuniform. Chance, 21(4):20– 24, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0035-1.

Croson:2008:PSS

[1037] Rachel Croson, Peter Fishman, and Devin G. Pope. Poker superstars: skill or luck? Similarities between golf — thought to be a game of skill — and poker. Chance, 21(4):25– 28, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0036-0.

Price:2008:VST

[1038] Phillip Price. The value of statistical thinking: A basketball gambling rookie conquers March madness. Chance, 21(4):29–36, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0037-z.

Schmotzer:2008:WLS

[1039] Brian Schmotzer. When is the lead safe in a college basketball game? Chance, 21(4):37–44, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-008-0038-y.

McGivney:2008:LIP

[1040] Katherine McGivney, Ray McGivney, and Ralph Zegarelli. Light it up: Predicting the winner of an NBA game before the end. Chance, 21(4):45– 50, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0039-x.

Glickman:2008:HYHc

[1041] Mark Glickman and Stephanie R.

Land. Here’s to your health: ‘we all survived’ and other failings of risk perception. Chance, 21(4):51– 55, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-008-
0040-4.

Wainer:2008:VRL

[1042] Howard Wainer and Paul Velleman. Visual revelations: Looking at blood sugar. Chance, 21(4):56–61, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00144-008-0041-
                                        3.

Anonymous:2008:BGD

[1043] Anonymous. Chance graphic display contest: Burtin’s antibiotic data. Chance, 21(4):62, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00144-008-0042-2.pdf.

Berkowitz:2008:GWTb

[1044] Jonathan Berkowitz. Goodness of wit test. Chance, 21(4):63–64, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00144-008-0043-1.

Anonymous:2008:HCd

[1045] Anonymous. Help & contacts. Chance, 21(4):??, 2008. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:EBa

[1046] Anonymous. Editorial Board. Chance, 22(1):2, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:AAa

[1047] Anonymous. About the authors. Chance, 22(1):3–4, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2009:ELa

[1048] Mike Larsen. Editor’s letter. Chance, 22(1):5, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s144-009-0011-4.pdf.

Keppel:2009:HPM

[1049] Kenneth G. Keppel and Jeffrey N.

Pearcy. Healthy people 2010: Measuring disparities in health. Chance, 22(1):6–9, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-009-0001-6.

Caffo:2009:OOS

[1050] Brian Caffo, Bruce Swihart, Ciprian Crainiceanu, Alison Laffan, and Naresh Punjabi. An overview of observational sleep research with application to sleep stage transitioning. Chance, 22(1):10–15, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-009-0002-5.

Slaven:2009:SMS

[1051] James E. Slaven, Michael E. Andrew, Anna Mnatsakanova, John M. Violanti, Cecil M. Burchfiel, and Bryan J. Vila. Statistical modeling of sleep. Chance, 22(1):16–21, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-009-0003-4.

Connolly:2009:DC

[1052] Michele Connolly. Disability: It’s complicated. Chance, 22(1):22–27, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s144-009-0004-3.

Hurley:2009:JJR

[1053] W. J. Hurley. Jussi Jokinen, regression to the mean, and the assessment of exceptional performance. Chance, 22(1):28–33, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s144-009-0005-2.

Clevenson:2009:GIW

[1054] M. Lawrence Clevenson and Jennifer Wright. Go for it: What to consider when making fourth-down decisions. Chance, 22(1):34–41, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s144-009-0006-1.

Cherkassky:2009:AML

[1055] Michael Cherkassky. Application of machine learning methods to medical diagnosis. Chance, 22(1):42–50, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s144-009-0007-0.

Wainer:2009:VRC

[1056] Howard Wainer. Visual revelations: A centenary celebration for Will Burtin: A pioneer of scientific visualization. Chance, 22(1):51–55, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://link.springer.com/
article/10.1007/s144-009-0008-z.

Glickman:2009:HYHa

[1057] Mark Glickman, Todd G. Nick, and Shannon N. Saldaña. Here’s to your health: Pharmacogenetics and pharmacogenomics: Statistical challenges in design and analysis. Chance, 22 (1):56–62, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s144-009-0009-y.

Berkowitz:2009:GWTa

[1058] Jonathan Berkowitz. Goodness of wit test. Chance, 22(1):63–64, March 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s144-009-0010-5.

Anonymous:2009:HCa

[1059] Anonymous. Help & contacts. Chance, 22(1):??, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:EBb

[1060] Anonymous. Editorial Board. Chance, 22(2):2, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:AAb

[1061] Anonymous. About the authors.

Chance, 22(2):3, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-009-0012-3.pdf.

Larsen:2009:ELb

[1062] Mike Larsen. Editor’s letter. Chance, 22(2):5, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-009-0013-2.pdf.

Aref:2009:LE

[1063] Susanne Aref and Ray Stefani. Letters to the Editor. Chance, 22(2):6–7, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00144-
009-0014-1.pdf.

Elliott:2009:HLA

[1064] Michael R. Elliott. Healthy for life: Accounting for transcription errors using multiple imputation. Chance, 22(2):7, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00144-
009-0022-1.pdf.

Albert:2009:RCW

[1065] Jim Albert. Is Roger Clemens’ WHIP trajectory unusual? Chance, 22(2):8–20, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0015-0.

Schmotzer:2009:NES

[1066] Brian Schmotzer, Patrick D. Kilgo, and Jeff Switchenko. ‘The Natural’? The effect of steroids on offensive performance in baseball. Chance, 22(2):21–32, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0016-z.

Chance:2009:WOA

[1067] Don M. Chance. What are the odds? Another look at DiMaggio’s streak. Chance, 22(2):33–42, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-009-
0017-y.

Larsen:2009:ABG

[1068] Mike Larsen. Announcing the Burtin graphic contest winners. Chance, 22(2):43–45, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0018-x.

Wainer:2009:VRP

[1069] Howard Wainer and Mike Larsen. Visual revelations: Pictures at an exhibition. Chance, 22(2):46–54, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-009-
0019-9.

Glickman:2009:HYHb

[1070] Mark Glickman, Xiaofei Wang, Herbert Pang, and Todd A. Schwartz. Here’s to your health: Building and validating high throughput lung cancer biomarkers. Chance, 22(2):55–62, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00144-009-
                                        0020-3.

Berkowitz:2009:GWTb

[1071] Jonathan Berkowitz. Goodness of wit test. Chance, 22(2):63–64, June 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00144-009-0021-2.

Anonymous:2009:HCb

[1072] Anonymous. Help & contacts. Chance, 22(2):??, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:EBc

[1073] Anonymous. Editorial Board. Chance, 22(3):2, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:AAc

[1074] Anonymous. About the authors. Chance, 22(3):3–4, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:ELa

[1075] Anonymous. Editor’s letter. Chance, 22(3):5, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2009:ELc

[1076] Mike Larsen. Editor’s letter. Chance, 22(3):5, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-009-0023-0.pdf.

Quade:2009:LE

[1077] Dana Quade. Letter to the Editor. Chance, 22(3):6–7, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00144-009-0024-z.

Khan:2009:ACC

[1078] Shahedul Ahsan Khan, Grace Chiu, and Joel A. Dubin. Atmospheric concentration of chlorofluorocarbons: Addressing the Global Concern with the longitudinal bent-cable model. Chance, 22(3):8–17, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0026-x.

Westreich:2009:CSI

[1079] Daniel Westreich. Comic: A statistical investigation of urban skylines in Illinois. Chance, 22(3):16, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaw:2009:MQC

[1080] Autar K. Kaw and Ali Yalcin. A metric to quantify college football’s topsy-turvy season. Chance, 22(3): 18–26, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0027-9.

Barnett:2009:GYW

[1081] Tristan Barnett. Gambling your way to new York: A story of expectation. Chance, 22(3):27–34, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-009-
0028-8.

Baccini:2009:HSJ

[1082] Alberto Baccini, Lucio Barabesi, and Marzia Marcheselli. How are statistical journals linked? A network analysis. Chance, 22(3):35–45, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-009-
0029-7.

Bilder:2009:HCG

[1083] Christopher R. Bilder. Human or cylon? Group testing on ‘Battlestar Galactica’. Chance, 22(3): 46–50, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0030-1.

Wainer:2009:VRL

[1084] Howard Wainer and Peter Baldwin. Visual revelations: A little ignorance: How statistics rescued a damsel in distress. Chance, 22(3):51–55, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-009-
0031-0.

Glickman:2009:HYHc

[1085] Mark Glickman and Scott Evans.

Here’s to your health: Noninferiority clinical trials. Chance, 22(3): 56–62, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0032-z.

Berkowitz:2009:GWTc

[1086] Jonathan Berkowitz. Goodness of wit test #5: Simple inference. Chance, 22 (3):63–64, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-009-0033-
y.

Anonymous:2009:HCc

[1087] Anonymous. Help & contacts. Chance, 22(3):??, September 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:EBd

[1088] Anonymous. Editorial Board. Chance, 22(4):2, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2009:AAd

[1089] Anonymous. About the authors.

Chance, 22(4):3–4, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-009-0034-

Anonymous:2009:ELb

[1090] Anonymous. Editor’s letter. Chance, 22 (4):5, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00144-
                                        009-0035-9.pdf.

Larsen:2009:ELd

[1091] Mike Larsen. Editor’s letter. Chance, 22(4):5, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Labby:2009:WDA

[1092] Zacariah Labby. Weldon’s dice, automated. Chance, 22(4):6–13, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00144-009-0036-
                                        8.

Min:2009:MAE

[1093] Wanli Min and Gang Luo. Medical applications of EEG wave classification. Chance, 22(4):14–20, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00144-009-0037-7.

Vars:2009:MWO

[1094] Fredrick E. Vars. Missing well: Optimal targeting of soccer shots. Chance, 22(4):21–28, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00144-009-0038-6.

Symanzik:2009:RCP

[1095] Jürgen Symanzik. A real Challenger of a puzzle. Chance, 22(4): 28, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00144-
009-0039-5.pdf. See [1128].

Schilling:2009:DME

[1096] Mark F. Schilling. Does momentum exist in competitive volleyball? Chance, 22(4):29–35, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0040-z.

Hurley:2009:COM

[1097] W. J. Hurley. Calculating the odds of the miracle at Brookline. Chance, 22(4):36–38, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0041-y.

Nguyen:2009:ESR

[1098] Andreas Nguyen and Shenghua K. Fan.

Ethics and stopping rules in a phase II clinical trial. Chance, 22(4):39– 44, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-009-
0042-x.

Haug:2009:EFM

[1099] Mark G. Haug and Heather Ardery.

Entangling finance, medicine, and law. Chance, 22(4):45–50, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0043-9.

Lewis:2009:PFD

[1100] Virginia Vimpeny Lewis. Pascal and Fermat divide the stakes. Chance, 22(4):51–54, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0044-8.

Wainer:2009:VRG

[1101] Howard Wainer. Visual revelations: A good table can beat a bad graph: It matters who plays Mozart. Chance, 22(4):55–57, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-009-0045-7.

Glickman:2009:HYHd

[1102] Mark Glickman, Douglas A. Noe, Ian M. Nelson, Shahla Mehdizadeh, and A. John Bailer. Here’s to your health: Will they stay or will they go? Predicting disenrollment from home care services to nursing homes using classification trees. Chance, 22(4):58– 62, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-009-
0046-6.

Berkowitz:2009:GWTd

[1103] Jonathan Berkowitz. Goodness of wit test #6: Standardizing conflict. Chance, 22(4):63–64, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/

article/10.1007/s00144-009-0047-
5.

Anonymous:2009:HCd

[1104] Anonymous. Help & contacts. Chance, 22(4):??, 2009. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:EBa

[1105] Anonymous. Editorial Board. Chance, 23(1):2, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:AAa

[1106] Anonymous. About the authors.

Chance, 23(1):3–4, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-010-0001-6.pdf.

Anonymous:2010:ELa

[1107] Anonymous. Editor’s letter. Chance, 23 (1):5, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00144-
010-0002-5.pdf.

Larsen:2010:ELa

[1108] Mike Larsen. Editor’s letter. Chance, 23(1):5, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mebane:2010:FPE

[1109] Walter R. Mebane, Jr. Fraud in the 2009 Presidential Election in Iran? Chance, 23(1):6–15, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0003-4.

Baccini:2010:MIA

[1110] Michela Baccini, Samantha Cook, Constantine E. Frangakis, Fan Li, Fabrizia Mealli, Donald B. Rubin, and Elizabeth R. Zell. Multiple imputation in the anthrax vaccine research program. Chance, 23(1):16–23, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0004-3.

Bell:2010:ATN

[1111] Robert M. Bell, Yehuda Koren, and Chris Volinsky. All together now: A perspective on the Netflix prize. Chance, 23(1):24, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-010-0005-2.pdf.

Lohr:2010:CSC

[1112] Steve Lohr. The contest that shaped careers and inspired research papers. Chance, 23(1):25–29, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0006-1.

Remund:2010:AED

[1113] Todd G. Remund. Analysis of engineering data: The case of rocket firings. Chance, 23(1):30–36, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0007-0.

Gelbrich:2010:GBP

[1114] Götz Gelbrich, Annegret Franke, Bianca Gelbrich, Susann Blüher, and Markus Löffler. Gummy bears and the pleasure of biostatistics lessons. Chance, 23(1):37–40, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0008-z.

Kvam:2010:EVP

[1115] Paul Kvam. Electoral voting and population distribution in the United States. Chance, 23(1):41–47, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0009-y.

Glickman:2010:HYHa

[1116] Mark Glickman. Here’s to your health: The virtual patient: Developing drugs using modeling and simulation. Chance, 23(1):48–52, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0010-5.

Wainer:2010:VRS

[1117] Howard Wainer. Visual revelations:

Schrödinger’s cat and the conception of probability in item response theory. Chance, 23(1):53–56, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0011-4.

Loly:2010:CPLa

[1118] Peter D. Loly and George P. H. Styan. Comments on 4 × 4 philatelic latin squares. Chance, 23(1):57–62, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00144-010-
                                        0012-3.

Berkowitz:2010:GWTa

[1119] Jonathan Berkowitz. Goodness of wit test #7: Figure it out. Chance, 23(1): 63–66, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00144-010-
                                        0013-2.

Anonymous:2010:HCa

[1120] Anonymous. Help & contacts. Chance, 23(1):??, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:EBb

[1121] Anonymous. Editorial Board. Chance, 23(2):2, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:AAb

[1122] Anonymous. About the authors. Chance, 23(2):3–4, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:ELb

[1123] Anonymous. Editor’s letter. Chance, 23(2):5, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-010-0015-0.pdf.

Larsen:2010:ELb

[1124] Mike Larsen. Editor’s letter. Chance, 23(2):5, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Asher:2010:CDC

[1125] Jana Asher. Collecting data in challenging settings: In the Global South — say, East Timor — data collection is not for the faint of heart or weak of stomach. Chance, 23(2):6–13, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0016-z.

Zheng:2010:LDD

[1126] Qi Zheng. The Luria–Delbrück distribution: Early statistical thinking about evolution. Chance, 23(2):15–18, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0017-y.

Finch:2010:UIR

[1127] Holmes Finch. Using item response theory to understand gender differences in opinions on women in politics. Chance, 23(2):19–24, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0018-x.

Symanzik:2010:RRC

[1128] Jürgen Symanzik, Stephanie Kovalchik, and Brad Thiessen. Results of “A Real Challenger of a Puzzle” graphics contest. Chance, 23(2):25–27, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-

0019-9. See [1095].

Dillard:2010:HWT

[1129] Bernard Dillard. Hey, who turned off the lights? A look at electricity consumption. Chance, 23(2):28–37, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0020-3.

Petras:2010:LSL

[1130] Ivo Petras and Igor Podlubny. Least squares or least circles? A comparison of classical regression and orthogonal regression. Chance, 23(2):38–42, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0021-2.

Bohannon:2010:CPD

[1131] John Bohannon, Robin Goldstein, and Alexis Herschkowitsch. Can people distinguish pâté from dog food? Chance, 23(2):43–46, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0022-1.

Wainer:2010:VRC

[1132] Howard Wainer. Visual revelations:

Commentary on the graphic displays in the 2008 National Healthcare Quality Report and State Snapshots. Chance, 23(2):47–53, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0023-0.

Smeltzer:2010:OHB

[1133] Ronald K. Smeltzer. One for the history books: An early time-line bar graph. Chance, 23(2):54–56, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0024-z.

Loly:2010:CPLb

[1134] Peter D. Loly and George P. H. Styan.

Comments on 5 × 5 philatelic latin squares. Chance, 23(2):57–62, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0025-y.

Berkowitz:2010:GWTb

[1135] Jonathan Berkowitz. Goodness of wit test #8: Employs magic. Chance, 23(2):63–64, June 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-010-0026-

Anonymous:2010:HCb

[1136] Anonymous. Help & contacts. Chance, 23(2):??, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:EBc

[1137] Anonymous. Editorial Board. Chance, 23(3):2, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:AAc

[1138] Anonymous. About the authors. Chance, 23(3):3–4, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mosley:2010:VME

[1139] Lawrence Mosley, Dianne Cook, Heike Hofmann, Chris Kielion, and Barret Schloerke. Visually monitoring the 2008 election. Chance, 23(3):4, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2010:ELc

[1140] Mike Larsen. Editor’s letter. Chance, 23(3):5, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rothschild:2010:DEA

[1141] David Rothschild. Debia$ed aggregated poll$ and prediction market price$. Chance, 23(3):6–7, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00144-010-0039-5.

Gelman:2010:VE

[1142] Andrew Gelman and Yu-Sung Su. Voting by education in 2008. Chance, 23(3):8, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-010-0040-z.pdf.

Stark:2010:RLV

[1143] Philip B. Stark. Risk-limiting votetabulation audits: The importance of cluster size. Chance, 23(3): 9–12, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0041-y.

Hobbs:2010:FCV

[1144] Jonathan Hobbs, Luke Fostvedt, Adam Pintar, David Rockoff, Eunice Kim, and Randy Griffiths. Fixed-cost vs. fixed-risk post-election audits in Iowa. Chance, 23(3):13–17, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0042-x.

Young:2010:EAF

[1145] Linda J. Young and Dan McCrea. Election audits in Florida. Chance, 23 (3):18–19, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-010-0043-
9.

Redelmeier:2010:RCN

[1146] Donald A. Redelmeier and Robert J. Tibshirani. Road crashes and the next U.S. Presidential Election. Chance, 23 (3):20–24, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0044-8.

McCarthy:2010:CPS

[1147] David McCarthy, David Groggel, and A. John Bailer. Career pitching statistics and the probability of throwing a no-hitter in MLB: A case-control study. Chance, 23(3):25–33, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0045-7.

Berthet:2010:BHW

[1148] Vincent Berthet. Best hand wins: How poker is governed by chance. Chance, 23(3):34–38, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0046-6.

Jones:2010:CBP

[1149] Martin L. Jones and Ryan J. Parker. The contribution of Star Players in the NBA: A story of misbehaving coefficients. Chance, 23(3):39–45, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0047-5.

Mulry:2010:OCC

[1150] Mary H. Mulry and Patrick J.

Cantwell. Overview of the 2010 Census Coverage Measurement Program and its evaluations. Chance, 23(3): 46–51, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0048-4.

Glickman:2010:HYHb

[1151] Mark Glickman and Arlene S. Swern.

Here’s to your health: A story of evidence-based medicine: Hormone replacement therapy and coronary heart disease in postmenopausal women. Chance, 23(3):52–56, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0049-3.

Wainer:2010:VRI

[1152] Howard Wainer and Jim Ramsay.

Visual revelations: Inside-out plots. Chance, 23(3):57–62, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://link.springer.com/
article/10.1007/s00144-010-0050-
x.

Berkowitz:2010:GWTc

[1153] Jonathan Berkowitz. Goodness of wit test #9: In one ear and out the other. Chance, 23(3):63–64, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00144-010-0051-9.

Anonymous:2010:HCc

[1154] Anonymous. Help & contacts. Chance, 23(3):??, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:EBd

[1155] Anonymous. Editorial Board. Chance, 23(4):2, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2010:AAd

[1156] Anonymous. About the authors. Chance, 23(4):3–4, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://link.springer.com/
                                        accesspage/article/10.1007/s00144-
                                        010-0027-9.

Anonymous:2010:ELc

[1157] Anonymous. Editor’s letter. Chance, 23(4):5, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00144-010-0028-8.pdf.

Larsen:2010:ELd

[1158] Mike Larsen. Editor’s letter. Chance, 23(4):5, 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Crowley:2010:WJG

[1159] Jason Crowley, Brenna Curley, and Dave Osthus. What is Jeopardy!? A graphical exploration from 1984– 2009. Chance, 23(4):6–14, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0029-7.

Doganaksoy:2010:VPR

[1160] Necip Doganaksoy, Gerald J. Hahn, and William Q. Meeker. Validating product reliability. Chance, 23 (4):15–21, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0030-1.

Lent:2010:SDE

[1161] Janice Lent. Statistics in a dynamic energy environment. Chance, 23(4): 22–28, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0031-0.

Allen:2010:IHE

[1162] Terry Allen, Amber Thom, and Glen Buckner. Infant homicides: An examination using multiple correspondence analysis. Chance, 23(4):29–34, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0032-z.

Glickman:2010:SCL

[1163] Mark Glickman, Rick Ittenbach, Todd G. Nick, Ralph O’Brien, Sarah J. Ratcliffe, and Justine Shults. Statistical consulting with limited resources: Applications to practice. Chance, 23 (4):35–42, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0033-y.

Stark:2010:NVC

[1164] Philip B. Stark. Null and vetoed: Chance coincidence? Governor Schwarzenegger’s acrostic veto. Chance, 23(4):43–46, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://link.springer.com/
article/10.1007/s00144-010-0034-
x.

Huber:2010:MAL

[1165] Michael Huber. Mick has (almost) left the building. Chance, 23(4): 47–51, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0035-9.

McCotter:2010:HSD

[1166] Trent McCotter. Hitting streaks don’t obey your rules: Evidence that hitting streaks are not just byproducts of random variation. Chance, 23 (4):52–57, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0036-8.

Wainer:2010:VRP

[1167] Howard Wainer. Visual revelations: Pies, spies, roses, lines, and symmetries. Chance, 23(4):58–61, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-010-
0037-7.

Berkowitz:2010:GWTd

[1168] Jonathan Berkowitz. Goodness of wit test #10: Once is enough. Chance, 23 (4):62–64, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-010-0038-6.

Anonymous:2010:HCd

[1169] Anonymous. Help & contacts. Chance, 23(4):??, September 2010. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:EBa

[1170] Anonymous. Editorial Board. Chance, 24(1):2, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:AAa

[1171] Anonymous. About the authors.

Chance, 24(1):3–4, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:ELa

[1172] Anonymous. Editor’s letter. Chance, 24 (1):5, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00144-
011-0001-1.pdf.

Behseta:2011:ELa

[1173] Sam Behseta. Editor’s letter. Chance, 24(1):5, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:LE

[1174] Anonymous. Letter to the Editor. Chance, 24(1):7, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Feitelson:2011:LE

[1175] Dror Feitelson. Letter to the editor. Chance, 24(1):7, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-011-0012-y.pdf.

Stigler:2011:GVB

[1176] Stephen M. Stigler. Galton visualizing Bayesian inference. Chance, 24(1):8– 10, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0002-0.

Wainer:2011:VAM

[1177] Howard Wainer. Value-added models to evaluate teachers: A Cry for Help. Chance, 24(1):11–13, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0003-z.

Rockoff:2011:JDD

[1178] David Rockoff and Philip Yates. Joe DiMaggio done it again … and again and again and again? Chance, 24(1): 14–18, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0004-y.

Davis:2011:AST

[1179] Harry Zvi Davis, Hershey H. Friedman, and Jianming Ye. An ancient sampling technique: Flawed, surprisingly good, or optimal? Chance, 24(1):19– 23, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0005-x.

Burks:2011:JBM

[1180] Robert E. Burks. Jelly belly mixing and uniform multinomial distribution. Chance, 24(1):24–28, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0006-9.

Wenz:2011:TCT

[1181] Michael Wenz and Joren Skugrud.

Tackling the chart: Two-point conversions and team differences in football. Chance, 24(1):29–35, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0007-8.

Liu:2011:ERW

[1182] Hongmei Liu, Jay Parker, and Wei Sun. Estimating rates at which books are mis-shelved. Chance, 24(1):36– 43, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0008-7.

Loyer:2011:CPE

[1183] Milton W. Loyer and Gene D. Sprechini. Can the probability of an event be larger or smaller than each of its component conditional probabilities? Chance, 24(1):44–53, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00144-011-0009-6.

Marks:2011:TCP

[1184] Stephen Marks and Gary Smith. The two-child paradox reborn? Chance, 24 (1):54–59, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://link.

                                        springer.com/article/10.1007/s00144-
                                        011-0010-0.

Berkowitz:2011:GWTa

[1185] Jonathan Berkowitz. Goodness of wit test #11: Definite article rejection! Chance, 24(1):61–63, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00144-011-0011-z.

Anonymous:2011:HCa

[1186] Anonymous. Help & contacts. Chance, 24(1):??, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:EBb

[1187] Anonymous. Editorial Board. Chance, 24(2):2, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:AAb

[1188] Anonymous. About the authors. Chance, 24(2):3–4, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:ELb

[1189] Anonymous. Editor’s letter. Chance, 24(2):5, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-011-0013-x.pdf.

Behseta:2011:ELb

[1190] Sam Behseta. Editor’s letter. Chance, 24(2):5, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Slavkovic:2011:PWAa

[1191] Aleksandra Slavkovic and Stephen E.

Fienberg. O privacy, where art thou? What should privacy mean to you? Chance, 24(2):6–9, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0014-9.

Fienberg:2011:PWA

[1192] Stephen E. Fienberg. O privacy, where art thou? Chance, 24(2):7–9, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barnett:2011:HMB

[1193] Tristan Barnett. How much to bet on video poker. Chance, 24(2): 10–14, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0015-8.

Deutsch:2011:LBS

[1194] Roland C. Deutsch. Looking back at South Africa: Analyzing and reviewing the 2010 FIFA World Cup. Chance, 24 (2):15–23, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0016-7.

Lackritz:2011:NAG

[1195] James R. Lackritz. A new analysis of the golf handicap system: Does the better player have an advantage? Chance, 24(2):24–34, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0017-6.

Rockoff:2011:HGY

[1196] David Rockoff and Heike Hofmann. How good is your eyeballing? Chance, 24(2):35–45, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0018-5.

Hurley:2011:TWB

[1197] W. J. Hurley and Andrey Pavlov. There will be blood: On the risk-return characteristics of a blackjack counting system. Chance, 24(2):47–52, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0019-4.

Glickman:2011:HYH

[1198] Mark Glickman and Meena Doshi.

Here’s to your health: Assessing the seasonality of influenza. Chance, 24 (2):53–59, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0020-y.

Wainer:2011:VRF

[1199] Howard Wainer. Visual revelations:

The first step toward wisdom. Chance, 24(2):60–61, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00144-011-0021-
x.

Berkowitz:2011:GWTb

[1200] Jonathan Berkowitz. Goodness of wit test #12: Straightforward! Chance, 24 (2):62–64, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0022-9.

Anonymous:2011:HCb

[1201] Anonymous. Help & contacts. Chance, 24(2):??, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:EBc

[1202] Anonymous. Editorial Board. Chance, 24(3):2, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:AAc

[1203] Anonymous. About the authors. Chance, 24(3):3–4, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:ELc

[1204] Anonymous. Editor’s letter. Chance, 24(3):5, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00144-011-0023-8.pdf.

Behesta:2011:EL

[1205] Sam Behesta. Editor’s letter. Chance, 24(3):5, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Trempel:2011:DBH

[1206] Rebecca E. Trempel, Sergey Y. Kyrychenko, and Matthew J. Moore. Døes banning hand-held cell phøne use while driving reduce cøllisiøns? Chance, 24(3):6–11, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://link.springer.com/
article/10.1007/s00144-011-0024-
7.

Rutter:2011:RWC

[1207] Michael A. Rutter. Ranking women’s college hockey teams: Ties and home ice. Chance, 24(3):12–22, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0025-6.

Guzman:2011:SSJ

[1208] Daniel Guzmán. Speaking stats to justice: Expert testimony in a Guatemalan human rights trial based on statistical sampling. Chance, 24 (3):23–29, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0026-5.

Ding:2011:PAD

[1209] Jimin Ding, Jürgen Symanzik, Abbass Sharif, Jia Wang, Stephen Duntley, and William D. Shannon. Powerful actigraphy data through functional representation. Chance, 24(3):30–36, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0027-4.

McCarthy:2011:EDS

[1210] David McCarthy. Estimating different sources of variation and predicting tournament outcomes in professional bowling. Chance, 24(3):37–46, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0028-3.

Bring:2011:TPW

[1211] Johan Bring and Marcus Thuresson.

Three points for a win in soccer: Is it fair? Chance, 24(3):47–53, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0029-2.

Wainer:2011:VRH

[1212] Howard Wainer. Visual revelations: How much is tenure worth? Chance, 24(3):54–57, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00144-011-0030-9.

Slavkovic:2011:PWAb

[1213] Aleksandra Slavkovic, John M. Abowd, and Lars Vilhuber. O privacy, where art thou? Science, confidentiality, and the public interest. Chance, 24 (3):58–62, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00144-011-0031-8.

Abowd:2011:PWA

[1214] John M. Abowd and Lars Vilhuber. O privacy, where art thou? Chance, 24 (3):59–62, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2011:GWTc

[1215] Jonathan Berkowitz. Goodness of wit test #13: Don’t be mean. Chance, 24 (3):63–64, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00144-011-0032-
                                        7.

Anonymous:2011:HCc

[1216] Anonymous. Help & contacts. Chance, 24(3):??, September 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:EBd

[1217] Anonymous. Editorial Board. Chance, 24(4):2, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:AAd

[1218] Anonymous. About the authors.

Chance, 24(4):3–4, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2011:ELd

[1219] Anonymous. Editor’s letter. Chance, 24 (4):5, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00144-
011-0033-6.pdf.

Behseta:2011:ELc

[1220] Sam Behseta. Editor’s letter. Chance, 24(4):5, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Keyes:2011:MRG

[1221] Tim K. Keyes. Model risk and the great recession. Chance, 24(4):6– 16, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0034-5.

Lampe:2011:RWC

[1222] Kristen A. Lampe. Reconciling wild cards and poker. Chance, 24(4):17– 24, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0035-4.

Alamar:2011:BPM

[1223] Benjamin Alamar and Keith Goldner. The Blindside Project: Measuring the impact of individual offensive linemen. Chance, 24(4):25–29, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0036-3.

Stone:2011:ONQ

[1224] Mervyn Stone. The owl and the nightingale: The Quetelet/nightingale nexus. Chance, 24(4):30–34, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://link.springer.com/
article/10.1007/s00144-011-0037-
2.

Mata:2011:APO

[1225] Fernando Mata and Sarah Watts. Analyzing the placement odds of favorite horses in the thoroughbred racing industry of the British Isles. Chance, 24(4):35–40, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0038-1.

Slavkovic:2011:PWAc

[1226] Aleksandra Slavkovic, Saki Kinney, and Alan Karr. O privacy, where art thou? Research access to restricteduse data. Chance, 24(4):41–45, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0039-0.

Jensen:2011:SRS

[1227] Shane Jensen. A statistician reads the sports pages: Evaluating general managers in baseball. Chance, 24(4):47– 49, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0040-7.

Gelman:2011:ESO

[1228] Andrew Gelman. Ethics and statistics open data and open methods. Chance, 24(4):51–53, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

//link.springer.com/article/10.
1007/s00144-011-0041-6.

Anonymous:2011:ES

[1229] Anonymous. Ethics and statistics.

Chance, 24(4):51–54, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2011:VRR

[1230] Howard Wainer and Lawrence Hubert.

Visual revelations: A remarkable horse: An inquiry into the accuracy of medical predictions. Chance, 24(4):55– 57, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00144-011-
0042-5.

Robert:2011:BRB

[1231] Christian Robert and Kenneth Lange.

Book reviews: Numerical Analysis for Statisticians. Chance, 24(4):58–59, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lange:2011:BR

[1232] Kenneth Lange, Shravan Vasishth, Michael Broe, and Neal Stephenson. Book reviews. Chance, 24(4):58–61, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00144-011-
                                        0043-4.

Vasishth:2011:FSS

[1233] Shravan Vasishth and Michael Broe. The foundations of statistics: A simulation-based approach. Chance, 24 (4):59–61, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stephenson:2011:A

[1234] Neal Stephenson. Anathem. Chance, 24 (4):61, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2011:GWTd

[1235] Jonathan Berkowitz. Goodness of wit test #14: Outcome analysis. Chance, 24(4):62–64, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00144-011-0044-3.

Anonymous:2011:HCd

[1236] Anonymous. Help & contacts. Chance, 24(4):??, 2011. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2012:ELa

[1237] Sam Behseta. Editor’s letter. Chance, 25(1):3, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Heiny:2012:YSR

[1238] Erik L. Heiny and Robert L. Heiny.

An 11-year study of the relative importance of driving accuracy and driving distance on scoring average on the PGA tour. Chance, 25(1):4–14, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2012:BP

[1239] Nicole Lazar. The big picture. Chance, 25(1):15–17, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2012:TCCa

[1240] Dalene Stangl, Mine Çetinkaya-Rundel, and Kari Lock Morgan. Taking a chance in the classroom: Uncounted votes: A case study in analyzing aggregated data. Chance, 25(1):18–23, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2012:SBM

[1241] Anonymous. Sharon Bertsch McGrayne answers questions about: ‘The Theory That Would Not Die’. Chance, 25(1): 24–29, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Koyama:2012:LBS

[1242] Tatsuki Koyama. Length of The Beatles’ songs. Chance, 25(1):30–33, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Slavkovic:2012:PWA

[1243] Aleksandra Slavkovic and Jerry Reiter. O privacy, where art thou?: Statistical disclosure limitation research and practice: Fascinating and growing areas of importance. Chance, 25(1):34–37, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jensen:2012:SRSa

[1244] Shane Jensen and Andrew C. Thomas. A statistician reads the sports pages: When can you blame your tiles in Scrabble? Chance, 25(1):38–42, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2012:VRM

[1245] Howard Wainer. Visual revelations: Moral statistics and the thematic maps of Joseph Fletcher. Chance, 25(1):43– 46, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2012:ESSa

[1246] Andrew Gelman and Eric Loken. Ethics and statistics: Statisticians: When we teach, we don’t practice what we preach. Chance, 25(1):47–48, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:BRS

[1247] Christian Robert. Book review: Sharon Bertsch McGrayne, The Theory That Would Not Die: Hardcover: 320 + xiv pages, Publisher: Yale University Press (first edition, May 2011), Language: English, ISBN-10: 0-300-16969-8. Chance, 25(1):49–50, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ziliak:2012:BRB

[1248] Stephen Ziliak and Deirdre McCloskey.

Book review: The Cult of Significance: Hardcover: 322 + xxiv pages, Publisher: The University of Michigan Press (First edition, Feb. 2008), Language: English, ISBN-13: 978-0-472-05007-9. Chance, 25(1):51–53, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Brooks:2012:BRB

[1249] Steve Brooks, Andrew Gelman, Galin Jones, and Xiao-Li Meng. Book review: Handbook of Markov Chain Monte Carlo: Hardcover: 619 pages, Publisher: Chapman and Hall/CRC Press (first edition, May 2011), Language: English, ISBN-10: 1-4200-7941-7. Chance, 25(1):53–55, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Karian:2012:BRB

[1250] Z. Karian and E. J. Dudewicz. Book review: Handbook of Fitting Statistical Distributions with R: Hardcover: xlv + 1672 pages + 1 CD-ROM, Publisher: CRC Press, Taylor and Francis Group, Chapman & Hall, Boca Raton (first edition, Oct. 2010), Language: English, ISBN-13: 978-1-58488-711-9. Chance, 25(1):56–57, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2012:GWTa

[1251] Jonathan Berkowitz. Goodness of wit test #15: Outcome analysis. Chance, 25(1):58–60, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

House:2012:LER

[1252] Dennis House, Carl F. Blackman, and Andrew Gelman. Letters to the Editor: Responses to Gelman. Chance, 25(1): 61–64, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2012:ELb

[1253] Sam Behseta. Editor’s letter. Chance, 25(2):3, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Leman:2012:IMM

[1254] Scotland C. Leman and Leanna House. Improving Mr. Miyagi’s coaching style: Teaching data analytics with interactive data visualizations. Chance, 25(2): 4–12, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kucharski:2012:FTE

[1255] Adam Kucharski. The family tree of an epidemic. Chance, 25(2):13–16, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gross:2012:EDA

[1256] Alden L. Gross. An exploratory data analysis of a graduate student (n=1). Chance, 25(2):17–26, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bar-Hillel:2012:WSP

[1257] Maya Bar-Hillel and Ro’i Zultan. We sing the praise of good displays: How gamblers bet in casino roulette. Chance, 25(2):27–30, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Taylor:2012:CRB

[1258] Laura Taylor. Competing risks in basketball … competing risks in basketball … competing risks in basketball . . .. Chance, 25(2):31–36, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2012:BPM

[1259] Nicole Lazar. The big picture: Multiplicity control in large data sets presents new challenges and opportunities. Chance, 25(2):37–40, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cetinkaya-Rundel:2012:TCC

[1260] Mine Çetinkaya-Rundel, Dalene Stangl, and Kari Lock Morgan. Taking a chance in the classroom: Exploring Google’s transparency report. Chance, 25(2):41–45, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jensen:2012:SRSb

[1261] Shane Jensen. A statistician reads the sports pages: Baseball playoffs are a coin toss. Chance, 25(2):46–48, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2012:VRWa

[1262] Howard Wainer. Visual revelations:

When nothing is not zero. Chance, 25 (2):49–51, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2012:ESEa

[1263] Andrew Gelman. Ethics and statistics: Ethics in medical trials: Where does statistics fit in? Chance, 25(2):52–54, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Iacobucci:2012:BRB

[1264] Alessandra Iacobucci and Christian Robert. Book review: The Art of R Programming: A Tour of Statistical Software Design, by Norman Matloff. Chance, 25(2):55–56, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:UCB

[1265] Christian Robert and William M. Boldstad. Understanding computational Bayesian statistics. Chance, 25(2):57– 58, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:BID

[1266] Christian Robert, Ronald Christensen, Wesley Johnson, Adam Branscum, and Timothy Hanson. Bayesian ideas and data analysis. Chance, 25(2):58–60, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:BMU

[1267] Christian Robert and Ioannis Ntzoufras. Bayesian modeling using Win-BUGS. Chance, 25(2):60–61, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2012:GWTb

[1268] Jonathan Berkowitz. Goodness of wit test #16: Logical shortcomings.

Chance, 25(2):62–64, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2012:ELc

[1269] Sam Behseta and Jo Hardin. Editor’s letter. Chance, 25(3):3, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ince:2012:PR

[1270] Darrel Ince. The problem of reproducibility. Chance, 25(3):4–7, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barr:2012:ECR

[1271] Christopher D. Barr and Jukka-Pekka Onnela. Establishing a culture of reproducibility and openness in medical research with an emphasis on the training years. Chance, 25(3):8–10, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zhao:2012:ASR

[1272] Xiaohui Zhao, Keith Dowd, and Cynthia A. Searcy. Assessing statistics and research methodology in the MCAT exam. Chance, 25(3):11–17, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kwasny:2012:CIP

[1273] Mary J. Kwasny and Constantine Daskalakis. Career issues pertaining to teaching introductory biostatistics in a medical center. Chance, 25(3):18–22, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2012:TIB

[1274] Dalene Stangl, Allison Dorogi, Kyu Seo Kim, Josh Latner, Daniel Li, Clara Starkweather, Cynthia Wang, Michelle Wu, Young In Yun, and Lisa Zhang. Transforming introductory biostatistics education: A decision-theoretic perspective using deliberate practice and team-based learning. Chance, 25 (3):23–30, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Brieger:2012:MSE

[1275] Katharine Brieger and Johanna Hardin. Medicine, statistics, and education: The inextricable link. Chance, 25(3):31–34, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hardin:2012:CAY

[1276] Johanna Hardin. Changes across 25 years of statistics in medicine. Chance, 25(3):35–37, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Glickman:2012:HYH

[1277] Mark Glickman, Sowmya R. Rao, Gary Shapiro, and Theresa Diaz. Here’s to your health: Use of GPS-enabled mobile devised to conduct health surveys: Child mortality in Sierra Leone. Chance, 25(3):38–42, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2012:ESSb

[1278] Andrew Gelman. Ethics and statistics: Statistics for cigarette sellers. Chance, 25(3):43–46, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2012:BPB

[1279] Nicole Lazar. The big picture: Big data hits the big time. Chance, 25(3):47–49, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jensen:2012:SRSc

[1280] Shane Jensen and Abraham Wyner. A statistician reads the sports pages: Can the skill level of a game of chance be measured? Chance, 25(3):50–53, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2012:VRC

[1281] Howard Wainer. Visual revelations:

Cheating: Some ways to detect it badly. Chance, 25(3):54–57, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:PAS

[1282] Christian Robert. Principles of applied statistics. Chance, 25(3):58–59, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:LSI

[1283] Christian Robert. Large-scale inference: Empirical Bayes methods for estimation, testing, and prediction. Chance, 25(3):59–61, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:WST

[1284] Christian Robert. A whistle-stop tour of statistics. Chance, 25(3):61, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:CBP

[1285] Christian Robert. Correlations between the physical and social sciences. Chance, 25(3):62, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2012:GWTc

[1286] Jonathan Berkowitz. Goodness of wit test #17. Chance, 25(3):63–64, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2012:ELd

[1287] Sam Behseta. Editor’s letter. Chance, 25(4):3, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hubele:2012:VSS

[1288] Norma Faris Hubele and Mark W. Arndt. Vehicle safety standard update: A case study in a regulatory debate using statistical models. Chance, 25(4):4–12, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gorroochurn:2012:SLP

[1289] Prakash Gorroochurn. Some laws and problems of classical probability and how Cardano anticipated them. Chance, 25(4):13–20, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:IPD

[1290] Christian Robert and Sam Behseta. Interview with: Persi Diaconis. Chance, 25(4):21–25, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Watnik:2012:IWB

[1291] Mitchell R. Watnik and Richard A.

Levine. If we build it, will they compete? Chance, 25(4):26–31, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wilson:2012:PMH

[1292] Jonathan L. Wilson. Prediction markets: How accurate are they? Chance, 25(4):32–38, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lane:2012:PWA

[1293] Julia Lane. O privacy, where art thou?: Protecting privacy and confidentiality in an era of big data access. Chance, 25(4):39–41, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2012:BPF

[1294] Nicole Lazar. The big picture: Functional magnetic resonance imaging — introduction to a neuroimaging modality. Chance, 25(4):42–45, 2012. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2012:TCCb

[1295] Dalene Stangl, Mine Çetinkaya-Rundel, and Kari Lock Morgan. Taking a chance in the classroom: Making the old new again. Chance, 25(4):46–49, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2012:VRWb

[1296] Howard Wainer. Visual revelations:

Waiting for Achilles. Chance, 25(4):50– 51, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2012:ESEb

[1297] Andrew Gelman. Ethics and statistics: Ethics and the statistical use of prior information. Chance, 25(4):52–54, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Donnet:2012:SMS

[1298] Sophie Donnet and Christian Robert. Stochastic modelling for systems biology (second edition). Chance, 25(4): 55–56, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Donnet:2012:STE

[1299] Sophie Donnet. Statistical thinking in epidemiology. Chance, 25(4):56–57, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:UZW

[1300] Christian Robert. The universe in zero words. Chance, 25(4):57–58, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:SN

[1301] Christian Robert. Simulating nature. Chance, 25(4):59–60, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:WN

[1302] Christian Robert. A wealth of numbers. Chance, 25(4):60, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2012:LES

[1303] Christian Robert. Logicomix: An epic search for truth. Chance, 25(4):61, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2012:GWTd

[1304] Jonathan Berkowitz. Goodness of wit test #18: Plus or minus. Chance, 25 (4):63–64, 2012. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2013:ELa

[1305] Sam Behseta. Editor’s letter. Chance, 26(1):3, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sullivan:2013:BBP

[1306] Lisa M. Sullivan, Marie Davidian, Anita L. DeStefano, and Roslyn A. Stone. Building the biostatistics pipeline: Summer institutes for training in biostatistics (SIBS). Chance, 26 (1):4–9, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robinson:2013:HBK

[1307] Stephen J. Robinson. How to beat kindergartners at battleship. Chance, 26(1):10–15, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rambharat:2013:SIU

[1308] Bhojnarine R. Rambharat. Statistical intelligence units. Chance, 26(1):16–21, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2013:SIG

[1309] Christian Robert. Special Issue on George Casella’s books. Chance, 26(1): 22–32, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2013:BPD

[1310] Nicole Lazar. The big picture: Data, data, everywhere . . .. Chance, 26(1): 33–37, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Slavkovic:2013:PWA

[1311] Aleksandra Slavkovic and Anne-Sophie Charest. O privacy, where art thou?: Privacy protection for the European Census Hub. Chance, 26(1):38–41, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2013:TCC

[1312] Dalene Stangl, Mine Çetinkaya-Rundel, and Kari Lock Morgan. Taking a chance in the classroom: The American Community Survey. Chance, 26(1): 42–46, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jensen:2013:SRSa

[1313] Shane Jensen. A statistician reads the sports pages: Salaries and wins in baseball. Chance, 26(1):47–52, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2013:VRS

[1314] Howard Wainer, Polina Harik, and John Neter. Visual revelations: Stigler’s Law of Eponymy and Marey’s train schedule: Did Serjev do it before Ibry, and what about Jules Petiet? Chance, 26(1):53–56, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2013:ESW

[1315] Andrew Gelman and Mark Palko.

Ethics and statistics: The war on data. Chance, 26(1):57–60, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2013:GWTa

[1316] Jonathan Berkowitz. Goodness of wit test #19: No real opposition. Chance, 26(1):61–63, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2013:ELb

[1317] Sam Behseta. Editor’s letter. Chance, 26(2):3, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Blyth:2013:WBR

[1318] Stephen Blyth. Of wranglers and bankers: Reflections on applied statistics from a career in finance. Chance, 26 (2):4–7, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Redelmeier:2013:CTD

[1319] Donald A. Redelmeier and Christopher J. Yarnell. Can tax deadlines cause fatal mistakes? Chance, 26(2): 8–14, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

DAbramo:2013:HDE

[1320] Germano D’Abramo. How to defuse Earth impact threat announcements. Chance, 26(2):17–23, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2013:DPU

[1321] Sam Behseta and Michelle Dunn. Discussing Principles of Uncertainty with Jay Kadane. Chance, 26(2):24–27, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2013:BPB

[1322] Nicole Lazar. The big picture: Big data computing. Chance, 26(2):28–32, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cetinkaya-Rundel:2013:TCCa

[1323] Mine Çetinkaya-Rundel, Kari Lock Morgan, and Dalene Stangl. Taking a chance in the classroom: Looking good on course evaluations. Chance, 26(2): 33–37, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cook:2013:VRP

[1324] Robert Cook and Howard Wainer. Visual revelations: Plotting evidence to affect social policy: Guns, murders, life, death, and ignorance in contemporary America. Chance, 26(2):38–44, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2013:ESTa

[1325] Andrew Gelman. Ethics and statistics: They’d rather be rigorous than right. Chance, 26(2):45–49, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2013:MMM

[1326] Christian Robert. Magical mathematics: The mathematical ideas that animate great magic tricks. Chance, 26(2): 50–51, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McCabe:2013:MEI

[1327] Brendan McCabe. Methods for estimation and inference in modern econometrics. Chance, 26(2):51–52, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2013:PSI

[1328] Christian Robert. Paradoxes in scientific inference. Chance, 26(2):52–54, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2013:PUE

[1329] Christian Robert. In pursuit of the unknown: 17 equations that changed the world. Chance, 26(2):54–58, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2013:GSW

[1330] Christian Robert. Guesstimation:

Solving the World’s problems on the back of a cocktail napkin and Guesstimation 2.0: Solving Today’s problems on the back of a napkin. Chance, 26(2): 58–59, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2013:GWTb

[1331] Jonathan Berkowitz. Goodness of wit test #20: Uncertainty. Chance, 26(2): 60–62, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Greenland:2013:SMA

[1332] Sander Greenland. Should a metaanalyst want the likelihood or the posterior from each study? Chance, 26(2): 63–64, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2013:ELc

[1333] Sam Behseta. Editor’s letter. Chance, 26(3):3, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Pagano:2013:MBD

[1334] Marcello Pagano and Sarah Anoke. Mommy’s baby, daddy’s maybe: A closer look at regression to the mean. Chance, 26(3):4–9, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stewart:2013:FFU

[1335] Gavin Stewart, Kerrie Mengersen, Georgina M. Mace, Jeff A. McNeely, Jegar Pitchforth, and Ben Collen. To fund or not to fund: Using Bayesian networks to make decisions about conserving our World’s endangered species. Chance, 26(3):10–17, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rosenbaum:2013:UDC

[1336] Paul R. Rosenbaum. Using differential comparisons in observational studies. Chance, 26(3):18–25, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Horowitz:2013:VWM

[1337] Ira Horowitz and Jim Lackritz. The value of s at which a major league player with multiple hitting streaks of at least s games might be termed ‘streak prone’. Chance, 26(3):26–33, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jensen:2013:SRSb

[1338] Shane Jensen. A statistician reads the sports pages: Measuring player contributions in hockey. Chance, 26(3):34– 38, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2013:BPS

[1339] Nicole Lazar. The big picture: Symbolic data analysis. Chance, 26(3):39– 42, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cetinkaya-Rundel:2013:TCCb

[1340] Mine Çetinkaya-Rundel and Dalene Stangl. Taking a chance in the classroom: A celebration of data. Chance, 26(3):43–46, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2013:VRH

[1341] Howard Wainer. Visual revelations:

How the rule of 72 can provide guidance to advance your wealth, weight, career, and gas mileage. Chance, 26(3):47–48, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2013:ESTb

[1342] Andrew Gelman. Ethics and statistics: It’s too hard to publish criticisms and obtain data for republication. Chance, 26(3):49–52, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jacob:2013:PSE

[1343] Pierre Jacob. Probability and statistics for engineers and scientists (9th edition). Chance, 26(3):53, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Saloman:2013:PSF

[1344] Jean-Bernard Saloman. Probability and statistics (fourth edition). Chance, 26(3):54, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Keller:2013:SLS

[1345] Merlin Keller. Statistics for the life sciences (4th edition). Chance, 26(3):54– 56, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoffmann:2013:SSP

[1346] Marc Hoffmann. Stationary stochastic processes. Chance, 26(3):56–57, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoffmann:2013:SMS

[1347] Marc Hoffmann. Statistical methods for stochastic differential equations.

Chance, 26(3):57–58, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2013:GWTc

[1348] Jonathan Berkowitz. Goodness of wit test #21: Cogitation. Chance, 26(3): 59–60, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Teoh:2013:BIM

[1349] Eric R. Teoh. Beware the Ides of March? Chance, 26(3):61–64, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2013:ELd

[1350] Sam Behseta. Editor’s letter. Chance, 26(4):3, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Slavkovic:2013:SMM

[1351] Aleksandra Slavkovic. Steve the matchmaker: The marriage of statistics and computer science in the world of data privacy. Chance, 26(4):4–7, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tanur:2013:SCS

[1352] Judith M. Tanur. Steve’s contributions to sample surveys, censuses, and Federal statistics. Chance, 26(4):8–11, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Airoldi:2013:SFC

[1353] Edoardo M. Airoldi. Stephen E.

Fienberg’s contributions to categorical data analysis and the social sciences.

Chance, 26(4):12–14, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stigler:2013:BRS

[1354] Stephen Stigler. Book review: Stephen Fienberg: Cause or Effect? Node or Edge? A Comparative Study. Chance, 26(4):15, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

George:2013:SB

[1355] Edward I. George. Steve the Bayesian. Chance, 26(4):16–17, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2013:ISF

[1356] Sam Behseta and Aleksandra Slavkovic. Interview with Steve Fienberg. Chance, 26(4):18–29, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2013:TCI

[1357] Howard Wainer. Taking a chance: An interview with William F. Eddy and Stephen E. Fienberg. Chance, 26(4): 30–34, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2013:NCC

[1358] Michael D. Larsen, Alexander F. Wakim, Reid Rottach, and Yang Cheng. Nonparametric comparisons of current population survey panel estimates. Chance, 26(4):35–41, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2013:BPA

[1359] Nicole Lazar. The big picture: The arts — digitized, quantified, and analyzed. Chance, 26(4):42–45, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schutt:2013:TCC

[1360] Rachel Schutt. Taking a chance in the classroom: Embracing the ambiguity and potential of data science. Chance, 26(4):46–51, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2013:ESI

[1361] Andrew Gelman. Ethics and statistics: Is it possible to be an ethicist without being mean to people? Chance, 26(4):52–55, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fouley:2013:BBP

[1362] Jean-Louis Fouley. The BUGS book:

A practical introduction to Bayesian analysis. Chance, 26(4):56–59, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Emmert-Streib:2013:BCA

[1363] Frank Emmert-Streib. Biostatistics: A computing approach. Chance, 26(4): 59–60, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Beck:2013:ABM

[1364] Benoît Beck. Advanced Bayesian methods for medical test accuracy. Chance, 26(4):60–61, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2013:BRB

[1365] Christian Robert. Book review: R For Dummies. Chance, 26(4):61, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2013:GWTd

[1366] Jonathan Berkowitz. Goodness of wit test #22: Judge for yourself. Chance, 26(4):62–64, 2013. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2014:ELa

[1367] Scott Evans. Editor’s letter. Chance, 27 (1):3, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rosenthal:2014:SOL

[1368] Jeffrey S. Rosenthal. Statistics and the Ontario lottery retailer scandal. Chance, 27(1):4–9, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Johnson:2014:CSC

[1369] H. Dean Johnson. The conduct of statistical consulting sessions between statisticians and the researchers with whom they interact: A survey of consulting clients from academic settings. Chance, 27(1):10–19, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaplan:2014:CYB

[1370] Andee Kaplan, Eric Hare, Heike Hofmann, and Dianne Cook. Can you buy a President? Politics after the Tillman Act. Chance, 27(1):20–30, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hurley:2014:EGC

[1371] W. J. Hurley, Jack Brimberg, and Richard Kohar. Examining a gambler’s claims: Probabilistic fact-checking and Don Johnson’s extraordinary blackjack winning streak. Chance, 27(1):31–37, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2014:BPF

[1372] Nicole Lazar. The big picture: Functional data analysis. Chance, 27(1):38– 40, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McClintock:2014:TCC

[1373] Shannon McClintock, Dalene Stangl, and Mine Çetinkaya-Rundel. Taking a chance in the classroom: The real secret to genius? Reading between the lines. Chance, 27(1):41–44, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2014:VRC

[1374] Howard Wainer. Visual revelations: On the crucial role of empathy in the design of communications: Genetic testing as an example. Chance, 27(1):45– 50, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2014:ESA

[1375] Andrew Gelman and Eric Loken.

Ethics and statistics: The AAA tranche of subprime science. Chance, 27(1):51–56, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:MHH

[1376] Christian Robert. The most human human. Chance, 27(1):57–58, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:NSS

[1377] Christian Robert. Naked statistics: Stripping the dread from data. Chance, 27(1):58–59, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sohier:2014:IPT

[1378] Julien Sohier. Introduction to probability with Texas hold ’em examples. Chance, 27(1):60–61, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:CIS

[1379] Christian Robert. The cartoon introduction to statistics. Chance, 27(1):61, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berkowitz:2014:GWT

[1380] Jonathan Berkowitz. Goodness of wit test #23: Times are changing: A farewell puzzle. Chance, 27(1):62–64, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2014:ELb

[1381] Scott Evans. Editor’s letter. Chance, 27 (2):3, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Trono:2014:HPO

[1382] John A. Trono and Philip Yates. How predictable is the overall voting pattern in the NCAA Men’s Basketball Post Tournament Poll? Chance, 27(2):4–12, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lopez:2014:YTC

[1383] Michael J. Lopez, Adrian Esparza, Michael Lavine, and Jenna Marquard. Your textbook can’t help you here: Applying traditional methods to atypical data. Chance, 27(2):13–20, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Young:2014:BJB

[1384] Derek S. Young. Bond, James Bond.

A statistical look at cinema’s most famous spy. Chance, 27(2):21–27, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mitchell:2014:OOO

[1385] Jason D. Mitchell and Li Lian Ong.

Optimism and the occult octopus: Favorites Lose, Underdogs Triumph, and Spain Finally Wins the World Cup. Chance, 27(2):28–38, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2014:IJB

[1386] Sam Behseta and Mine Çetinkaya-Rundel. Interview with Jim Berger. Chance, 27(2):39–43, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morgan:2014:TCC

[1387] Kari Lock Morgan, Mine Çetinkaya-Rundel, and Dalene Stangl. Taking a chance in the classroom: Speed dating: Exploring initial romantic attraction. Chance, 27(2):44–49, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2014:BPM

[1388] Nicole Lazar. The big picture: On mice, badges for your CV, and the big data backlash . . .. Chance, 27(2):50– 52, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cook:2014:VGE

[1389] Di Cook. Visiphilia: A graphical expedition into a statistics gradebook. Chance, 27(2):53–55, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2014:VRL

[1390] Howard Wainer. Visual revelations: Life follows art: Gaming the Missing Data Algorithm. Chance, 27(2):56–57, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2014:ESC

[1391] Andrew Gelman and Phillip N. Price. Ethics and statistics: The Commissar for Traffic presents the latest Five-Year Plan. Chance, 27(2):58–60, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:SMC

[1392] Christian Robert. Statistical Modeling and Computation. Chance, 27(2):61– 62, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:MLP

[1393] Christian Robert. Machine Learning, a Probabilistic Perspective. Chance, 27 (2):62–63, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:SST

[1394] Christian Robert. Statistics for Spatio-Temporal Data. Chance, 27(2):64, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2014:ELc

[1395] Scott Evans. Editor’s letter. Chance, 27 (3):3, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Albert:2014:SHR

[1396] Jim Albert. Streakiness in home run hitting. Chance, 27(3):4–9, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fischer:2014:SIS

[1397] Krista Fischer and Donald A. Berry.

Statisticians introduce science to International Doping Agency: The Andrus Veerpalu case. Chance, 27(3):10–16, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Morris:2014:ICM

[1398] Carl N. Morris. Interview with Carl Morris. Chance, 27(3):17–24, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Smith:2014:QTB

[1399] Richard L. Smith, Scott Powers, and Jessi Cisewski. Qualifying times for the Boston Marathon. Chance, 27(3):25– 33, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wilson:2014:CIQ

[1400] Jason Wilson and Jarvis Greiner. A curveball index: Quantification of Breaking Balls for Pitchers. Chance, 27(3):34–40, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Baumer:2014:SRS

[1401] Benjamin S. Baumer and Gregory J. Matthews. A statistician reads the sports pages: There is no avoiding WAR. Chance, 27(3):41–44, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2014:BPT

[1402] Nicole Lazar. The big picture: Take me out to the ball game. Chance, 27(3): 45–48, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cook:2014:VCN

[1403] Di Cook. Visiphilia: Can Nick Kyrgios be a tennis champion? Chance, 27(3): 49–52, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bray:2014:TCC

[1404] Andrew Bray, Mine Çetinkaya-Rundel, and Dalene Stangl. Taking a chance in the classroom: Five concrete reasons your students should be learning to analyze data in the reproducible paradigm. Chance, 27(3):53–56, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:MIU

[1405] Christian Robert. Medical illuminations: Using evidence, visualization and statistical thinking to improve healthcare. Chance, 27(3):57–58, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2014:VRM

[1406] Howard Wainer. Visual revelations:

Musing about changes in the SAT: Is the College Board getting rid of the bulldog? Chance, 27(3):59–63, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2014:ELd

[1407] Scott Evans. Editor’s letter. Chance, 27 (4):3, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hicks:2014:WWS

[1408] Stephanie C. Hicks. When women in statistics come to know their power. Chance, 27(4):4–5, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2014:KYP

[1409] Dalene Stangl. Know your power: A panel discussion by past and future ASA Presidents. Chance, 27(4):6–14, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2014:IJW

[1410] Scott Evans. An interview with Janet Wittes, President of Statistics Collaborative. Chance, 27(4):15–22, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sellers:2014:SGS

[1411] Kimberly F. Sellers, K. Nicole Meyer, Maria A. Terres, Samantha Tyner, and Kaitlin Woo. Surviving graduate school: What happens in session, stays in session. Chance, 27(4):23–25, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lindborg:2014:IES

[1412] Stacy R. Lindborg. Influencing — even without statistics. Chance, 27(4):26– 27, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mejdal:2014:MMS

[1413] Arati Mejdal. Making the most of social media. Chance, 27(4):28–30, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wendelberger:2014:WIT

[1414] Joanne Wendelberger, Alyson Wilson, Sandra Stinnett, and Brenda Gaydos. Working in interdisciplinary teams. Chance, 27(4):31–33, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Harvill:2014:TPC

[1415] Jane L. Harvill and Kristen Tecson. Two perspectives on celebrating women in statistics. Chance, 27(4):34– 36, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kafadar:2014:CWS

[1416] Karen Kafadar. Contributions of women in science: Ingredients for success? Chance, 27(4):37–43, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Huzurbazar:2014:OSN

[1417] Snehalata Huzurbazar, Jessi Cisewski, Bailey Fosdick, and Xia Wang. Opportunities at SAMSI and NISS. Chance, 27(4):44–47, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cook:2014:VGG

[1418] Di Cook. Visiphilia: The gender gap in math is not universal. Chance, 27(4): 48–52, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Pileggi:2014:TCC

[1419] Shannon McClintock Pileggi, Mine Çetinkaya-Rundel, and Dalene Stangl. Taking a chance in the classroom: A classroom project for Generation Z. Chance, 27(4):53–55, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Baaaath:2014:CGS

[1420] Rasmus Bååth. The cartoon guide to statistics. Chance, 27(4):56, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Baaaath:2014:MGS

[1421] Rasmus Bååth. The Manga guide to statistics. Chance, 27(4):56–57, 2014.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Baaaath:2014:CIS

[1422] Rasmus Bååth. The cartoon introduction to statistics. Chance, 27(4):57, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Baaaath:2014:ISG

[1423] Rasmus Bååth. Introducing statistics, a graphic guide. Chance, 27(4):57–58, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:SSU

[1424] Christian Robert. Straightforward statistics: Understanding the tools of research. Chance, 27(4):58–59, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2014:FSA

[1425] Christian Robert. Foundations of statistical algorithms. Chance, 27(4):59– 60, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2014:VRH

[1426] Howard Wainer. Visual revelations: Happiness and causal inference. Chance, 27(4):61–64, 2014. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2015:ELa

[1427] Scott Evans. Editor’s letter. Chance, 28 (1):3, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dorum:2015:UF

[1428] Guro Dørum and Mariam Mjærum Bouzga. Urns and forensics. Chance, 28(1):4–11, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2015:IHW

[1429] Sam Behseta. An interview with Howard Wainer. Chance, 28(1):12–20, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hurley:2015:MDT

[1430] W. J. Hurley. Mandatory drug testing in the Canadian workplace: A note on the recent Supreme Court decision and its effect on the misclassification rate of non-drug users. Chance, 28(1):21–26, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Polson:2015:BVJ

[1431] Nick Polson and Jan Hendrik Witte.

A Bellman view of Jesse Livermore. Chance, 28(1):27–31, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2015:PM

[1432] Sam Behseta and Michelle Dunn. A peek into MoMath. Chance, 28(1):32– 38, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2015:BPB

[1433] Nicole Lazar. The big picture: Big data and privacy. Chance, 28(1):39–42, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Thomas:2015:SRS

[1434] A. C. Thomas. A statistician reads the sports pages: Pucksberry: What we can learn from hexagonal plots of NHL data. Chance, 28(1):43–50, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2015:OJP

[1435] Mary W. Gray. The odds of justice: P -values: What do they prove? Chance, 28(1):51–53, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mulder:2015:CSR

[1436] Kees Mulder. Circular statistics in R. Chance, 28(1):54, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2015:ESD

[1437] Andrew Gelman. Ethics and statistics: Disagreements about the strength of evidence. Chance, 28(1):55–59, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoaglin:2015:LE

[1438] David C. Hoaglin, Jarvis Greiner, and Jason Wilson. Letter to the Editor. Chance, 28(1):60–64, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2015:ELb

[1439] Scott Evans. Editor’s letter. Chance, 28 (2):3, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zhang:2015:SSB

[1440] Xinping Zhang, Byran J. Smucker, and Jay Woffington. Statistics and show business: Shakespeare meets predictive analytics. Chance, 28(2):4–12, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vesper:2015:PHH

[1441] Andrew Vesper. Putting the “hot hand” on ice. Chance, 28(2):13–18, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Radman:2015:CAS

[1442] Tom Radman, Erica Rosemond, and Michelle Dunn. Calling all statisticians for the next wave of biomedical big data discoveries. Chance, 28(2):19–22, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Goldman:2015:DBM

[1443] Robert Goldman. Development of a biostatistics major: The Simmons College story. Chance, 28(2):23–25, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hurley:2015:HNB

[1444] W. J. Hurley and Tyler Sauerbrei. Handicapping net best-ball team matches in golf. Chance, 28(2):26–30, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2015:ESH

[1445] Andrew Gelman and David Madigan.

Ethics and statistics: How is ethics like logistic regression? Ethics decisions, like statistical inferences, are informative only if they’re not too easy or too hard. Chance, 28(2):31–33, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2015:BPN

[1446] Nicole Lazar. The big picture: Now trending on Twitter. . .. Chance, 28(2): 34–36, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Slavkovic:2015:PWA

[1447] Aleksandra Slavkovic and Fei Yu. O privacy, where art thou?: Genomics and privacy. Chance, 28(2):37–39, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Horton:2015:TCC

[1448] Nicholas J. Horton, Benjamin S. Baumer, and Hadley Wickham. Taking a chance in the classroom: Setting the stage for data science: Integration of data management skills in introductory and second courses in statistics. Chance, 28(2):40–50, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2015:OJW

[1449] Mary W. Gray. The odds of justice: Where is the Reverend Bayes when needed? Chance, 28(2):51–53, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Craig:2015:SRS

[1450] Bruce A. Craig, Zachary Hass, and Sean McCabe. A statistician reads the sports pages: Easy to criticize, harder to verify: Fourth and goal in post-2011 NFL overtime. Chance, 28(2):54–57, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2015:VRC

[1451] Howard Wainer and Donald B. Rubin. Visual revelations: Causal inference and death. Chance, 28(2):58–64, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2015:ELc

[1452] Scott Evans. Editor’s letter. Chance, 28(3):3, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
09332480.2015.1099360.

Usiskin:2015:RBS

[1453] Zalman Usiskin and Kathy Hall. The relationships between statistics and other subjects in the K–12 curriculum. Chance, 28(3):4–18, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/09332480.2015.1099361.

Bargagliotti:2015:SET

[1454] Anna Bargagliotti and Christine Franklin. The statistical education of teachers: Preparing teachers to teach statistics. Chance, 28 (3):19–27, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
09332480.2015.1099362.

Starnes:2015:ASE

[1455] Daren Starnes and Anna Martin.

The AP statistics exam: An insider’s guide to its distinctive features.

Chance, 28(3):28–37, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/09332480.2015.1099363.

Hedrick:2015:ASC

[1456] Benjamin Hedrick. The AP statistics curriculum: Past, present, and future. Chance, 28(3):38–43, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/09332480.2015.1099364.

Zieffler:2015:CCH

[1457] Andrew Zieffler and Michael D. Huberty. A catalyst for change in the high school math curriculum. Chance, 28(3):44–49, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/09332480.2015.1099365.

Endriss:2015:DS

[1458] Kay Endriss and Michael MacWilliams. Data surfing. Chance, 28(3):50– 55, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/abs/10.1080/
09332480.2015.1099367.

Lee:2015:TCC

[1459] Hollylynne Lee and Dalene Stangl. Taking a chance in the classroom: Professional development MOOCs for teachers of statistics in K–12. Chance, 28(3):56–63, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/09332480.2015.1099368.

Evans:2015:ENS

[1460] Scott Evans. Editor’s note: Staytuned for Part II of “Nurturing statistical thinking before college”. Chance, 28(3):64, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http:/

/www.tandfonline.com/doi/abs/10.
1080/09332480.2015.1099371.

Evans:2015:ELd

[1461] Scott Evans. Editor’s letter. Chance, 28 (4):3, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Simoneau:2015:SDC

[1462] Eric J. Simoneau. STATS4STEM:

Data, computing, and assessment resources for high-school statistics students. Chance, 28(4):4–11, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Turner:2015:CMS

[1463] Peter R. Turner, Rachel Levy, and Kathleen Fowler. Collaboration in the mathematical sciences community on mathematical modeling across the curriculum. Chance, 28(4):12–18, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hayden:2015:SCP

[1464] Robert (Bob) W. Hayden. Statistics: From college to pre-college. Chance, 28(4):19–25, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kwasny:2015:SKE

[1465] Mary J. Kwasny. Statistics in K–12:

Educators, students, and us: Challenges and resources for K–12 educators and students. Chance, 28(4):26– 29, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Holcomb:2015:CCS

[1466] John Holcomb. Consider a career in statistics. Chance, 28(4):30–33, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2015:CAS

[1467] Anonymous. A conversation about statistics before college: Barriers and opportunities for statistics in secondary education. Chance, 28(4):34–42, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2015:BPP

[1468] Nicole Lazar and Christine Franklin. The big picture: Preparing students for a data-centric world. Chance, 28(4):43– 45, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kaplan:2015:TCC

[1469] Daniel Kaplan. Taking a chance in the classroom: Modeling statistical thinking. Chance, 28(4):46–51, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2015:OJL

[1470] Mary W. Gray. The odds of justice: Leveling the playing field. Chance, 28 (4):52–54, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Roey:2015:PWA

[1471] Stephen “Shep” Roey, Tom Krenzke, and Robert Perkins. O privacy, where art thou?: Can ‘dirty’ data be your friend? considering the impact on disclosure risk as illustrated in the education data context. Chance, 28(4):55–59, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Layton:2015:IRR

[1472] Richard Layton. Influences of reproducible reporting on work flow. Chance, 28(4):60–64, 2015. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2016:ELa

[1473] Scott Evans. Editor’s letter. Chance, 29 (1):3, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Pierson:2016:SFS

[1474] Steve Pierson and Karen Kafadar.

Statisticians and forensic science: A perfect match. Chance, 29(1):4–8, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Neumann:2016:FEF

[1475] Cedric Neumann and Hal Stern. Forensic examination of fingerprints: Past, present, and future. Chance, 29(1):9– 16, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Curran:2016:MMD

[1476] James M. Curran and Bruce S. Weir.

Modern methods of DNA interpretation. Chance, 29(1):17–26, 2016.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cheng:2016:BLB

[1477] Edward K. Cheng. A Bayesian look at the Baby Annie case. Chance, 29(1): 27–31, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zabell:2016:VBB

[1478] Sandy Zabell. The virtues of being blind. Chance, 29(1):32–36, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Neumann:2016:PQQ

[1479] Cedric Neumann, David Kaye, Graham Jackson, Valerie Reyna, and Anjali Ranadive. Presenting quantitative and qualitative information on forensic science evidence in the courtroom. Chance, 29(1):37–43, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lancaster:2016:ESI

[1480] Mark J. Lancaster. An exploration of statistical issues in forensic science related to document examination. Chance, 29(1):44–48, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Starr:2016:OJA

[1481] Sonja Starr. The odds of justice: Actuarial risk prediction and the criminal justice system. Mary W. Gray, Column Editor. Chance, 29(1):49–51, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Nathenson:2016:SRS

[1482] Robert A. Nathenson. A statistician reads the sports pages: A historical tracking of parity in baseball. Shane Jensen, Column Editor. Chance, 29(1):52–60, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2016:VRD

[1483] Howard Wainer. Visual revelations:

Defeating deception: Escaping the shackles of truthiness by learning to think like a data scientist. Chance, 29 (1):61–64, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2016:ELb

[1484] Scott Evans. Editor’s letter. Chance, 29 (2):3, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lavine:2016:STC

[1485] Michael Lavine, J. Roger Brothers, Kenneth J. Lohmann, and Isaac Lavine. Sea turtles: A case of animal magnetism. Chance, 29(2):4–9, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Buckley:2016:DEP

[1486] Hannah L. Buckley, Bradley S. Case, Ronny Vallejos, J. Julio Camarero, Emilia Gutiérrez, Eryuan Liang, Yafeng Wang, and Aaron M. Ellison. Detecting ecological patterns along environmental gradients: Alpine treeline ecotones. Chance, 29(2):10–15, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ogle:2016:PEM

[1487] Kiona Ogle and Jarrett J. Barber. Plant and ecosystem memory. Chance, 29(2):16–22, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wikle:2016:MEM

[1488] Christopher K. Wikle, William B. Leeds, and Mevin B. Hooten. Models for ecological models: Ocean primary productivity. Chance, 29(2):23–30, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Record:2016:MSR

[1489] Sydne Record and Noah Charney. Modeling species ranges. Chance, 29 (2):31–37, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gotelli:2016:CMS

[1490] Nicholas J. Gotelli. Checkerboards and missing species combinations: Are ecological communities assembled by chance? Chance, 29(2):38–45, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dey:2016:DHL

[1491] Tanujit Dey, Walter Hickey, Bimal Parakkal, and Wojbor A. Woyczynski. Deepwater horizon: Locating submerged oil in ocean water using the k-means clustering technique. Chance, 29(2):46–52, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rundel:2016:TCC

[1492] Colin Rundel and Mine Çetinkaya-Rundel. Taking a chance in the classroom: La Quinta is Spanish for “next to Denny’s”. Chance, 29(2):53–57, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2016:OJS

[1493] Mary W. Gray. The odds of justice: Statistics go to school and then to court. Chance, 29(2):58–60, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2016:BPI

[1494] Nicole Lazar. The big picture: Imagining genetics: A tale of two modalities. Chance, 29(2):61–63, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Steinberg:2016:VRV

[1495] Linda Steinberg and Howard Wainer.

Visual revelations: VAMs and baseball. Chance, 29(2):64, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2016:ELc

[1496] Scott Evans. Editor’s letter. Chance, 29(3):3–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tyner:2016:URP

[1497] Sam Tyner. Using the R package geomnet: Visualizing trans-Atlantic slave trade of Africans, 1514–1866. Chance, 29(3):4–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Smith:2016:ZHW

[1498] Abigail Smith. Zika is here, and we need statistics. Chance, 29(3):17–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Huber:2016:WBR

[1499] Michael Huber. Why Babe Ruth never hit for the cycle, and other hitting for the cycle tidbits. Chance, 29(3):27–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Proschan:2016:FAC

[1500] Michael Proschan. FDA advisory committees: The role of statisticians. Chance, 29(3):31–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stanek:2016:PUS

[1501] Edward J. Stanek III. Predicting the unknown — sampling, smoke, and mirrors. Chance, 29(3):41–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2016:CPS

[1502] Scott Evans, Cyrus Mehta, and Nitin Patel. Creators of pioneering software and clinical services company look back and to the future. Chance, 29(3):51–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2016:SDW

[1503] Christian Robert. Book review: Statistics Done Wrong: The Woefully Complete Guide. Chance, 29(3):58–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2016:MSE

[1504] Christian Robert. Book review: Measuring Statistical Evidence Using Relative Belief. Chance, 29(3):59–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2016:VRE

[1505] Howard Wainer. Visual revelations:

Evidence in the 2016 Presidential campaign. Chance, 29(3):62–??, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2016:ELd

[1506] Scott Evans. Editor’s letter. Chance, 29 (4):3, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Friedenberg:2016:MPH

[1507] David A. Friedenberg and Michael A.

Schwemmer. Moving a paralyzed hand — a biomedical big data success story. Chance, 29(4):4–13, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bosse:2016:SBB

[1508] Michael Bossé, Eric Marland, Gregory Rhoads, and Michael Rudziewicz. Searching for the black box: Misconceptions of linearity. Chance, 29(4):14– 23, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mulholland:2016:PDN

[1509] Jason Mulholland and Shane T.

Jensen. Projecting the draft and NFL performance of wide receiver and tight end prospects. Chance, 29(4):24–31, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2016:SBR

[1510] Anonymous. Statistics books to read for pleasure. Chance, 29(4):32–33, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2016:CSA

[1511] Scott Evans, Dennis Lock, and Stephanie Kovalchik. Careers in sports analytics-highlights of an ASA webinar. Chance, 29(4):34–46, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2016:OJW

[1512] Mary W. Gray. The odds of justice: Who gets to vote? Chance, 29(4):47– 51, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Freilich:2016:SRS

[1513] Brian Freilich. A statistician reads the sports pages: The value of coaching tenure in the NBA. Chance, 29(4):52– 60, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2016:BBH

[1514] Scott Evans. Brief bits of humor in statistics: “a statistician is … ”. Chance, 29(4):61, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2016:IPS

[1515] Anonymous. International Prize in Statistics awarded to Sir David Cox for survival analysis model: Cox model applied in medicine, science, and engineering. Chance, 29(4):62–63, 2016. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2017:ELa

[1516] Scott Evans. Editor’s letter. Chance, 30 (1):3, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

ONeill:2017:LGV

[1517] Robert T. O’Neill and Marcia A. Testa.

Let’s go to the videotape: Promoting the visibility and understanding of regulatory science by using case studies from publicly available FDA information. Chance, 30(1):4–14, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hamasaki:2017:IPG

[1518] Toshimitsu Hamasaki and Scott Evans.

Interview with Professor Geert Molenberghs. Chance, 30(1):16–23, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eddy:2017:RSF

[1519] William F. Eddy. Remembering Steve Fienberg. Chance, 30(1):24–25, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Albert:2017:DSM

[1520] Jim Albert and Stephanie Kovalchik.

The do’s and dont’s of sports metrics: The Tennis ATP Leaderboard. Chance, 30(1):26–34, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ball:2017:AIC

[1521] Richard Ball. Anti-isomorphisms in classical hypothesis testing and children’s literature. Chance, 30(1):35, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2017:ESH

[1522] Andrew Gelman. Ethics and statistics: Honesty and transparency are not enough. Chance, 30(1):37–39, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2017:VBV

[1523] Christian Robert. Une vie brève (one hundred twenty-one days). Chance, 30 (1):40, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2017:SR

[1524] Christian Robert. Statistical rethinking. Chance, 30(1):40–42, 2017. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2017:SPD

[1525] Christian Robert. Superintelligence: Paths, dangers, strategies. Chance, 30 (1):42–43, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2017:VRG

[1526] Howard Wainer. Visual revelations: The Grabovsky curve. Chance, 30(1): 44–48, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2017:ELb

[1527] Scott Evans. Editor’s letter. Chance, 30 (2):3, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Brogan:2017:CSD

[1528] Donna Brogan. Challenging sex discrimination: Reflections over seven decades. Chance, 30(2):4–18, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Furlong:2017:LDT

[1529] Cathy Furlong. Leadership development through volunteer activities. Chance, 30(2):19–21, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wendelberger:2017:SAD

[1530] Joanne R. Wendelberger. Scaling up to address data science challenges. Chance, 30(2):22–25, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Datta:2017:AOD

[1531] Susmita Datta. Advancing omics data analysis: A call for participation by a statistician in the field. Chance, 30(2):26–29, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wiley:2017:CEC

[1532] Erin Wiley. Cause and effective communication. Chance, 30(2):30–37, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sellers:2017:AIB

[1533] Kimberly F. Sellers, Emma K. T. Benn, Maria Garcia, and Meghan Kellam. Addressing implicit bias among women statisticians and data scientists. Chance, 30(2):38–41, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

An:2017:ACF

[1534] Ming-Wen An and Donna LaLonde. ASA’s Committee on Funded Research: How can it help You? Chance, 30(2): 42–44, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2017:TDT

[1535] Dalene Stangl. From ‘truth or dare’ to ‘truth and dare’. Chance, 30(2):45–53, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stangl:2017:TCC

[1536] Dalene Stangl and Mine Çetinkaya-Rundel. Taking a chance in the classroom: An interview with Mine Çetinkaya-Rundel. Chance, 30(2):54– 59, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2017:OJA

[1537] Mary W. Gray. The odds of justice: Adjusting the odds. Chance, 30(2):60– 64, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2017:ELc

[1538] Scott Evans. Editor’s letter. Chance, 30 (3):3, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bales:2017:USS

[1539] Kevin Bales. Unlocking the statistics of slavery. Chance, 30(3):4–12, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zhang:2017:SEG

[1540] Sheldon X. Zhang and Kyle Vincent.

Strategies to estimate global prevalence of trafficking for sexual exploitation. Chance, 30(3):13–20, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2017:MVE

[1541] Jacqueline Joudo Larsen and Davina P.

Durgana. Measuring vulnerability and estimating prevalence of modern slavery. Chance, 30(3):21–29, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Larsen:2017:USE

[1542] Jacqueline Joudo Larsen and Pablo Diego-Rosell. Using surveys to estimate the national prevalences of modern slavery: Experience and lessons learned. Chance, 30(3):30–35, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Durgana:2017:PGD

[1543] Davina P. Durgana and Graham K.

Brown. The politics of global data reporting: A triangulated solution for estimating modern slavery. Chance, 30 (3):36–40, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cruyff:2017:CCV

[1544] Maarten Cruyff, Jan van Dijk, and Peter G. M. van der Heijden. The challenge of counting victims of human trafficking: Not on the record: A multiple systems estimation of the numbers of human trafficking victims in the Netherlands in 2010–2015 by year, age, gender, and type of exploitation. Chance, 30(3):41–49, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Durgana:2017:FST

[1545] Davina P. Durgana and Paul L. Zador. Fighting slavery through statistics: A discussion of five promising methods to estimate prevalence in the United States. Chance, 30(3):50–53, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

David:2017:MSS

[1546] Fiona David. Modern slavery — from statistics to prevention. Chance, 30(3): 54–60, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2017:VRB

[1547] Howard Wainer. Visual revelations: The birth of statistical graphics and their European childhood: On the historical development of W. E. B. Du Bois’s graphical narrative of a people. Chance, 30(3):61–67, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lagoze:2017:PWA

[1548] Carl Lagoze and Lars Vilhuber. O privacy, where art thou? Making confidential data part of reproducible research. Chance, 30(3):68–72, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2017:ELd

[1549] Scott Evans. Editor’s letter. Chance, 30 (4):3, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Craigmile:2017:ERS

[1550] Peter F. Craigmile. Editorial: The role of statistics in climate research. Chance, 30(4):4–5, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Guttorp:2017:HWK

[1551] Peter Guttorp. How we know that the Earth is warming. Chance, 30(4):6–11, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Craigmile:2017:IPS

[1552] Peter F. Craigmile and Bo Li. Instruments, proxies, and simulations: Exploring imperfect measures of climate. Chance, 30(4):12–18, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jun:2017:CMI

[1553] Mikyoung Jun. Climate model intercomparison. Chance, 30(4):19–25, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hammerling:2017:CCD

[1554] Dorit Hammerling. Climate change detection and attribution: Letting go of the null? Chance, 30(4):26–29, 2017.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gilleland:2017:QRE

[1555] Eric Gilleland, Richard W. Katz, and Philippe Naveau. Quantifying the risk of extreme events under climate change. Chance, 30(4):30–36, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Haran:2017:SFA

[1556] Murali Haran, Won Chang, Klaus Keller, Robert Nicholas, and David Pollard. Statistics and the future of the Antarctic ice sheet. Chance, 30(4): 37–44, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wikle:2017:EIC

[1557] Christopher K. Wikle. Ecological impacts of climate change: the importance of temporal and spatial synchrony. Chance, 30(4):45–54, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Chang:2017:PHI

[1558] Howard H. Chang, Stefanie Ebelt Sarnat, and Yang Liu. Projecting health impacts of climate change: Embracing an uncertain future. Chance, 30(4):55– 61, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Javaid:2017:TCC

[1559] Azka Javaid, Xiaofei Wang, and Nicholas J. Horton. Taking a chance in the classroom: Using a “study of studies” to help statistics students assess research findings. Chance, 30(4):62–66, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schwab-McCoy:2017:TSH

[1560] Aimee Schwab-McCoy. Teaching statistics in the health sciences: Life on an island: Using peer consulting in applied statistics courses. Chance, 30(4):67–72, 2017. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2018:ELa

[1561] Scott Evans. Editor’s letter. Chance, 31 (1):3, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mejia:2018:WWS

[1562] Robin Mejia and Megan Price. Working where statistics and human rights meet. Chance, 31(1):4–6, 2018. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hemenway:2018:FDO

[1563] David Hemenway. Firearms data, and an ode to data systems. Chance, 31(1): 7–11, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jans:2018:MAS

[1564] Matt Jans, Bianca D. M. Wilson, and Jody L. Herman. Measuring aspects of sexuality and gender: A sexual human rights challenge for science and official statistics. Chance, 31(1):12–20, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kendall:2018:MRC

[1565] Arthur Kendall. Maryland rent courts fail to protect human rights to due process and fair trials. Chance, 31(1):21– 28, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Pham:2018:HRM

[1566] Phuong Pham and Patrick Vinck. Human rights and mixed methods. Chance, 31(1):29–37, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ball:2018:SG

[1567] Patrick Ball and Megan Price. The statistics of genocide. Chance, 31(1): 38–45, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vehkalahti:2018:RLI

[1568] Kimmo Vehkalahti, Juuso Koponen, and Hanna Kuusi. From Russia with love to Infovis: Graphic train schedules reflect the history of Finland. Chance, 31(1):46–53, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2018:OJS

[1569] Mary W. Gray (Column Editor). The odds of justice: Is statistics down for the count? Chance, 31(1):54–56, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2018:BPD

[1570] Nicole Lazar (Column Editor). The big picture: Divide and combine to conquer big data. Chance, 31(1):57–59, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2018:TRC

[1571] Christian Robert. Testing R code.

Chance, 31(1):60, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2018:TTD

[1572] Christian Robert. Truth or truthiness: Distinguishing fact from fiction by learning to think like a data scientist. Chance, 31(1):60–61, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2018:SPS

[1573] Christian Robert. The seven pillars of statistical wisdom. Chance, 31(1):61– 62, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2018:EBL

[1574] Christian Robert. Errors, blunders, and lies. Chance, 31(1):62–63, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2018:WMV

[1575] Christian Robert. What makes variables random: Probability for the applied researcher. Chance, 31(1):63–64, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2018:ELb

[1576] Scott Evans. Editor’s letter. Chance, 31 (2):3, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Freeman:2018:IWW

[1577] Laura J. Freeman and Catherine Warner. Informing the warfighter — why statistical methods matter in defense testing. Chance, 31(2):4–11, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Fricker:2018:DSD

[1578] Ronald D. Fricker, Jr. and Steven E. Rigdon. Disease surveillance: Detecting and tracking outbreaks using statistics. Chance, 31(2):12–22, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Banks:2018:ARA

[1579] David Banks. Adversarial risk analysis. Chance, 31(2):23–29, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ray:2018:ATD

[1580] Douglas M. Ray and Paul A. Roediger. Adaptive testing of DoD systems with binary response. Chance, 31(2):30–37, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Buttrey:2018:DSM

[1581] Samuel E. Buttrey, Lyn R. Whitaker, and Jonathan K. Alt. Developments in the statistical modeling of military recruiting. Chance, 31(2):38–44, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sanchez:2018:DFR

[1582] Susan M. Sanchez. Data farming: Reaping insights from simulation experiments. Chance, 31(2):45–52, 2018.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wood:2018:UGU

[1583] Beverly L. Wood, Megan Mocko, Michelle Everson, Nicholas J. Horton, and Paul Velleman. Updated guidelines, updated curriculum: The GAISE College Report and introductory statistics for the modern student. Chance, 31(2):53–59, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2018:OJW

[1584] Mary W. Gray (Column Editor). The odds of justice: Whom shall we kill? How shall we kill them? Chance, 31(2): 60–61, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2018:VRA

[1585] Howard Wainer (Column Editor) and Michael Friendly. Visual revelations: Ancient visualizations. Chance, 31(2): 62–64, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2018:ELc

[1586] Scott Evans. Editor’s letter. Chance, 31 (3):3, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Stigler:2018:LST

[1587] Stephen M. Stigler and Margaret L.

Stigler. Luck and skill in tournament golf. Chance, 31(3):4–13, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cupingood:2018:WRK

[1588] Leonard Cupingood. Why a 1-for-45 record in the Kentucky Derby does not necessarily equate to underachievement. Chance, 31(3):14–19, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cottrell:2018:AFG

[1589] David Cottrell and Michael C. Herron. All in the family: German twins’ finishing times in the 2016 Women’s Olympic Marathon. Chance, 31(3):20–28, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Sackrowitz:2018:PAT

[1590] Harold Sackrowitz. The point(s)-after-touchdown decision revisited. Chance, 31(3):29–36, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Chandler:2018:AFR

[1591] Gabriel Chandler and Simon Rosenbaum. An analysis of the first round of the MLB first-year player draft. Chance, 31(3):37–43, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Heiny:2018:OLR

[1592] Erik L. Heiny and Cody C. Frisby. An ordinal logistic regression model for the masters golf tournament. Chance, 31 (3):44–58, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Heard:2018:DOH

[1593] Craig A. Heard and A. John Bailer. Defying the odds: How likely are we to see another team pull a ‘Leicester’ and win the EPL? Chance, 31(3):59–66, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wexler:2018:OJC

[1594] Rebecca Wexler. The odds of justice:

Code of silence: How private companies hide flaws in the software that governments use to decide who goes to prison and who gets out. Chance, 31(3):67–72, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2018:ELd

[1595] Scott Evans. Editor’s letter. Chance, 31 (4):3, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barnett:2018:EFP

[1596] Arnold Barnett. Epic fail? The polls and the 2016 Presidential Election. Chance, 31(4):4–8, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gill:2018:EST

[1597] Richard D. Gill, Piet Groeneboom, and Peter de Jong. Elementary statistics on trial — the case of Lucia de Berk. Chance, 31(4):9–15, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zubizarreta:2018:SEI

[1598] Jose R. Zubizarreta, Dylan S. Small, and Paul R. Rosenbaum. A simple example of isolation in building a natural experiment. Chance, 31(4):16–23, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Mowrey:2018:W

[1599] Daniel H. Mowrey, Jonathan Potts, Susan Spruill, Walter Stroup, and Michiko I. Wolcott. “With great power comes great responsibility”: Harnessing the extraordinary power of statistics. Chance, 31(4):24–28, 2018. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cameron:2018:MAV

[1600] Briana Cameron, 23andMe Research Team, and Robert Gentleman. Mobile apps versus Web browsers: A comparison of self-administered survey platforms. Chance, 31(4):29–36, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Redelmeier:2018:LMM

[1601] Donald A. Redelmeier and Eldar Shafir. The lunacy of motorcycle mortality. Chance, 31(4):37–44, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dawson:2018:MAC

[1602] Deborah V. Dawson and Derek R. Blanchette. Meta-analysis: Can we extract gold from the biomedical literature? Chance, 31(4):45–50, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Rohlfs:2018:TCC

[1603] Chris Rohlfs. Taking a chance in the classroom: Top 10 reasons not to drop insignificant regressors: A statistical listicle. Chance, 31(4):51–56, 2018. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2019:ELa

[1604] Scott Evans. Editor’s letter. Chance, 32 (1):3, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ioannidis:2019:WMP

[1605] John P. A. Ioannidis. Why most published research findings are false. Chance, 32(1):4–13, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Koyama:2019:LBS

[1606] Tatsuki Koyama. Length of the Beatles’ songs. Chance, 32(1):14–17, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bales:2019:USS

[1607] Kevin Bales. Unlocking the statistics of slavery. Chance, 32(1):18–26, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Young:2019:BJB

[1608] Derek S. Young. Bond. James Bond.

A statistical look at cinema’s most famous spy. Chance, 32(1):27–35, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Guttorp:2019:HWK

[1609] Peter Guttorp. How we know that the Earth is warming. Chance, 32(1):36– 41, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hernan:2019:SCG

[1610] Miguel A. Hernán, John Hsu, and Brian Healy. A second chance to get causal inference right: A classification of data science tasks. Chance, 32(1): 42–49, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Adams:2019:BEI

[1611] Tom Adams. Book excerpt: Improving Your NCAA(R) Bracket with Statistics. Chance, 32(1):50–54, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2019:BPB

[1612] Nicole Lazar. The big picture: Big data and privacy. Chance, 32(1):55–58, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:BRP

[1613] Christian Robert. Book review: Pragmatics of Uncertainty. Chance, 32(1): 59, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:BRG

[1614] Christian Robert. Book review: 10 Great Ideas About Chance. Chance, 32 (1):60–62, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:BRI

[1615] Christian Robert. Book review: Independent Random Sampling Methods. Chance, 32(1):62–63, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:BRC

[1616] Christian Robert. Book review: Computational Methods for Numerical Analysis with R. Chance, 32(1):63, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2019:ELb

[1617] Scott Evans. Editor’s letter. Chance, 32 (2):3, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoaglin:2019:ALM

[1618] David C. Hoaglin. Another look at meta-analysis. Chance, 32(2):4–5, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wapner:2019:FEC

[1619] Leonard M. Wapner. Fair and efficient by chance. Chance, 32(2):6–10, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kvam:2019:DCQ

[1620] Paul Kvam. Demonstrating the consequences of quota sampling in graduate student admissions. Chance, 32(2): 11–16, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Krenzke:2019:RPD

[1621] Tom Krenzke and Jianzhu Li. Replicating published data tables to assess sensitivity in subsequent analyses and mapping. Chance, 32(2):17–27, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Singh:2019:BPV

[1622] Sarjinder Singh. Box plot versus human face. Chance, 32(2):28–30, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hu:2019:TCC

[1623] Jingchen (Monika) Hu. Taking a chance in the classroom: Teaching upper-level undergraduate statistics through a shared/hybrid model. Chance, 32(2):31–36, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cetinkaya-Rundel:2019:TSH

[1624] Mine Çetinkaya-Rundel. Teaching statistics in the health sciences: Teaching to, and learning from, the masses. Chance, 32(2):37–42, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2019:BPC

[1625] Nicole Lazar. The big picture: Crowdsourcing your way to big data. Chance, 32(2):43–46, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:BMC

[1626] Christian Robert. The beauty of mathematics in computer science. Chance, 32(2):47–48, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:A

[1627] Christian Robert. AIQ. Chance, 32(2): 48–49, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:LES

[1628] Christian Robert. Let the evidence speak. Chance, 32(2):49–50, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:BN

[1629] Christian Robert. Is that a big number? Chance, 32(2):50–51, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2019:SPS

[1630] Christian Robert. Surprises in probability-seventeen short stories. Chance, 32(2):51–52, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2019:VRS

[1631] Howard Wainer. Visual revelations: A statistician reads the obituaries: On the relative effects of race and crime on employment. Chance, 32(2):53–56, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2019:ELc

[1632] Scott Evans. Editor’s letter. Chance, 32 (3):3, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cisewski-Kehe:2019:SIA

[1633] Jessi Cisewski-Kehe and Chad Schafer.

Special issue on astrostatistics. Chance, 32(3):4–5, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Green:2019:TOC

[1634] Sheridan B. Green, Abby Mintz, Xin Xu, and Jessi Cisewski-Kehe. Topology of our cosmology with persistent homology. Chance, 32(3):6–13, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Politsch:2019:MLS

[1635] Collin A. Politsch and Rupert A. C. Croft. Mapping the large-scale universe through intergalactic silhouettes. Chance, 32(3):14–19, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Freeman:2019:JHF

[1636] Peter E. Freeman. Just how far away is that galaxy, anyway? Estimating Galaxy Distances Using Low-Resolution Photometric Data. Chance, 32(3):20–26, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eadie:2019:SSS

[1637] Gwendolyn Eadie. Statistics for stellar systems: From globular clusters to clusters of galaxies. Chance, 32(3):27– 34, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Feigelson:2019:ASH

[1638] Eric D. Feigelson. ARIMA for the stars: How statistics finds exoplanets. Chance, 32(3):35–40, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Segal:2019:IMW

[1639] Mark R. Segal and Jacob W. Segal. Identifying Milky Way open clusters with extreme kinematics using PRIM. Chance, 32(3):41–49, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Corliss:2019:TSC

[1640] David J. Corliss. Time series clustering methods for analysis of astronomical data. Chance, 32(3):50–58, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2019:ADI

[1641] Mary W. Gray (Column Editor). Alexa did it! Chance, 32(3):59–61, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hoaglin:2019:RHF

[1642] David C. Hoaglin. Reconsidering the human face as boxplot. Chance, 32(3): 62–63, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Evans:2019:ELd

[1643] Scott Evans. Editor’s letter. Chance, 32 (4):3, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bretz:2019:RRM

[1644] Frank Bretz, Willi Maurer, and Dong Xi. Replicability, reproducibility, and multiplicity in drug development. Chance, 32(4):4–11, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Akacha:2019:CMT

[1645] Mouna Akacha. Choosing measures of treatment benefit: Estimands and beyond. Chance, 32(4):12–17, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jiang:2019:RCR

[1646] Yunyun Jiang and Diane Uschner. Romantic comedies and responsive adaptive randomization in clinical trials. Chance, 32(4):18–22, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Halabi:2019:APO

[1647] Susan Halabi and Donna Niedzwiecki. Advancing precision oncology through biomarker-driven trials: Theory vs. practice. Chance, 32(4):23–31, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lai:2019:CTG

[1648] I-Chun Lai, Chin-Fu Hsiao, Lih jiuan Hsu, and Churn-Shiouh Gau. Clinical trials in global drug developments. Chance, 32(4):32–39, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hering:2019:YSN

[1649] Amanda S. Hering and Daniel Cooley. 20 years of statistics at the National Center for Atmospheric Research. Chance, 32(4):40–43, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Campbell:2019:WDH

[1650] R. B. Campbell. When did history occur? Chance, 32(4):44–45, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Pelechrinis:2019:RCN

[1651] Konstantinos Pelechrinis, Ronald Yurko, and Sam Ventura. Reducing concussions in the NFL: A data-driven approach. Chance, 32(4):46–56, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McDonald:2019:DSI

[1652] Sarah McDonald and Nicholas J. Horton. Data scraping, ingestation, and modeling: Bringing data from cars.com into the intro stats class. Chance, 32(4):57–64, 2019. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2020:ELa

[1653] Amanda Peterson-Plunkett. Editor’s letter. Chance, 33(1):3, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Chai:2020:IDC

[1654] Christine P. Chai. The importance of data cleaning: Three visualization examples. Chance, 33(1):4–9, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zaman:2020:MPG

[1655] Zia Zaman and Hong Ming Tan. A modification to pull the goalie that takes the state of play into account: Coach Markov returns. Chance, 33(1): 10–15, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Eck:2020:CNP

[1656] Daniel J. Eck. Challenging nostalgia and performance metrics in baseball. Chance, 33(1):16–25, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Farebrother:2020:NAM

[1657] Richard William Farebrother. Neopythagorean approaches to measures of central tendency and dispersion. Chance, 33(1):26–29, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kim:2020:MEV

[1658] Allen Kim and Steven Skiena. Maximizing the expected value of a lottery ticket: How to sell and when to buy. Chance, 33(1):30–37, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2020:SHC

[1659] Amanda Peterson-Plunkett. Statistics history in a crossword. Chance, 33(1): 38–39, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2020:BCE

[1660] Christian Robert. Bayesian costeffectiveness analysis of medical treatments. Chance, 33(1):40–41, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2020:PDS

[1661] Christian Robert. The 9 pitfalls of data science. Chance, 33(1):41–42, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2020:PS

[1662] Christian Robert. Prime suspects. Chance, 33(1):43, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2020:GBP

[1663] Howard Wainer and Michael Friendly.

The graphical birth of plate tectonics. Chance, 33(1):44–47, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2020:CPS

[1664] Amanda Peterson-Plunkett. Crossword puzzle solution. Chance, 33(1):48, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2020:ELb

[1665] Amanda Peterson-Plunkett. Editor’s letter. Chance, 33(2):3, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Behseta:2020:MCR

[1666] Sam Behseta and Michelle Dunn. Meet Cynthia Rudin a champion of interpretable machine learning. Chance, 33 (2):4–6, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Holmes:2020:WWF

[1667] David Holmes and Ferris Samara. Was the wild frontiersman a prolific penman? A stylometric investigation into the works of Davy Crockett. Chance, 33(2):7–18, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2020:WS

[1668] Amanda Peterson-Plunkett. Word search. Chance, 33(2):19, 2020. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Trott:2020:DMS

[1669] David Trott. Deceiving machines: Sabotaging machine learning. Chance, 33 (2):20–24, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2020:SSI

[1670] Andrew Gelman and Alexey Guzey. Statistics as squid ink: How prominent researchers can get away with misrepresenting data. Chance, 33(2):25–27, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2020:CMF

[1671] Mary Gray. Call Mr. Fisher. Chance, 33(2):28–30, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

xcetinkaya-Rundel:2020:DFT

[1672] Mine Çetinkaya-Rundel and Maria Tackett. From drab to fab: Teaching visualization via incremental improvements. Chance, 33(2):31–41, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Jacobs:2020:TGT

[1673] Justin Jacobs. Three is greater than two? Not so fast . . .. Chance, 33(2): 42–46, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Campbell:2020:MDR

[1674] Paul J. Campbell. Messy data in the real world. Chance, 33(2):47, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2020:WSS

[1675] Anonymous. Word search solution.

Chance, 33(2):48, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2020:ELc

[1676] Amanda Peterson-Plunkett. Editor’s letter. Chance, 33(3):3, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ewing:2020:UCT

[1677] E. Thomas Ewing, Steven E. Rigdon, and Ronald D. Fricker Jr. Understanding COVID-19 in 2020 through the lens of the 1918 “Spanish Flu” epidemic. Chance, 33(3):4–21, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Yin:2020:OS

[1678] Xiaoyan Yin, Toshimitsu Hamasaki, Dean Follmann, and Scott R. Evans. OutSMARTing superbugs. Chance, 33 (3):22–30, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Berg:2020:SAE

[1679] Arthur Berg. Statistical analysis of the Elo rating system in chess. Chance, 33(3):31–38, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zhou:2020:EMO

[1680] Tianjian Zhou and Yuan Ji. Emerging methods for oncology clinical trials. Chance, 33(3):39–48, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wang:2020:PRP

[1681] Dong Wang and Fang Liu. Privacy risk and preservation in contact tracing of COVID-19. Chance, 33(3):49–55, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2020:DM

[1682] Mary W. Gray. Dangerous models? Chance, 33(3):56–57, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2020:AEV

[1683] Andrew Gelman. Assessing evidence vs. truth in the coronavirus pandemic. Chance, 33(3):58–60, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2020:EPT

[1684] Christian Robert. Essentials of probability theory for statisticians. Chance, 33(3):61, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2020:CAS

[1685] Christian Robert. A computational approach to statistical learning. Chance, 33(3):61–62, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tackett:2020:CDC

[1686] Maria Tackett and Mine Çetinkaya-Rundel. COVID-19 data in the classroom. Chance, 33(3):63–66, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Boka:2020:HCW

[1687] Debra M. Boka and Howard Wainer.

How can we estimate the death toll from COVID-19? Chance, 33(3):67–72, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2020:ELd

[1688] Amanda Peterson-Plunkett. Editor’s letter. Chance, 33(4):3, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Liu:2020:SIN

[1689] Fang Liu, Saki Kinney, and Aleksandra (Sesa) Slavkovic’. Special issue: A new generation of statisticians tackles data privacy. Chance, 33(4):4–5, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Snoke:2020:HSS

[1690] Joshua Snoke and Claire McKay Bowen. How statisticians should grapple with privacy in a changing data landscape. Chance, 33(4):6–13, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Nixon:2020:DDD

[1691] Michelle Pistner Nixon, Steven R.

Nixon, and Roberto Molinari. Data of the defense and the defense of data. Chance, 33(4):14–21, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Zhang:2020:PPA

[1692] Zhe Zhang and Linjun Zhang. Privacypreserving algorithms: the gain and the loss. Chance, 33(4):22–28, 2020.

CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Hu:2020:RWD

[1693] Jingchen (Monika) Hu, Terrance Savitsky, and Matthew Williams. Riskweighted data synthesizers for microdata dissemination. Chance, 33(4):29– 36, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Avella-Medina:2020:RRS

[1694] Marco Avella-Medina. The role of robust statistics in private data analysis. Chance, 33(4):37–42, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Awan:2020:FPM

[1695] Jordan Awan, Matthew Reimherr, and Aleksandra (Sesa) Slavković. Formal privacy for modern nonparametric statistics. Chance, 33(4):43–49, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Burnham:2020:LLR

[1696] Ella Burnham and Erin Blankenship. Lessons learned: Revising an online introductory course. Chance, 33(4):50– 55, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bilder:2020:NFD

[1697] Christopher R. Bilder. In or out? The new flagstick dilemma for putting in golf. Chance, 33(4):56–61, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Reynolds:2020:OOE

[1698] Evangeline Reynolds. Outcomeoriented educational tasks with flipbooks. Chance, 33(4):62–64, 2020. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2021:ELa

[1699] Amanda Peterson-Plunkett. Editor’s letter. Chance, 34(1):3, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bhaduri:2021:DRB

[1700] Moinak Bhaduri, Joe Dery, Dominique Haughton, Tao Li, Piaomu Liu, and Wenxiu Nan. Dynamics of rewardbased crowdfunding: Kickstarter rock music projects. Chance, 34(1):4–13, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Anonymous:2021:CEB

[1701] Anonymous. A conversation with Emery Brown. Chance, 34(1):14–24, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dickens:2021:DSP

[1702] James C. Dickens. The data science public schools outreach project: An initiative for diversity and inclusion in statistics. Chance, 34(1):25–27, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Walker:2021:MPP

[1703] Ernie Walker. Monty Python and the probability of winning the Holy Grail. Chance, 34(1):28–38, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Smeeton:2021:ESF

[1704] Nigel Smeeton. Early statistical findings and authorship misattribution: An unsystematic review of the literature. Chance, 34(1):39–43, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ehsani:2021:CRR

[1705] Johnathon P. Ehsani and Brian Tefft. Crash risk and roadway familiarity. Chance, 34(1):44–48, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2021:W

[1706] Mary W. Gray. And the winner is . . .. Chance, 34(1):49–51, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wiorkowski:2021:DG

[1707] Robert Christian and John J. Wiorkowski. Data games. Chance, 34(1):52, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2021:UET

[1708] Christian Robert. Understanding elections through statistics: Polling, prediction, and testing. Chance, 34(1):52– 53, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2021:BRP

[1709] Christian Robert. Book review: Principles of Uncertainty (Second Edition). Chance, 34(1):54–55, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2021:BRS

[1710] Christian Robert. Book review: Statistics and Analysis of Scientific Data (Second Edition). Chance, 34(1):55–56, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2021:ELb

[1711] Amanda Peterson-Plunkett. Editor’s letter. Chance, 34(2):3, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

T:2021:MES

[1712] Jon T., Nicholas B., and Thomas Middleton. Modeling the economic and societal impact of non-pharmaceutical interventions during the COVID-19 pandemic. Chance, 34(2):4–17, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

McIntire:2021:SCS

[1713] Brenda McIntire. The secret career of Solomon Kullback. Chance, 34(2):18– 23, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Allen:2021:POK

[1714] Terry Allen. Police officers killed in the line of duty: A correspondence analysis of circumstances and time of day. Chance, 34(2):24–30, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Marchand:2021:BLC

[1715] Chase Marchand and Dalton Maahs.

Benford’s Law and COVID-19 data. Chance, 34(2):31–38, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Parkey:2021:WAD

[1716] Charna Parkey. Who is accountable for data bias? Chance, 34(2):39–43, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Palayew:2021:TRP

[1717] Adam Palayew, Sam Harper, and James A. Hanley. Toward reducing the possibility of false-positive results in epidemiologic studies of traffic crashes. Chance, 34(2):44–52, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Nelson:2021:DSI

[1718] Valerie Nelson and Jason Crea. The data science instructional escape room — a successful experiment. Chance, 34 (2):53–58, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lazar:2021:STT

[1719] Nicole Lazar and Hyunnam Ryu. The shape of things: Topological data analysis. Chance, 34(2):59–64, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2021:LRH

[1720] Howard Wainer and Richard Feinberg. Looking at reported hate crimes. Chance, 34(2):65–72, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2021:ELc

[1721] Amanda Peterson-Plunkett. Editor’s letter. Chance, 34(3):3, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Augustin:2021:SIS

[1722] Caitlin Augustin, Matt Brems, and Davina P. Durgana. Special issue on statistics and data science for good. Chance, 34(3):4–5, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Andersen:2021:LAC

[1723] Christian Andersen, Uyen Kim Huynh, Andrés Ochoa Toasa, Christine Wells, and Melissa Wong. Lessons from applying the community rapid assessment method to COVID-19 protective measures in three countries. Chance, 34(3): 6–12, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Olubusoye:2021:TEA

[1724] Olusanya E. Olubusoye, Olalekan J.

Akintande, and Eric A. Vance. Transforming evidence to action: The case of election participation in Nigeria. Chance, 34(3):13–23, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Augustin:2021:PIT

[1725] Caitlin Augustin, Isaac Holeman, Erika Salomon, Helen Olsen, Phil Azar, and Mitali Ayyangar. Pathways to increasing trust in public health data. Chance, 34(3):24–32, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Unangst:2021:SPG

[1726] Jennifer Unangst. Surveys for the public good: Examples from statistics without borders. Chance, 34(3):33–37, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Vance:2021:BSD

[1727] Eric A. Vance and Kim Love. Building statistics and data science capacity for development. Chance, 34(3):38–46, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gentry:2021:ILM

[1728] Blake Gentry, Marianne Richardson, Diego Piña Lopez, and Joseph Watkins. Indigenous language migration along the U.S. Southwestern Border — the view from Arizona. Chance, 34(3):47–55, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Cheng:2021:KHM

[1729] Samantha Cheng and Caitlin Augustin. Keep a human in the machine and other lessons learned from deploying and maintaining Colandr. Chance, 34(3):56–60, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Czapski:2021:MLA

[1730] Michal Czapski, Stephen Godfrey, Joshua Derenski, and Isaac Khader. A machine learning approach to helping small businesses find pandemic economic-impact relief. Chance, 34(3): 61–68, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tackett:2021:TCF

[1731] Maria Tackett, Kendra Burbank, Judith Canner, and Mine Çetinkaya-Rundel. Teaching courses focused on social good. Chance, 34(3):69–72, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Roy:2021:CMF

[1732] Preetam Debasish Saha Roy and Sangeeta Jayadevan. COVID monitoring framework for Indian cities. Chance, 34(3):W73–W81, 2021. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Balinskaite:2021:ACC

[1733] Violeta Balinskaite. An assessment of the change in co-worker well-being at Sydenham Garden. Chance, 34(3): W82–W85, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kinsella:2021:DSS

[1734] Benjamin Kinsella. Data science for social good volunteer motivations and limitations: an exploratory survey. Chance, 34(3):W86–W95, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2021:ELd

[1735] Amanda Peterson-Plunkett. Editor’s letter. Chance, 34(4):3, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Orkin:2021:GCG

[1736] Michael Orkin. Games of chance and games of skill. Chance, 34(4):4–9, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Benjamini:2021:APT

[1737] Yoav Benjamini, Richard D. De Veaux, Bradley Efron, Scott Evans, Mark Glickman, Barry I. Graubard, Xuming He, Xiao-Li Meng, Nancy M. Reid, Stephen M. Stigler, Stephen B. Vardeman, Christopher K. Wikle, Tommy Wright, Linda J. Young, and Karen Kafadar. ASA President’s Task Force statement on statistical significance and replicability. Chance, 34(4):10–11, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tong:2021:SEA

[1738] Christopher Tong. The statistical endeavors of Andrew Carnegie. Chance, 34(4):12–17, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Weitzman:2021:DRS

[1739] R. A. Weitzman. From divining rods to statistics: a forensic analysis of the misuse of statistics in the estimation of environmental impact. Chance, 34(4):18–20, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gelman:2021:ERR

[1740] Andrew Gelman. Ethical requirements of a research assistant who is concerned about the behavior of a supervisor. Chance, 34(4):21–22, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Gray:2021:LDL

[1741] Mary Gray and Nimai Mehta. Liars, damned liars, and . . .. Chance, 34(4): 23–27, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wainer:2021:STW

[1742] Howard Wainer. Some thoughts on a world without tests. Chance, 34(4):28– 29, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bigelow:2021:TSH

[1743] Carol Bigelow and Amy S. Nowacki.

Teaching statistics in the Health Sciences Resources Portal. Chance, 34(4): 30–33, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2021:QC

[1744] Christian P. Robert. Quick(er) calculations. Chance, 34(4):34, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2021:ET

[1745] Christian P. Robert. The error of truth.

Chance, 34(4):34–36, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Robert:2021:PSP

[1746] Christian P. Robert. Poems that solve puzzles. Chance, 34(4):36, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wagenmakers:2021:BF

[1747] Eric-Jan Wagenmakers. Bernoulli’s fallacy. Chance, 34(4):37–38, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Wilkinson:2021:HDV

[1748] Leland Wilkinson. A history of data visualization and graphic communication. Chance, 34(4):38–40, 2021. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2022:ELa

[1749] Amanda Peterson-Plunkett. Editor’s letter. Chance, 35(1):3, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barnett:2022:SFG

[1750] Arnold Barnett, Pengchen Han, and Gege Zhang. A simple fix for gerrymandering? Chance, 35(1):4–10, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Tijms:2022:MAG

[1751] Steven Tijms. The mathematical anatomy of the gambler’s fallacy. Chance, 35(1):11–17, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Schutt:2022:PWI

[1752] Randy Schutt. Probit and wealth inequality-how random events and the laws of probability are partially responsible for wealth inequality. Chance, 35 (1):18–25, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Bhaduri:2022:CC

[1753] Moinak Bhaduri. Contrary currents. Chance, 35(1):26–33, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Kim:2022:ADD

[1754] Steven B. Kim and Joonghak Lee. Approaches to a dilemma during the pandemic: Sequential successes and simultaneous successes. Chance, 35(1):34– 37, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Lee:2022:LAS

[1755] Hollylynne S. Lee, Zachary T. Vaskalis, David J. Stokes, and Taylor R. Harrison. A look into the AP Statistics Classroom: Who teaches it and what aspects of statistics do they emphasize? Chance, 35(1):38–47, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Paller:2022:WSP

[1756] Alan Paller. Why some people don’t listen to statisticians. Chance, 35(1): 48–50, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Dasgupta:2022:SDD

[1757] Nairanjana Dasgupta. Skewed distributions in data science. Chance, 35(1): 51–55, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Patil:2022:BTS

[1758] Sujata Patil and Jaya Satagopan.

Building and teaching a statistics curriculum for post-doctoral biomedical scientists at a free-standing cancer center. Chance, 35(1):56–64, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Peterson-Plunkett:2022:ELb

[1759] Amanda Peterson-Plunkett. Editor’s letter. Chance, 35(2):3–4, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2022.
2066407.

Groharing:2022:BLC

[1760] Brooks Groharing and David Mc-Cune. Benford’s Law and countylevel votes in US Presidential elections. Chance, 35(2):4–10, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2066408.

Velleman:2022:ECD

[1761] Paul Velleman and Howard Wainer. Exploring COVID data with Benford’s and Zipf’s laws. Chance, 35(2):11–15, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2022.
2066410.

Lines:2022:PPP

[1762] Nicholas A. Lines. The past, problems, and potential of readability analysis. Chance, 35(2):16–24, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2066411.

Orkin:2022:MML

[1763] Michael Orkin. The Mega Millions Lottery and hypothesis testing. Chance, 35(2):25–28, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2066412.

Gray:2022:SEP

[1764] Mary W. Gray. The struggle for equal pay, the lament of a female statistician. Chance, 35(2):29–31, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2066413.

Horton:2022:SFW

[1765] Nicholas J. Horton, Jie Chao, William Finzer, and Phebe Palmer. Spam four ways: Making sense of text data. Chance, 35(2):32–40, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2066414.

Gelman:2022:HSS

[1766] Andrew Gelman. How should scientific journals handle ‘Big If True’ submissions? Chance, 35(2):41–43, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2022.
2066415.

Robert:2022:MAM

[1767] Christian Robert. Measuring abundance: Methods for the estimation of population size and species richness. Chance, 35(2):44, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2066417.

Robert:2022:WCW

[1768] Christian Robert. What are the chances? (Why we believe in luck). Chance, 35(2):45, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2066418.

Robert:2022:LBR

[1769] Christian Robert. Learning Base R (2nd edition). Chance, 35(2):46–47, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2022.
                                        2066419.

Velleman:2022:MLW

[1770] Paul Velleman and Howard Wainer. In memoriam: Leland Wilkinson (1944–2021). Chance, 35(2):48–49, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2022.
                                        2066422.

Pham:2022:CAV

[1771] Nguyen-Khang Pham, Jean-Hugues Chauchat, and Jean Dumais. Correspondence analysis visualization. Chance, 35(2):50–52, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2066423.

Peterson-Plunkett:2022:ELc

[1772] Amanda Peterson-Plunkett. Editor’s letter. Chance, 35(3):3, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2123144.

Messner:2022:SGS

[1773] Michael Messner and Claire McKay Bowen. Spotlight on the government statistics section of the American Statistical Association. Chance, 35(3):4–8, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2022.
2123147.

Eltinge:2022:CHH

[1774] John Eltinge and Wendy Martinez. A conversation with Howard Hogan, former Chief Demographer at the United States Census Bureau. Chance, 35(3):9–16, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2123150.

LaLonde:2022:RGN

[1775] Donna LaLonde. R Govys: A new community that wants You. Chance, 35(3):17–18, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2123151.

Molfino:2022:TGF

[1776] Emily Molfino. Tips for getting a Federal statistics job. Chance, 35(3):19–22, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2022.
2123152.

Gray:2022:DCD

[1777] Simone Gray. An in-depth conversation with Dionne Price, 2023 President of the American Statistical Association. Chance, 35(3):23–26, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2123153.

Wainer:2022:TMC

[1778] Howard Wainer. Transforming medical cucumbers into sunbeams. Chance, 35(3):27–32, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2123156.

Sedgwick:2022:TTT

[1779] Philip Sedgwick. Trials and tribulations of teaching null hypothesis significance testing in the health sciences. Chance, 35(3):33–41, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2123159.

Fox:2022:ABD

[1780] Andrea Fox, Marica Manisera, Marco Sandri, and Paola Zuccolotto. Analyzing basketball data with BasketballAnalyzeR. Chance, 35(3):42–56, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2022.
                                        2123161.

Peterson-Plunkett:2022:ELd

[1781] Amanda Peterson-Plunkett. Editor’s letter. Chance, 35(4):3, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2145147.

Tijms:2022:MHL

[1782] Steven Tijms. Monty Hall and “the Leibniz Illusion”. Chance, 35(4):4–14, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2022.
2145124.

Orkin:2022:RE

[1783] Michael Orkin. Random entanglement.

Chance, 35(4):15–17, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2145126.

Behseta:2022:IHS

[1784] Sam Behseta. An interview with Hal Stern — from academics to administration. Chance, 35(4):18–26, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2022.
2145129.

Mitra:2022:TSD

[1785] Sinjini Mitra. Teaching statistics and data science to business students. Chance, 35(4):27–35, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2022.2145133.

Gray:2022:R

[1786] Mary Gray. Risk. Chance, 35(4):36–39, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2022.
                                        2145134.

Veaux:2022:HFA

[1787] Richard De Veaux, Anna Plantinga, and Elizabeth Upton. Are the handicaps fair? Age and participation effects in the Dipsea Race. Chance, 35(4):40–49, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2145138.

Robert:2022:BR

[1788] Christian Robert. Bayes rules! Chance, 35(4):50–51, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2145140.

Robert:2022:AL

[1789] Christian Robert. Amy’s luck. Chance, 35(4):51, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2022.2145142.

Liu:2022:QAD

[1790] Fang Liu. A Q&A about differential privacy. Chance, 35(4):52–56, 2022. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2022.
                                        2145146.

Martinez:2023:ELa

[1791] Wendy Martinez and Donna LaLonde. Editors’ letter. Chance, 36(1):3, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2179255.

Ensor:2023:IDE

[1792] Katherine Bennett Ensor and Donna LaLonde. The influencing discovery exploration and action forum: An initiative of the American Statistical Association. Chance, 36(1):4–8, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2179258.

Cressie:2023:ASS

[1793] Noel Cressie. Adapting statistical science for a fast-changing climate. Chance, 36(1):9–13, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2179263.

Ensor:2023:HSM

[1794] Katherine Bennett Ensor and Rich Sorkin. Helping society mitigate effects of climate change. Chance, 36 (1):14–16, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/
09332480.2023.2179268.

Sain:2023:CRA

[1795] Steve Sain. Climate risk analytics.

Chance, 36(1):17–22, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2179270.

Ensor:2023:ENC

[1796] Katherine Bennett Ensor and Jennifer Grancio. Engine No. 1 CEO reflects on the future of climate change. Chance, 36(1):23–24, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2023.2179271.

Li:2023:RIF

[1797] Bo Li and Douglas Simpson. Reflections on the IDEA forum-statistics, climate change, and sustainability. Chance, 36(1):25–30, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2023.2179273.

Risser:2023:ACC

[1798] Mark Risser and Steve Pierson. Advisory Committee on Climate Change Policy: A committee of the American Statistical Association. Chance, 36(1):31–34, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2023.2179274.

Tractenberg:2023:SRI

[1799] Rochelle Tractenberg. 10 simple rules for integrating ethics training into any quantitative course. Chance, 36(1):35–39, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2023.2179275.

Tackett:2023:ARD

[1800] Maria Tackett and Mine Çetinkaya-Rundel. Analyzing and recreating data visualizations of W. E. B. Du Bois. Chance, 36(1):40–47, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2179279.

Wainer:2023:WTW

[1801] Howard Wainer and Daniel Robinson.

Why testing? Why should it cost you? Chance, 36(1):48–52, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2179281.

Martinez:2023:ELb

[1802] Wendy Martinez and Donna LaLonde.

Editors’ letter. Chance, 36(2):3, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2203642.

Lund:2023:CMC

[1803] Robert Lund and Xueheng Shi.

Changepoint methods in climatology. Chance, 36(2):4–8, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2203643.

TumasianIii:2023:TEC

[1804] Robert A. Tumasian Iii. Toward enhancing clinical trial diversity. Chance, 36(2):9–13, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2203644.

Eltinge:2023:RRD

[1805] John L. Eltinge. Re-reading Deming’s 14 points in the era of data science: 14 questions for vibrant management of statistical information systems. Chance, 36(2):14–22, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2203645.

LaLonde:2023:RSE

[1806] Donna LaLonde. Remembering Samuel “Ed” Nevius. Chance, 36(2):23–27, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2203646.

Martinez:2023:PPA

[1807] Wendy Martinez and Donna LaLonde. Pixel by pixel: The art of CHANCE. Chance, 36(2):28–30, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2203648.

Robert:2023:CL

[1808] Christian P. Robert. Casanova’s lottery. Chance, 36(2):31–32, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2203649.

Robert:2023:BFF

[1809] Christian P. Robert. Bayes factors for forensic decision analyses with R. Chance, 36(2):32–33, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2203651.

Robert:2023:BPB

[1810] Christian P. Robert. Bayesian probability for babies. Chance, 36(2):33, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2203652.

Reynolds:2023:FP

[1811] Penny S. Reynolds and Chaitra Nagaraja. Fresh perspective. Chance, 36 (2):34–36, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/
09332480.2023.2203655.

Wasserstein:2023:OTA

[1812] Ron Wasserstein. One thing about . . ..

Chance, 36(2):37–38, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2203657.

Ellis:2023:IWR

[1813] Amanda Ellis. Incorporating workforce readiness skills to increase how students perceive value of biostatistics coursework. Chance, 36(2):39–44, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2203659.

Wainer:2023:TWA

[1814] Howard Wainer. Three welcome arrivals for 2023: 1. Florence Nightingale. Chance, 36(2):45–48, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2203662.

Martinez:2023:ELc

[1815] Wendy Martinez and Donna LaLonde. Editors’ letter. Chance, 36(3):3, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2264727.

Muoio:2023:ARC

[1816] Patricia Muoio. Assessing risk in cybersecurity: How sound data science can raise the bar. Chance, 36(3):4–8, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2264732.

Baumer:2023:MMD

[1817] Benjamin S. Baumer. More math for data scientists, not less. Chance, 36 (3):9–11, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/
09332480.2023.2264734.

Bhaduri:2023:BUG

[1818] Moinak Bhaduri. Bringing unrelated games closer: Excitements unified with changes picked through a careful parallax. Chance, 36(3):12–18, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2264728.

Banks:2023:FIS

[1819] David Banks. The future of industrial statistics. Chance, 36(3):19–24, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2264738.

Wu:2023:PMP

[1820] JunBo Wu and Nathaniel C. Comfort. Precision medicine for the population — the hope and hype of public health genomics. Chance, 36(3):25–27, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2264730.

Durso:2023:PSW

[1821] John Durso and Howard Wainer.

A physicist and a statistician walk into a bar. Chance, 36(3):28–30, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2264731.

Bailer:2023:WWC

[1822] A. John Bailer and Rosemary Pennington. Winner, winner, chicken dinner … well, maybe a podcast appearance and some swag: Using contests to engage learner communities.

Chance, 36(3):31–34, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2023.2264733.

Govindarajulu:2023:CRV

[1823] Usha Govindarajulu. COVID-19 research and p-values. Chance, 36 (3):35–38, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        09332480.2023.2264736.

Nagaraja:2023:RS

[1824] Chaitra H. Nagaraja. Refracted selves. Chance, 36(3):39–43, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2023.2264737.

Andrews:2023:TTJ

[1825] R. J. Andrews. Timeless temples — a journey through Emma Willard’s chronographic masterpieces. Chance, 36(3):44–48, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2023.2264735.

Martinez:2023:ELd

[1826] Wendy Martinez and Donna LaLonde. Editors’ letter. Chance, 36(4):3–4, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2023.
                                        2290961.

Hagan:2023:FSH

[1827] Nancy Hagan and Margaret Henderson. Four strategies for human trafficking research. Chance, 36(4):5–11, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290942.

Banks:2023:MSE

[1828] David Banks and Emmanuel Mokel.

Multiple systems estimation and human trafficking. Chance, 36(4):12–14, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290944.

Kunke:2023:RDS

[1829] Jessica P. Kunke, Adam Visokay, and Tyler H. McCormick. Respondentdriven sampling: an overview in the context of human trafficking. Chance, 36(4):15–21, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2290949.

Seymour:2023:HCW

[1830] Rowland G. Seymour and Bernard W.

Silverman. How can we estimate modern slavery globally? Chance, 36 (4):22–29, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/
09332480.2023.2290950.

Waller:2023:CRW

[1831] Lance A. Waller. A conversation with Ray Waller. Chance, 36(4):30–36, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290952.

Emeritus:2023:CSC

[1832] A. John Bailer Professor Emeritus and Rosemary Pennington. I can see clearly now, the data’s shown: Data visualization insights from podcast contest. Chance, 36(4):37–41, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2023.2290953.

Shipp:2023:MED

[1833] Stephanie Shipp, Donna LaLonde, and Wendy Martinez. Making ethical decisions is hard! Chance, 36(4):42–50, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290955.

Lorenzo-Arribas:2023:SSS

[1834] Altea Lorenzo-Arribas, Penny S. Reynolds, and Chaitra H. Nagaraja. Suffrage, statistics, and spurious correlations. Chance, 36(4):51–54, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290956.

Ziliak:2023:OTA

[1835] Stephen T. Ziliak and Ron Wasserstein. One thing about … the rhetoric of statistics. Chance, 36(4):55–56, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290957.

Andrews:2023:GM

[1836] Rj Andrews, Georges Hattab, Étienne-Jules Marey, and Howard Wainer. The graphic method. Chance, 36(4):57–61, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290959.

Robert:2023:PN

[1837] Christian Robert. Probabilistic numerics. Chance, 36(4):62–64, 2023. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2023.
2290960.

Martinez:2024:ELa

[1838] Wendy Martinez and Donna LaLonde.

Editors’ letter. Chance, 37(1):3, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2024.
2322914.

Price:2024:IFC

[1839] Dionne Price and Donna LaLonde. The 2023 IDEA Forum: Celebrating the year of open science. Chance, 37(1):4– 5, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2024.
2324658.

Cannon:2024:PVO

[1840] Matt Cannon. A publisher’s view of open research. Chance, 37(1):6–9, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2024.
                                        2324659.

Molfino:2024:OCP

[1841] Emily Molfino, Michael Walsh, and Robert Sienkiewicz. Open census: Promoting scientific integrity at the US Census Bureau. Chance, 37(1):10–13, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2024.
                                        2324660.

Ensor:2024:IPH

[1842] Kathy Ensor and Donna LaLonde. Interview with Peter Hotez, influencing discovery exploration and action forum plenary presenter. Chance, 37 (1):14–16, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

                                        tandfonline.com/doi/full/10.1080/
                                        09332480.2024.2324661.

LaLonde:2024:DSS

[1843] Donna LaLonde. Data Science in Science: an open-access journal. Chance, 37(1):17–21, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2024.2324663.

Lancaster:2024:CHL

[1844] Vicki Lancaster, Cesar Montalvo, Joseph Salvo, and Stephanie Shipp.

Creating a household living budget using publicly available data: a 21st century census curated data enterprise use case. Chance, 37(1):22–32, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2024.
2324664.

Ziliak:2024:LYS

[1845] Stephen T. Ziliak. Letter to a young statistician: On ‘Student’ and the Lanarkshire Milk Experiment. Chance, 37(1):33–36, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2024.2324665.

Martinez:2024:RNS

[1846] Wendy L. Martinez. Remembering Nozer Singpurwalla. Chance, 37(1): 37–40, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/
09332480.2024.2324666.

Author:2024:IPR

[1847] Penny S. Reynolds and Chaitra H. Nagaraja. An interview with Professor Rosemary Bailey. Chance, 37(1):41–54, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

URL http://www.tandfonline.com/
doi/full/10.1080/09332480.2024.
2324667.

Ethics:2024:EEL

[1848] Members of the Asa Committee on Professional Ethics and Ron Wasserstein. Exploring the ethical landscape of AI. Chance, 37(1):55–56, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2024.
                                        2324668.

Martinez:2024:ELb

[1849] Wendy Martinez and Donna LaLonde. Editors’ letter. Chance, 37(2):3, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2024.
                                        2348950.

LaLonde:2024:PPA

[1850] Donna LaLonde. Pixel by pixel: Artistic achievements. Chance, 37(2):4–5, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2024.
                                        2348951.

Manzione:2024:IAT

[1851] Katherine L. Manzione, Bailey K. Fosdick, Ryan Elmore, and Connor Gibbs. The impact of altitude training on NCAA Division I female swimmers’ performance. Chance, 37(2):6–12, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2024.
                                        2348952.

Andersson:2024:CIB

[1852] Per Gösta Andersson. Confidence interval for a binomial p: Recommend the Wilson interval to students. Chance, 37(2):13–15, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2024.2348953.

Wasserstein:2024:RKW

[1853] Ron Wasserstein. Remembering Katherine Wallman. Chance, 37(2): 16–20, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/
09332480.2024.2348955.

Baker:2024:MDC

[1854] Stuart G. Baker and Karen S. Lindeman. Multiple discoveries in causal inference: LATE for the party. Chance, 37(2):21–25, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2024.2348956.

Nagaraja:2024:RCR

[1855] Chaitra H. Nagaraja, Penny S.

Reynolds, and Chaitra H. Nagaraja. Remembering C. R. Rao (1920–2023). Chance, 37(2):26–28, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2024.2348970.

Loyer:2024:OTA

[1856] Milton W. Loyer and Ron Wasserstein. One thing about … strange expectations: a dickens of a problem. Chance, 37(2):29–33, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

www.tandfonline.com/doi/full/10.
1080/09332480.2024.2348971.

Thiesmeier:2024:TSH

[1857] Robert Thiesmeier, Nicola Orsini, Edward Gracely, and Bob Oster. Teaching statistics in health sciences: The potential of simulations in public health. Chance, 37(2):34–39, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2024.2348972.

Robert:2024:SCP

[1858] Christian Robert. A second course in probability. Chance, 37(2):40–41, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

                                        URL http://www.tandfonline.com/
                                        doi/full/10.1080/09332480.2024.
                                        2348976.

Robert:2024:E

[1859] Christian Robert. Exhalation. Chance, 37(2):41, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2024.2348977.

Hardin:2024:CCU

[1860] Jo Hardin, Mine Çetinkaya-Rundel, and Maria Tackett. CURV: Connecting, uplifting, and recognizing voices. Chance, 37(2):42–48, 2024. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic). URL http://

                                        www.tandfonline.com/doi/full/10.
                                        1080/09332480.2024.2348978.

deVries:2012:RD

[1861] Andrie de Vries and Joris Meys. R for dummies. For dummies. Wiley, New York, NY, USA, 2012. ISBN 1-119-96284-6 (paperback), 1-119-96312-5 (e-book), 1-119-96313-3 (e-book), 1-119-96314-1 (e-book). xviii + 387 pp. LCCN QA276.45.R3 V75 2012.

Evans:2015:MSE

[1862] Michael Evans. Measuring Statistical Evidence Using Relative Belief, volume 144 of Monographs on statistics and applied probability. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2015. ISBN 1-4822-4279-6. xvii + 232 pp. LCCN QA277.5 .E93 2015.

Karian:2011:HFS

[1863] Zaven A. Karian and Edward J.

Dudewicz. Handbook of fitting statistical distributions with R. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-58488-711-7 (hardcover), 1-58488-712-5 (e-book: PDF). xlv + 1672 pp. LCCN QA273.6 .K374 2011.

URL http://marc.crcnetbase.com/
isbn/9781584887126. Includes CD-

ROM.

Matloff:2011:ARP

[1864] Norman S. Matloff. The art of R programming: tour of statistical software design. No Starch Press, San Francisco, CA, USA, 2011. ISBN 1-59327-384-3. xxiii + 373 pp. LCCN QA276.4

.M2925 2011. URL http://www.loc.
gov/catdir/enhancements/fy1205/
2011025598-b.html;   http://www.
loc.gov/catdir/enhancements/fy1205/
2011025598-d.html;   http://www.
loc.gov/catdir/enhancements/fy1205/
2011025598-t.html.

McGrayne:2011:TWH

[1865] Sharon Bertsch McGrayne. The theory that would not die: how Bayes’ rule cracked the Enigma code, hunted down Russian submarines, and emerged triumphant from two centuries of controversy. Yale University Press, New Haven, CT, USA, 2011. ISBN 0-300-16969-8. xiii + 320 pp. LCCN QA279.5 .M415 2011.

Reinhart:2015:SDW

[1866] Alex Reinhart. Statistics done wrong: the woefully complete guide. No Starch Press, San Francisco, CA, USA, 2015. ISBN 1-59327-620-6 (paperback). xviii + 152 pp. LCCN QA276 .R396 2015.

Ziliak:2008:CSS

[1867] Stephen Thomas Ziliak and Deirdre N. McCloskey. The Cult of Statistical Significance: How the Standard Error Costs Us Jobs, Justice, and Lives. Economics, cognition, and society. University of Michigan Press, Ann Arbor, MI, USA, 2008. ISBN 0-472-07007-X (cloth), 0-472-05007-9 (paperback). xxiii + 321 pp. LCCN HB137 .Z55 2008. URL http://

                                        www.loc.gov/catdir/enhancements/
                                        fy0809/2007035401-d.html.

Gelman:2011:HMC

[1868] Andrew Gelman, Xiao-Li Meng, Steve Brooks, and Galin L. Jones, editors. Handbook of Markov Chain Monte Carlo. Chapman and Hall/CRC handbooks of modern statistical methods. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4200-7941-7 (hardcover), 1-4200-7942-5 (e-book). xxv + 592 pp. LCCN QA274.7 .H36 2011.